A Complete Bibliography of Publications in SIAM Journal on Mathematical Analysis for 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

16 October 2025 Version 1.12

Title word cross-reference

−Δu = u^{−γ} [MMS25]. 1/3 < K < 1/2 [Oli21]. 2 [FMN21, LXZ23]. 2 + 1 [DM22]. 2 ≥ p\frac{5}{3} [CMMS24]. 3 [HL23a, HJLZ21, LX23, LZ24a, SH22]. 8π [GH21a]. ^{∘} [AM22]. α [BIFL20]. α → 0 [BIFL20]. b [LP22]. BD [CFV20]. BV [GJ21]. BV^{s} [BCGJ22]. C^{2} [GJ21]. d [Chi22]. d ≥ 1 [KNP21]. Ḃ_{∞,∞}^{1} [CN23]. ϵ ∈ [1, 2] [DK23]. g [Ran23]. Γ [BK23b, FMP22, MPC22, MW23]. γ = 3 [Gol23]. γ^{2} > 1/4 [ZZ23a]. H(p) [MPZ22]. H^{2} [HOV25]. H_{n}^{k} [Den20]. k [PU24]. L^{1} [CN24, GHW21]. L_{k}^{1} ∩ L_{k}^{p} [DSU24]. L^{2} [BGK21, BHG25, Dod21, JMQ23, TX24]. L^{∞} [BHI24, FV25, PW21]. L^{p} [CY21a, Har20, LXZ25]. L^{r} [Xia20]. L_{p} [DY24]. L_{q}(L_{p}) [CKP24]. L_{v}^{p}L_{T}^{∞}L_{x}^{∞} [Li22b]. m [LZ20, MS21]. R [BCFP23, DM20a]. R^{1+3} [Tes20]. R^{2} [HY22a, NS20b, CiDRdTY23]. R_{+}^{2} [WW24]. R^{3} [BMM22, CDLZ21, Kat20, LLWZ24, MB20, DSU24]. R_{+}^{3} [Wan23]. R^{d} [BK23a, CLLZ23, FMS21]. R^{d} × T^{n} [HLS24]. R^{N} [WHQ25, JM23, SWX25, MS21]. R × T^{3} [YYZ21]. T [HOV25]. Z_{2} [CCJT23]. A [KV21]. N [PZ23, XMN23]. p [CP24, EIN22, EFP^{+}24, Lia22, MN24a, NP20, dDUFJK25, Wic23]. P 1 [Wan21]. p = Kρ [Oli21]. p → 1^{+} [CP24]. Q [LW20a]. R^{n} [Fos21]. SU(3) [WZ20b]. P_{I}^{2} [DL24a]. ε [DW20]. [GDS24]. |k| [KSS23].

-Approximation [Wan21]. -based [FV25]. -bounds [PW21]. -Convergence [FMP22, GDS24, GHW21, MPC22, MW23]. -Convex [LZ20]. -Convexity [Ran23].

-Critical [TX24]. -Critical-Mass [GH21a]. -D [LX23, HL23a, HJLZ21, LZ24a, XMN23]. -Dimensional [LXZ23, SH22, PZ23]. -Elastic [NP20]. -Equivariant [CCJT23]. -Euler [BIFL20]. -Family [LP22]. -Gradient [CN24]. -Initial [LXZ25]. -Laplace [CP24, EFP^{+}24, Wic23]. -Laplace-Type [EIN22]. -Laplacian [Lia22, dDUFJK25]. -Limit [BK23b]. -Plane [Chi22, PU24]. -Quasiconvexity [KV21]. -Regularity [DW20]. -Solutions [HOV25]. -spaces [CY21a]. -System [MN24a]. -Tensor [MS21, LW20a]. -Torus [FMN21]. -Vector [BGK21].

1 [DN22, WZ20b]. 1-d [YY22]. 1D [CMVP23, Fei25, Gib24, MN24a, RS25].

2D [BHI24, Bob20, Cai23, CW22, CMX24, HNX25, IPTT22, JWX22, KL24b, KM24, LWZ21a, Ma22, Zha23].

3 [Cia20]. 3D [BHMCR24, BDF24, BHI24, BW25, CWZ24, DGWZ23, GHM23, LPS22, LWZ23, MRR20, MSW24b, WJW22, WYZ23, WW23, ZZZ22a]. 3D-Grain [MSW24b]. 3dimensional [WO22].

above [Dod21]. Absolute [MPZ22]. Absorption [CGU23, RZ21, SZ22]. Acceleration [AMMT22]. Accounting [DDG24]. Accretions [Liu23]. Accumulation [CRS21]. Accurate [GLS23]. Acid [CQ25]. Acoustic [EFHS21, EI20, FHS24, MS23b]. Acoustics [BSH24, KN22, UZ23]. Acting [Ger24]. Action [DGMP20]. Activation [LSW23]. Activations [JLL^{+}23]. Active [Abd24, AO23, BBES22, HW22a]. Adaptive [BB23]. Additive [ACS23]. Adjacent [AKKY23, gLZ25]. Adjoint [AT20, Liu24]. Adler [AHLW21, LW20b]. Admissible [AW21, AKK^{+}20, FG22]. Advection [FMTY21, IV22, KW23, LLPZ21, Liu24]. Advective [WXZ24]. Adverse [WZ20a]. Age [FS22, PR20]. Age- [FS22]. Age-Structured [PR20]. Agents [BHS25]. Aggregation [Shu21]. Air [DKLT25]. Airy [AP24, CDM24a, HZ22b, Nem21]. Alarm [JWW23]. Alarm-Taxis [JWW23]. Alfvén [RWZ21]. Algorithm [EN22]. Alignment [CFP25, Lea23]. All-time [Ou21]. Allee [KW23]. Allen [AF23, AM22, BV20b, FLS20, Gen25, HL23b, HMNW20, MMN23]. Allied [ND20]. Alloys [Sim21]. Almost [BN24, GG24, LX23, Mar22, MP20b, MP21b, PS23a, RS20a]. Almost-Sure [BN24]. Along [BBD^{+}23]. Alternating [CNT23]. Amalgam [BT21]. Ambient [KM25, RS24]. Ambrosio [BCR21]. Among [Pru24]. Ampere [CHX24b, Ham25, Lew25, TT22]. Amplitude [Li22b, LG21, Sin21]. Analogue [GL24a, MN24b]. Analyses [CJ24]. Analysis [ALV23, ACL22, BK20a, BBBH25, BMR20, CJL21, CWSGS20, CQ25, CCM21, CC23, Chi22, DLL20, DN20, DL21, DJ20, EOE20, FHKM22, FSW24, HJJ20, HM25, HXH22, Kat20, Kat23, KLSW24, KA24, LZ20, LXZ23, LP23, LZ23b, MPV23, MS20c, MS23b, NS21b, TW22a, ZHY24, vM22]. Analytic [FMMS22, Fen25, HMS24, MS20a, Ngu20, Wan20]. Analytical [ED22, Pan24]. Analyticity [ALW23, Dec24, DZZ23, KN23]. Anchoring [BG24a]. Anderson [HY21, KNP21]. Angle [AB20, AM22, MPMS23, MW20, NT25, RS20a]. Anisotropic [CCM21, CC23, CKP24, HL23a, HW24, Kim21, KSS24, LWZ21a, LZ22, LLX21, RZZ21, UW24, Yan21a]. Annihilation [vMP24]. Anomalous [CDX21, CL21c, DP25]. Ansatz [DLX21]. Anti [Hog21]. Anti-Plane [Hog21]. Antisymmetric [DDG24]. Any [KS21, KS24a, WdlL20]. Appearing [Tet22]. Appendix [GG20]. Application [Aik23, AB21, ACC^{+}21, ACC^{+}23, BHI24, BDF25, BCD24, GLSW23, Hei23, LYZZ24, MRB22, VY23]. Applications [ARS21, ADJS24, BLPS22, BCFP23, CDLZ21, CMP25, CNW20, CWSGS20, CJ24, DPS24, DH24, DMS22c, FLRX24, JL21, KLM24, KN22, KLLT25, KPS20, KM24, KV21, KV20, Pru24, QRW24, Ran23, WWZ23a, WHY25, Yan25, dHKLN24]. Applied [Kim21]. Approach [AK23a, AMMT24, AK23b, BHMCR24, BDF24, BF23, CL20, CS23, DSU24, DG23, GK20, HH23, HL23b, JKZ21, Jin20, RR25, RS23b, SW22c, SUT22]. Approaches [LXZ23]. Approximate [Zha21]. Approximation [BB23, BP20, BJ23, BJL24, BNP22, Cia20, CLR24, DIS25, DST23, FHM20, FOS20, JJL21, JLZ25, JLL^{+}23, KN22, LLS25, LSYZ21, Ros20, Wan21, WW22, Wri24]. Approximations [BDMA22, GMŚ21b, KvM21, LWX24, QB24, XC22]. Arbitrary [DL22b, Lew25]. Arc [CNT23]. Area [DLW22]. Arguments [Pru24]. Arising [BK23b, CCJT23, DR20a, Jui20, LSW23, LNRV21, UZ23, Wan21]. Arnold [KT22a]. Arrays [ADH23]. Aspects [Sim21]. Assimilation [BP21]. Associated [BBBH25, BHTY25, CH21b, Eit22]. Assumption [LY21]. Astrophysics [Liu23]. Asymmetric [CF25, MS24a]. Asymptotic [AV24b, AS21, BCK25, BM24, CNT23, CaDZ24, CM23c, CL21b, CHXZ21, CJ24, CMMS24, DN20, DR20b, Dun24, EK20, EY23, FR23, Fen25, GLW23, GR23, GLM20, GBS20, GM20b, GOW20, GLP23, JZZ22, KMY20, KM24, LPW22, MN24a, ND20, Pan23, PR20, YZ21, ZZ21a]. Asymptotically [CWZ24, CMVP23, IKL20, WKM23]. Asymptotics [ANWS22, CL24, DL24a, GRHSZ20, HZ22b, Li20, LPW21, LYZZ24, LX25, LLS20, LLPZ21, OP20, Ste25b, YZ24b]. Atmospheric [DKLT25]. Atomistic [ALP21a]. Attractive [BCT24, NR24]. Attractive-Repulsive [BCT24]. Attractors [CWZ24, KM24, WKM23]. Aurora [FMN21]. Automata [Pro20]. Autonomous [CWZ24, CDG23, WKM23]. Autonomy [KM24]. Autoparametric [BN24]. Autophagy [LX21]. Average [DLW22]. Averaging [BDOZ21, CL23, Gao21, HNS20, KLSZ22, Zhu21]. Avoidance [CFP25]. Axially [DW21]. Axisymmetric [BHI24, Boc20, BW25, Hu22, LX23, WXX21].

B [LLW21, Pan24, WW24]. Back [Gar24, WZ20a]. Background [BLW20, FP20, LWZ21a, Zha23]. Backward [ACC^{+}21, ACC^{+}23, CFS22, LZ23b, WJW22]. Bacterial [EY23]. Balance [CMP25, CL21c, DR20b, FR23]. Balanced [CR21]. Ball [BCT24, FS20]. Banach [Kor22]. Band [CW25]. Bandlimited [GIT20]. Barenblatt [GHW21]. Barotropic [GMN22, HKK23, HL22a]. Barrier [vM22]. Barron [CLLZ23]. Barycenters [Buz25, CEK21, FMS21]. Based [ABR22, ATZ24, CDGH20, CHQ22, FLS20, HL20a, IMS22, LWZZ22, ST20a, FV25]. Basic [ZZ21a]. Basis [EB23]. Bathtub [MRB22]. Bayesian [MSW24a]. BBM [CWZ24]. Beale [LZ24b]. Beam [BP20, CDG23, Wan24]. Beam-Hardening [Wan24]. Behavior [AO24, BK24, CL21b, DS22, DG20, EY23, Fen25, FS22, GLW23, GLM20, GM20b, GOW20, GLP23, HKK23, JLN23, KSS23, KSS24, Lea23, LLW20, LM21, MN24a, MSM24, MB20, MSW24b, MPS21, Pan21, Pan23, PW21, PR20, RWZ21, WZZ20, XC22, ZZ21a]. Bellman [GMŚ21b, JMQ23]. Belousov [HKlT24a]. Below [KL24b]. Beltrami [GN23]. Bence [BV20b]. Benjamin [Gas23, HZ22a, Wri24]. Benney [MS20b, NO23]. Bernoulli [CF20]. Besov [DK23]. Bessel [Dun24, FLPS24].

Between [AKKY23, JR21, LLhW23, TW23, gLZ25]. Beyond [BJL24]. BGK [BY20, BKPY23, BY24, HRY22]. Bidomain [HHS23, MMN23]. Bifluid [KKNN22]. Bifurcation [BDF25, CM22, Gar24, Var20]. Bifurcations [ADVW20]. Biharmonic [Duc23, HLS24, LW24, Yan21a]. Bilaplacian [BK22]. Bilayer [BZ25]. Bilinear [CMU23, MRB22]. Bimolecular [LLN20]. Binary [RS24]. Biomembranes [GHPZ24]. Biot [Gah25, Ger24, HS23]. Bistable [HKlT24a, KMY20, KM23]. Bitemperature [ADBP21]. Black [PP24]. Blaschke [FL21]. Blood [ZZZ24]. Blow [AB20, Bar23, BCJ20, CW24, GL24a, He23, HJ11, HR23, KLS23, KS21, KS24a, LZ20, TX24, Tré24]. Blow-Up [Bar23, BCJ20, GL24a, He23, HJ11, HR23, KLS23, KS21, KS24a, LZ20, TX24, Tré24, AB20, CW24]. Blowup [DK23, IMO20, LS21a, LXZ23, LY23, Miz20]. Blowups [Liu23]. Bodies [NS20b]. Body [BNSS23, BBG^{+}24, Eit22, JV24a, KO24, NR24]. Bogolyubov [CL23]. Boltzmann [AČ24, AMP22, AL23, AL25, BBBH25, CHJ24, CDL24, CPW24, Che22, Den20, Den21, DSU24, Eva21, Ger23, GZW21, HLlZ23, HW22b, HW22c, HY21, IMS20, Jia20b, Jun25, KL24b, LUY21, Li22b, LW23a, LLW20, TSMO24, WWZ21, WY24, WO22, WZGX25]. Boltzmann/Landau [CDL24]. Bose [BDS24, CJL21, CH21b, GLP23, LLWZ24, LLWZ25, NR24, PWZZ24]. Both [Fue20]. Bound [BDS24, Cai23, Chu23, FSW22, HZ25, HKSW24, KL24a]. Boundaries [BZGZ24, CLWX20, Cla23, CDF^{+}22, DNS25, ELM21]. Boundary [AC21, AM22, ACL22, AB21, ANPS21, ABFV22, BK20a, BZGZ24, BG24b, BC24b, BGMW20, BDLL25, BPSS22, BK22, BIFL20, CYY23, CNT23, CHaW24, CK23b, CHS22, CCS22, Che22, CMZZ22, CLMZ23, CHZ24, CDL22, CFS22, CF20, DFV25, DS21a, DMS22b, DLX21, DLZ25, DH24, DW20, DGOY24, DN22, EFHS21, EFP^{+}24, FK25, FHM^{+}24a, GDS24, GR23, GK20, GZ20a, GHM22, Gin21, GV21, GBS20, GZ22, HdRS20, HM25, HXH22, HMS24, HSV20, HSV21, HSVW25, HC21, HW22c, IKL20, JK22, KL21, KMY20, Kil22, KUW21, KKNN22, Kre21, KL22, KS24b, LZ20, LPW21, LWZ21b, LY21, LXZ23, Liu23, LY23, LNRV21, LZ21, LZ24b, MW23, MSW24b, Ngu20, NWG24, NT25, OP20, Ou21, Ric23, SSZ20, SH22, SD24, SD25, SZ22, SUT22, ST21, TW25, TWZ20, Taw24, UZ23, VY23, WZ20a, Wan20, WWZ23b, XX23a, XMN23]. Boundary [Yan21a, YZ25, ZRW22, ZiDRdW24, vM22]. Boundary-Field [HSVW25]. Boundary-Layer [vM22]. Boundaryless [dN23]. Bounded [Abd24, ADJS24, Bad25, CJ21, CDN21, GK22, Ger21, HY21, LY23, Mar22, MW23, Mou20, OY22, PVY25]. Boundedness [CMT21]. Bounds [CWSGS20, CFS23, Don22, IMS20, IV22, KMW24, LNOY24, Sne20, PW21]. Bourgain [DDG24, HSSY24]. Boussinesq [BHI24, CMX22, CMX24, CL24, CGS23, DST20, FL24, LOWX20, LLW24, SX20a, ZZ23a, Zil23]. Brakke [NT25, Tak21]. Branching [Sim21]. Break [JLL^{+}23]. Breakage [AGL24, GL21]. Breakdown [Gin21, LZ24b]. Breaking [SW22b, SSWZ24]. Brezis [DDG24]. Broadening [Hog21]. Brödinger [Bar20]. Brownian [BBES22]. Bubbles [ABLW24, AFH^{+}20]. Bubbling [WZ20b]. Bulk [BBC20]. Buoyancy [Hu24b]. Burgers [CMVP23, DGWZ23, KV20, Yan21b]. BV [KMW20]. BV-Continuous [KMW20].

Cahn [AM22, AF23, BV20b, CEGP23, CFS22, CGSS24, CMM20, FLS20, GK20, Gen25, GLRS22, HL23b, HMNW20, MMN23, RW23, SW20a]. Calcium [EOE20].

Calculus [GLSW23]. Calderón [BCR25, BS24a, Cov22, GH24b, GV24, Har23, KLW22b, LLR20, LS20, Rül21, UW24]. Calogero [Bad24, Wri24]. Camassa [LP22, MP21a]. Camouflaging [PR25]. Canards [ADVW20]. Cancer [CQ25, DL22a]. Canonical [BDS24]. Capacity [CL25]. Capillarity [BH22b, CL21b, LZ23a]. Capillary [DS21b, IPTT22, LZ23a, MS24a, Set24, Ste25a]. Capillary-Gravity [MS24a]. Capture [FJ22]. Caputo [TR20]. Carleman [IY20, KA24]. Carpet [ALV23]. Case [Boc20, CFG25, Gar24, GMS21c, GLM20, JM23, JR21, Jui20, KT22a, LXZ23, Mou20, RW23, VX25]. Cases [PWZZ24]. Casimir [AK23b]. Category [CT23]. Cattaneo [BSH24]. Cauchy [BDMA22, CH21b, GM20a, HL22a, LLW21, MN24a, PZ23]. Cavitation [MS23b]. Cavity [Kwe23]. Cell [CMM20, LNRV21]. Cells [CCS23]. Cellular [IZ23]. Center [CMMS24, NK21]. Centers [WHQ25]. Central [LSW23]. Cercignani [Che22]. Certain [DK23, DS21a, Ste20]. Chain [ZQ24]. Chains [AKV20, HJM25, HS24, Zil23]. Changing [BC24b]. Channel [BBG^{+}24, CW24]. Channels [DW21]. Chaos [HRZ24, LNOY24, LLN20, Xie25]. Character [KZ21b]. Characterization [BGK21, CEK21, CCE24]. Charged [DMS22a]. Charges [HY22a]. Chemical [LLN20]. Chemotactic [Hu24b]. Chemotaxis [CHaW24, CBHS23, EY23, He23, Hei23, HKLT24b, KS21, KS24a, LS21a, LT24, LPW21, NP23, DL22b]. Chemotaxis-Fluid [He23]. Chemotaxis-Navier [Hei23, DL22b]. Choquard [XZ24, YCRS22]. Christian [GG20]. Circle [GNNCR20]. Circular [AKKY23, Dai21a]. Circulation [Ma22]. Class [AO24, ALP21a, AG24, BT23, BDGdN24, BM24, BCGJ22, BY23, CT23, Dai21b, FMW23, GP24, GL24b, GK25, JLL^{+}23, KL21, KL24b, Li23, LW23b, RRTT24, Saf20, SW20a, Wan22]. Classical [ALM23, CHLW25, DLY21, HL22a, LY23, Mou20, Ou21, Per20, WZZ24]. Classification [BMR20, DLQ21, DQ23, MMS25, Pen23]. Cloaking [BLPS22, Ngu22]. Close [AM22, MPMS23]. Closed [BDF25]. Closely [LX25]. Closure [DS23]. Clouds [CTL22, Liu23]. CLT [CEK21, KNP21]. Coagulation [FLNV24, FHP24]. Coarse [HNS20, HS24]. Coarse-Graining [HS24]. Coated [HSVW25]. Codimension [Lew25]. Coefficient [BH22b, KM20, KSS23, LLS20, LSY24]. Coefficients [AIN25, BS24a, BR23, CL23, CDX22, Cia20, Don25, FHS24, HL23a, KRS21, Mou20, PYZ24, Pin21, RZ21, ST24, WZ22, Xia20, XX23b, Zhu21]. Coercive [RS20b]. Cohesive [NS21a]. Coils [GRR23]. Collaboration [GG20]. Collapse [HY22b, JLS25]. Collective [BDMA22, NP23]. Collision [AGL24, CFP25, DW21, GL21, Jia20b, LUY21]. Collision-Induced [AGL24, GL21]. Collisions [CPW24]. Colloidal [DR20a, GRZ24]. Colored [CWZ24, HKY23]. Combined [BDF24, DM20b, HSV21, PS22]. Combined-Type [HSV21]. Combustion [HK23]. Coming [Rös21]. Communities [CL21a]. Compact [AW21, Com22, CDM24b, HKY23]. Compactness [ARS21]. Comparison [Ham25]. Comparisons [GV24]. Compatibility [KL22]. Competition [DNS25]. Complementarity [KLSW24]. Complementary [Ngu22]. Complementing [FN23]. Completely [BG24c]. Completeness [FN23]. Complex [BK20a, BG24a, BD22, CNW20, CEW21, CCJT23, CCM21, DCWF^{+}22, Duc23, FF25, Gao21, GH21b, NZ24, QWYY25, SS22b, Ste25b, Xia20]. Complex-Scaled [DCWF^{+}22]. Component [BKPY23, LZ24a, PWZZ24, YL22]. Composite [HKK23, HWWW24, YY22]. Composites [Bel22]. Compressible [AW21, BFHZ22, CCW23, CLZ22, CL22a, CLWX20, CL21b, CCZ21, CHS22, CHWY25, CDS20, DKLT25, FK25, GHW21, Ger21, HR23, HL22a, Hua24, JM23, JJL21, KKNN22, Kwe23, LS21b, LS23c, LX24, LLW21, LY23, MS20d, Ou21, Pan24, PZ23, RvW20, TW22b, TW23, WZ22, Wan23, WZ25, WYZ23, WW23, YZ21, YY22, ZZZ22a, Zen23, ZLZ20, ZZZ22b, Zho22]. Computation [DJLR24]. Concatenation [Bob23]. Concave [AL25, Jui20]. Concentrating [HZ25]. Concentration [AKKY23, BF23, HH24, Li20, gLZ25]. Concentric [CKW20]. Concentric-Ring [CKW20]. Concerning [VX21]. Condensate [CJL21, CH21b]. Condensates [GLP23, LLWZ24, LLWZ25, PWZZ24]. Condition [AM22, CFS22, DGOY24, EYZ22, GZ22, JJZ24, JK22, JW23, KNP21, Kre21, LWZ21b, NT25, Rös21, TWZ20, Taw24]. Conditional [FK25]. Conditions [AB21, BK20a, BFLP21, BC24b, BK22, BIFL20, CNT23, CHaW24, CY21a, EFHS21, FN23, GK20, GHM22, HMS24, HSV21, HW22c, KS24b, LY22, MW23, Ngu20, OP20, Ric23, SD24, SD25, ZRW22, ZiDRdW24]. Conductance [CaDZ24]. Conducting [BNSS23, BMPS21, DLX21, JW23, LZZ21, Wan23]. Conductive [EB23]. Conductivity [ABFV22, BFV21, CL22a, CNW20, CRTZ24, FFV20, GV24, Li20, PZ23, ZZ23b]. Conductors [HSVW25]. Conduit [JP20]. Cone [HLZ24, Yos21]. Cones [CKZ21]. Configurations [ACC22, BBG^{+}24]. Confined [CKP^{+}22]. Confinement [HJ24, Laf20]. Conformally [DQ23]. Congestion [RS23a]. Conic [Buz25].

Conical [CKZ21, DFV25, LS23a].

Conjectures [CL25, Nem21]. Connected [GHM23]. Connection [LLhW23, Tet22]. Connections [HNS20]. Conormal [CDL22]. Consensus [BST22, CHQ22]. Consensus-Based [CHQ22]. Conservation [CKP20, DJLR24, DMS22c, FP20, GJ21, GJP24, GK21b, HL22b, JR20, KV20, LZ23b, Pan20, RS23a, Yua23]. Conservative [DLW24, GLW22, GLW23]. Conserved [HMNW20]. Consistency [WT24]. Constant [FLNV24, GIT20, HL23b, KL22, MPMS23, Mon20, PSW25]. Constrained [Buz25, GJP24, KV21, MW20]. Constraint [ABHN21, CFS21, KTT20]. Constraints [AMMT22, AdSST24, BHM24]. Construction [HL22b, Kil22]. Contact [AB20, AM22, MP20b, MW20, YZ21, YZ25]. Container [Web20]. Content [Li22a]. Continuation [BCF24, DFV25, KMW20]. Continued [HY22b]. Continuity [BWF24, Lia22]. Continuous [BC24a, BP21, CMP25, CCZ21, NS21b, SW22a, KMW20]. Continuum [BBC20, CL21d, CC22, KvM21]. Contour [AHK24]. Contrast [Bel22, CCi23]. Control [AMMT22, ACS23, BHTY25, CMU23, CC25, CGSS24, FMTY21, JJZ24, KLSW24, MRB22, Web20, dW23]. Controlled [RS20a]. Controls [LT22]. Convection [BEM21, CGG20, ZRW22]. Convective [Gar24, RW23]. Convergence [AM22, ADH23, BLZ24, BMZ23, Bla23, BST22, CHJ24, CHaW24, CC22, DJLR24, DL21, Dod21, EN22, FHM20, FFV21, FLS20, FMP22, FW21, GMN22, GDS24, GHW21, Ger23, HRZ24, HL23b, IM21a, JQSY21, KL24b, LT22, MWX23, MPC22, MW23, MS24b, Ric23, RW23, Saf20, SAB21, Wan23, WWZ23b, ZZZ24, MPC22]. Converging [JLS25]. Converging/Diverging [JLS25]. Convex [AL25, BDN22, BGL23, CN24, Che22, EIN22, Har23, HL21, Lew25, LZ20, NS20b, PX22, Pru24, WO22]. Convexity [BDD^{+}20, BHM24, CJL21, CM23b, GKLP21, Ran23]. Convolutional [LLS25]. Cooperative [FLRX24]. Core [ABLW24]. Corner [KSS24, MW20]. Cornered [PX22]. Corners [JRS24]. Corotational [DS23]. Correct [BY24]. Correlated [BNP22]. Corresponding [CKP20, MN24b]. Corrigendum [BCJ20]. Cosmology [BLPS22]. Cost [BGL23, Buz25, GSGM24]. Costs [PVJ21]. Couette [ZZZ22a, ZZ23a]. Coulomb [BY23, CFS21, GGL24, HLlZ23, MMR^{+}23]. Coupled [AKKR23, BMZ23, BC25, CR21, CLR24, DGB23, EOE20, JZZ22, JLX23, Win20]. Coupling [ABR22, DKLT25, HJJ20]. Couplings [CFP25, DM22, SWX25]. Crack [MM23]. Cracks [GV24]. Craig [HSSY24]. Credit [BDLL25]. Criteria [DW20, LZ24a]. Criterion [AB20, LZ24b, Miz20]. Critical [AV24a, AS21, BSW23, CiDRdTY23, CHL23, DK23, Dod21, Gar24, GLM20, GH21a, HZ25, KT20, Ngu20, NR24, PVY25, TT22, TX24, Tio23, Var20, Wan22, WZ20b, XZ24, YZZ22, YCRS22, FX21]. Cross [DbMS24, DJ20, HPS23, HJ11, Jac23, TW22a, BCJ20]. Cross-Diffusion [DbMS24, DJ20]. Cross-Diffusive [TW22a]. Cross-Interaction [HPS23]. Crossing [KMV21]. Crystal [CHZ24, CHX24a, GHL22, GRZ24, HHJ24, HJLZ21]. Crystal-Colloidal [GRZ24]. Crystalline [FFV21, KP21]. Crystals [CCS24, Kim21, RS21]. Cubic [BLZ24, KMV21, LP23, LLS20, SW20b, Tes20, YYZ21, ZZ21b]. Cubic-Quintic [KMV21]. Curl [BCD24, GMS23, KbP21]. Current [FFV20, LM20]. Currents [Ger24]. Curse [JLL^{+}23]. Curvature [AM22, ACL22, CGSS24, FLS20, KL22, Mon20, Tak21]. Curvatures [BL21, IM21a]. Curve [AB20, MB20]. Curved [BBD^{+}23, FGW21, HLZ24, Sil23]. Curves [DK21c, LL21, PS23a]. Cut [FFV21]. Cutoff [DSU24, IMS20, WZGX25, Den20, Den21]. Cyclic [PVJ21]. Cylinder [CGZ20, Dun21, GK23, GWX21, Ma22, Nem21, Zha21]. Cylinders [AKKY23]. Cylindrical [WWZ23b]. Cylindrically [GMS23].

D [LX23, YY22, HL23a, HJLZ21, LZ24a, XMN23]. Daido [HJJ20]. Damage [BEMZ23, CR21, CLR24]. Damped [AS20, GG20, HH21, JWX22, Kof23, MMSH21]. Damping [AS20, BMM22, CCZ22, GLM20, GHW21, HNX25, JM23, Jia20a, JJZ22, JK22, JLX23, LSY24, MN24a, Zen23]. Darcy [GLRS22, GBS20]. Data [ANPS21, BKS23, BRR25, BP21, BPSS22, BIFL20, CTL22, CLZ22, CHaW24, CGU23, CCZ22, CHS22, CN23, CHWY25, DM20a, DH21, EFP^{+}24, FK25, Fei25, FMN21, GDS24, Ger23, GM20b, GNR22, HW22b, HW24, HJLZ21, HL22a, HL22b, IM21b, JR20, Kat20, KR21, KKNN22, KO24, LLZ24, LNOY24, LS23b, LLW24, LSY24, LXZ25, Mar22, MSW24a, Ngu20, Ngu22, OY22, RZ20, Sel24, TX24, Tes20, Wan20, WZ23]. Davey [EGN22, KSS23]. DB [Kre21]. Debye [WY24]. Debye-Yukawa [WY24]. Decay [CLYaY23, JM23, Kof23, KSW24, LS23c, LX24, LLS20, Pan20, Ste25b, Wan21]. Decaying [FSZ24, HW22b]. Deck [DGV22]. Decomposition [BGK21, BHG25, CDLLSG20]. Decompositions [GLSW23]. Decoupled [YZ24a]. Decreases [dN23]. Deep [JLL^{+}23, LLS25, LSYZ21]. Defect [ADH23, CCi23]. Defects [GHL22]. Deficient [DS22]. Defocusing [CGZ20, LL22, LP23, ZZ21b]. Degeneracies [CW25]. Degeneracy [CP24]. Degenerate [BDGdN22, BMZ23, BHTY25, BDGdN24, BH22b, CMU23, CLZ22, CHaW24, CLMZ23, CHX24b, DbMS24, DS21a, DFGV25, DG25, EOE20, EMT20, EM22, HS23, HWZ22, JLN23, JV24b, QRW24, Tio23, WZ25]. Degenerated [CY22]. Degradation [DL22b]. Degree [NS21b]. Delay [BK20b, BFLP21, DG20, Fen25, HXH22, LWC22, MMSH21, RH20, YGD21]. Delayed [ADVW20, HKlT24a, PR20]. Delta [KNP21, TX24]. Denoising [IM21a]. Density [BLN24b, BH22b, CL21b, DFGV25, LN20, LX24, LZZ21, LM20, Sin21]. Density-dependent [CL21b, LZZ21]. Dependence [FHP24, FP20]. Dependent [BK20b, BR23, BS24b, CL22a, CCFO21, Don25, Fen25, GLM20, GHW21, HJK25, JM23, JKZ21, KMW20, KMM24, KRS21, LLZ24, LSY24, MN24a, MP20b, PP24, RR20, SCB17, Tet22, YGD21, Zho22, CL21b, Li21, LZZ21, SCB20]. Depending [MPC22, WZ22]. Deposition [RW23]. Depth [DS21b, IPTT22]. Derivation [ALP21a, AMP22, BPIR25, CPW24, Eme21, Gah25, GMS21c, GL23b, LP20, Mén24, RS25, TT24, YZ24a]. Derivative [Bad24, FSW24, Ger21, HOV25, LX25]. Derivatives [FLPS24]. Describing [AČ24, CDG23, KT22b]. Description [MS21]. Design [EB23]. Desingularization [CWZ20]. Destabilizing [dN23]. Destination [DMS22c]. Destruction [RWZ21]. Determinant [YZ24b]. Determinantal [BY23]. Determination [ABFV22, BK21, BFV21, KB23, RZ21, RSS21]. Determined [HKLT24b, KRS21]. Determining [DLZ25]. Deterministic [AMMT22, AMMT24, BK24, HWXZ25, JJZ24, MS24d, RS23a]. Detonation [BW24]. Developments [Liu23]. Deviation [GL23a, HLL21]. Deviations [WZ24, XCV24]. Diagram [FL21]. Diameter [ABHN21]. Dielectric [HSVW25]. Difference [BCR21, BFLP21]. Difference-Delay [BFLP21]. Different [AF23]. Differentiability [EYZ22, Fos21]. Differential [BK20b, BGMW20, BBES22, CL20, CCS23, HKY23, HNS20, HKSW24, NS21b, OS25, RRTT24, WHY25, XC22, XCV24, YGD21, BK23a]. Differentiation [KT22b]. Diffraction [CFHX20]. Diffuse [JK22]. Diffusion [AB20, AV24a, ADK23, BMZ23, BHTY25, BEM21, Bob20, BCJ20, BMPS21, CKW20, CMT21, DbMS24, DM20a, DS22, DJ20, DN22, DLW24, DNS25, DG20, EOE20, EMT20, FHKM22, FMTY21, GR23, HSZ20, HH20, HJ11, HWZ22, IMS22, Jac23, JJL21, JV24b, KMY20, KLS23, KM20, KST23, LWX24, LLN20, LLW21, LPW22, Liu24, MV23, MS24b, MSY23, PSW25, RH20, Shu21, ST20a, TT24, TR20, WW24, Yam22, ZZ21a, ZZZ24, dN23]. Diffusion-Driven [KLS23]. Diffusions [CJK24]. Diffusive [Bla23, CS24, GHM22, HT23, HW22c, Jun25, PW21, TW22a, WZGX25]. Diffusivity [dN23]. Dilute [DS23, EFK20, FSW24, GL23b]. Dimension [BH22b, CHX24a, KL21, KNP21, KvM21, Lew25, Pan23, Pro20, TX24, vMP24]. Dimensional [ATZ24, AG24, ADL20, BLPS22, BFV21, Bl23, BP21, BLW20, BCJ20, BCD24, CW21, CWZ20, CMX22, CNW21, CKZ21, CLYaY23, CGZ20, CL22b, CKP^{+}22, Dai21b, DP25, DS21b, DL24b, Dod21, DST23, DM24, DN22, DM22, FOS20, GL24a, GMŚ21b, GH24b, GMS21c, GK22, GRR23, HJ11, HJK25, HC21, Hu22, HK23, Hu24a, HL22a, IKL20, JR20, JJZ22, KS21, KS24a, LS23a, LS21b, LXZ23, LZZ21, LP23, LLS20, MP20a, Mar22, MSM24, NR24, Nov20, PX22, PZ23, SH22, Sun21, Wan20, WWZ21, WWZ23a, WXX21, Win20, XY21, Yam22, ZLZ20, ZZ21b, CLZ22, CHS22, DW20, GLY20, LG21, Ou21, WZ20a]. Dimensionality [JLL^{+}23]. Dimensions [AKK^{+}20, BCDH25, BST22, DK21b, FGW21, FMP22, GJ21, GL20, HTY21, HWXZ25, JTY23, LL22, LYZZ24, LLWZ25, MS20b, PL23, QX25, Ste25a]. Diophantine [JLZ25]. Dirac [BY20, CW25, CHL23, DGR21, KSW24, LWW20, SH22, Tes20].

Directed [GK23]. Direction [LWZ23]. Directionally [BDN22]. Dirichlet [AO24, BK20a, BK21, BFV21, BPSS22, CDL22, DMS22b, DLW24, FL21, GDS24, GZ22, Ham25, HZ25, Jin20, KB23, KS24b, Li22a, RWZZ20, XX23b]. Dirichlet-Conormal [CDL22]. Dirichlet-to [BFV21]. Dirichlet-to-Neumann [BK21, KB23, XX23b]. Disclinations [CDM24a]. Discontinuity [ATZ24, FLS25]. Discontinuous [ANPS21, BH22b, CKKP20, FMTY21, GJP24]. Discrete [AGL24, AKV20, BB23, BBC20, BC24a, BY23, CL21d, GKM20, Kat20, LY22, RH20, dW23, vMP24]. Discrete-to-Continuum [BBC20]. Discreteness [CN21]. Discretization [GMN22, Ste20]. Discretizations [BCR21]. Disjoint [LNOY24]. Disks [Mon20, VX25]. Dislocation [vMP24]. Dislocations [CDM24a, DPS24, FGW21, GMS21c]. Disorder [CT21]. Dispersal [CL21a, FLRX24, WXZ24]. Dispersion [Bad24, CJL23, Eme21, Gas23, PP24]. Dispersive [BJ23, BP23, COS24, DST20, EGN22, KSW24, SX20a, Zha22]. Displacement [DN20, MM24]. Dissipation [CM23b, CL21c, CMT21, DP25, Del23, Den20, He23, Hol22, IZ23, LWZ23, WdlL20]. Dissipative [AV24a, BT23, BO24, CbGwG24, CH21a, GK22, HJM25, Hua24, TSMO24]. Distance [DLW22, FP21, FMS21, KB23, SS20]. Distribution [BY20, JV24a, LN20]. Distributional [PU24]. Divergence [BCD24, Cio21, DY24, DM22, PS23a]. Divergence-Curl [BCD24]. Divergence-Free [PS23a]. Divergent [Lai20, LS23a, NWG24, WXX21, XY21]. Diverging [SS22b, JLS25]. DMO [CDX22]. Domain [CW21, CK23a, CDLLSG20, Che22, CGJL24, HSV20, Jun25, KSW24, LWW20, MW20, MS23b, NWG24, RRV22, SS22b, Sen23]. Domains [Abd24, ATZ24, AKLM22, ABD22, ADL20, BLN24a, BGT22, BCDM21, EB23, FF25, GV21, Hal22, Har20, IM21a, JJZ22, Jin20, KTT20, KSS24, Mar22, MW23, MR24, OP20, OY22, PVY25, SH22, Sil23, ST24, WKM23, WZZ24, WO22, Xia20]. Dominated [Liu23]. Doping [CHWY25]. Double [ABLW24, DFGV25, DK21c, KZ21b, LN20, ZRW22, ZiDRdW24]. Double-Integral [KZ21b]. Double-Well [DFGV25, LN20]. Doubly [CS24, GHM23, JLN23, Ngu22, Set24]. Drag [ELM21]. Drift [BMZ23, BK23a, Chu23, JV24b, KT20, MS24d]. Drift-Diffusion [JV24b]. Drifts [WHY25]. Driven [CWZ24, DJ20, KLS23, KM24, WZ24, TW21]. Drop [EFK20, SAB21]. Dual [PU24]. Duct [Lai20]. Ducts [GLY20]. During [KZ21b]. Dyadic [LW23b]. Dynamic [CFS22, ELM21, GR23, GK20, MSW24a, NS21a, QWYY25, Tet22, dDUFJK25]. Dynamical [BBBH25, HH23, JMQ23, WWZ23a]. Dynamics [AHK24, BBD^{+}23, BIL^{+}23, BHS25, BDMA22, CNW21, CHW24, CDS20, CBHS23, DM20a, DK21a, DKLT25, DNS25, GGBS25, GSS23, HZ22a, HWZ22, JKL24, KT24, KV21, Kwe23, MMSH21, Mar22, MS20c, RH20, ST20a, ST20b, Tan21, XCV24, YZZ22, ZQ24, ZZ21b, vMP24].

Echo [Zil23]. Echoes [GNR22]. Eckhaus [HdRS20]. Ecosystem [HNNW23]. Eddy [FFV20, KT24]. Edge [BBD^{+}23, CDM24a, CXZ22, Cia20, GV21, YZ24b]. Edges [DGMP20, Kat23]. Effect [AO23, AL25, CKP^{+}22, EMT20, GJ21, KW23, LZ23a, WY24, dN23]. Effective [FV25, GL23b, JTY23, NS20a, SW22a]. Effects [CGSS24, Mén24, Wan24]. Eigenfunctions [BL21, CDLZ21, CCi23, FN23, GRHSZ20]. Eigenproblem [LS20]. Eigenvalue [FL21, KL24a, LP23, LLPZ21, MPV23, ST24]. Eigenvalues [AO24, ABHN21, BHM24, CN21, CRS21, CM20b, FN23, GMS21a, Gas23, Hal23, LPS22, Liu24, MRB22, SS22b]. Einstein [AK23b, CH21b, GLP23, LLWZ24, LLWZ25, NS20a, PWZZ24]. Elastic [BMPS21, CGJL24, DJLR24, DLL20, EI20, HWXZ25, KO24, LLW22, MS22, NP20, SS22a, WW23, Yos21]. Elasticity [ALP21a, CDLLSG20, FGW21, FLS25, GKLP21, KM25, LX25]. Elasto [CR21]. Elasto-plastic [CR21]. Elastodynamics [CW21, Cai23, LWZ21b, RSS21]. Elastomers [HHJ24]. Elastoplastic [CLR24]. Electric [CKP^{+}22, Li20, WY21]. Electrical [CDGH20, GHK20]. Electrically [BNSS23, DLX21]. Electro [OP20]. Electro-Magneto-Statics [OP20]. Electrodiffusion [ALW23]. Electrohydrodynamic [WY21]. Electromagnetic [CCM21, Cia20, CM20b, GH21b, HSVW25, Kre21]. Electrons [HNX25]. Electrorheological [KR23]. Electrostatic [CL21d, FMW23, RS21]. Element [GLSW23]. Elements [BB23, Cia20]. Ellipsoids [AKLM22, MMR^{+}23]. Elliptic [AC21, BDGdN22, BO21, BLS22, CT23, DU20, DM21, DK21b, DR24, GL24b, IK22, JZZ22, KZ21a, KT20, KS21, KS24a, Pen23, Sen23, ST24]. Elliptical [Ma22]. Embedded [LPS22]. Emergence [CCFO21, CW25]. Emergent [ST20a]. Enclosures [FLPS24]. Endpoint [PWZZ24]. Energies [ALP21a, BCT24, GDS24, MMR^{+}23, Tio23]. Energy [ABR22, AG23, AZ24, BSW23, BHMCR24, BDS24, BT21, BK23b, Cai23, CM23c, CLWX20, CL21c, CJ24, CL21d, CS23, DMS22b, Don22, FHKM22, GNNCR20, GR23, GV24, GMS21c, HL23b, IKL20, Kof23, KMM24, LL22, LLX21, MMR^{+}23, Miz20, NS21a, SAB21, YZZ22]. Energy-Based [ABR22]. Energy-Critical [BSW23]. Energy-Dependent [KMM24]. Energy-Supercritical [LL22]. Enhanced [Del23, He23]. Enhancement [IZ23]. Ensemble [DL21]. Ensembles [BY23, HH24]. Entropic [ADdR22, ANWS22, CEK21, FMP22]. Entropic-Wasserstein [CEK21]. Entropies [FLS20]. Entropy [BDN22, BCGJ22, CJ21, CDN21, CHW24, CHLL22, CDS20, DLW24, Eva21, GJP24, Hol22, JSW23, LX24, LY23, Pan20, QRW24, RS23a]. Entropy-Bounded [LY23]. Entry [HR21]. Entry-Exit [HR21]. Envelope [DST23]. Envelopes [EZ20]. Environment [CC25, PW21]. Environments [WXZ24]. Epidemic [PW21, PSW25]. Epitaxially [KP21]. Equality [CLWX20]. Equation [AM22, AG23, AGL24, AKV20, AHLW21, AS20, ACC^{+}21, ACC^{+}23, Bad24, BSW23, BK20a, BK20b, BH22a, BIL^{+}23, BLZ24, BHG25, BBBH25, BCFP23, BNP22, Ben24, BPR24, BEM21, BJP24, BL20, BCGJ22, CGY21, CMU23, CHJ24, CS22, CPW24, CNT23, CJL21, CCS23, CbGwG24, CNW20, CGU23, CWSGS20, CL24, Cha23, CF22, CEW21, Che22, CCJT23, CGS23, CGZ20, CMVP23, CHXZ21, CJL23, Cla23, CFS22, CGSS24, CCLM23, CMMS24, DT24, DU20, Den21, DIS25, DGWZ23, DH21, DN22, DSU24, DG20, DG25, ES24, Eva21, EY23, FX21, FLNV24, FLS20, FOS20, FHP24, FS22, Gao21, GLW22, GLW23, Gas23, GG20, Gib24, GL21, GGL24, GWZ21, GM20b, GL20, GZ22, HSZ20, HdRS20, HLS24, HH21, HL20a, HW22b, HMNW20, HJ24, HW22c, HY21, IMS20, IMO20, JK22, Jun25, KMY20, KT22a]. Equation [KM20, KM23, Kof23, KLW22b, KV20, Laf20, LLR20, LL20, LUY21, LLZ24, LZ22, LS23b, LL22, Li22b, LW24, LWX24, LP23, LLW20, LR21, LLS20, LW20b, LSY24, MP21a, MS24a, MMN23, MSY23, MP21b, MPS21, NZ24, PS22, PW25, QZ22, RZZ21, RZ21, RS25, RZ20, RW23, Saf20, SSWZ24, SS22a, SW22c, Sne20, SW20b, SZ22, SUT22, Ste25b, Tes20, Tet22, WWZ21, WZ23, WY24, WO22, XZ24, YGD21, Yan21b, YY24, EGN22]. Equations [Abd24, ALM23, AV24a, AW21, ADK23, AL23, AHM20, AKKR23, AIN25, BDGdN22, BFLP21, BR23, BCF24, Bar23, BCDH25, BMZ23, BHTY25, BC24a, BD22, BDN22, BHI24, BP21, BDS21, BK23a, BF23, BMR20, BT21, BW25, BP23, BS24b, BIFL20, BCM25, CEGP23, CCW23, CN21, CJ21, CLZ22, CW22, CM23b, CH21a, CK23b, CCZ21, CMZZ22, CiDRdTY23, CLLZ23, CWZ24, CHW24, CDZ24, CGJL24, CHWY25, CHX24b, CL21c, CJ24, CKP24, CD20, Com22, CDS20, CKP20, COS24, CF25, CS23, DLQ21, DGT20, DFV25, DM20a, DL24b, DK21a, DST23, DM24, DW20, DM21, Don22, DZZ23, DY24, DR24, DGMP20, DLW24, Eme21, EIN22, EM22, EYZ22, EZ20, FSZ24, FF25, Fei25, FMW24, FHS24, FMP22, FP21, GMS23, GL24a, GMŚ21b, GLM20, GHW21, GG24, Gin21, GP24, Gol23, GK25, HdRS20, HH21]. Equations [Hal22, HH23, HH20, HKY23, HKK23, HNS20, HW22a, HMS24, HHS23, HKSW24, HL21, HR20, HC21, HR23, HWWW24, IS21, IY20, JLN23, JQSY21, JWX22, JM23, JJZ22, JMQ23, JLX23, JW23, KL21, KLM24, KN22, KST23, KT20, KTT20, KL24b, KM24, KSS23, KV22, KT24, KLW22a, KSW24, Kwe23, LP22, LS20, LP20, LOWX20, LW21, LX24, LY22, LLW24, LWC22, LWZZ22, LX23, LY23, LZ24a, LZ21, LZ24b, MS20a, MMS21, MS24d, Nov20, NS21b, Oli21, OP20, Ou21, OY22, PYZ24, PX22, PZ23, PVY25, Pin21, QRW24, Ran23, RWZZ20, RvW20, Ros20, SW22b, SW20a, SH22, Shu21, Sil23, SCB17, SUT22, ST21, Ste25a, Sun21, TWZ20, TR20, WZ22, WW22, WJW22, WKM23, WWZ23b, WHY25, WZ25, Win20, Wri24, WYZ23, WW23, WZ24, Xia20, XC22, XCV24, Yam22, YL22, YZ24a, YZ21, YCRS22, YZ25]. Equations [YYZ24, YY22, ZZZ22a, Zen23, ZLZ20, Zha22, ZHY24, ZZ21b, Zha23, Zhu21, Zil23, dN23, SCB20]. Equilibration [DS23, Shu21]. Equilibria [DMS22a, HW22a, Hog21]. Equilibrium [BMZ23, BBG^{+}24, GGL24, HNX25, HRY22, LWC22]. Equipartition [AZ24]. Equivalent [Har23]. Equivariant [CCJT23, dW23]. Eradicate [ALV23]. Ergodic [CCKW21]. Ergodicity [DGT20, FZ21, KNP21]. Ericksen [CHZ24, CHX24a, FHM20, HJLZ21, LW20a]. Erratum [ACC^{+}23, CMX24, KS24a, Mef24, SCB20]. ES-BGK [BY24]. Essential [AS21]. Estimate [BHL23, IKL20, IY20, KvM21, KA24, QRW24]. Estimates [AKKY23, ALP21b, AIN25, Bar23, BDS21, BGMW20, CNW20, CL22b, CDX22, CJ24, CDM24b, COS24, CRTZ24, DFGV25, DY24, GGL24, HH23, Jia20b, KLLT25, KA24, KSW24, LZ20, LX25, Lia24, LR21, LZ21, LXZ25, MW20, VY23, WZ20b, Xia20, Xie25, Zha22, dHKLN24]. Estimation [RS24]. Euclidean [Chi22]. Euler [BIFL20, AW21, AKK^{+}20, ADBP21, BPIR25, BHI24, BW25, BV20b, BCM25, CCW23, CWZ20, CJ21, CW22, CYY23, CbGwG24, CCZ21, CMZZ22, CHW24, CHWY25, CD20, Din23, DL24b, DYZ21, FG22, GK22, GLM20, GHW21, Gin21, Gol23, HC21, HK23, HY22b, IS21, JM23, JJL21, KL24b, KM24, Lea23, LS23c, LW23a, Liu23, LZ24b, PX22, PL23, TT22, XY21, XMN23, YL22, Zen23]. Eulerian [RS23b, Tan21]. Evacuations [AG24]. Even [CMMS24, KPS20, TX24, WdlL20]. Eventual [Hei23]. Eventually [GIT20]. Every [Mar22]. Everywhere [PS23a]. Evolution [AT20, CL21a, Fei25, HR20, JLX23, KbP21, KT22b, LWC22, MMS21, WXZ24]. Evolutionary [FMP22, WO22]. Evolving [CEGP23, DJLR24]. Exact [HW24, Pan21, QZ22]. Exceeds [DK23].

Exceptional [ADH^{+}22, FMW23]. Excessively [Tak25]. Exchange [DDG24]. Exchanger [Tré24]. Excitable [CNW21]. Exciton [CH21b]. Exist [IPTT22]. Existence [AG24, ADBP21, AIN25, Bar20, BK23a, BC25, BJ25, BPSS22, BCDM21, CS24, CGG20, CYY23, CP24, Cha23, CCZ21, CQ25, CY21b, CT23, CKP20, CF25, DS23, DS21b, DM21, EGN22, Fen25, FHKM22, FFV20, FLS25, GL24a, GH24a, GRHSZ20, GRZ24, GG24, GLRS22, GL21, GLP23, HLS24, Hal23, HY22a, HR23, Hua24, Jac23, JSW23, KR23, KT20, KW23, KLW22a, Lea23, LS23c, LLW24, LX21, LX23, MRR20, MV23, Mar22, NS21a, NS21b, PST23, QX25, SX20a, Sil23, TW22a, Tré24, Tsu22, WZ22, WY24, WZ25]. Existence-Uniqueness [BK23a]. Exit [Bob23, HR21, LN20]. Expanding [CCW23]. Expansion [AV24b, BLN24b, Jun25]. Expansions [AP24, AS21, CNT23, Dun24, JL21, LYZ23, ND20]. Explicit [CFS23, Gib24, GZ20b]. Exponential [BMZ23, BST22, CiDRdTY23, Coo23, FR23, JLL^{+}23, NZ24]. Exponentially [DQ23, GLS23]. Extended [AP24, BPR24, MMS21, ZHY24, dHKLN24]. Extension [PU24]. Extensions [CDLLSG20, GRR23]. Exterior [ADL20, GWX21, Hal22, Jun25, OP20, TW25, WW23]. External [CHLV25, Tak25]. Extinction [CKKP20, Gir25, MSY23]. Extrapolation [BG24c, GH21b]. Extreme [CDGH20].

Faber [JL21]. Families [Hog21]. Family [LP22]. Far [BHL23, CLZ22, HL20b, LYZZ24, MWX23]. Far-Field [HL20b, LYZZ24]. Fast [BDOZ21, HR21, LX24, NY24, SX20b]. Fast-Slow [NY24]. Fatemi [MN24b]. Feasibility [BG24c, GH21b]. Feller [Bob23]. Fermi [BY20, HJM25]. Ferromagnetic [ZQ24]. Feynman [CDX21]. Fibered [Bel22]. Field [AMMT22, AMMT24, ACS23, Bar20, BK24, BZ25, BHL23, BHTY25, BC24a, BZGZ24, BLW20, BDMA22, CLZ22, CS22, CCS23, CFP25, CJK24, DLZ25, GSS23, HSVW25, HL20b, IMS22, KLSW24, Kim21, KA24, LWZ21a, LT22, Li20, LYZZ24, LLWZ25, MWX23, Mén24, MM24, MSM24, MS24c, MSW24b, NP23, OS25, PW25, Reg21, Ric23, RR25, WY21, Zha23, BDOZ21, DL21, MS20c]. Fields [BGK21, BHG25, Cia20, CKP^{+}22, FSW22, GRR23, HBK24, KN23, MPC22, MS21, PS23a]. Fierz [LP20]. Filament [Aik23]. Film [DT24, GK25]. Films [KP21]. Filter [BNP22]. Finely [CFS23]. Fingering [PTE25]. Finite [Aik23, ADH23, BB23, BCR21, BST22, CKKP20, Cia20, CW24, DS21b, GL24a, GLSW23, GMŚ21b, GK25, IPTT22, KS21, KS24a, LLW20, Miz20, WT24]. Finite-Difference [BCR21]. Finite-Time [GL24a, KS21, KS24a, Miz20]. Finitely [Har23]. Finiteness [VX21, VX25]. Finsler [EIN22]. Finslerian [HPS23]. First [BST22, CCS22, FL21, GY24, HL22b, KM20, LN20, LLhW23, MSM24, SH22]. First- [KM20]. First-Kind [SH22]. First-Order [BST22, HL22b, MSM24]. First-Passage [CCS22]. Fisher [HBK24, DN22, Gir25, KLS23]. FitzHugh [HHS23, BF23]. Fixed [CFG25, MPMS23, RS20a, UW24]. Flat [AKLM22, FG22]. Flatness [ED22]. Floating [Boc20]. Flocks [Lea23]. Floquet [SW22a]. Flow [AB20, AM22, AKKY23, BBG^{+}24, BLW20, BCMW21, BCD24, BV20b, CF21, CN24, CFHX20, CHZ24, CH23, CHLL22, CW24, Del23, DK21c, Eit22, FXX20, FLS20, FMP22, GMN22, GGBS25, GK20, Ger21, GBS20, Hol22, HLZ24, Hu24b, LXZ23, LKPR20, LLX21, Ma22, MB20, NT25, NP20, RvW20, TW25, Tak21, Tsu22, WZ20a, ZZZ22a, ZZ23a, ZZZ24, gLZ25]. Flowing [RWZ21]. Flows [ADL20, BO24, BY24, CS24, CWZ20, CYY23, CKZ21, CXZ22, CLYaY23, CHX24a, CHLW25, Coo23, CCLM23, DL24b, DMS22c, FG22, GG23, GHL22, GLY20, GK22, HT23, Hu22, HK23, Hua24, IZ23, Jia20a, JKL24, KR23, Lai20, LS23a, MWX23, NWG24, SSZ20, Wan22, WO23, XY21]. Fluctuations [GL20, GK23]. Fluid [BT23, BNSS23, CJ21, CDLLSG20, CL21b, CD21, DJ20, He23, JP20, KR23, KO24, MRR20, MS22, MS20d, PYZ24, PL23, Taw24, TW23, WZZ20, Zha23]. Fluid-Particle [CJ21]. Fluid-Rigid [BNSS23]. Fluid-Structure [MRR20, Taw24]. Fluids [ALW23, BG24a, BMR20, BMPS21, CC25, CLWX20, CDS20, DGB23, DP25, DS23, DLX21, FV25, Gal23, GMN22, GY25, KKNN22, LZ23a, Oli21]. Fluorescence [ACC^{+}21, ACC^{+}23]. Flux [FLNV24, GJP24, Hei23]. Fluxes [GJ21]. Focusing [DU20, KT22a, LW22, RS25, YYZ21]. Fokker [AIN25, Bla23, DY24, DLY21, BP20, BR23, FMP22, JQSY21, Laf20, LWZZ22, QRW24, Zhu21]. Fontana [CY21b]. Force [Laf20]. Forced [BO24, NZ24, WdlL20]. Forces [Gao21, Shu21, Tak25]. Form [BHI24, DY24, DM22]. Formally [KRS21]. Formation [BW24, BH23, CCW23, CCZ21, CHX24a, CMVP23, CGSS24, DR20a, GHPZ24, HL22b, LT24, TW21, Yan21b]. Forms [RWZZ20, IM21b]. Formula [DDG24, KLM24, NS20a, PS23b, Shu20]. Formulation [Cio21, EFHS21, GN23, HSVW25, KLSW24, RZZ21]. Formulations [Buz25, CCM21, GDS24, HSV21]. Fornberg [SSWZ24]. Forward [CFS22, GHK20, LZ23b]. Forward-Backward [CFS22]. Fourier [GZW21, BPS23, BFN20, Don25, Ger23, Pec20]. Fourth [IMO20, YYZ24]. Fourth-Order [IMO20, YYZ24]. Fractal [CMVP23]. Fraction [CJL21]. Fractional [AC21, ARS21, BCR25, BS24a, BLN24a, CL22b, Cov22, CRTZ24, DPS24, DFV25, DN20, FMMS22, FSW24, HL20a, KST23, KLW22b, Laf20, LLR20, Lea23, Li21, LW22, LWC22, NO23, RRV22, Rül21, ST20a, ST20b, TR20, WT24, XCV24, KM20]. Fractional-Order [BLN24a]. Fracture [KLSW24]. Fractured [LKPR20]. Fragmentation [CGY21]. Frame [WO23]. Framework [KKK20, LYZ23]. Free [BDS24, BDLL25, CCS22, CD21, CK22, CF20, DS21a, DN22, DNS25, FLS25, Gin21, GBS20, HM25, HXH22, KMY20, LPW21, LWZ21b, LLWZ24, LY23, LNRV21, LZ21, LZ24b, NWG24, Ou21, PS23a, SSZ20, ST21, YZ25]. Free-Boundary [LZ21, LZ24b]. Freidlin [GL23a, HLL21]. Frequencies [JLZ25]. Frequency [BJ23, CFS23, COS24, GLSW23, GLS23, KN23, OP20, RS24, UW24, WdlL20, BJL24, CWSGS20]. Frequency-Explicit [CFS23]. Frequency-Restricted [COS24]. Friction [TWZ20, Tsu22]. Frictional [MP20b]. Front [DM20a, Gar24, HK23, MMN23]. Front-Like [DM20a]. Fronts [AS21, BD22, CMX22, CMX24, HJM25, JTY23]. Froude [Ou21]. Fuglede [Pru24]. Fuglede-Type [Pru24]. Fujita [Tré24]. Fujita-Type [Tré24]. Full [AKK^{+}20, CL22a, CHZ24, CHX24a, Din23, DLZ25, DYZ21, Eme21, HC21, Hua24, LS21b, LXZ25, PX22, Yam22, YY22]. Full/Partial [DLZ25]. Fully [BGMW20, CJK24, EGN22, IK22, KL21, KO24, LLS25, TW22a]. Function [BBBH25, CDM24a, DK21b, GH21b, HR21, KL24a, LY21, LY24, LYZZ24, LR21, RZZ21, SD24]. Functional [BB23, BCR21, DR20a, FS20, GLSW23, Hei23, LL21, Mon20]. Functionals [BM24, CFV20, CN24, FFV21, MPC22, RS20b]. Functions [AP24, BG24c, CDN21, DPS24, Dun21, Dun24, ED22, FLPS24, Fos21, GIT20, JLZ25, Kat23, LSYZ21, ND20, Nem21]. Fundamental [AIN25]. Future [Oli21].

Gårding [KV21]. Gain [Jia20b]. Galbrun [HH21]. Game [DLZ25, OS25]. Games [AMMT22, AMMT24, ALP21b, BK24, BC24a, CS22, CJK24, KA24, LT22, MM24, MSM24, MS24c, Ric23]. Gamma [LL21]. Gamma-Limit [LL21]. Gap [CGY21]. Gaps [SW22a]. Gas [AČ24, BDS24, BY23, CJL21, CHW24, CHLW25, JZZ22, LZZ21, Tsu22, Wan23]. Gaseous [HY22b]. Gases [AMP22, BKPY23, CFS21, HRY22, NR24]. Gaussian [IMS20, Sne20]. General [ALP21a, BM24, BDMA22, CNW20, CCZ22, DM21, DR24, Fos21, Ger23, HM25, KR23, KKNN22, KM25, Laf20, LL20, LWC22, MW23, Pen23, VX25, Zhu21, HL20a]. Generalization [LSW23]. Generalized [AHLW21, ARD20, BCC^{+}22, BC24b, CS24, CDN21, CPW24, CGS23, CF25, DL22b, FN23, GHW21, Ham25, HOV25, HT23, Kat20, LOWX20, LR21, LZ24b, NT25, SSWZ24, WHQ25, WZ24]. Generated [LLX21, Ran23]. Generation [CHLV25, HMNW20]. Generators [AT20]. Generic [GG23]. Gennes [LLX21]. Gent [KT24]. Geodesics [BDF25]. Geometric [BHM24, CDLZ21, ES24, JRS24, KL24a, Kil22, Sai23]. Geometrical [QWYY25]. Geometrically [GJP24]. Geometries [Dai21a, Yan21a]. Geometry [KN23]. Geometry/Frequency [KN23]. Geostrophic [GHM23, Hu24a, JKL24, Nov20, Ros20]. Gevrey [Dec24, LY21, LPZ23]. Gibbs [BLN24b, LW22, RS25, ST20b]. Ginzburg [BK20a, BD22, CCJT23, Dai21a, FHM20, Gao21, GR23, IKL20, LZ22, NZ24, QX25]. Global [ALW23, AV24a, AHK24, ADBP21, BY20, BKPY23, BSW23, BIL^{+}23, BG24a, BZGZ24, Bl23, BC25, BJ25, BPSS22, CW21, CGG20, CJ21, CLZ22, CMX22, CCZ21, CNW21, CCZ22, CHS22, CLYaY23, CHWY25, CL23, CHLL22, CHLW25, CBHS23, DL22a, DS23, Din23, Don22, DGWZ23, DY24, DLY21, DM22, FHKM22, FMN21, Fue20, GM20a, GRZ24, GG24, GGL24, GZW21, GWX21, HOV25, HHS23, Hog21, HR23, HK23, Hu24a, HJLZ21, HL22a, IKL20, IS21, JWX22, JJZ22, JWW23, Jun25, KT22b, KV22, KT24, Lai20, Lea23, LS21b, LS23c, LW23a, LWZ23, LLW24, Mar22, MS20d, Oli21, PYZ24, PX22, PW21, PL23, Ran23, TT22, Tré24, Tsu22, Var20, Wan21, WZ22, Wan22, WZZ24, WZ25, WYZ23, XZZ24, XY21, YYZ21, YYZ24, YY24, ZQ24, Zen23, ZLZ20, ZZ23b, Zho22, CMX24]. Global-in-Time [Bl23, CGG20, CLYaY23, LS21b]. Globally [GK22, Pan24]. Gordon [Don22, BLZ24, Com22, CMMS24, DGR21, DM22, LLS20]. Govern [BMR20]. Grade [PYZ24]. Gradient [ALP21b, BEMZ23, BDGdN24, BGMW20, CS24, CN24, CDX22, DJ20, DG23, FMP22, GK20, HH23, HT23, IMO20, LZ20, LY24, LX25, LLX21, NP20, WZ20a, WO23]. Grain [BZGZ24, ELM21, MSW24b]. Graining [HNS20, HS24]. Grand [BDS24]. Graph [BV20b, CTL22, WT24]. Graph-Laplacian [WT24]. Graphene [BZ25]. Graphons [BH23]. Graphs [Bob23, DM24, DGMP20, HPS23, PS22, Yos21]. Gravitational [HY22b, PTE25]. Gravity [DS21b, FG22, GGBS25, MS24a, Set24, Ste25a, Zen23]. Gravity-Capillary [DS21b, Set24, Ste25a]. Green [DK21b, KLLT25, LYZZ24, SD25]. Greenberg [CHLL22]. Grid [CCS23]. Griffith [BB23]. Gross [AHLW21, BDS24, BLN24b, GZ22, LW20b]. Ground [BLS22, CJL21, Dod21, GLP23, HLS24, LLWZ25, PWZZ24, PS22]. Group [CMP25]. Growing [WHY25]. Growth [BGMW20, CGY21, CiDRdTY23, DbMS24, GMS21c, HXH22, HZ25, KL23, LX21, MP21a, RS20b, RR25]. Growth-Fragmentation [CGY21]. Gyroscopic [Mén24].

H [Sim21]. H-Measures [Sim21]. Hadamard [AP24]. Half [BPS23, BL20, BIFL20, CM23c, DCWF^{+}22, DYZ21, HW22c, JZZ22, KLLT25, KM20, KMM24, MMS25, WW23]. Half-Line [BPS23, BL20, CM23c, KMM24]. Half-order [KM20]. Half-Plane [BIFL20]. Half-Space [DCWF^{+}22, DYZ21, HW22c, WW23]. Hall [Dai21b]. Hamilton [BDN22, DM24, EIN22, EYZ22, EZ20, GMŚ21b, HH23, JMQ23, KLM24, KTT20]. Hamiltonian [BCC^{+}22, BDN22, BDF25, CCLM23, DL24a, GY24, LP20, MPZ22, Nei20]. Hamiltonians [SW22a]. Haptotaxis [DL22a]. Hard [AMP22, Cha23, Den21, HM25, LZ20, Sne20, WZGX25, YZ24b]. Hard-Sphere [AMP22]. Hardening [Wan24]. Hardy [CP24]. Harmonic [BHL23, CHLV25, CC23, GRR23, Hal22, HWXZ25, LLWZ24, OP20, CCM21, HH21]. Harnack [BDS21]. Harris [CGY21]. Hartree [DLQ21, CHL23, Saf20, Tes20]. Hartree-Type [CHL23]. Hasimoto [Aik23]. Heat [ACC^{+}21, ACC^{+}23, BMPS21, CL22a, DG25, GHM22, GL20, IV22, Li22a, LZZ21, PZ23, RWZZ20, RZ20, SL22, Tré24, Wan23]. Heat-Conducting [BMPS21, LZZ21]. Heat-Reflection [SL22]. Heated [JV24a]. Heisenberg [CMP25]. Hele [Chu23]. Helical [KLW22a]. Hellinger [Buz25, FMS21]. Helmholtz [BHG25, CWSGS20, CEW21, Cla23, GMS21a, GLSW23, GRHSZ20]. Hemivariational [Han20, MP20b]. Herglotz [VX21, VX25]. Hermitian [Tet22]. Heterogeneous [BEMZ23, CF21, DJLR24, JR21, PW21]. Hidden [CM23b, GKLP21]. Hierarchies [CDS20]. Hierarchy [HZ22b]. High [ACL22, AL23, BJ23, BJL24, Bel22, CCi23, CWSGS20, DN22, Fep21, FJ22, GLSW23, GLS23, LL22, Pan23, UW24].

High-Contrast [CCi23]. High-Curvature [ACL22]. High-Frequency [BJ23, GLSW23, BJL24, CWSGS20]. High-Velocity [AL23]. Higher [ATZ24, CPW24, CCS22, GJP24, GL20, HZ22b, JLN23, NS21a, NS21b, RRTT24]. Higher-Dimensional [ATZ24]. Higher-Order [CPW24, CCS22]. Highest [Cai23]. Highest-Order [Cai23]. Highly [CGG23, GL24b]. Hilbert [BPS23, KV20, Yan21b]. Hilliard [GLRS22, CEGP23, CFS22, CGSS24, CMM20, GK20, RW23, SW20a]. Hilliard-Type [RW23, SW20a]. History [MP20b]. History-Dependent [MP20b]. Hoff [Bel22]. Hölder [BDS21, CMP25, CD20, JLL^{+}23, Lia24, Zhu21]. Hölder’s [Cio21]. Holes [ABD22]. Holm [LP22, MP21a]. Holomorphy [DH24]. Homoclinic [LY22]. Homogeneous [Fue20, RRTT24, WXZ24, WY24, XX23b, CHXZ21]. Homogenization [Bel22, BEM21, CGG23, CFS23, CCKW21, CFG25, DGB23, Fep21, FFV21, GZ20a, GL24b, Jin20, JR21, JRS24, KZ21a, RvW20, Sen23, Zha21]. Homogenized [FJ22]. Honeycomb [AFH^{+}20]. Honeycomb-Lattice [AFH^{+}20]. Hookean [BHMCR24, DS23]. Hopf [Dai21b, Shu20]. Horizontal [FG22]. Horizontally [JJZ22]. Hot [Web20]. Hotspots [TW21]. Hull [HL21]. Hump [DS21b]. Hunter [GLW22, GLW23]. Hybrid [NP23]. Hydrodynamic [CLMZ23, CY21b, FHM^{+}24a, FHM^{+}24b, WWZ21, WMZZ21, WO22]. Hydrodynamics [FW21, HHJ24, Kim21, Wan21, WZZ24, ZZ23b]. Hydrostatic [LPZ23, PYZ24, WW22]. Hyperbolic [BCF24, CMU23, CY22, CK22, CBHS23, FR23, FP20, GK21b, HHJ24, HJLZ21, HL22b, JZZ22, Kil22, KRS21, LPZ23, LPW22, LSY24, MP20a, Per20]. Hyperbolic-Elliptic [JZZ22]. Hyperbolic-Parabolic [CBHS23, LPW22].

Hyperbolicity [BCDH25]. Hyperbolized [HR23]. Hyperelastic [DMS22a]. Hypersonic [HLZ24]. Hypersurface [CMMS24]. Hypocoercivity [Eva21]. Hypodissipative [CD20]. Hypoelliptic [MS24c]. Hypothesis [FL24]. Hysteresis [ATZ24, EK20].

Ideal [ALW23, ADL20, CD21, Gal23, JJZ22, JW23, TW22b]. Identifiability [GSGM24]. Identification [HJK25, RZ20, Wan24]. Identities [SD25]. Identity [NS21a]. Idiosyncratic [CS22]. II [JM23, CJL21, HL20a, KSS23, MPC22, SD25]. Ill [BHI24, DGV22, OY22, WdlL20, Zha23]. Ill-Posed [WdlL20]. Ill-Posedness [BHI24, DGV22, Zha23]. Ill-Prepared [OY22]. Imaging [LM20]. Immediate [LY23]. Immersed [BBG^{+}24, KO24]. Impedance [BC24b, CDGH20, CC23, GHK20, LM20]. Impenetrable [vM22]. Impermeable [CL21b]. Implicit [BV20b, ZRW22, ZiDRdW24]. Improved [FW21, HH24]. Incidence [PSW25]. Incident [VX21]. Inclined [Ste25a]. Inclusion [AdSST24]. Inclusions [ACL22, ABFV22, BFV21, BWF24, CDGH20, LZ20, LX25, OS25]. Incompatible [KM25]. Incompressible [BCK25, BPIR25, BNSS23, BMR20, BMPS21, CW21, Cai23, CW22, CYY23, CLYaY23, CD21, CH23, DP25, DM20b, FV25, FMW24, Gin21, GZW21, GWX21, HL21, JJZ22, JW23, KM23, KL24b, KO24, LZ23a, LWZ23, LX21, LZ21, MS20a, MS22, OY22, SSZ20, ST21, WX21]. Increasing [DQ23, EI20, KUW21]. Indefinite [MPV23]. Independent [CR21, CLR24]. Index [CCLM23, KS24b]. Indices [BCC^{+}22, SL22]. Indirect [EMT20, JLX23, Li23]. Induced [AGL24, CQ25, GL21]. Inelastic [AL23]. Inequalities [ABD22, BDS21, BGMW20, Han20, Har20, Hei23, KV21, Ngu22].

Inequality [Cio21, DFGV25, LW21, MP20b, Zha21]. Inertia [CM22, CHXZ21, CHQ22, HSZ20]. Inertial [KLSZ22, LW20a]. Infection [TW21]. Inference [DM24]. Infinite [BST22, DM24, GH24b, IKL20, Lai20, LS21a, MMN23, WWZ23a]. Infinite-Dimensional [DM24, GH24b, WWZ23a]. Infinite-time [LS21a]. Infinitely [LPS22, XY21]. Infinity [CF20, LLhW23]. Inflow [WXZ24]. Influence [HL23a]. Inhomogeneities [KSS24, VX25]. Inhomogeneous [CM23a, CHJ24, DK21a, FK25, GWX21, HL20b, QX25, SSZ20, Sne20]. Initial [BZGZ24, BH22b, BIFL20, CHaW24, CCZ22, CHS22, CN23, CHZ24, CHWY25, DM20a, DH21, EYZ22, Fei25, Ger23, GHM22, GM20b, HW22b, HL22a, HL22b, JJZ24, KKK20, KNP21, KST23, LXZ25, Mar22, Miz20, RRV22, WZ23, YL22, ZZZ22b]. Initial-Boundary [BZGZ24, CHZ24]. Initial-Boundary-Value [CHS22]. Initial-Value [YL22]. Injection [CD21]. Injectivity [MS21]. Inputs [HJJ20]. Instabilities [BG24b]. Instability [BCR25, BJP24, CW25, JP20, KMW24, LP22, LOWX20, Yan25]. Instantaneous [LX24, Pin21]. Insulating [ACNT24]. Integrability [AČ24]. Integrable [AKV20, KT22a]. Integral [BDD^{+}20, CFV20, DH24, EFHS21, FMMS22, FHS24, HSV20, HSV21, KZ21b, MS21, PS23a, RZZ21, RS20b, SH22, SUT22]. Integrals [DL24a, GZ20a]. Integrated [Kof23]. Integration [Lew25]. Integro [BK23a, BBES22]. Integro-Differential [BBES22, BK23a]. Intensity [PP24]. Intensity-Dependent [PP24]. Interacting [BCMW21, CaDZ24, DIS25, DGR21, FOS20, KvM21, LRW20, MS22, SS22a, Xie25]. Interaction [ALM23, BDF24, BNSS23, BM24, CDF^{+}22, DR20a, GK21a, Ger21, HPS23, LT22, MRR20, PW25, Shu21, Sol24, Taw24, TW23]. Interactions [BBC20, BK24, CCFO21, DDG24, GRZ24, HRZ24, NR24, ST20a]. Intercritical [CM23a]. Interface [AF23, ACC22, DT24, DST23, GGBS25, GHPZ24, GN23, HL23b, HMNW20, JR21, NS21a, QB24, RR25, SS20]. Interfaces [BBD^{+}23, SL22, TW22b]. Interfacial [WY21]. Interior [LS20, XMN23]. Intermediate [MPS21]. Intermittency [ED22]. Internal [Set24]. Interpolation [Har20]. Interval [Bad25, HY21]. Invader [DNS25]. Invariance [BCDH25, BWF24]. Invariant [AKV20, DQ23, GL23a, KMW24, KLW22a, WWZ23a]. Invariants [QB24]. Invasion [CQ25, DL22a]. Invasive [ALV23]. Inverse [BHL23, BKS23, BL21, CDLZ21, CNW20, CGU23, CCS22, CM20b, CRTZ24, EI20, EFP^{+}24, EZ20, GV24, GSGM24, HL20b, HLZ24, IY20, KM20, KMM24, KUW21, KRS21, LLZ24, LNOY24, Li21, LLW22, LW24, LSY24, MPMS23, Mef21, Mef24, RS20a, Sel24, SW22c, Ste20, SZ22, UZ23, Yan21a, YZZ22]. Inversion [HL20a, QWYY25]. Invertibility [PU24]. Inverting [Kat20]. Investigation [SCB17, SCB20]. Inviscid [DbMS24, GY25, HWWW24, Jia20a, Ngu20, VY23, Wan20]. Invisibility [BLPS22]. Involving [BCGJ22, CP24, HZ25, Li21, Lia22, MPZ22]. Ioffe [LLWZ25]. Ion [EOE20]. Ionized [LZZ21]. Irrational [SW20b]. Irregularly [Liu23]. Irrotational [Ma22, MWX23]. Isentropic [CHW24, Gol23, OY22, Tsu22, YZ25, ZZZ22a]. Island [RWZ21]. Isoperimetric [ABLW24, Pru24]. Isothermal [CKZ21, CHW24, FG22, WZ25]. Isotropic [BJ25]. Issuing [DW21]. Iteration [LWZZ22]. Itself [CDM24b].

Jacobi [GMŚ21b, JMQ23, BDN22, DM24, EIN22, EYZ22, EZ20, HH23, HH24, KLM24, KTT20]. Jacobian [Ran23]. Jet [CH23, LS23a, NWG24, PX22]. Jets [DW21]. Joint [KN23]. Josephson [PWZZ24]. Jump [ATZ24]. Junction [Gen25, Hol22, PWZZ24]. Junctions [ELM21]. Justification [LW20a, LWZZ22, Ros20].

Kac [CPW24, CDX21]. Kadomtsev [EGN22]. Kalman [DL21]. KAM [GL23a, HSSY24]. Kantorovich [Buz25, FMS21, SHM25]. Kato [LZ24b]. Kawasaki [Mou20]. KdV [BL20, CM23c, HdRS20, LW21]. Keller [BCJ20, CW24, GH24a, GH21a, CKW20, CBHS23, HJ11, MP20a, Miz20, Win20]. Kepler [BDF21]. Kernel [HKLT24b, LUY21, RZZ21]. Kernels [CJ24, CGK23, DN20]. Kimura [CM23b]. Kind [SH22]. Kinematic [HJLZ21]. Kinetic [Bad25, BDMA22, CaDZ24, CY21b, CJ24, DY24, HKLT24b, JK22, Per20, QRW24, ZHY24, Zhu21]. Kinetics [TT24]. Kirchhoff [BH22a, GG24, QZ22]. Klar [CHLL22]. Klein [DGR21, BLZ24, Com22, CMMS24, Don22, DM22, GG20, LLS20]. Knowledge [JJZ24]. Kohn [SCB20, SCB17]. Kolmogorov [AIN25, DY24]. Kompaneets [BIL^{+}23]. Korn [Har20]. Korteweg [LM21, BH22b, CL21b, GM20b, MS20d, RvW20, YZ24a]. KPP [DN22, Gir25, KLS23]. Kramers [GH21b]. Krein [KPS20]. Kronig [GH21b]. Krylov [BGMW20]. Kuramoto [CM22, CHXZ21, HSZ20, HJJ20]. Kuznetsov [Dun21, LR21, YZ24a]. KWC [AMS24].

Label [MS20c]. Labels [WT24]. Lack [CHLV25]. Lagrangian [AMMT24, GMN22, KR21]. Laguerre [HH24]. Lamé [Fen25, LZ20, MMSH21]. Lamellar [ACC22]. Lamination [HL21]. Lampis [Che22]. Landau [BK20a, BD22, BJ25, CDL24, Cha23, CCJT23, Dai21a, DGOY24, FHM20, Gao21, GR23, GGL24, HLlZ23, HNX25, IKL20, LZ22, LLX21, NZ24, QX25, Sne20]. Langevin [CF22]. Laplace [CP24, Dun21, EIN22, EFP^{+}24, GN23, Wic23]. Laplacian [AO24, BGT22, CDLZ21, CK23a, CF20, FMMS22, GV21, Lia22, dDUFJK25, WT24, Yam22]. Laplacians [CTL22, CL22b]. Large [ADH23, BSW23, BBBH25, CLZ22, CHS22, CLYaY23, CHWY25, DR20b, EFP^{+}24, FMN21, GL23a, HJJ20, HKK23, HLL21, Hu24b, HL22a, KSS23, KW23, KT24, LT22, LS23b, Li22b, LLW20, LLPZ21, LG21, Liu24, MSW24b, ND20, Pan21, RS23b, Sin21, WZZ24, WZ25, WZZ20, WZ24, XCV24, ZLZ20, ZZ23a, ZZZ22b]. Large-Amplitude [Sin21]. Large-Parameter [ND20]. Large-Scale [KT24]. Large-Time [DR20b, HKK23, MSW24b, WZZ20]. Lattice [AFH^{+}20, ELM21]. Lattices [LRW20, Wri24]. Laval [Tsu22]. Law [FLNV24, FN23, GK21b, HL22b, KV20, Yam22]. Laws [Bob23, CMP25, CKP20, DMS22c, DR20b, FR23, FP20, GJ21, GJP24, JR20, LZ23b, Pan20, RS23a, Yua23]. Lax [Gas23, LP23, PS23b, Shu20]. Lax-Oleinik [PS23b]. Layer [ACNT24, DLX21, Gah25, GLS23, HW22c, JL21, Kor22, LY21, Sin21, WZ20a, WWZ23b, vM22]. Layered [BFV21, CFS23, KN23, LYZZ24]. Layers [AHK24, BG24b, Bob20, CHaW24, Var20]. Leading [Kil22]. Leaky [FS22]. Leapfrogging [BCM25]. Learning [WT24]. Least [Buz25, DG23]. Least-Cost [Buz25]. Lebesgue [Pec20]. Lefever [BPR24]. Legendre [ND20]. Length [Aik23]. Lens [NS20b]. Leray [Dai21b, WJW22]. Leslie [CHZ24, CHX24a, FHM20, HJLZ21, LW20a]. Level [IM21a]. Lévy [WZ24, XCV24]. Liénard [CCJT23]. Lifespan [BH22a, HL23a]. Lighthill [CFHX20]. Like [DM20a, CCKW21, GM20b]. Limit [AF23, AHM20, AKKR23, Bad24, BLZ24, Ben24, BSH24, Bla23, Bob23, BDS24, BK23b, BDMA22, BH22b, BIFL20, BCM25, CM20a, CCS23, CaDZ24, CP24, CLMZ23, CHW24, CY21b, CHLL22, Chu23, CHQ22, CC22, CBHS23, DbMS24, DL21, DM20b, EFK20, EB23, FMW24, Gas23, GHPZ24, GZW21, GZ22, HSZ20, HL23b, HT23, HWWW24, JJL21, JW23, KM23, KV22, LS23b, LW20a, LL21, LW23a, LX21, LSW23, LW23b, MWX23, Mén24, NP23, Ngu20, Ou21, OY22, PL23, RR25, TWZ20, VY23, Wan20, WWZ21, WWZ23b, WW24, WO22, WZGX25, ZZ24, Zho22, vMP24]. Limiting [BPSS22, Lea23]. Limits [BBC20, CFP25, Din23, FHM^{+}24a, FHM^{+}24b, Gal23, GHM22, GKM20, GP24, GY25, IMS22, SX20b, WX21]. Line [BPS23, BL20, CM23c, CXZ22, GNNCR20, GMS21c, JZZ22, Jui20, KMM24, MS20b]. Line-Tension [GMS21c]. Linear [ALP21a, Bad25, BHTY25, BST22, BLS22, Eva21, FLS25, GL24b, HXH22, HSSY24, HSV20, HSV21, IY20, Jia20a, LPZ23, LX25, LLPZ21, LPW22, Ma22, MP21b, Oli21, RRTT24, RS20b, Ste20, ZZZ22a, Zha22, BLN24a]. Linearity [Sai23]. Linearization [KM25]. Linearized [BMM22, Den20, FGW21, GH24b, Ger21, SS24, Zil23]. Liouville [BCFP23, Ben24, KT22a]. Lipschitz [ACC^{+}21, ACC^{+}23, ABFV22, ALP21b, BGK21, BFV21, BLN24a, CCS24, CTL22, CKZ21, DL24b, DG25, GH24b, GKK22, HBK24, KM20, KA24, Mef21, Mef24, OP20, SH22, TW25, Xia20]. Liquid [CCS24, CGG20, CHZ24, CHX24a, EFK20, GHL22, GRZ24, HHJ24, HJLZ21, Kim21, RvW20, RRV22, SAB21]. Liquid-Vapor [RvW20]. Listmode [MSW24a]. Local [ABR22, AG23, ABFV22, ADVW20, BSH24, Boc20, BT21, CL22a, CK23b, Cha23, CJK24, COS24, DGOY24, DKLT25, HJJ20, HW22b, Kof23, LT24, Lia22, LY22, LR21, MS24d, NS20a, SD24, SD25, Sil23, Zha22, DR20a]. Local/Nonlocal [DR20a]. Localised [CCi23]. Localization [DST23]. Localized [ACL22, Bar23, KW23, XX23b]. Locally [HLL21, Sen23]. Located [LX25]. Location [LY24]. Logarithmic [CS23, GHK20, KST23, MSY23, SW20a]. Logistic [KS21, KS24a, MSY23]. Long [AKKR23, BBC20, BK24, DK21a, DNS25, FS22, GGBS25, GM20b, JLN23, LPW21, LM21, MSM24, MB20, MPS21, PW21, RWZ21, SX20a, Ste25b, Sun21, XY21, XC22, ZZ21b]. Long-Range [BBC20]. Long-Term [Sun21, MB20]. Long-Time [AKKR23, BK24, DNS25, GM20b, JLN23, LPW21, PW21]. Longer [BH22a]. Longtime [Liu23]. Loss [BIL^{+}23, CFHX20, KZ21b]. Lotka [PR20]. Low [AG23, AHM20, BK23b, CCM21, DGMP20, Gal23, JJL21, MWX23, OP20, Ou21, Pec20, Sel24, WHY25]. Low-Energy [BK23b]. Low-regularity [CCM21]. Lower [FSW22, HH23, IMS20, YCRS22]. Lowest [KL24a]. Lugiato [BPR24]. Luke [MS20b]. Lusternik [CT23].

Mach [AHM20, Gal23, JJL21, MWX23, Ou21]. Machlup [LL21]. Macro [CFP25]. Macroscopic [CF21, DS23, HKLT24b, Hol22]. Magnetic [BZ25, BHL23, Ben24, BLW20, FSW22, GV21, KL24a, Kim21, LWZ21a, Li21, LLWZ25, Reg21, RWZ21, Sel24, Zha23]. Magneto [OP20, Zha23]. Magneto-Micropolar [Zha23]. Magnetoelasticity [BK23b]. Magnetohydrodynamic [BLW20, Hua24, JW23, LWZ21a, CH21a]. Magnetohydrodynamics [LZ21]. Magnetorheological [DGB23]. Majda [BW24, LZ24b]. Malik [GP24]. Managed [CJL23]. Manifold [CDM24b, RWZZ20]. Manifolds [BLPS22, JJZ24, KLSZ22, KL22, LNOY24, NK21, SS24, WWZ23a, YYZ24]. Many [AKKR23, BH22a, Har23, LPS22].

Many-Server [AKKR23]. Map [BFV21, GHK20, KB23, XX23b]. Mapping [Mon20]. Maps [CDM24b, MR24]. Marginal [CL20, PVJ21]. Marginals [GSGM24]. Markov [HS24]. Markovian [FZ21]. Masks [RS24]. Maslov [BCC^{+}22, CCLM23]. Mass [AGL24, BL20, BGL23, CMT21, Dod21, FX21, FMTY21, GLRS22, GZ20b, GH21a, HLS24, HMNW20, LS21a, Liu23, MSW24a, TX24, Win20]. Mass-Critical [Dod21, FX21]. Mass-Subcritical [HLS24]. Massless [HNX25]. Master [BC24a, BDS21, CS22]. Matched [GLS23]. Matching [CY22, DCWF^{+}22]. Material [EB23, Ger21]. Materials [AMS24, BEMZ23]. Mathematical [BMR20, CQ25, CK22, EOE20]. Matrices [BLN24b]. Matrix [BV20a, Cio21, DS22, KPS20]. Matrix-Valued [BV20a]. Matter [BG24a, EFK20]. Maximal [LR21, MRR20]. Maximal-in-Time [MRR20]. Maximization [ABHN21]. Maximizing [CL25]. Maximum [LY24]. Maxwell [DGOY24, WZGX25, CN21, CC23, DLQ21, DST23, Hal23, HJT22, LS20, LP20, OP20, Xia20, YZ21, dHKLN24]. Maxwellian [BKPY23, CDL24]. May [TW21]. McKean [CDF^{+}22, HRZ24, LWX24]. McWilliams [KT24]. Mean [AM22, AMMT22, AMMT24, ACS23, BK24, BHTY25, BC24a, BDMA22, CS22, CCS23, CFP25, CJK24, DN20, DL21, DLZ25, FLS20, IMS22, KA24, LT22, Mén24, MM24, MSM24, MS24c, MS20c, NP23, OS25, Ric23, Tak21, BDOZ21]. Mean-Field [ACS23, BDMA22, CFP25, Mén24, NP23, DL21, MS20c]. Measure [AKKR23, BHS25, CbGwG24, Gal23, GK21b, Hua24, LW22, ST20b, WY24]. Measure-Valued [AKKR23, BHS25, CbGwG24, Gal23, GK21b, Hua24]. Measurement [DLZ25, Fei25, Rül21]. Measurements [ABFV22, CM20a].

Measures [AKV20, ARD20, BV20a, BCD24, Cio21, GMŚ21b, GL23a, RS25, Sim21]. Mechanism [PSW25]. Media [ALW23, BJ23, CCi23, CNW21, CCKW21, CCM21, CC23, DGT20, EK20, GMN22, Gir25, GBS20, HL21, JR21, KN23, KM23, LKPR20, Ngu22, RvW20, WZ24]. Medians [CCE24]. Medium [BFV21, CHLV25, CGU23, DIS25, Fep21, GK21a, HL20b, LYZZ24]. Meissner [KbP21, QX25]. Mellin [BPS23]. Membrane [Bob20]. Memory [DN20, FS22, JLX23, Sim21]. Memory-Damping [JLX23]. Memory-Structured [FS22]. Memristors [JV24b]. Menten [TT24]. Merriman [BV20b]. Mesh [BCR21]. Mesh-Size [BCR21]. Mesoscopic [BF23]. Metamaterial [CK22]. Metamaterials [ADH^{+}22]. Method [ADL20, CWZ20, CDM24a, Cla23, DCWF^{+}22, FW21, GLSW23, HSSY24, YGD21, ZHY24]. Methods [HSV20, RS23a]. Metric [BDN22, BDF25, CCS24, CDN21, HBK24, PS22]. MHD [CHS22, Dai21b, FK25, JWX22, JJZ22, LS21b, LY21, LWZ23, LYZ23, TW22b, WW22, YZ25]. Michaelis [TT24]. Micro [CFP25]. Micro-to-Macro [CFP25]. MicroBubbles [MS23b]. Microlocal [Chi22]. Micromagnetics [DMRS20]. Microphysics [DKLT25]. Micropolar [KR23, Zha23]. Microscopic [CF21, KP21, RS25]. Microscopy [ACC^{+}21, ACC^{+}23]. Microstructure [MM23]. Microstructures [Sim21]. Microswimmers [GL23b]. Migration [BDLL25]. Mild [Zha23]. Mildly [RS23a]. Mills [Pec20]. Minimal [CT21, GNNCR20, TX24]. Minimization [BHMCR24]. Minimizer [CJL23]. Minimizers [DMS22b, FS20, MPC22, MMR^{+}23, MPZ22]. Minimizing [DPS24]. Minnaert [AFH^{+}20]. Mironescu [DDG24]. Mirrors [Pin21].

Misfit [DPS24]. Misorientations [ELM21]. Mitochondrial [EOE20]. Mixed [AB21, AL25, CDL22, DQ23, EFHS21, HMS24, HC21, LWZ21b, OP20, Pen23, Saf20, SWX25, ZRW22]. Mixed-type [HC21]. Mixing [CM22, Coo23, GG23, Gao21, KMW24, NZ24]. Mixture [AMP22, AL23, KKNN22]. Mixtures [AČ24, DJ20]. mKdV [HZ22b, Ste25b]. Mobilities [HPS23]. Mobility [HL23b]. Mode [BN24]. Model [ACC22, ATZ24, ADBP21, AKKR23, BY20, BKPY23, BCK25, BSH24, BZGZ24, Bla23, BJ25, BK23b, BY24, BBES22, CGG20, CPW24, CaDZ24, CKW20, CLMZ23, CHZ24, CQ25, CM22, CY21b, CHLL22, CBHS23, CLR24, CMM20, DL22a, DGB23, DbMS24, DGV22, DFGV25, DKLT25, DMS22c, DNS25, EOE20, EFK20, FHM^{+}24a, FHM^{+}24b, FL24, FSW24, Fue20, GM20a, GHM23, GMS21c, GRZ24, GHPZ24, GBS20, GSS23, HJJ20, HM25, HJ11, HJLZ21, HY21, HRY22, IOPR21, JJL21, KR23, KLS23, KLSW24, KNP21, KT24, KKNN22, KP21, KO24, KS24b, LS21a, LT24, LLN20, LLW21, LX21, LPW22, LNRV21, MRR20, MSW24a, Mén24, MN24b, MS20d, NS21a, Pan24, PSW25, PTE25, SAB21, TW21, Wan21, WW24, WMZZ21, WZZ24, WZZ20, ZZ23b, ZZZ24, vMP24, BCJ20]. Model-the [CQ25]. Modeling [Bob20, CCS24, CCS23, CDM24a, FMN21, HXH22]. Models [ALW23, ABR22, AG24, BEMZ23, BJP24, BW24, BDMA22, CF21, CCFO21, CHaW24, CL21b, CMZZ22, CL21d, DS23, Eme21, FJ22, Hol22, IMS22, JKZ21, KW23, KS21, KS24a, LW23b, NP23, PR20, RR25, XMN23]. Moderately [SX20b]. Modes [ADH23]. Modified [BRR25, CM20b, GM20b, Hal23, Ros20, YL22]. Modular [MMS21]. Modulated [CJ24]. Modulation [HdRS20]. Modulational [BJP24, JP20]. Modulus [IKL20]. Moisture [DKLT25]. Molecular [Liu23]. Moment [BCDH25, MS21]. Moments [CHJ24, LS23b, Reg21, ZHY24]. Momentum [BKS23]. Monge [TT22, CHX24b, Ham25, Jui20, Lew25]. Monge-Ampère [Ham25]. Monokinetic [Mén24]. Monostable [KM23]. Monotone [BG24c, CS22, CL23, HLL21, Jia20a, MV23]. Monotonic [BG24b]. Monotonicity [CDGH20, CJL21, GP24, HL20a, MM24]. Monotonicity-Based [CDGH20, HL20a]. Morawetz [BDF24]. Morrey [WJW22]. Morse [DIS25, KS24b]. Moser [AHLW21, Bad24, LW20b, Wri24]. Motility [CMM20]. Motion [BZGZ24, BMR20, CGG20, ELM21, FGW21, MSW24b]. Motions [NP23]. Motivated [KbP21]. Moving [DLX21, TW25, TW22b]. Multi [CL20, DS21b, PVJ21]. Multi-Hump [DS21b]. Multi-Marginal [CL20, PVJ21]. Multibody [BBC20]. Multibump [XZ24]. Multidimensional [FHM^{+}24b, Fue20, HR21, HR23, HL22b, MWX23, Pan20, Yua23]. Multidimensions [Tan21]. Multilane [CHLL22]. Multilayer [HXH22, ST21]. Multilayered [BCMW21]. Multimarginal [Buz25, GZ20b]. Multiphase [Kil22]. Multiple [AHLW21, BNP22, CYY23, CT23, FMP22, GMS23, GY24, Var20]. Multiplicative [DGT20, DGWZ23, HH20, KM24, Zha22]. Multiplicity [DM21]. Multipole [HW22a]. Multipopulation [MS20c]. Multiscale [ARD20, CW22, HLL21, LWX24, MN24b]. Multispecies [AL23, CW24, CL21a]. Multistable [GKK22]. Multivalued [ZRW22, ZiDRdW24]. Multivortex [LW20b]. Murat [ARS21].

Nabarro [vMP24]. Nagumo [BF23, HHS23]. Nanoparticles [DLL20]. Navier [CW24, Dec24, GH24a, GH21a, LX23, ZZ24, AF23, AB21, AHM20, Bar23, BBG^{+}24, BP21, BT21, BO24, BFHZ22, BFN20, BH22b, CLZ22, CL22a, CK23b, CDZ24, CL21c, CDX22, CD20, DK23, DL22b, DM20b, DW20, Don25, FMW24, FOS20, GL24a, Ger23, GZW21, GWZ21, GWX21, HKK23, HL23a, HMS24, Hei23, HL23b, HR23, HL22a, HWWW24, KLW22a, Kwe23, LS23c, LPZ23, LX24, LY23, LZ24a, MS20a, Ngu20, Ou21, OY22, PZ23, RvW20, SS22a, ST21, Sun21, TW25, Tak25, TWZ20, Taw24, VY23, WZ22, WJW22, WKM23, WZ23, WWZ23b, WZ25, Win20, WYZ23, WW23, Yam22, YZ21, YY22, ZZZ22a, ZLZ20, ZZ23a, ZZZ22b, Zho22, gLZ25]. Navier-Slip [Ngu20, Taw24]. Navier-Stokes [HL23b]. Navier-Stokes/Allen [HL23b]. Navier-type [AB21]. Near [AS21, BKPY23, BHL23, CDL24, CK23b, GGL24, GNR22, HdRS20, HRY22, Jia20a, Liu23, Wan20, Zha23, BY20, Fue20, LWZ21a]. Near-Boundary [Liu23]. Negative [LLhW23]. Nematic [CCS24, CHZ24, CHX24a, GRZ24]. Neo [BHMCR24]. Neo-Hookean [BHMCR24]. Neohookean [IOPR21]. Nernst [GWZ21, HY22a, IS21, ZZ24]. Nested [KTT20]. Network [LSYZ21]. Networks [AMMT24, BFLP21, BDOZ21, BH23, DMS22c, JLL^{+}23, Kor22, LLS25, LWZZ22, NP20, PS23b]. Neumann [AKLM22, BK21, BFV21, BHM24, DH21, IK22, KL24a, KL21, KB23, LPS22, MPV23, PST23, Ric23, XX23b]. Neural [JLL^{+}23, Kor22, LLS25, LWZZ22, LSW23]. Neuronal [FS22]. Neurons [CaDZ24]. Neutral [CLMZ23, PL23, ZZ24]. Newtonian [CC25, Din23, MS22]. Nield [Dun21]. NLS [Bad24, BDF24, CM23a, Dod21, DGMP20, HOV25, KT22a, KMV21, LW22, PS22, ST20b, TX24, YZZ22, YYZ21]. No [BIFL20, DP25, IPTT22, LT24, RRV22, Fos21]. No-Slip [BIFL20]. Node [GM20a]. Noise [AV24a, ACS23, BK20a, BFHZ22, CS22, CWZ24, DGT20, DGWZ23, DLW24, HH20, HKY23, KM24, LLW24, LWC22, LXZ25, RS24, WZ24, XCV24, Zha22]. Noisy [BN24, CM20a, HW24]. Non [BCC^{+}22, BDOZ21, BST22, CDG23, Den20, Den21, DSU24, GKK22, GL23c, Liu24]. Non-Autonomous [CDG23]. Non-Cutoff [DSU24, Den20, Den21]. Non-Hamiltonian [BCC^{+}22]. Non-Lipschitz [GKK22]. Non-mean-field [BDOZ21]. Non-Self-Adjoint [Liu24]. Non-Symmetric [BST22]. Non-Uniqueness [GL23c]. Nonautonomous [WWZ23a]. Noncentral [LSW23]. Noncompact [PS22]. Nonconservative [ADBP21]. Nonconvergence [JKL24]. Nonconvex [MR24]. Noncooperative [Gir25]. Noncutoff [CHJ24]. NonDegeneracy [FS20]. Nondifferentiable [ED22]. Nondiffusive [CMX24, CMX22]. Nondivergence [DK21b]. Nondominated [CC22]. Nonequilibrium [CH21b, JJL21]. Nonequilibrium-Diffusion [JJL21]. Nonexistence [CP24, GJ21, GL21]. Nonexpanding [FP21]. NonGradient [ADdR22]. Nonhexagonal [LRW20]. Nonhonest [Bob23]. Noninteracting [KKNN22]. Nonisentropic [CHLW25, LY23, Lai20]. Nonlinear [AM22, AKV20, BSW23, BR23, BJL24, BN24, BMZ23, BNP22, BSH24, BG24b, BK23a, BW24, BGMW20, BK23b, BLS22, CS24, CHLV25, CEW21, CWZ24, CGZ20, CHL23, CJL23, CFP25, CJK24, Com22, CDG23, COS24, CCLM23, CMMS24, DU20, DGR21, DK21a, Don25, DLW24, ES24, FX21, FSZ24, Gar24, GMS21c, GK21a, GRHSZ20, GKLP21, GGL24, GK21b, GSGM24, GL20, HLS24, HZ25, HPS23, HJ24, HNX25, IK22, KL21, KN22, KMY20, KKK20, Kil22, LUY21, LLZ24, LL22, LG21, MMS21, PSW25, PW25, QRW24, RS25, SWX25, SW20b, TW23, UZ23, Wan24, YZ25, YY24, ZZZ24, ZZ21b, LP23, Bad25, DGB23]. Nonlinearities [BDF24, GKK22, LLS20, PS22]. Nonlinearity [AC21, CKKP20, DQ23, ES24, Fen25, IMO20, KMY20, Pen23, Tes20]. Nonlinearly [AKKR23]. Nonlocal [ALM23, ACC22, ABR22, ABLW24, ADK23, ADJS24, BCFP23, BP23, BS24b, CKP24, CJK24, CKP20, DQ23, DbMS24, DDG24, DT24, DIS25, DFGV25, DZZ23, DR24, DN22, DNS25, DG20, EB23, FLRX24, GDS24, GHPZ24, HPS23, HKSW24, HR20, KLM24, KKK20, KLS23, KT22b, KV20, MW23, Mou20, RS23a, RvW20, SD24, SD25, Sol24, XC22, ZiDRdW24, CL22b, Cla23, DR20a]. Nonlocal-Interaction [Sol24]. Nonmonotone [BK24]. Nonmonotonicity [Liu24, FP22]. Nonnegative [DLQ21]. Nonorientable [BK21]. Nonperiodic [GL24b]. Nonpower [FLNV24]. Nonquadratic [HPS23]. Nonradial [BDF24, DK21a, SWX25]. Nonrelativistic [BLZ24]. NonResistive [LS21b]. Nonreversible [HNS20]. Nonscattering [CHL23, VX21, VX25]. Nonsmooth [BGT22, FFV20, KR21]. Nonsonic [FHM^{+}24a]. Nonstationary [KLLT25]. NonStrictly [FP20]. Nontrivial [CHX24b, LWC22]. Nonunique [Dai21b, AW21]. Nonuniqueness [AKK^{+}20, BCFP23, CDZ24, MS24d, Nov20, PS23a, Yam22]. Nonvanishing [WYZ23]. Nordström [DLY21]. Normal [FXX20, HBK24, PYZ24]. Normalized [CiDRdTY23, CT23, GY24, HLS24, PVY25, YCRS22]. Norton [Bel22]. Novel [CDLZ21, Kat23]. Novikov [AG23]. Nowak [TW21]. Nowhere [Rös21]. Nozzle [CHLW25, FG22, LS23a, Tsu22, WXX21]. Nozzles [MWX23, NWG24, Wan22, XY21]. Nuclear [EFK20]. Nucleation [MM23]. Null [BGK21]. Null-Spaces [BGK21]. Number [AK23b, AHM20, BY24, Gal23, GK23, JJL21, MWX23, Ou21, QZ22, ZZ23a]. Numbers [VX21]. Numerical [CFS21, CWSGS20, LZ23b]. Numerics [CCE24]. Nutrients [KL23].

Oblique [CD21]. Observability [AT20, LW21]. Observational [CM20a]. Obstacle [CH23, Li22a, MWX23, NT25, TW25, Tak21, Yos21, ZRW22, ZiDRdW24]. Obstacles [CFS23]. Obtained [VX25]. Ocean [KT24]. Odd [FV25]. ODE [BCD24, EOE20]. Oldroyd [LLW21, Pan24, WW24]. Oldroyd-B [LLW21, Pan24, WW24]. Oleinik [PS23b]. One [AG24, BCF24, Bl23, BH22b, CNW21, CL22b, DS21a, Dod21, Fei25, GL24a, HNNW23, HJK25, HK23, IM21b, JR20, KL21, KvM21, LZZ21, LP23, LWZ23, LLS20, LG21, LZ24a, MSM24, NR24, RRV22, TX24, WZGX25, Zha21, vMP24, CLZ22, Ou21]. One-Cylinder [Zha21]. One-Dimensional [AG24, Bl23, CNW21, CL22b, Dod21, GL24a, HJK25, HK23, JR20, LZZ21, LP23, LLS20, MSM24, NR24, LG21, CLZ22, Ou21]. One-forms [IM21b]. One-Phase [RRV22, DS21a]. One-Species [WZGX25]. Only [LWZ23]. Ono [Gas23, HZ22a, Wri24]. Onsager [LL21]. Onset [BG24b]. Open [CNT23]. Operator [AKLM22, BK21, BHL23, BKS23, Den20, GMS21a, Ger24, GZZ22, Jia20b, LS20, LPS22, LP23, LLPZ21, MRB22, SH22]. Operators [AV24b, BS24a, BLN24a, BS24b, CGG23, CDL22, CKP24, DH24, DK21b, DR24, Duc23, EIN22, JL21, KPS20, LYZ24, Liu24, LWW20, RRTT24, Sel24, SS22b, Yan21a, Zhu21]. Optical [Mef21, Mef24, PP24]. Optics [BLPS22, ES24, HKSW24, Kil22, QWYY25]. Optimal [AK23a, ACNT24, ANWS22, ACS23, BCR21, BV20a, BGS22, BCDM21, CCS24, CC25, CDL22, CGK23, Cio21, CGSS24, CDM24b, DG23, EN22, EB23, GKM20, GSGM24, JJZ24, JM23, KLSW24, KLW22b, LS23c, MRB22, MN24b, MR24, PVJ21, PR25, SS24, Shu20, ST24, WWZ23b, Web20, WZZ20, XX23a]. Optimality [KUW21]. Optimization [BHM24, BGL23, CHQ22, Har23]. Optimize [Li22a]. Optimized [Cla23]. Optimizers [MPV23]. Orbital [DU20, HJ24]. Orbits [BDF25]. Order [BLN24a, BST22, BLN24b, Cai23, CGG23, CPW24, CCS22, CGSS24, DQ23, Fep21, FJ22, GY24, HJK25, HZ22b, HL22b, IMO20, JLN23, JLX23, MSM24, NY24, NS21b, NP20, OS25, Pen23, RRTT24, RW23, SW20a, YYZ24, KM20]. Ordinary [YGD21]. Organized [CDS20]. Oscillating [BD22, CGG23, GL24b]. Oscillations [HH21, Hal22, HR21]. Oscillatory [GZ20a]. Oseen [Eit22]. Oseen-Type [Eit22]. Osher [MN24b, BV20b]. Ostrovsky [BJP24]. Outflow [Wan23, WXZ24]. Overdamping [JM23].

P1 [JJL21]. Packets [CM20a, DST23]. Painlevé [HZ22b, LLhW23]. Pantograph [LWC22]. Parabolic [AK23a, AMS24, BS24a, BDD^{+}20, BDGdN24, CWSGS20, CDL22, CKP24, CJK24, CBHS23, DZZ23, DR24, Dun21, EM22, HSZ20, IMO20, JLN23, KL21, KN22, KbP21, KNP21, KV22, KS21, KS24a, LLR20, Li21, LLPZ21, LPW22, LNRV21, Miz20, Zha21]. Parabolic-parabolic [Miz20]. Parameter [LL20, ND20]. Parameterization [YGD21]. Parameters [WZ20b]. Parametric [DH24]. Parametrization [KT24]. Parity [ADH^{+}22]. Part [Cia20, DN22, WZ20b]. Partial [BKS23, CGU23, DLZ25, GWZ21, HKY23, HKSW24, HJ24, IM21b, LLZ24, LSY24, Ngu22, OS25, Sel24, XC22, XCV24]. Partially [HWZ22, RS20b, TSMO24]. Particle [AK23b, CJ21, CHQ22, CDF^{+}22, FOS20, FW21, IMS22, KvM21, LXZ23, RS23a, Xie25]. Particle-Based [IMS22]. Particles [BBES22, DIS25, gLZ25, vM22]. Partition [AdSST24]. Passage [CCS22]. Passing [CL21d]. Passive [Fei25, IV22]. Past [CKZ21, CXZ22, CH23, DL24b, HLZ24, Eit22]. Pasta [HJM25]. Patch [CF25, HW22a, Hu22]. Patches [AHK24, CW22, KL23]. Path [RR20]. Path-Dependent [RR20]. Patlak [CW24, GH24a, GH21a]. Pattern [BH23, CGSS24, DR20a, GHPZ24, HL20b, LT24]. Patterns [BHL23, KW23]. Pauli [LP20]. Paveri [CY21b]. Payoff [XX23a]. PDE [CEK21, CCE24, EOE20, Ger21, KT22b, KV21, KRS21]. PDEs [BLS22, CS24, IK22, KT22a, Per20, RR20, WdlL20]. Peaked [DGMP20, MP21a]. Peakons [LP22]. Pearcey [YZ24b]. Pedestrian [AG24]. Peierls [vMP24]. Pekar [FS20]. Penalization [DMS22b, DLW22]. Penalty [GDS24]. Pencil [BP20]. Pencil-Beam [BP20]. Pennes [BSH24]. Penrose [GNR22, HNX25]. Penrose-Stable [HNX25]. Perfect [Li20, Oli21]. Perfectly [GLS23, JW23]. Perfectly-Matched-Layer [GLS23]. Perforated [Jin20]. Perimeter [DLW22, FL21]. Perimeter-to-Area [DLW22]. Perimeters [MW23]. Periodic [AHK24, AV24b, BDF21, BP23, CGG23, CT21, Del23, DM20a, Eit22, FLRX24, Fep21, GHM23, GZ20a, Gir25, GZZ22, HWZ22, JQSY21, JJZ22, JLZ25, JP20, KMW20, KN23, KZ21a, LY22, LLPZ21, MP21a, MS22, MP21b, Nei20, NS21b, RS21, Sen23, Set24, Ste25a, YY24, YY22, Yua23, ZQ24, ZZ21a, Zha21]. Periodically [WdlL20]. Periodicity [BL20]. Permanent [HY22a]. Permittivity [GH21b]. Perona [GP24]. Persistence [FXX20, Gir25]. Perspective [LWZZ22]. Perturbated [Gao21]. Perturbation [CDL24, JKZ21, YGD21, YCRS22, ZZZ22b]. Perturbations [Cov22, GH24b, Han20, MP21a, RS20a, VX25, YY22, Yua23]. Perturbative [AK23a]. Perturbed [AO24, BDF21, CK23a, CKZ21, DMS22b, HZ25, MMR^{+}23, WXX21]. Peskin [CN23]. PET [MSW24a]. Petviashvilii [EGN22]. Phan [CLYaY23]. Phan-Thien [CLYaY23]. Phase [BBBH25, BG24a, BZGZ24, BC25, CHXZ21, CFG25, FLPS24, GM20a, GSS23, GL23c, HM25, KLSW24, MSW24b, RR25, RRV22, SSZ20, WZZ20, Yan25, ZRW22, ZiDRdW24, DS21a]. Phase-Field [BZGZ24, KLSW24, MSW24b, RR25]. Phase-homogeneous [CHXZ21]. Phenomena [BF23, FMN21]. Phenomenon [ADK23, LWZ21a]. Phi [Eva21]. Phi-Entropy [Eva21]. Photon [BIL^{+}23, SZ22]. Physical [CHaW24, CLWX20, Cla23, DLX21, Zen23]. Physics [BG24a]. Piecewise [CDX22, GN23, KV20, XX23b]. Piezoelectricity [XX23b]. Pinning [BPR24]. Piston [Din23, HK23]. Pitaevskii [AHLW21, BDS24, BLN24b, GZ22, LW20b]. Placement [Li22a]. Plan [Shu20]. Planar [BBG^{+}24, CYY23, DK21c, MR24, NP20, WWZ21, Wan23, Yua23]. Planck [GWZ21, IS21, ZZ24, AIN25, BP20, BR23, Bla23, DY24, DLY21, FMP22, HY22a, JQSY21, Laf20, LWZZ22, QRW24, Zhu21]. Plane [BIFL20, CF22, Chi22, Hog21, MPZ22, PU24, VX21]. Planning [MSM24]. Plasma [Bad25, CHWY25, GNR22, RWZ21, Web20]. Plasmas [Pan21]. Plasmonic [Ngu22]. plastic [CR21]. Plate [Gah25, MS22, SS22a, TW23]. Plates [BK23b, IOPR21]. Poincaré [AKLM22, LPS22, WKM23, YL22]. Point [CCFO21, GNNCR20, LN20, Mar22, PW25, Ros20]. Points [ADH^{+}22, ACL22, CW25, DFV25, GZZ22, LW21, Tio23]. Poiseuille [CHZ24, CW24, Del23]. Poisson [Bla23, CDL24, GZW21, HLlZ23, LW23a, BMM22, BPIR25, BRR25, CNT23, CbGwG24, CCZ22, CMZZ22, CHWY25, DM20b, DYZ21, GZW21, GWZ21, HY22a, HY22b, HNX25, KR21, LW23a, LWX24, Liu23, Pan23, PL23, Reg21, Saf20, Sun21, WW23, XMN23, YZ24a, ZZZ22b]. Polariton [DLL20]. Polaritons [CH21b]. Polarization [LNRV21]. Polarized [BJL24]. Polaron [CCFO21]. Poles [LLhW23]. Pollution [QWYY25]. Polyatomic [AČ24, HRY22]. Polyconvex [Tio23]. Polygon [JTY23]. Polygonal [BFV21]. Polygons [FMMS22, HMS24, MS20a]. Polyharmonic [BKS23]. Polyhedral [ABFV22]. Polymer [FSW24, GK23]. Polymeric [CLYaY23, DS23]. Polynomial [CDL24, KPS20, KT22b]. Polynomially [HW22b]. Polynomials [AHLW21, JL21, LW20b]. Population [FS22, Fue20, KW23, LS21a, LT22, PSW25]. Poroelastic [Gah25]. Poroelasticity [BCMW21]. Poromechanics [BC25]. Porous [ALW23, CGU23, DGT20, DIS25, EK20, Fep21, GMN22, GK21a, GBS20, HL21, KM23, LKPR20, RvW20, RS23b, WZ24]. Posed [WdlL20]. Posedness [ALW23, AG23, AV24a, AGL24, AMS24, BCK25, BT23, BSH24, BHI24, BBES22, CW21, CGS23, CGSS24, COS24, DGV22, DST20, DGWZ23, DGOY24, DKLT25, FV25, GGL24, HW22b, HHS23, HK23, HL22a, JWX22, KT24, LWZ21b, LR21, LW23b, MS20d, PYZ24, Pec20, TW22b, Wan21, XZZ24, YYZ24, YY24, Zha23, Boc20, HH21, LY21, WY21, YYZ21]. Positions [FG22]. Positive [Bl23, KMY20, Li23, QZ22]. Positivity [LX24]. Possibly [CGU23, WdlL20]. Potential [BGK21, Bar23, BRR25, CFHX20, Den20, Den21, DIS25, DFGV25, FXX20, HLlZ23, HW22b, HZ22a, JL21, KV22, LN20, Li21, LLW22, LW23a, LW24, LLX21, LSY24, MP21b, RS21, SW20a, TX24, Tet22, WY24, YZZ22]. Potentials [AB21, BCDM21, BY23, CGG23, CHJ24, CP24, Cha23, Duc23, GMS23, GL24b, GZZ22, HL20a, KMM24, KL22, Li22b, LYZ24, Sel24, SS22b, Sne20, SHM25, WO23, WZGX25]. Prandtl [BY24, DLX21, LYZ23, WZ20a, XZZ24].

Predator [Fue20, HNNW23, Li23, TW22a]. Predator- [Fue20]. Predator-Prey [TW22a]. Prepared [GHM22, OY22]. Presence [LZ20, TT24]. Pressure [AB21, Bar20, CDLLSG20, Don25, DJ20, WZ20a]. Pressureless [HY22b]. Preventing [BCJ20, HJ11]. Prevention [KS21, KS24a]. Prey [Fue20, HNNW23, Li23, TW22a]. Prey-Predator [Li23]. Prey-Taxis [Fue20, Li23]. Primitive [KT24]. Principal [FLRX24, LLPZ21, Liu24, MPV23]. Principle [CL23, Gao21, GL23a, Ham25, HLL21, KMW20, MRB22, SS20]. Principles [DMRS20, GIT20, SD24]. Priori [MW20, LZ21]. Pritchard [LLWZ25]. Probabilistic [WO23]. Probabilities [BHS25]. Probability [GMŚ21b, JJZ24, JQSY21, WKM23]. Problem [AC21, AK23a, AKK^{+}20, ABLW24, AB21, BCR25, BS24a, Bar20, BHL23, BKS23, BZGZ24, BO21, BDF21, BDLL25, BGL23, CGU23, CFHX20, CL21b, CHS22, CCS22, CN23, CJL23, CH21b, CDG23, Cov22, CRTZ24, CF20, Din23, Don25, DLW22, DLW24, DYZ21, DG23, EFHS21, Eit22, EZ20, FMW23, GM20a, GH24b, GV24, GK21a, GKLP21, GZ20b, GN23, Ham25, Har23, HXH22, HC21, HK23, HLZ24, HL22a, IY20, Jin20, Jui20, JL21, KMY20, KUW21, KLW22b, KRS21, LLR20, LNOY24, Li20, Li21, LWZ21b, LZ23a, Lia22, Lia24, LLW21, LY23, LSY24, LNRV21, LLhW23, LZ24b, MN24a, MRB22, Mef21, Mef24, MSM24, MW20, NWG24, NT25, Ou21, PVJ21, PZ23, PST23, PR25, Ric23, RRV22, Rül21, SSZ20, SW22c, SZ22, Tak21, Taw24, TW23, UZ23, UW24, Wan20, Wan23, Web20, WXX21, YL22, Zen23]. Problems [AO24, ACL22, AdSST24, ANPS21, BL21, CDLZ21, CNW20, CDLLSG20, CP24, CHZ24, CCM21, CC23, CDL22, CT23, DS21a, EI20, EFP^{+}24, FHS24, FLS25, GMS23, GZ20a, HZ25, HL20b, IK22, KL21, KKK20, KM20, Kil22, KMM24, KV21, LLZ24, MPV23, MP20b, Pru24, RRTT24, SD24, SD25, Sel24, Sen23, ST24, Ste20, WW23, Yan21a, ZRW22, ZiDRdW24]. Process [CHLV25]. Processes [Bob23, CDX21, CCKW21, KMW20]. Product [YYZ21, ZZ21b]. Profile [CXZ22, CHWY25, Kil22]. Profiles [FP22, KMY20, MS24b, PSW25]. Project [FFV21]. Proof [FLS20, Sai23]. Proofs [Nem21]. Propagating [KMY20, PTE25]. Propagation [ADK23, AČ24, BJ23, BJL24, CHJ24, DT24, GK25, HRZ24, HWZ22, KZ21a, LS23b, LLN20, LLW20, Reg21, WZ23]. Proper [BWF24]. Properties [BCR25, BCF24, DT24, DMS22b, FSZ24, Ger24, Gir25, GP24, Mon20, RRTT24]. Property [CY22, HKY23, LLS25, MRB22, Ngu22]. Protein [GSS23]. Prototypical [FP20]. Proudman [FL24]. Pseudo [AMS24]. Pseudo-Parabolic [AMS24]. Pseudomodes [Duc23]. Pseudospectra [AS20]. Pullback [KM24]. Pulled [AS21, Gar24]. Pulse [CRS21]. Pulses [MV23]. Pure [IPTT22, PST23]. Purely [BMPS21]. Pursuit [EB23]. Pushed [vM22]. Pyragas [dW23].

Qian [LW20a]. Quadrant [IK22]. Quadratic [BG24a, BGMW20, CW25, CKW20, DIS25, GMS21c]. Quadrature [ZHY24]. Qualitative [TW22a]. Quantifying [ADK23, IZ23]. Quantitative [AKKY23, Bar23, Bla23, EN22, HNS20, HS24, JR21, KvM21, LLN20, LXZ25, MRB22, MPZ22, Sim21]. Quantities [HKLT24b]. Quantum [AHM20, BY20, BT23, CKP^{+}22, FSW22, HKSW24, LNOY24]. Quasi [BLN24a, CLMZ23, EIN22, FFV21, GHM23, Hu24a, JKL24, LN20, LPZ23, LPW22, Nov20, PL23, Ros20, Sai23, WdlL20, ZZ24]. Quasi-Convex [EIN22].

Quasi-Crystalline [FFV21].

Quasi-Geostrophic

[GHM23, Hu24a, JKL24, Nov20, Ros20]. Quasi-Linear [LPZ23, LPW22, BLN24a]. Quasi-Linearity [Sai23]. Quasi-Neutral [CLMZ23, PL23, ZZ24]. Quasi-Periodically [WdlL20]. Quasi-stationary [LN20]. Quasiconvexity [KV21]. Quasilinear [DM21, EFP^{+}24, JLN23, KN22, KbP21, Kil22, LT24]. Quasilocal [Cov22]. Quasineutral [DM20b]. Quasiperiodic [JLZ25]. Queueing [BK20b]. Queues [AKKR23]. Quintic [CCJT23, CGZ20, KMV21, RS25, YY24].

R2 [Kat23]. R3 [FMW23]. Radial [DN22, LZ22, TT22, Tan21, XMN23]. Radially [CCW23]. Radiating [JZZ22]. Radiation [CCZ22, JV24a, JJL21, Wan21, WZZ24, ZZ23b]. Radiationless [BL21]. Radiative [GHM22, RZZ21]. Radon [Kat20, Kat23, WHQ25]. Ramified [XX23a]. Random [BCR21, BH23, CTL22, CCi23, CWZ24, Gao21, HJJ20, HNNW23, HT23, HWXZ25, LLW22, LW24, MT24, WKM23]. Randomized [RR20]. Range [BBC20, BM24, MS21, ST20a]. Rank [DS22]. Rank-Deficient [DS22]. Rao [HBK24]. Rapidly [BO24]. Rarefaction [HWWW24, WWZ21, YZ21, ZZZ22b]. Rascle [CHLL22]. Rate [BBBH25, BMPS21, CHJ24, CR21, CLR24, IV22, Liu24, Wan23, WWZ23b]. Rate-Independent [CR21, CLR24]. Rate-Type [BMPS21]. Rates [ADH23, BLZ24, Bar23, CLYaY23, CC22, DK23, FLS20, HL23b, JM23, LS23c, MWX23, RW23, Wan21]. Rating [BDLL25]. Ratio [DLW22, ZZ21a]. Ratios [CL25]. Ray [BKS23, Mon20, IM21b, MS23a, Wan24]. Rayleigh [GLY20, KMW24, LZ23a]. Reacting [LS23a]. Reaction [AV24a, CMT21, DM20a, EMT20, FHKM22, FMTY21, GR23, HH20, HKlT24a, HWZ22, IMS22, Jac23, KM23, LLN20, MV23, MS24b, MSY23, PSW25, RH20, ZZ21a, dN23]. Real [GNNCR20, Jui20, Nem21]. Receptor [ATZ24]. Receptor-Based [ATZ24]. Reconstructing [ACL22]. Reconstruction [CDGH20, Fei25, GH24b, GV24, HW24, JKZ21, LL20, LUY21, MSW24a, Mef21, Mef24, SL22]. Recovering [DH21, XX23b]. Recovery [ES24, KST23, Tet22]. Recursive [FP22]. Redi [KT24]. Reduced [BLN24b, BK23b]. Reduction [Nei20]. Reflecting [CDF^{+}22]. Reflection [DGOY24, HW22c, JLS25, KSS23, SL22]. Reformulated [KPS20]. Regime [BLN24b, HJJ20]. Regimes [FJ22]. Regular [Bl23, CLZ22, DK21c, FF25]. Regularity [Abd24, AG23, AHK24, AL25, BK20a, BDGdN24, CTL22, CMP25, CEK21, CMX22, CMX24, CH21a, CK23b, CFHX20, CCS22, Che22, CLLZ23, CHL23, CDL22, CKP24, CD20, DK23, Dec24, Den21, DW20, DGWZ23, DMS22c, DG25, DG23, FMMS22, FK25, FFV20, GLW22, GHK20, GJP24, Gol23, GWZ21, HMS24, HJLZ21, KT20, KT22b, Kwe23, LLX21, LZ24a, MS20a, MPZ22, MSM24, MR24, NS21a, Pec20, Ran23, Reg21, SD25, Sel24, Sun21, TT22, WHY25, Wic23, XZZ24, ZZ23b, Zhu21, CCM21]. Regularization [ADdR22, CEGP23, GGL24, Mon20, OS25, WT24]. Regularized [CS23, EN22]. Regularizing [GJ21, Jia20b]. Regulated [JR20]. Related [BJP24, BO21, CJL23, GMS23, LS20]. Relation [CKP20]. Relations [GH21b, Mon20]. Relative [FLS20, HL23b]. Relativistic [BDF21, Din23, DL24b, DGOY24, HY21, HRY22, KR21, LS23b, LLW20, Oli21]. Relativity [HM25]. Relaxation [BHMCR24, BSH24, CHLL22, CBHS23, Eva21, FHM^{+}24a, FHM^{+}24b, HR21, KZ21b, RS20b]. ReLU [JLL^{+}23]. ReLU-Sine-Exponential [JLL^{+}23].

Remark [Yos21]. Remodeling [DL22a]. Reneging [AKKR23]. Renormalization [GL24a]. Renormalized [MS24c]. Repartition [LN20]. Repelling [vM22]. Replication [CRS21]. Representation [CFV20, KLM24]. Representations [JMQ23]. Reproduction [ZZ21a]. Repulsive [BRR25, BCT24, DIS25]. Requirements [Fos21]. Resolution [Kat20, Kat23]. Resolvent [AS21, Eit22, dHKLN24]. Resonance [DLL20, KMM24]. Resonances [HKSW24, LYZ24]. Resonating [MS23b]. Resonator [ADH23]. Resonators [CK23a]. Respect [EYZ22]. Response [BMPS21, CHLV25, WdlL20]. Restricted [COS24]. Result [ARS21, BC25, BCD24, CDLLSG20, DL22a, FW21]. Results [Bla23, BGT22, CS24, CFS21, DM21, DMS22c, Gen25, Gol23, HLS24, HNS20, KT20, VX21, VX25]. Retardation [Pan24]. Retrieval [GL23c]. Retrospective [KA24]. Revisited [DMRS20, HZ22a, KLSZ22, Kre21, SW22b, SW22c]. Revolution [WHQ25]. Rheology [GL23b]. Richardson [ZZ23a]. Riemann [AKK^{+}20, DR20b, ED22, FR23, KB23]. Riemannian [CDM24b, JJZ24]. Riesz [CL25]. Right [NT25]. Right-Angle [NT25]. Rigid [BNSS23, CGG20, LXZ23, RSS21, gLZ25]. Rigidity [CDM24b, Gen25, MT24, NS20b, Sim21]. Rigorous [AMP22, DST23, Eme21, LW20a, LKPR20, LWZZ22, RR25, TT24]. Ring [CKW20]. Rings [AHLW21, BCM25]. RN [CT23]. RNA [GSS23]. RNA-Protein [GSS23]. Robin [AM22, BGT22, BK22, GZ20a]. Robin-Laplacian [BGT22]. Robustness [CCE24, CWZ24, WKM23]. Roll [Ste25a]. Rolling [ALV23]. Root [DG25, HBK24]. Root-Lipschitz [DG25]. Roots [BGS22, KT22b]. Rotating [Eit22, GLP23, JSW23, JKL24]. Rotational [BCDH25, BY23, Hei23, Lai20]. Rough [AIN25, GG23, HWXZ25, Kat23, LLW24, NK21, ST24, WZ23, Zhu21]. Rudin [MN24b]. Run [EY23].

Saint [KL22]. Saint-Venant [KL22]. Sakaguchi [CHXZ21, HSZ20]. Sampled [GL23c]. Sampler [DL21]. Sampling [MS23a, Ste20]. Santaló [FL21]. Savart [Ger24]. Saxton [GLW22, GLW23]. Scalar [Abd24, Bar20, BGK21, DMS22c, DR20b, GJ21, HW22a, JR20, KV20, PW25, RS23a, Yua23]. Scale [FFV21, KMW24, KT24, SX20b]. Scaled [DCWF^{+}22]. Scales [GY25]. Scaling [BCM25, GKM20, RRTT24]. Scattered [KN23]. Scatterer [RSS21]. Scatterers [HSVW25]. Scattering [BDF24, BRR25, BL21, CFS23, CEW21, CGS23, CGZ20, CHL23, CM20b, HSVW25, HWXZ25, KMV21, KSS24, Kre21, LLW22, LW24, LYZ24, LYZZ24, LSW23, MPMS23, NS20b, Pan23, RS20a, Tes20, YYZ24]. Schauder [CL22b]. Scheme [BV20b, GP24, QWYY25]. Schnirelmann [CT23]. Schödinger [HL20a]. Schrödinger [LYZ24, AKV20, AV24b, BSW23, BCF24, BJL24, BHL23, CGG23, CL20, CF22, CiDRdTY23, CLLZ23, CGZ20, CJL23, Com22, CCLM23, CS23, DLQ21, DU20, DK21a, FX21, FSZ24, GMS23, Gib24, GZZ22, HLS24, HJ24, LLZ24, LP20, LL22, LP23, LY22, LM21, MRB22, MP21b, PVY25, RS25, SWX25, Sel24, SS22b, SW20b, Tet22, YYZ24, YY24, ZZ21b]. Schrödinger-Type [LYZ24]. Schwarz [Cla23]. Schwarzschild [BDF25]. Science [AMS24]. SDE [CDF^{+}22]. SDEs [HRZ24, QRW24]. Second [BLN24b, CHLV25, CGG23, Har20, JLX23, LX25, NY24, NP20, OS25, PYZ24, CL23]. Second-Grade [PYZ24]. Second-Harmonic [CHLV25].

Second-Order [CGG23, JLX23, NY24, OS25]. Sectional [KL22]. Sector [CL24]. Segel [BCJ20, CW24, GH24a, GH21a, CKW20, CBHS23, HJ11, MP20a, Miz20, Win20]. Selection [GL23a]. Self [CMVP23, CDS20, DG20, JLS25, Liu24, MS24b, WJW22]. Self-Organized [CDS20]. Self-Similar [CMVP23, DG20, JLS25, MS24b, WJW22]. Semi [Lai20]. Semi-Infinite [Lai20]. Semiclassical [Ben24, GZ22, LS23b, Ste20]. Semiconcave [EZ20]. Semiconductor [CMZZ22, XMN23]. Semiconductors [CLMZ23, FHM^{+}24a, FHM^{+}24b, WMZZ21]. Semidefinite [Har23]. Semidiscrete [ANWS22, CDM24a, GP24, BV20b]. Semigeostrophic [Sil23]. Semigroup [Den20, Ger21]. Semigroups [FZ21]. Semilinear [HS23, RZ21]. Seminorms [DPS24]. Semipermeable [Bob20]. Semisupervised [WT24]. Sensing [LT24]. Sensitivity [ACL22, Cia20, HJJ20, KS21, KS24a]. Separability [RZZ21]. Separated [Bob20]. Separation [CEGP23, CFG25]. Sequential [Dod21]. Series [JL21]. Server [AKKR23]. Sets [BWF24, Fos21, IM21a]. Setting [KS21, KS24a]. Several [GJ21, PL23]. Shallow [BCDH25, Boc20, Eme21, Ste25a]. Sham [SCB20, SCB17]. Shape [ACNT24, DH24, Sim21]. Shaped [Liu23]. Shapes [BGS22, JTY23, Pru24]. Sharp [AO24, AF23, ACC22, CHJ24, CLYaY23, CDZ24, COS24, GHPZ24, Har20, HL23b, Jia20b, LPW21, Mon20, RR25]. Shaw [Chu23]. Shear [Coo23, GG23, Hog21, Jia20a, WWZ23b, Zho22]. Shearing [TSMO24]. Sheath [Bad25]. Sheet [BIFL20]. Shells [ABLW24, MM23]. Sheng [LW20a]. Shigesada [Mou20]. Shock [CFHX20, CMVP23, Din23, FXX20, HK23, HWWW24, HL22b, JLS25, LG21, PX22, WXX21, Yan21b]. Shocks [CKZ21, DR20b, FG22, FR23, GLY20, HKK23, YY22, Yua23].

Short [BM24, FLS20, ST20a]. Short-Range [BM24, ST20a]. Shortening [MB20]. Sign [BC24b, GIT20]. Sign-Changing [BC24b]. Signature [EFP^{+}24]. Signorini [HS23]. Silent [BHG25]. Similar [CMVP23, DG20, JLS25, MS24b, WJW22]. Simple [CJL21, MS23a]. Simplified [CaDZ24, FGW21, Zho22]. Sine [JLL^{+}23]. Single [BN24, HL20b, Rül21]. Singular [BSH24, CGU23, CL22b, CJ24, FF25, GMS23, Han20, HRZ24, HR20, HY22a, Hu24a, KL21, KS21, KS24a, Lia22, LLX21, MPV23, Pin21, RS23a, SX20b, Tak25, TW21]. Singularities [BW24, CCZ21, DK21c, JRS24]. Singularity [CCW23, CHX24a, Hu24b]. Singularly [AO24, HZ25]. Sinkhorn [EN22]. Sixth [CGSS24, RW23, SW20a]. Sixth-Order [CGSS24, SW20a]. Size [BCR21, Xie25]. Skeleton [FHS24]. Skew [AT20]. Skew-Adjoint [AT20]. SKT [KS24b]. Slightly [PST23]. Slip [BIFL20, Ngu20, Taw24]. Slope [DPS24]. Slot [CD21]. Slow [CL21a, HR21, NY24, TT24]. Slow-Fast [HR21]. Slowly [HZ22a]. Small [BRR25, CK23a, GY25, HJLZ21, KO24, LLW24, RW23, Tes20, Win20]. Small-Mass [Win20]. Smallness [KZ21a]. Smoluchowski [FLNV24, FHP24]. Smooth [ADBP21, FMN21, GV21, GN23, HL22b, KV20, LSYZ21, MMSH21, Sil23, Wan22, WMZZ21]. Smoothed [FW21]. Smoothing [COS24, WY24, Zha22]. Smoothness [Hei23]. Sobolev [DFGV25, DR24, DG23, JWX22, LYZ23, MPZ22, MR24, PVY25, PS23a, YCRS22]. Soft [BG24a, Buz25, CHJ24, Den20, HW22b, Li22b, LW23a]. Solid [CGG20, SL22]. Solidification [CGG20]. Solids [DMS22a, RS23b]. Solitary [EGN22, IPTT22, LOWX20, MS20b, Sin21]. Soliton [CMMS24, HZ22a]. Solitons [CM23c, PP24]. Solution [AKK^{+}20, AL25, BN24, CHXZ21, DL24a, Dod21, Don22, Don25, GMS21a, GHW21, Gen25, Gib24, GZ20b, GZW21, HW22c, Jui20, MRR20, Tsu22, Wan23, WY24, WZ25, Zen23]. Solutions [Abd24, AW21, AGL24, AK23b, AG24, ADBP21, AIN25, BCK25, BSW23, BH22a, BT23, BPIR25, BG24a, BZGZ24, Bl23, BRR25, BDF21, BCGJ22, BT21, BGMW20, BDLL25, BFN20, BW25, BS24b, BPSS22, CGG20, CM23a, CJ21, CLZ22, CL22a, CMP25, CS22, CNT23, CbGwG24, CH21a, CP24, CF22, CFHX20, CKW20, CL21b, CEW21, CCZ21, CCZ22, CHS22, CiDRdTY23, CLYaY23, CHW24, CDZ24, CHWY25, CHX24b, CCM21, CHLL22, CHLW25, CT23, CD20, Com22, CKP20, CR21, CMMS24, DLQ21, Dai21b, DQ23, DS23, Dec24, DU20, DL22b, Din23, DK21a, DGMP20, DLW24, DLY21, DM22, DG23, EK20, FLNV24, FMW24, FP20, FMN21, Fue20, GMS23, GL24a, GLSW23, GLW22, GLW23, GL23a, GH24a, GM20a, GHM23, GK20, GLM20, GRZ24, GG24, GJP24, GKK22, Gin21, GLRS22, GP24, GWZ21, GK21b, GWX21, GY24]. Solutions [HH23, HKY23, HOV25, HMS24, Hei23, HMNW20, HR20, HY22a, HC21, HK23, Hu24a, HL22a, Hua24, HY21, IKL20, IS21, Jac23, JQSY21, JZZ22, JJZ22, KLM24, KR23, KMY20, KR21, KLW22a, KKNN22, KV20, LZ22, LS21b, Li22b, LPZ23, Li23, Lia22, LY22, LWZ23, LM21, LWC22, LX23, LY23, MS24a, MMN23, MN24a, MS24c, MS22, MSY23, MMS25, MPS21, NS21a, Nei20, Nov20, NS21b, Ou21, Pan24, Pan23, Pan20, Pen23, Per20, PVY25, PST23, QZ22, QX25, QWYY25, RS23a, RR20, RW23, SWX25, SD25, Sil23, ST21, Ste25a, Tak25, TX24, TR20, WdlL20, Wan21, WZ22, WJW22, WZZ24, Wic23, Win20, WZ20b, XZ24, XZZ24, YZZ22, YCRS22, ZLZ20, ZZZ24]. Solvability [DL22a, JWW23, LZ23b, Zho22]. Some [ALW23, Abd24, CFS21, Eme21, FP21, Liu24, MW23, SSZ20]. Sonic [CMZZ22, FHM^{+}24a, HC21, Hu22, Wan22, XMN23]. Sonic-Supersonic [Wan22, HC21]. Sorting [CC22]. Sound [MPMS23]. Source [CY21a, DMS22c, EI20, GLRS22, HL20b, IY20, JKZ21, KM20, KST23, KS21, KS24a, RZ20, SW22c]. Source-Destination [DMS22c]. Sources [BHG25, BL21]. Space [AW21, BS24a, BC24a, BHS25, BS24b, CM23c, Chi22, CK22, DCWF^{+}22, DYZ21, DM22, GJ21, Gir25, HW22c, JWX22, JMQ23, JKZ21, KLLT25, KLSW24, Kof23, Kor22, KRS21, LLR20, LL21, LY21, LWX24, LXZ25, MP20a, MN24a, MMS25, NZ24, PL23, RRV22, SUT22, TR20, WWZ21, WO23, WZZ20, WW23, YYZ21, ZZ21b]. Space-Dependent [BS24b, MN24a]. Space-Fractional [RRV22, TR20]. Space-Time [DM22, Gir25, KLSW24, LLR20, SUT22]. Space-Time-Dependent [JKZ21]. Space-Times [Kof23]. Spaces [BGK21, BWF24, BT21, CLLZ23, CKP24, DK23, DR24, GMŚ21b, GH24b, KT20, KM25, MT24, MMS21, Pec20, SD24, WJW22, CY21a]. Sparse [ACS23, DM24, RZ20]. Spatial [KNP21, KLSZ22, PSW25]. Spatially [HJK25]. Spatially-Dependent [HJK25]. Spatio [ADVW20]. Spatio-Temporal [ADVW20]. SPDEs [CL23, CL22b, HLL21, LXZ25]. Species [ALV23, DL22a, LRW20, WZGX25]. Spectra [DS22]. Spectral [AKLM22, AdSST24, BZ25, CGY21, CLLZ23, CCLM23, FLRX24, LP22]. Spectrum [AV24b, AS21, CK23a, Com22]. Specular [DGOY24]. Speed [GK25, LLW20]. Speeds [MPMS23]. Sphere [AMP22, Ngu22, Zha21]. Spherical [BMPS21, CHWY25, Dai21a]. Spherically [AK23b, KR21, Pan21]. Spin [AKV20, LLWZ24, LLWZ25, ZQ24]. Spin-1 [LLWZ24, LLWZ25]. Spiral [Dai21a, DS22, DK21c]. Split [LS21a]. Spray [Mén24]. Spreading [DNS25, Gir25, MSY23]. SQG [CF25]. Square [HBK24, YZZ22]. Squared [DN20]. Stability [ACC22, ACC^{+}21, ACC^{+}23, AS21, Bad25, BCK25, BFLP21, BN24, BHL23, BD22, BW24, BCT24, BDLL25, BFN20, BGT22, CM23c, CKZ21, CLYaY23, CGJL24, CM22, CHXZ21, CRTZ24, CCLM23, CMMS24, DU20, Din23, Don25, DR20b, EN22, EK20, EI20, FR23, GLY20, GR23, GH24b, Gar24, HLS24, Hal23, HH20, HTY21, HM25, HL20a, HXH22, HSSY24, HJ24, Hua24, JLN23, JZZ22, JWW23, JLX23, KM20, KL23, KA24, KW23, KB23, KUW21, KLW22b, KRS21, KV20, LZ22, LWZ21b, LG21, LPW22, LWC22, LM20, Ma22, MMN23, MMR^{+}23, MRB22, Mef21, Mef24, MS20b, Oli21, PP24, Pru24, Rül21, SW20b, Wan23, WXX21, WYZ23, XMN23, YZ21, YZ25, ZZZ22a, ZZ23a, ZZZ22b, ZHY24]. Stabilization [BLW20]. Stabilizing [AO23, LWZ21a, LZ23a]. Stable [ABFV22, BFV21, CCKW21, DR24, EFHS21, GNR22, HNX25, KST23, RSS21]. Stable-like [CCKW21]. Stacking [CT21]. Standing [CCLM23, HJ24]. Star [HY22b, KPS20]. Star-Even [KPS20]. Stark [CKP^{+}22]. Stars [JSW23]. State [AMMT22, BBD^{+}23, BK20b, BC24a, BLN24b, CJL21, CLMZ23, Dod21, KMW20, KST23, KTT20, Li23, Oli21, YGD21]. State-Constraint [KTT20]. State-Dependent [BK20b, KMW20, YGD21]. States [ATZ24, BLS22, CMZZ22, DGR21, FHM^{+}24a, FHM^{+}24b, Fue20, GV21, GLP23, HLS24, HM25, HZ25, HKSW24, KbP21, KS24b, LLWZ25, PWZZ24, PS22, Saf20, WMZZ21, WYZ23, XMN23]. Static [AK23b, CLLZ23, DLQ21, KTT20]. Statics [KV21, OP20]. Stationarity [KNP21].

Stationary [AGL24, BP20, BBG^{+}24, BO21, BY24, CKW20, CEW21, CCKW21, DGR21, DLZ25, GLY20, HL21, HY21, Kof23, OS25, Wan23, LN20]. Statistical [DM24, FMW24]. Steady [ATZ24, CWZ20, CW22, CKZ21, CMZZ22, CLMZ23, DLX21, DL24b, DW21, FXX20, FG22, FHM^{+}24a, FHM^{+}24b, Fue20, GY25, HM25, Hu22, KR23, KLW22a, KS24b, Li23, LYZ23, PX22, WMZZ21, XMN23]. Steady-State [CLMZ23, Li23]. Steady-States [CMZZ22, KS24b, XMN23]. Stefan [HJT22, Lia22, Lia24, RRV22]. Steklov [ABHN21, Hal23, LS20]. Stellar [HH21, Hal22]. Stellarator [GRR23]. Step [GM20b]. Step-like [GM20b]. Stewartson [EGN22, KSS23]. STFT [GL23c]. Stochastic [BK20a, BEM21, CCS23, CC25, CDZ24, CS23, DGT20, DM20b, DGWZ23, FX21, FOS20, FL24, GK25, GL20, HKY23, HNS20, HHS23, IMS22, JRS24, KM24, LLW24, LWC22, LW23b, MS24d, NY24, RWZZ20, Taw24, WKM23, WHY25, Wic23, WZ24, XC22, XCV24, Zha22]. Stokes [AF23, BFN20, BH22b, DM20b, Don25, Ger23, GZW21, GWZ21, LS23c, RvW20, Sun21, WW23, YZ21, ZZZ22b, AKKY23, AB21, AHM20, Bar23, BBG^{+}24, BP21, BCMW21, BT21, BO24, BFHZ22, CLZ22, CL22a, CK23b, CDZ24, CL21c, CDX22, CD20, CW24, DK23, Dec24, DL22b, DW20, Fep21, FJ22, FMW24, FOS20, GL24a, GGBS25, GH24a, GH21a, GWX21, HKK23, HL23a, HMS24, Hei23, HR23, HL22a, HWWW24, KLLT25, KLW22a, Kwe23, LPZ23, LXZ23, LX24, LX23, LY23, LZ24a, MS20a, Ngu20, Ou21, OY22, PZ23, SS22a, ST21, TW25, Tak25, TWZ20, VY23, WZ22, WJW22, WKM23, WZ23, WWZ23b, WZ25, Win20, WYZ23, Yam22, YY22, ZZZ22a, ZLZ20, ZZ23a, ZZ24, Zho22, gLZ25]. Stokes/Allen [AF23, HL23b]. Strain [BPSS22, CY22]. Strain-Limiting [BPSS22]. Strained [KP21]. Strains [RS23b]. Strange [Rös21]. Strategy [ALV23]. Stratified [JKL24, Sin21]. Stream [CD21]. Stress [AKKY23, BMPS21, CDM24a, KL22, LXZ23, LLW21, PYZ24, gLZ25]. Stress-Diffusion [BMPS21]. Stretching [HR20]. Strichartz [Zha22]. Strict [Ran23]. Strictly [Jui20, NS20b]. Strip [Del23, MMN23, WW22]. Strong [BHI24, Bla23, BFN20, CL22a, CHS22, CKP^{+}22, DS23, DM22, EM22, FSW22, FLS25, GMS21a, GLRS22, GZW21, GWX21, Ham25, HRZ24, HHS23, HJT22, KW23, KR21, MRR20, PX22, Saf20, SS22a, Tak25, WdlL20, Wan21, Wic23]. Strongly [DGB23, SX20a]. Structural [LY21, WXX21, XMN23]. Structure [AKLM22, BMM22, FXX20, GLW22, Ger21, HPS23, MRR20, Ngu22, Per20, Taw24]. Structured [FS22, FP21, PR20]. Structures [Boc20, CDLZ21, FMP22, LX23]. Study [CM23c, CK22, ED22, Gar24, JRS24, KT22a]. Subcritical [HLS24]. Subdiffusion [HJK25, JKZ21]. Subdomains [FMW23]. Subject [DIS25]. Sublinear [FZ21]. Subsonic [CXZ22, FHM^{+}24a, FHM^{+}24b, NWG24, PX22, XMN23]. Subwavelength [ADH^{+}22]. Sufficient [BFLP21]. Suitable [GWZ21]. Superconductors [KbP21]. Supercritical [CFG25, LL22, PST23]. Superlinear [AV24a, DL22b]. Superposition [YZ21]. Superquadratic [LY22]. Supersonic [CKZ21, DL24b, Hu22, Lai20, LS23a, Wan22, XY21, HC21]. Supersonic-Sonic [Hu22]. Support [AW21, HKY23]. Suppression [CW24, He23, Hu24b]. Sure [BN24]. Surface [BBC20, BHG25, BK21, CEGP23, DLL20, DS21b, Ger24, Gin21, KB23, LZ21, LZ24b, MW20, Pro20, Ros20, TW22b, WX21, WHQ25]. Surfaces [BGK21, CT21, CY22, GN23, HBK24, HWXZ25, MS23a, WHQ25, dN23]. Suspending [LXZ23]. Suspension [AO23].

Suspensions [GL23b, NS20a]. Swarm [CHQ22]. Sweeping [KMW20]. Swelling [EOE20, RS23b]. Swimming [AO23]. Switching [HNNW23, MS20c]. Symmetric [ADH^{+}22, AK23b, ABD22, BBG^{+}24, BST22, BY23, CCW23, CCKW21, DW21, GMS23, KR21, Pan21, RS20a, SWX25, Yos21]. Symmetry [BP23, BLS22, CHWY25, DMS22b, DN22, Nei20, TT22, Tan21, WWZ23b]. Synchronizing [Pro20]. System [AF23, AKK^{+}20, AČ24, AMP22, AK23b, AMS24, BN24, BPIR25, BRR25, BO21, BFHZ22, BFN20, BH22b, CCS24, CJ21, CL22a, CDL24, CL20, CMX22, CMX24, CCZ22, CHS22, CCJT23, CLYaY23, CHX24a, CHL23, CDF^{+}22, CBHS23, CR21, CW24, Dai21b, DbMS24, Dec24, DLL20, DL22b, Din23, DLZ25, DST20, DM20b, DGOY24, Don25, DKLT25, DLY21, DYZ21, EK20, FP22, FK25, FHM20, Fen25, Fep21, FFV20, FP20, FMN21, Gah25, GH24a, GK20, Gen25, GHM22, GLRS22, GH21a, GZW21, GWX21, HL23a, He23, HLlZ23, Hei23, HL23b, HS23, HY22a, HY22b, HL22a, HNX25, JZZ22, JWW23, JLX23, KbP21, KLLT25, KA24, KR21, LWZ21a, Lea23, Lew25, LS21b, LY21, LS23c, LW23a, LPZ23, LWZ23, LM21, LG21, LRW20, MP20a, MN24a, MS24c, Miz20, MSW24b, Pan23, Pec20, PW21, Pen23, PL23, QX25, Reg21, Saf20]. System [Sai23, SX20a, TT22, Tak25, TW22a, Tré24, Win20, WZGX25, XZZ24, XMN23, YZ24a, ZQ24, ZZ23a, ZZZ22b, ZZ24, Zho22]. Systems [ADdR22, AV24a, AT20, ACS23, BBC20, BCC^{+}22, BT23, BMZ23, BNP22, BDD^{+}20, BDGdN24, BDF25, BST22, BCGJ22, CCS23, CDM24a, CDX22, CMT21, DQ23, DR20a, DGR21, DJ20, DM22, EMT20, FR23, FLRX24, FHKM22, FMTY21, FOS20, GNNCR20, Gir25, GY24, HH20, HPS23, HR20, HR21, Hu24a, HWZ22, HJT22, Jac23, JMQ23, JV24b, KvM21, LZ20, Li23, MMSH21, MV23, MM24, MS24b, Mou20, Nei20, NO23, NY24, PP24, Per20, RH20, SWX25, TSMO24, WdlL20, WWZ23a, Wic23, WZ20b, Xie25, ZZ21a, dHKLN24]. Szego [ABD22, GG20].

Tails [AL23]. Tangential [WY21]. Tanner [CLYaY23]. Taxis [Fue20, JWW23, Li23, TW21]. Taxis-driven [TW21]. Taylor [FL24, KMW24, LZ23a]. Teichmüller [KB23]. Telegrapher [BFLP21]. Temperature [Bl23, CL22a, CMX22, CMX24, Don25, JV24a, LX24, WZ22, Zho22]. Temperature-Dependent [CL22a, Don25, Zho22]. Temperature-Depending [WZ22]. Tempered [LWC22]. Temporal [ADVW20, Wic23]. Tension [GMS21c, LZ21, LZ24b, MW20, TW22b, WX21]. Tensor [KLLT25, LW20a, MS21]. Tensors [CY22]. Teramoto [Mou20]. Term [CGU23, CCZ22, Jia20b, KM23, RZ20, Sun21, Tsu22, MB20]. Terminal [DGMP20]. Terms [DGR21, KT20, KB23]. Terrace [GKK22, KMY20, PTE25]. Tessellations [HT23]. Textiles [GOW20]. TFDW [SAB21]. Their [BHTY25, BMR20, CNW20, Hei23, JL21, dHKLN24, CJ24, Fei25, FLPS24, JMQ23]. Theorem [CGY21, FX21, KT22a, KV22, Yan25, CL23]. Theorems [AZ24, Bob23, CL25, LSW23, LW23b]. Theoretical [CFS21, SCB17, SCB20]. Theory [ADVW20, BZ25, BM24, BDMA22, CH21a, CLLZ23, CKP24, CM20b, FLRX24, GL23c, KM24, LW20a, MMS21, MN24b, NS21b, TW22a, WWZ23a]. Thermal [ZZ23b]. Thermistor [BO21]. Thermo [BSH24]. Thermo-Acoustics [BSH24]. Thermoelastic [TW23]. Thermoelasticity [Bl23, HSV20, HSV21]. Thickness [MM23]. Thien [CLYaY23]. Thin [ACNT24, AKLM22, Bob20, CW21, CDLLSG20, DT24, Gah25, GK25, Har20, WW22]. Thin-Film [DT24, GK25]. Three [ALM23, BLPS22, BP21, CW21, CKZ21, CLYaY23, Dai21b, DS21b, DW20, FJ22, FGW21, GMS21c, GL20, Hu22, Hu24a, HWXZ25, JJZ22, LS23a, LLWZ25, Ngu22, NR24, QX25, SX20b, Wan20, WWZ21, WXX21, XY21, ZLZ20, ZZ21b]. Three-Body [NR24]. Three-Dimensional [BP21, CW21, CKZ21, CLYaY23, Dai21b, DS21b, GMS21c, Hu22, Hu24a, JJZ22, LS23a, Wan20, WXX21, XY21, DW20]. Three-Scale [SX20b]. Three-Sphere [Ngu22]. Three-Wave [ALM23]. ThreeFold [GZZ22]. Threshold [ADK23, CM23a, Del23, GH21a, KMV21, LS21a, TT22, YZZ22, ZZ23a]. Through-the-Thickness [MM23]. Tiles [Pro20]. Time [ADH^{+}22, AKKR23, BS24a, BFLP21, BR23, BCF24, BK24, BDOZ21, BBBH25, BHL23, BSH24, Bl23, BL20, CGG20, CCFO21, CLYaY23, CGJL24, CCM21, CC23, CKKP20, CD20, DM20a, DK21a, DZZ23, DNS25, DM22, DR20b, DG20, Eit22, FLRX24, FHM^{+}24a, FHM^{+}24b, Fen25, FS22, FSW24, GL24a, GGBS25, GHM23, GLM20, GHW21, Gir25, GM20b, HH21, Hal22, HKK23, HXH22, HSV20, Hu24a, HWXZ25, HWZ22, JLN23, JM23, JKZ21, KbP21, KM20, KLSW24, KST23, KRS21, KS21, KS24a, LLR20, LLZ24, Li21, LS21b, LW21, LPW21, LS23c, LLW20, LM21, LLPZ21, LZ24a, LSY24, MRR20, MSM24, MS22, MSW24b, MS23b, MPS21, NS21a, OP20, Pan21, Pin21, RR20, RWZ21, RS24, SX20a, SCB17, SCB20, SUT22, Ste25b, Tet22, WZZ20, Xie25, XC22, ZZ21b, LS21a, Miz20, Ou21, PW21]. Time-Decay [LS23c]. Time-Dependent [BR23, CCFO21, Fen25, GLM20, GHW21, JM23, LLZ24, LSY24, SCB17, Tet22, Li21, SCB20]. Time-Domain [CGJL24, HSV20, MS23b].

Time-Fractional [FSW24, KM20].

Time-Frequency [RS24]. Time-Harmonic [CC23, Hal22, HWXZ25, OP20, CCM21, HH21]. Time-Periodic [DM20a, Eit22, FLRX24, HWZ22, LLPZ21, MS22]. Time-Periodicity [BL20]. Time-Singular [Pin21]. Time-Symmetric [ADH^{+}22]. Time-Varying [LZ24a, BFLP21]. Times [Kof23]. Timescale [BNP22]. Timoshenko [CDG23, TSMO24]. Tissue [DL22a, DbMS24, LX21]. Toda [WZ20b]. Tomography [CDGH20, GHK20, IM21b, Mef21, Mef24, Wan24]. Toner [BCK25]. Toner-Tu [BCK25]. Topological [QB24]. Topologically [ST20a]. Tori [HSSY24]. Torsion [KL24a, LY24]. Torsional [MT24]. Tortorelli [BCR21]. Torus [Eva21, FMN21, Gas23, LW22, SW20b, YZ24a]. Total [CDN21, DL24a, HW24, IM21a, PSW25]. Trace [Gol23]. Traceability [QWYY25]. Traces [DH21, EM22, Fos21]. Traffic [CF21, CY21b, CHLL22, DMS22c, Hol22]. Trailing [CXZ22]. Transcendent [LLhW23]. Transfer [AGL24, GHM22, IV22, RZZ21, Shu20]. Transform [CDX21, Chi22, Kat20, Kat23, LSW23, Mon20, MS23a, PU24]. Transformation [Aik23, BLPS22]. Transforms [BKS23, BPS23, Dun21, MS21, WHQ25]. Transient [HSV21, HSVW25]. Transition [Del23]. Transitions [BBBH25]. Transmission [ACL22, BL21, CN21, CM20b, EFHS21, FHS24, FN23, JL21]. Transonic [CMZZ22, FG22, GLY20, Hu22, PX22, WMZZ21, WXX21]. Transport [AK23a, AV24a, ANWS22, BL20, BCGJ22, BFHZ22, BV20a, CCS24, CGK23, Cio21, CKP20, Don25, DG23, EN22, FP21, GKM20, GSGM24, HHS23, HJLZ21, JK22, Jui20, LL20, LLW24, LXZ25, MN24b, MS24d, MR24, RZ21, SS20, SS24, SZ22, Zhu21]. Transportation [GZ20b, PVJ21, XX23a]. Transversally [Yan21a]. Trapped [LLWZ24]. Trapping [CWSGS20, GMS21a]. Traveling [BD22, BDLL25, BP23, CQ25, FP22, HH20, LW20b, MS24a, ST21, Zil23]. Treatment [Hal22]. Tree [BGS22]. Triangular [BCGJ22, Mou20]. Triple [DGV22, ELM21, Gen25]. Tritronquée [DL24a]. Truncated [GMS21a]. Truncation [GLS23]. Truncations [SS22b]. Tsingou [HJM25]. Tu [BCK25]. Tubes [PTE25]. Tug [ALP21b]. Tug-of-War [ALP21b]. Tumble [EY23]. Tumbling [HKLT24b]. Tumor [HXH22, KL23, RR25]. Tunneling [FSW22]. Turing [Gar24]. Turnpike [MRB22]. Twisted [BZ25]. Two [AKK^{+}20, ADL20, BKPY23, BCDH25, BFV21, BG24a, BC25, BCJ20, BCD24, CWZ20, CMX22, CHX24a, CGZ20, CKP^{+}22, DL22a, DPS24, DP25, DL24b, DST23, DK21b, DW21, FFV21, GLY20, GM20a, GK22, GRR23, HTY21, Hei23, HNNW23, HJ11, HC21, HL22a, IKL20, JTY23, JR21, KA24, Kor22, KKNN22, LS21b, LW21, LYZZ24, LLWZ25, LRW20, MP20a, Mar22, Nem21, NR24, PX22, PL23, PWZZ24, PTE25, RRTT24, SSZ20, Sin21, SZ22, Ste25a, Sun21, WZ20a, Win20, WZZ20, Yam22, YL22, Yan25, YY22, Zha21, gLZ25, CHS22]. Two-Body [NR24]. Two-Component [BKPY23, PWZZ24, YL22]. Two-Dimensional [ADL20, BFV21, BCJ20, BCD24, CWZ20, CMX22, CGZ20, CKP^{+}22, DP25, DL24b, GK22, GRR23, HJ11, HC21, HL22a, IKL20, LS21b, Mar22, PX22, Sun21, Win20, Yam22, GLY20, WZ20a, CHS22]. Two-Dimensional-Hyperbolic [MP20a]. Two-Fluid [PL23]. Two-Layer [Kor22, Sin21]. Two-Layered [LYZZ24]. Two-Phase [BG24a, BC25, GM20a, SSZ20, WZZ20, Yan25]. Two-Scale-Cut-and-Project [FFV21]. Two-Sphere [Zha21]. Two-Tubes [PTE25]. Two-Well [RRTT24]. Type [ABD22, ALP21b, BHTY25, BGMW20, BMPS21, CL21b, CHL23, CY21b, DLQ21, DST20, EGN22, Eit22, EIN22, FMN21, GP24, HLL21, HSV21, JWX22, KLS23, LYZ24, MS20d, NO23, Pru24, RW23, SW22b, SW20a, TW21, Tes20, TR20, Tré24, dHKLN24, AB21, HC21].

Ulam [HJM25]. Unbalanced [dN23]. Unbounded [AS20, BHTY25, BCDM21, GH24b, HWXZ25, IM21a, LYZ24, NWG24, WKM23, WZZ24]. Unboundedness [LX24]. Uncertainty [GIT20, SS20]. Unconditional [Gol23]. Undergoing [Gar24]. Uniaxial [LW20a]. Unidirectional [Lea23]. Unified [Jin20]. Uniform [BDOZ21, Cai23, CMT21, Don22, Dun24, FLPS24, FOS20, IKL20, LX24, LYZZ24, Reg21, Xie25]. Uniform-in-Time [Xie25]. Uniformly [BDN22]. Unimodal [FP22]. Unions [NS20b]. Unique [BCF24, Bl23, DFV25, RZ21, SL22]. Uniqueness [AIN25, BK23a, BLS22, BPSS22, CNW20, CHX24b, CJL23, CKP20, Cov22, DS23, DG25, FS20, GL23c, HJ24, HWWW24, HJT22, KT20, KA24, KLW22b, KV20, MRR20, Mef21, SS22a, SHM25, XX23b, Mef24]. Unit [GNNCR20]. Universal [LLS25]. Unknowns [Har23]. Unnormalized [CGK23]. Unsaturated [LKPR20]. Unsteady [FXX20, WZ20a]. Upper [AIN25, BDS24, Chu23]. Upscaling [LKPR20]. Using [BH23, GV24, GH21b, JL21, MS23b].

V [CL24]. Vacuum [CLZ22, CHS22, HL22a, LY23, Ou21, WZ25, WYZ23, Zen23]. Validation [KP21]. Validity [LYZ23]. Value [AC21, ANPS21, BZGZ24, CHS22, CHZ24, CDL22, DK23, GZ20a, HC21, KKK20, Kil22, KUW21, UZ23, YL22, ZRW22]. Valued [AKKR23, BHS25, BV20a, CbGwG24, Gal23, GK21b, Hua24, WY24]. Values [GSGM24, Kor22, RWZZ20]. Vanishing [BSH24, BC24b, BH22b, CFS22, DNS25, EB23, FMW24, PYZ24, WX21, WWZ23b, WW24, Zho22]. Vapor [RvW20]. Variable [BS24a, Chu23, HJK25, JSW23, LLS20]. Variation [CDN21, HW24, IM21a]. Variational [CFV20, CY21a, CCM21, CJL23, CDG23, DMRS20, DFGV25, GHPZ24, Han20, IM21a, KP21, MMS21, MP20b, SD24]. Variational-Hemivariational [Han20, MP20b]. Variations [ADJS24]. Varying [FHS24, HZ22a, LZ24a, BFLP21]. Vasculogenesis [CBHS23, LPW22]. Vector [AB21, BGK21, BHG25, Cio21, GRR23, MPC22, PS23a]. Vectors [WdlL20]. Velocity [AL23, CFP25, FV25, FHP24, JJZ22, Zhu21]. Venant [KL22]. Venttsel [ANPS21]. Verification [DLX21]. Version [ARS21, NS20a]. Versus [HR23, WXZ24]. Vertex [Rös21]. Vertical [FG22]. Very [HW22b, VX25]. Veselov [AG23]. via [AS21, BGK21, BK21, CS24, CGY21, CWZ20, CDM24a, CHZ24, CT23, CFS22, CW24, DIS25, DFGV25, EIN22, FFV21, GLSW23, GL23a, HSSY24, HT23, Hu24b, IV22, JKZ21, KL24a, KA24, KLSZ22, MMS21, NS21b, QWYY25, RS23a, RWZZ20, RH20, Sim21]. Viability [BWF24]. Vicsek [BDMA22]. Virial [AZ24, KMV21, Yan25]. Virus [TW21]. Viscoelastic [BMPS21, NS21a, dHKLN24]. Viscoelasticity [BPSS22, IY20, LG21, TSMO24]. Viscoelastodynamics [RS23b]. Viscoelastoplastic [GK21a]. Viscoplastic [CLR24]. Viscosities [AF23]. Viscosity [Bel22, BS24b, BH22b, CL22a, CL21b, CFS22, CR21, FV25, FMW24, FP20, HH23, KLM24, LZZ21, NS20a, RR20, TWZ20, TR20, WX21, WZ22, WWZ23b, WZ25, ZZ23b, Zho22]. Viscous [BG24b, CMX22, CMX24, HKK23, HL23a, Hu24b, JP20, KO24, LZ23a, LZZ21, LG21, Ste25a, TW23, WX21, YZ21, YZ25, YY22, Yua23]. Vlasov [AK23b, BMM22, BPIR25, BRR25, Bla23, CJ21, CDL24, CCZ22, CDF^{+}22, Dec24, DGOY24, DLY21, GZW21, HRZ24, HLlZ23, HNX25, KR21, LS23b, LW23a, LWX24, Nei20, Pan23, Reg21, Saf20, WZGX25, YZ24a]. Voltage [CaDZ24]. Voltage-Conductance [CaDZ24]. Volterra [FF25, PR20]. Volume [AdSST24, FL21]. Vortex [Aik23, AHK24, AHLW21, ADL20, BIFL20, BCM25, CW22, CXZ22, CF25, HW22a, Mar22, Ros20]. Vortexes [CYY23]. Vortexless [IKL20]. Vortices [CWZ20, Sol24]. Vorticity [BHI24, CWZ20, HR20, KL24b, Set24, VY23]. Vries [YZ24a, GM20b, LM21]. Vries/Zakharov [YZ24a]. vs [BW24, Tré24].

Walk [MT24]. Walker [HTY21]. Walks [HT23]. Wall [CL21b, HTY21, JW23, KSW24]. Walls [LWW20]. War [ALP21b]. Wasserstein [BWF24, CS24, CEK21, CCE24, HBK24, JMQ23, LWX24, SS20, WO23]. Water [AZ24, BCDH25, Boc20, CL24, DS21b, Eme21, MW20, QWYY25, Sai23, Sin21, Ste25a, Yan25]. Water-Wave [Yan25]. Wave [ALM23, AS20, BJ23, BJL24, BMM22, BDLL25, CCW23, CCS24, CM20a, CGJL24, CQ25, DS22, DST23, DH21, DM22, EFHS21, FP22, HWWW24, HL22b, JWX22, KN22, Kof23, KLW22b, LW24, LYZZ24, MS24a, MPS21, Pin21, SW22b, SSWZ24, SW22c, SUT22, ST21, Ste25a, VX21, WWZ21, Yan25, YZ21, YZ25, YY22]. Wave-Damping [BMM22]. Wave-Klein [DM22]. Wavefronts [HKlT24a]. Waveguide [YYZ24]. Waves [AZ24, BP23, CL24, CF22, CCLM23, Dai21a, DS21b, EGN22, Eme21, HH20, HKK23, HJ24, IPTT22, JLS25, JP20, LOWX20, LW20b, LG21, LPW22, MP21a, MW20, MS20b, RWZ21, Sai23, Set24, Sin21, Var20, VX21, VX25, WX21, WY21, YZ21, ZZZ22b, ZZZ24, Zil23, dW23]. Wayne [HSSY24]. Weak [AW21, AKK^{+}20, AT20, BPSS22, CGG20, CJ21, CHS22, CGK23, Dai21b, DS23, GL23a, GM20a, GK20, GRZ24, GWZ21, GK21b, Hu24a, HJT22, KT20, KM24, KKNN22, LS21b, Lia22, MS24c, MS22, MMS25, Nov20, OP20, SS22a, Shu20, Shu21, WZ22, XZZ24]. Weak-Strong [DS23, HJT22, SS22a]. Weaker [Yam22]. Weakly [AV24a, BC25, ES24, Kil22, MMSH21]. Wedge [CDM24a, PX22]. Wedges [DL24b]. Weighted [ABLW24, DR24, DG23, FMMS22, MPV23]. Weights [BDGdN22]. Weinberger [ABD22]. Well [ALW23, AG23, AV24a, AGL24, AMS24, BCK25, BT23, BSH24, Boc20, BBES22, CW21, CGS23, CGSS24, COS24, DST20, DFGV25, DGWZ23, DGOY24, DKLT25, FV25, GHM22, GGL24, HH21, HW22b, HHS23, HK23, HL22a, JWX22, KT24, LN20, LWZ21b, LY21, LR21, LW23b, MS20d, PYZ24, Pec20, RRTT24, TW22b, Wan21, WY21, XZZ24, YYZ21, YYZ24, YY24]. Well-Posedness [ALW23, AG23, AV24a, AGL24, AMS24, BCK25, BT23, BBES22, CW21, CGS23, CGSS24, COS24, DST20, DGWZ23, DGOY24, DKLT25, FV25, GGL24, HW22b, HHS23, HK23, HL22a, JWX22, KT24, LWZ21b, LR21, MS20d, PYZ24, Pec20, TW22b, Wan21, XZZ24, YYZ24, YY24, Boc20, HH21, LY21, WY21, YYZ21]. Well-Prepared [GHM22]. Wellposedness [Ger21]. Wells [CFG25]. Wentzell [GL23a, HLL21]. Westervelt [BSH24, ES24]. Weyl [FN23, GZZ22]. where [Oli21]. White [ACS23, BK20a, NZ24]. White-Forced [NZ24]. White-Noise [BK20a]. Whitham [DST20, MS24a, SW22b, SSWZ24]. Whitham-Type [SW22b]. Whole [CT23, NZ24, WZZ20]. Whose [DK23].

Widely [BDGdN24]. Wiener [BT21]. Wild [BW25]. Winding [GK23]. Wires [DJLR24]. Without [LLW21, Pan24, FN23, HJLZ21, IMS20, LY21, PZ23]. Witting [HY21]. Wong [FX21, XC22]. Work [CL21c].

X [IM21b, Mon20, MS23a, Wan24]. X-Ray [Mon20, IM21b, MS23a, Wan24].

Yang [Pec20]. Young [ARD20]. Yudovich [BPIR25, KL24b]. Yukawa [WY24].

Zakai [FX21, XC22]. Zakharov [YZ24a, Don22, LR21]. Zero [AG23, Bad24, CHLL22, CHQ22, Gas23, TWZ20]. Zero-Dispersion [Bad24, Gas23]. Zero-Inertia [CHQ22]. Zero-Viscosity [TWZ20]. Zeros [Dun24, FLPS24, Nem21]. Zhabotinsky [HKlT24a]. ZND [BW24].

References

Abels:2020:BCC

[AB20] Helmut Abels and Julia Butz. A blow-up criterion for the curve diffusion flow with a contact angle. SIAM Journal on Mathematical Analysis, 52(3): 2592–2623, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amrouche:2021:VPM

[AB21] Chérif Amrouche and Imane Boussetouan. Vector potentials with mixed boundary conditions. application to the Stokes problem with pressure and Navier-type boundary conditions. SIAM Journal on Mathematical Analysis, 53(2):1745– 1784, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anoop:2022:SWT

[ABD22] T. V. Anoop, Vladimir Bobkov, and Pavel Drábek. Szegő– Weinberger type inequalities for symmetric domains with holes. SIAM Journal on Mathematical Analysis, 54 (1):389–422, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1407227.

Abdo:2024:RSS

[Abd24] Elie Abdo. On the regularity of solutions to some active scalar equations on bounded domains. SIAM Journal on Mathematical Analysis, 56(6):7306–7326, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1614705.

Aspri:2022:LSD

[ABFV22] Andrea Aspri, Elena Beretta, Elisa Francini, and Sergio Vessella. Lipschitz stable determination of polyhedral conductivity inclusions from local boundary measurements. SIAM Journal on Mathematical Analysis, 54(5):5182–5222, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480550.

AlSayed:2021:MSE

[ABHN21] Abdelkader Al Sayed, Beniamin Bogosel, Antoine Henrot, and Florent Nacry. Maximization of the Steklov eigenvalues with a diameter constraint. SIAM Journal on Mathematical Analysis, 53(1):710–729, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alama:2024:CSD

[ABLW24] Stanley Alama, Lia Bronsard, Xinyang Lu, and Chong Wang. Core shells and double bubbles in a weighted nonlocal isoperimetric problem. SIAM Journal on Mathematical Analysis, 56(2):2357–2394, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Acosta:2022:LNE

[ABR22] Gabriel Acosta, Francisco Bersetche, and Julio D. Rossi. Local and nonlocal energy-based coupling models. SIAM Journal on Mathematical Analysis, 54(6):6288–6322, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1431977.

Abatangelo:2021:EBV

[AC21] Nicola Abatangelo and Matteo Cozzi. An elliptic boundary value problem with fractional nonlinearity. SIAM Journal on Mathematical Analysis, 53(3): 3577–3601, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alonso:2024:IPB

[AČ24] Ricardo Alonso and Milana Čolić. Integrability propagation for a Boltzmann system describing polyatomic gas mixtures. SIAM Journal on Mathematical Analysis, 56(1):1459– 1494, February 2024. CODEN SJMAAH. ISSN 1095-7154.

Arratia:2021:LSB

[ACC^{+}21] Pablo Arratia, Matías Courdurier, Evelyn Cueva, Axel Osses, and Benjamín Palacios. Lipschitz stability for backward heat equation with application to fluorescence microscopy. SIAM Journal on Mathematical Analysis, 53(5): 5948–5978, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Acerbi:2022:SLC

[ACC22] Emilio Acerbi, Chao-Nien Chen, and Yung-Sze Choi. Stability of lamellar configurations in a nonlocal sharp interface model. SIAM Journal on Mathematical Analysis, 54(1):558–594, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1394321.

Arratia:2023:ELS

[ACC^{+}23] Pablo Arratia, Matías Courdurier, Evelyn Cueva, Axel Osses, and Benjamín Palacios. Erratum: Lipschitz stability for backward heat equation with application to fluorescence microscopy. SIAM Journal on Mathematical Analysis, 55(3):2467–2473, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1534882.

Ammari:2022:LSA

[ACL22] Habib Ammari, Yat Tin Chow, and Hongyu Liu. Localized sensitivity analysis at high-curvature boundary points of reconstructing inclusions in transmission problems. SIAM Journal on Mathematical Analysis, 54(2):1543–1592, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1323576.

Acampora:2024:OST

[ACNT24] P. Acampora, E. Cristoforoni, C. Nitsch, and C. Trombetti. On the optimal shape of a thin insulating layer. SIAM Journal on Mathematical Analysis, 56 (3):3509–3536, May 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Ascione:2023:MFS

[ACS23] Giacomo Ascione, Daniele Castorina, and Francesco Solombrino. Mean-field sparse optimal control of systems with additive white noise. SIAM Journal on Mathematical Analysis, 55(6):6965–6990, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Aregba-Driollet:2021:GES

[ADBP21] Denise Aregba-Driollet, Stéphane Brull, and Yue-Jun Peng. Global existence of smooth solutions for a nonconservative bitemperature Euler model. SIAM Journal on Mathematical Analysis, 53(2):1886–1907, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adams:2022:ERN

[ADdR22] Daniel Adams, Manh Hong Duong, and Gonçalo dos Reis. Entropic regularization of Non-Gradient systems. SIAM Journal on Mathematical Analysis, 54(4):4495–4535, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1414668.

Ammari:2022:EPP

[ADH^{+}22] Habib Ammari, Bryn Davies, Erik Orvehed Hiltunen, Hyundae Lee, and Sanghyeon Yu. Exceptional points in parity– time-symmetric subwavelength metamaterials. SIAM Journal on Mathematical Analysis, 54(6):6223–6253, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469821.

Ammari:2023:CRD

[ADH23] Habib Ammari, Bryn Davies, and Erik Orvehed Hiltunen. Convergence rates for defect modes in large finite resonator arrays. SIAM Journal on Mathematical Analysis, 55(6): 7616–7634, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Antil:2024:NBV

[ADJS24] Harbir Antil, Hugo Díaz, Tian Jing, and Armin Schikorra. Nonlocal bounded variations with applications. SIAM Journal on Mathematical Analysis, 56(2):1903–1935, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Alfaro:2023:QTP

[ADK23] Matthieu Alfaro, Arnaud Ducrot, and Hao Kang. Quantifying the threshold phenomenon for propagation in nonlocal diffusion equations. SIAM Journal on Mathematical Analysis, 55(3):1596–1630, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1479099.

Arsenio:2020:VMT

[ADL20] Diogo Arsénio, Emmanuel Dormy, and Christophe Lacave. The vortex method for two-dimensional ideal flows in exterior domains. SIAM Journal on Mathematical Analysis, 52(4):3881–3961, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andrade:2024:SPP

[AdSST24] Pêdra D. S. Andrade, Ederson Moreira dos Santos, Makson S. Santos, and Hugo Tavares. Spectral partition problems with volume and inclusion constraints. SIAM Journal on Mathematical Analysis, 56(6):7136–7169, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M161553X.

Avitabile:2020:LTS

[ADVW20] Daniele Avitabile, Mathieu Desroches, Romain Veltz, and Martin Wechselberger. Local theory for spatio-temporal canards and delayed bifurcations. SIAM Journal on Mathematical Analysis, 52(6):5703–5747, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abels:2023:SIL

[AF23] Helmut Abels and Mingwen Fei. Sharp interface limit for a Navier–Stokes/Allen– Cahn system with different viscosities. SIAM Journal on Mathematical Analysis, 55 (4):4039–4088, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1523698.

Ammari:2020:HLM

[AFH^{+}20] Habib Ammari, Brian Fitzpatrick, Erik Orvehed Hiltunen, Hyundae Lee, and Sanghyeon Yu. Honeycomb-lattice Minnaert bubbles. SIAM Journal on Mathematical Analysis, 52 (6):5441–5466, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adams:2023:LRL

[AG23] Joseph Adams and Axel Grünrock. Low regularity local well-posedness for the zero energy Novikov–Veselov equation. SIAM Journal on Mathematical Analysis, 55(1):19–35, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458065.

Andreianov:2024:ESC

[AG24] Boris Andreianov and Theo Girard. Existence of solutions for a class of one-dimensional models of pedestrian evacuations. SIAM Journal on Mathematical Analysis, 56(3):3386– 3411, May 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Ali:2024:DCI

[AGL24] Mashkoor Ali, Ankik Kumar Giri, and Philippe Laurençot. The discrete collisioninduced breakage equation with mass transfer: Wellposedness and stationary solutions. SIAM Journal on Mathematical Analysis, 56(3):2915– 2937, May 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Ambrose:2024:CDG

[AHK24] David M. Ambrose, Fazel Hadadifard, and James P. Kelliher. Contour dynamics and global regularity for periodic vortex patches and layers. SIAM Journal on Mathematical Analysis, 56(2):2286– 2311, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Ao:2021:GAM

[AHLW21] Weiwei Ao, Yehui Huang, Yong Liu, and Juncheng Wei. Generalized Adler–Moser polynomials and multiple vortex rings for the Gross–Pitaevskii equation. SIAM Journal on Mathematical Analysis, 53(6):6959– 6992, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antonelli:2020:LMN

[AHM20] Paolo Antonelli, Lars Eric Hientzsch, and Pierangelo Marcati. On the low Mach number limit for quantum Navier–Stokes equations.

SIAM Journal on Mathematical Analysis, 52(6):6105–6139, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aiki:2023:HTF

[Aik23]  Masashi Aiki.  The Hasi-

moto transformation for a finite length vortex filament and its application. SIAM Journal on Mathematical Analysis, 55(6):7273–7295, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Auscher:2025:FSK

[AIN25] Pascal Auscher, Cyril Imbert, and Lukas Niebel. Fundamental solutions to Kolmogorov– Fokker–Planck equations with rough coefficients: Existence, uniqueness, upper estimates. SIAM Journal on Mathematical Analysis, 57(2):2114–2137, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1649241.

Abedin:2023:PAP

[AK23a] Farhan Abedin and Jun Kitagawa. A perturbative approach to the parabolic optimal transport problem. SIAM Journal on Mathematical Analysis, 55(6):6740–6763, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Andreasson:2023:SSS

[AK23b] Håkan Andreasson and Markus Kunze. Static solutions to the spherically symmetric Einstein–Vlasov system: a particle number–Casimir approach. SIAM Journal on Mathematical Analysis, 55 (5):4843–4879, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1522887.

AlBaba:2020:NAW

[AKK^{+}20] Hind Al Baba, Christian Klingenberg, Ondrej Kreml, Václav Mácha, and Simon Markfelder. Nonuniqueness of admissible weak solution to the Riemann problem for the full Euler system in two dimensions. SIAM Journal on Mathematical Analysis, 52(2):1729–1760, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Atar:2023:LTL

[AKKR23] Rami Atar, Weining Kang, Haya Kaspi, and Kavita Ramanan. Long-time limit of nonlinearly coupled measurevalued equations that model many-server queues with reneging. SIAM Journal on Mathematical Analysis, 55(6):7189– 7239, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Ammari:2023:QES

[AKKY23] Habib Ammari, Hyeonbae Kang, Do Wan Kim, and Sanghyeon Yu. Quantitative estimates for stress concentration of the Stokes flow between adjacent circular cylinders. SIAM Journal on Mathematical Analysis, 55(4):3755–3806, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1523972.

Ando:2022:SSN

[AKLM22] Kazunori Ando, Hyeonbae Kang, Sanghyuk Lee, and Yoshihisa Miyanishi. Spectral structure of the Neumann– Poincaré operator on thin ellipsoids and flat domains. SIAM Journal on Mathematical Analysis, 54(6):6164–6185, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1452275.

Angelopoulos:2020:IMI

[AKV20] Yannis Angelopoulos, Rowan Killip, and Monica Visan. Invariant measures for integrable spin chains and an integrable discrete nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 52(1): 135–163, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

An:2023:HVT

[AL23] Gayoung An and Donghyun Lee. High-velocity tails of the inelastic and the multispecies mixture Boltzmann equations. SIAM Journal on Mathematical Analysis, 55(5):4297–4336, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1529749.

An:2025:MCC

[AL25] Gayoung An and Donghyun Lee. The mixed convex– concave effect on the regularity of a Boltzmann solution. SIAM Journal on Mathematical Analysis, 57(1):1086–1136, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1670354.

Ablowitz:2023:TWI

[ALM23] Mark J. Ablowitz, Xu-Dan Luo, and Ziad H. Musslimani. Three-wave interaction equations: Classical and nonlocal. SIAM Journal on Mathematical Analysis, 55(4):4089–4139, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1488880.

Alicandro:2021:DLE

[ALP21a] Roberto Alicandro, Giuliano Lazzaroni, and Mariapia Palombaro. Derivation of linear elasticity for a general class of atomistic energies. SIAM Journal on Mathematical Analysis, 53(5):5060–5093, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Attouchi:2021:GLE

[ALP21b] Amal Attouchi, Hannes Luiro, and Mikko Parviainen. Gradient and Lipschitz estimates for tug-of-war type games. SIAM Journal on Mathematical Analysis, 53(2):1295–1319, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Almeida:2023:ARC

[ALV23] Luis Almeida, Alexis Léculier, and Nicolas Vauchelet. Analysis of the rolling carpet strategy to eradicate an invasive species. SIAM Journal on Mathematical Analysis, 55(1):275–309, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1427243.

Abdo:2023:GWP

[ALW23] Elie Abdo, Fizay-Noah Lee, and Weinan Wang. On the global well-posedness and analyticity of some electrodiffusion models in ideal fluids and porous media. SIAM Journal on Mathematical Analysis, 55(6):6838–6866, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Abels:2022:CAC

[AM22] Helmut Abels and Maximilian Moser. Convergence of the Allen–Cahn equation with a nonlinear Robin boundary condition to mean curvature flow with contact angle close to 90^{∘}. SIAM Journal on Mathematical Analysis, 54(1):114–172, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1424925.

Achdou:2022:DMF

[AMMT22] Yves Achdou, Paola Mannucci, Claudio Marchi, and Nicoletta Tchou. Deterministic mean field games with control on the acceleration and state constraints. SIAM Journal on Mathematical Analysis, 54(3):3757–3788, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1415492.

Achdou:2024:DMF

[AMMT24] Yves Achdou, Paola Mannucci, Claudio Marchi, and Nicoletta Tchou. Deterministic mean field games on networks: a Lagrangian approach. SIAM Journal on Mathematical Analysis, 56(5):6689–6730, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1615073.

Ampatzoglou:2022:RDB

[AMP22] Ioakeim Ampatzoglou, Joseph K.

Miller, and Natasa Pavlović. A rigorous derivation of a Boltzmann system for a mixture of hard-sphere gases. SIAM Journal on Mathematical Analysis, 54(2):2320–2372, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1424779.

Antil:2024:WPP

[AMS24] Harbir Antil, Daiki Mizuno, and Ken Shirakawa. Wellposedness of a pseudo-parabolic KWC system in materials science. SIAM Journal on Mathematical Analysis, 56 (5):6422–6445, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M163952X.

Apushkinskaya:2021:VBV

[ANPS21] Darya E. Apushkinskaya, Alexander I. Nazarov, Dian K. Palagachev, and Lubomira G. Softova. Venttsel boundary value problems with discontinuous data. SIAM Journal on Mathematical Analysis, 53 (1):221–252, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Altschuler:2022:ASE

[ANWS22] Jason M. Altschuler, Jonathan Niles-Weed, and Austin J. Stromme. Asymptotics for semidiscrete entropic optimal transport. SIAM Journal on Mathematical Analysis, 54 (2):1718–1741, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1440165.

Albritton:2023:SES

[AO23] Dallas Albritton and Laurel Ohm. On the stabilizing effect of swimming in an active suspension. SIAM Journal on Mathematical Analysis, 55(6):6093–6132, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Abatangelo:2024:SBD

[AO24] Laura Abatangelo and Roberto Ognibene. Sharp behavior of Dirichlet–Laplacian eigenvalues for a class of singularly perturbed problems. SIAM Journal on Mathematical Analysis, 56(1):474–500, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Alvarez-Perez:2024:EHE

[AP24] Jose Luis Alvarez-Perez. Extended Hadamard expansions for the Airy functions. SIAM Journal on Mathematical Analysis, 56(3):3537–3558, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Arroyo-Rabasa:2020:GMY

[ARD20] Adolfo Arroyo-Rabasa and Johannes Diermeier. Generalized multiscale Young measures. SIAM Journal on Mathematical Analysis, 52(4):3252– 3300, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antil:2021:FVM

[ARS21] Harbir Antil, Carlos N. Rautenberg, and Armin Schikorra. On a fractional version of a murat compactness result and applications. SIAM Journal on Mathematical Analysis, 53(3): 3158–3187, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arifoski:2020:PDW

[AS20] Ahmet Arifoski and Petr Siegl. Pseudospectra of the damped wave equation with unbounded damping. SIAM Journal on Mathematical Analysis, 52(2): 1343–1362, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Avery:2021:ASC

[AS21] Montie Avery and Arnd Scheel. Asymptotic stability of critical pulled fronts via resolvent expansions near the essential spectrum. SIAM Journal on Mathematical Analysis, 53(2): 2206–2242, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2020:WOE

[AT20] Kaïs Ammari and Faouzi Triki. On weak observability for evolution systems with skew-adjoint generators.

SIAM Journal on Mathematical Analysis, 52(2):1884–1902, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Akagi:2024:SSJ

[ATZ24] Goro Akagi, Izumi Takagi, and Conghui Zhang. Steady states with jump discontinuity in a receptor-based model with hysteresis in higher-dimensional domains. SIAM Journal on Mathematical Analysis, 56(2): 1996–2033, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Agresti:2024:RDE

[AV24a] Antonio Agresti and Mark Veraar. Reaction–diffusion equations with transport noise and critical superlinear diffusion: Global well-posedness of weakly dissipative systems. SIAM Journal on Mathematical Analysis, 56(4):4870–4927, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1562482.

Armstrong:2024:AES

[AV24b] Scott Armstrong and Raghavendra Venkatraman. Asymptotic expansion of the spectrum for periodic Schrödinger operators. SIAM Journal on Mathematical Analysis, 56(2): 1770–1808, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Akramov:2021:NAW

[AW21] Ibrokhimbek Akramov and Emil Wiedemann. NonUnique admissible weak solutions of the compressible Euler equations with compact support in space. SIAM Journal on Mathematical Analysis, 53(1): 795–812, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alazard:2024:VTE

[AZ24] Thomas Alazard and Claude Zuily. Virial theorems and equipartition of energy for water waves. SIAM Journal on Mathematical Analysis, 56 (4):5285–5329, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1574312.

Badreddine:2024:ZDL

[Bad24]  Rana Badreddine.  Zero-

dispersion limit of the Calogero– Moser derivative NLS equation. SIAM Journal on Mathematical Analysis, 56 (6):7228–7249, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1646935.

Badsi:2025:LNS

[Bad25] Mehdi Badsi. Linear and NonLinear stability for the kinetic plasma sheath on a bounded interval. SIAM Journal on Mathematical Analysis, 57(5):4953–4994, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1678660.

Baradat:2020:ESP

[Bar20] Aymeric Baradat. On the existence of a scalar pressure field in the Brödinger problem. SIAM Journal on Mathematical Analysis, 52(1):370– 401, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barker:2023:LQE

[Bar23] Tobias Barker. Localized quantitative estimates and potential blow-up rates for the Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 55(5):5221–5259, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1527179.

Babadjian:2023:DAG

[BB23] Jean-Francois Babadjian and Elise Bonhomme. Discrete approximation of the Griffith functional by adaptive finite elements. SIAM Journal on Mathematical Analysis, 55(6):6778–6837, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Basile:2025:LTA

[BBBH25] Giada Basile, Dario Benedetto, Lorenzo Bertini, and Daniel Heydecker. Large time analysis of the rate function associated to the Boltzmann equation: Dynamical phase transitions. SIAM Journal on Mathematical Analysis, 57(5):5746–5770, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1719165.

Bach:2020:DCL

[BBC20] Annika Bach, Andrea Braides, and Marco Cicalese. Discreteto-continuum limits of multibody systems with bulk and surface long-range interactions. SIAM Journal on Mathematical Analysis, 52(4):3600–3665, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:2023:ESD

[BBD^{+}23] Guillaume Bal, Simon Becker, Alexis Drouot, Clotilde Fermanian Kammerer, Jianfeng Lu, and Alexander B. Watson. Edge state dynamics along curved interfaces. SIAM Journal on Mathematical Analysis, 55(5):4219–4254, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1489708.

Bruna:2022:WPI

[BBES22] Maria Bruna, Martin Burger, Antonio Esposito, and Simon Schulz. Well-posedness of an integro-differential model for active Brownian particles. SIAM Journal on Mathematical Analysis, 54(5):5662–5697, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462039.

Berchio:2024:ECS

[BBG^{+}24] Elvise Berchio, Denis Bonheure, Giovanni P. Galdi, Filippo Gazzola, and Simona Perotto. Equilibrium configurations of a symmetric body immersed in a stationary Navier– Stokes flow in a planar channel. SIAM Journal on Mathematical Analysis, 56(3):3759–3801, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1568752.

Bertucci:2024:MFG

[BC24a] Charles Bertucci and Alekos Cecchin. Mean field games master equations: From discrete to continuous state space. SIAM Journal on Mathematical Analysis, 56(2):2569–2610, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Bourgeois:2024:GIB

[BC24b] Laurent Bourgeois and Lucas Chesnel. Generalized impedance boundary conditions with vanishing or signchanging impedance. SIAM Journal on Mathematical Analysis, 56(3):4223–4251, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1604217.

Both:2025:GER

[BC25] Jakub W. Both and Clément Cancès. A global existence result for weakly coupled two-phase poromechanics. SIAM Journal on Mathematical Analysis, 57(5):4995–5038, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1688783.

Baird:2022:GMI

[BCC^{+}22] Thomas J. Baird, Paul Cornwell, Graham Cox, Christopher Jones, and Robert Marangell. Generalized Maslov indices for non-Hamiltonian systems. SIAM Journal on Mathematical Analysis, 54 (2):1623–1668, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1381319.

Briane:2024:NDC

[BCD24] Marc Briane and Juan Casado-Díaz. A new divergence-curl result for measures. application to the two-dimensional ODE’s flow. SIAM Journal on Mathematical Analysis, 56 (5):6398–6421, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1617539.

Bauerle:2025:RIH

[BCDH25] Matthew Bauerle, Andrew J.

Christlieb, Mingchang Ding, and Juntao Huang. On the rotational invariance and hyperbolicity of shallow water moment equations in two dimensions. SIAM Journal on Mathematical Analysis, 57 (1):1039–1085, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1579789.

Buttazzo:2021:EOP

[BCDM21] Giuseppe Buttazzo, Juan Casado-Díaz, and Faustino Maestre. On the existence of optimal potentials on unbounded domains. SIAM Journal on Mathematical Analysis, 53(1):1088–1121, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barcelo:2024:UCP

[BCF24] Juan A. Barceló, Biagio Cassano, and Luca Fanelli. Unique continuation properties from one time for hyperbolic Schrödinger equations.

SIAM Journal on Mathematical Analysis, 56(6):7417–7438, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1578218.

Battaglia:2023:NNL

[BCFP23] Luca Battaglia, Matteo Cozzi, Antonio J. Fernández, and Angela Pistoia. Nonuniqueness for the nonlocal Liouville equation in R and applications. SIAM Journal on Mathematical Analysis, 55(5):4816–4842, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1538004.

Bourdarias:2022:ESC

[BCGJ22] Christian Bourdarias, Anupam Pal Choudhury, Billel Guelmame, and Stéphane Junca. Entropy solutions in BV^{s} for a class of triangular systems involving a transport equation. SIAM Journal on Mathematical Analysis, 54(1):791–817, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1351783.

Braukhoff:2020:CCD

[BCJ20] Marcel Braukhoff, Xiuqing Chen, and Ansgar Jüngel. Corrigendum: Cross Diffusion Preventing Blow-Up in the Two-Dimensional Keller– Segel Model. SIAM Journal on Mathematical Analysis, 52 (2):2198–2200, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bae:2025:WPA

[BCK25] Hantaek Bae, Young-Pil Choi, and Kyungkeun Kang. Wellposedness and asymptotic stability of solutions for the incompressible toner-Tu model. SIAM Journal on Mathematical Analysis, 57(1):637–660, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1599756.

Butta:2025:LVR

[BCM25] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Leapfrogging vortex rings as scaling limit of Euler equations. SIAM Journal on Mathematical Analysis, 57 (1):789–824, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1642391.

Bociu:2021:MPI

[BCMW21] Lorena Bociu, Suncica Canic, Boris Muha, and Justin T. Webster. Multilayered poroelasticity interacting with Stokes flow. SIAM Journal on Mathematical Analysis, 53(6):6243– 6279, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bach:2021:RFD

[BCR21] Annika Bach, Marco Cicalese, and Matthias Ruf. Random finite-difference discretizations of the Ambrosio–Tortorelli functional with optimal meshsize. SIAM Journal on Mathematical Analysis, 53(2):2275– 2318, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baers:2025:IPF

[BCR25] Hendrik Baers, Giovanni Covi, and Angkana Rüland. On instability properties of the fractional Calderón problem. SIAM Journal on Mathematical Analysis, 57(5):5454–5493, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M167010X.

Bonacini:2024:SBA

[BCT24] Marco Bonacini, Riccardo Cristoferi, and Ihsan Topaloglu. Stability of the ball for attractive-repulsive energies. SIAM Journal on Mathematical Analysis, 56(1):588–615, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Beyn:2022:STO

[BD22] Wolf-Jürgen Beyn and Christian Döding. Stability of traveling oscillating fronts in complex Ginzburg Landau equations. SIAM Journal on Mathematical Analysis, 54 (5):5447–5488, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1455760.

Bogelein:2020:ICP

[BDD^{+}20] Verena Bögelein, Bernard Dacorogna, Frank Duzaar, Paolo Marcellini, and Christoph Scheven. Integral convexity and parabolic systems. SIAM Journal on Mathematical Analysis, 52(2):1489–1525, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boscaggin:2021:PSP

[BDF21] Alberto Boscaggin, Walter Dambrosio, and Guglielmo Feltrin. Periodic solutions to a perturbed relativistic Kepler problem. SIAM Journal on Mathematical Analysis, 53(5): 5813–5834, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellazzini:2024:SNN

[BDF24] Jacopo Bellazzini, Van Duong Dinh, and Luigi Forcella. Scattering for nonradial 3D NLS with combined nonlinearities: The interaction morawetz approach. SIAM Journal on Mathematical Analysis, 56(3): 3110–3143, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Boscaggin:2025:BCO

[BDF25] Alberto Boscaggin, Walter Dambrosio, and Guglielmo Feltrin. Bifurcation of closed orbits of Hamiltonian systems with application to geodesics of the Schwarzschild metric. SIAM Journal on Mathematical Analysis, 57(5):5531–5569, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1711200.

Balci:2022:EED

[BDGdN22] Anna Kh. Balci, Lars Diening, Raffaella Giova, and Antonia Passarelli di Napoli. Elliptic equations with degenerate weights. SIAM Journal on Mathematical Analysis, 54(2):2373–2412, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1412529.

Bogelein:2024:GRC

[BDGdN24] Verena Bögelein, Frank Duzaar, Raffaella Giova, and Antonia Passarelli di Napoli. Gradient regularity for a class of widely degenerate parabolic systems. SIAM Journal on Mathematical Analysis, 56 (4):5017–5078, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1589232.

Brauner:2025:STW

[BDLL25] Claude-Michel Brauner, Yuchao Dong, Jin Liang, and Luca Lorenzi. Stability of traveling wave solutions in a credit rating migration free boundary problem. SIAM Journal on Mathematical Analysis, 57(1):563–600, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1628061.

Briant:2022:CTG

[BDMA22] Marc Briant, Antoine Diez, and Sara Merino-Aceituno. Cauchy theory for general kinetic Vicsek models in collective dynamics and meanfield limit approximations. SIAM Journal on Mathematical Analysis, 54(1):1131–1168, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1405885.

Bianchini:2022:MEH

[BDN22] Stefano Bianchini, Prerona Dutta, and Khai T. Nguyen. Metric entropy for Hamilton– Jacobi equations with uniformly directionally convex Hamiltonian. SIAM Journal on Mathematical Analysis, 54(5):5551–5575, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1475430.

Barre:2021:FNM

[BDOZ21] Julien Barré, Paul Dobson, Michela Ottobre, and Ewelina Zatorska. Fast non-mean-field networks: Uniform in time averaging. SIAM Journal on Mathematical Analysis, 53(1): 937–972, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biswas:2021:HIH

[BDS21] Animesh Biswas, Marta De León-Contreras, and Pablo Raúl Stinga. Harnack inequalities and Hölder estimates for master equations. SIAM Journal on Mathematical Analysis, 53 (2):2319–2348, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boccato:2024:UBG

[BDS24] Chiara Boccato, Andreas Deuchert, and David Stocker. Upper bound for the grand canonical free energy of the Bose gas in the Gross– Pitaevskii limit. SIAM Journal on Mathematical Analysis, 56(2):2611–2660, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Bellieud:2022:HNH

[Bel22] Michel Bellieud. Homogenization of Norton–Hoff fibered composites with high viscosity contrast. SIAM Journal on Mathematical Analysis, 54(1):649–692, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1345785.

Bessaih:2021:SHC

[BEM21] Hakima Bessaih, Yalchin Efendiev, and Razvan Florian Maris. Stochastic homogenization of a convection–diffusion equation. SIAM Journal on Mathematical Analysis, 53(3): 2718–2745, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bach:2023:GDM

[BEMZ23] Annika Bach, Teresa Esposito, Roberta Marziani, and Caterina Ida Zeppieri. Gradient damage models for heterogeneous materials. SIAM Journal on Mathematical Analysis, 55(4):3567–3601, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1499145.

BenPorat:2024:MLE

[Ben24] Immanuel Ben Porat. The magnetic Liouville equation as a semiclassical limit. SIAM Journal on Mathematical Analysis, 56(1):955–992, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Blaustein:2023:CPF

[BF23] Alain Blaustein and Francis Filbet. Concentration phenomena in Fitzhugh–Nagumo equations: a mesoscopic approach. SIAM Journal on Mathematical Analysis, 55(1):367–404, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1469961.

Breit:2022:CNS

[BFHZ22] Dominic Breit, Eduard Feireisl, Martina Hofmanová, and Ewelina Zatorska. Compressible Navier–Stokes system with transport noise. SIAM Journal on Mathematical Analysis, 54(4):4465–4494, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1464701.

Baratchart:2021:SSC

[BFLP21] L. Baratchart, S. Fueyo, G. Lebeau, and J.-B. Pomet. Sufficient stability conditions for time-varying networks of telegrapher’s equations or difference-delay equations. SIAM Journal on Mathematical Analysis, 53(2):1831–1856, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brezina:2020:SSS

[BFN20] Jan Brezina, Eduard Feireisl, and Antonín Novotný. Stability of strong solutions to the Navier–Stokes–Fourier system. SIAM Journal on Mathematical Analysis, 52(2):1761–1785, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beretta:2021:LSD

[BFV21] Elena Beretta, Elisa Francini, and Sergio Vessella. Lipschitz stable determination of polygonal conductivity inclusions in a two-dimensional layered medium from the Dirichlet-to– Neumann map. SIAM Journal on Mathematical Analysis, 53 (4):4303–4327, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bevilacqua:2024:GST

[BG24a] Giulia Bevilacqua and Andrea Giorgini. Global solutions for two-phase complex fluids with quadratic anchoring in soft matter physics. SIAM Journal on Mathematical Analysis, 56(5):6057–6120, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1608902.

Bian:2024:ONI

[BG24b] D. Bian and E. Grenier. Onset of nonlinear instabilities in monotonic viscous boundary layers. SIAM Journal on Mathematical Analysis, 56 (3):3703–3719, ???? 2024.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1505773.

Brown:2024:FEC

[BG24c] Henry J. Brown and Yury Grabovsky. On feasibility of extrapolation of completely monotone functions. SIAM Journal on Mathematical Analysis, 56(6):7713–7747, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1652325.

Baratchart:2021:DVF

[BGK21] L. Baratchart, C. Gerhards, and A. Kegeles. Decomposition of L^{2}-vector fields on Lipschitz surfaces: Characterization via null-spaces of the scalar potential. SIAM Journal on Mathematical Analysis, 53(4):4096– 4117, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buttazzo:2023:MOP

[BGL23] Giuseppe Buttazzo, Maria Stella Gelli, and Danka Lucić. Mass optimization problem with convex cost. SIAM Journal on Mathematical Analysis, 55 (5):5617–5642, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1493525.

Braga:2020:KBG

[BGMW20] J. Ederson M. Braga, Diego E. M. Gomes, Diego Moreira, and Lihe Wang. Krylov’s boundary gradient type estimates for solutions to fully nonlinear differential inequalities with quadratic growth on the gradient. SIAM Journal on Mathematical Analysis, 52(5): 4469–4505, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:2022:OST

[BGS22] Alberto Bressan, Sondre T.

Galtung, and Qing Sun. Optimal shapes for tree roots. SIAM Journal on Mathematical Analysis, 54(4):4757–4784, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1440281.

Bucur:2022:SRR

[BGT22] Dorin Bucur, Alessandro Giacomini, and Paola Trebeschi. Stability results for the Robin-Laplacian on nonsmooth domains. SIAM Journal on Mathematical Analysis, 54 (4):4591–4624, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1471250.

Baldi:2022:LLM

[BH22a] Pietro Baldi and Emanuele Haus. Longer lifespan for many solutions of the Kirchhoff equation. SIAM Journal on Mathematical Analysis, 54(1):306–342, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1351515.

Burtea:2022:VCL

[BH22b] Cosmin Burtea and Boris Haspot. Vanishing capillarity limit of the Navier–Stokes– Korteweg system in one dimension with degenerate viscosity coefficient and discontinuous initial density. SIAM Journal on Mathematical Analysis, 54(2):1428–1469, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1428686.

Bramburger:2023:PFR

[BH23] Jason Bramburger and Matt Holzer. Pattern formation in random networks using graphons. SIAM Journal on Mathematical Analysis, 55 (3):2150–2185, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1455875.

Baratchart:2025:SSS

[BHG25] L. Baratchart, H. Haddar, and C. Villalobos Guillén. Silent sources on a surface for the Helmholtz equation and decomposition of L^{2} vector fields. SIAM Journal on Mathematical Analysis, 57(1):682–713, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1626578.

Bianchini:2024:SIP

[BHI24] Roberta Bianchini, Lars Eric Hientzsch, and Felice Iandoli. Strong ill-posedness in L^{∞} of the 2D Boussinesq equations in vorticity form and application to the 3D axisymmetric Euler equations. SIAM Journal on Mathematical Analysis, 56(5):5915–5968, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M159384X.

Bellassoued:2023:SEI

[BHL23] Mourad Bellassoued, Houssem Haddar, and Amal Labidi. Stability estimate for an inverse problem for the time harmonic magnetic Schrödinger operator from the near and far field patterns. SIAM Journal on Mathematical Analysis, 55(4):2475–2504, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1481956.

Bogosel:2024:ONE

[BHM24] Beniamin Bogosel, Antoine Henrot, and Marco Michetti. Optimization of Neumann eigenvalues under convexity and geometric constraints. SIAM Journal on Mathematical Analysis, 56(6):7327–7349, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1641099.

Barchiesi:2024:RAM

[BHMCR24] Marco Barchiesi, Duvan Henao, Carlos Mora-Corral, and Rémy Rodiac. A relaxation approach to the minimization of the neo-Hookean energy in 3D. SIAM Journal on Mathematical Analysis, 56(6):7830–7845, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1614547.

Borghi:2025:DMV

[BHS25] Giacomo Borghi, Michael Herty, and Andrey Stavitskiy. Dynamics of measurevalued agents in the space of probabilities. SIAM Journal on Mathematical Analysis, 57(5):5107–5134, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1675515.

Bensoussan:2025:DMF

[BHTY25] Alain Bensoussan, Ziyu Huang, Shanjian Tang, and Sheung Chi Phillip Yam. Degenerate mean field type control with linear and unbounded diffusion, and their associated equations. SIAM Journal on Mathematical Analysis, 57(5):5494–5530, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1617904.

Busuioc:2020:LEE

[BIFL20] Adriana V. Busuioc, Dragos Iftimie, Milton D. Lopes Filho, and Helena J. Nussenzveig Lopes. The limit α → 0 of the α-euler equations in the half-plane with no-slip boundary conditions and vortex sheet initial data. SIAM Journal on Mathematical Analysis, 52(5): 5257–5286, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ballew:2023:GDP

[BIL^{+}23] Joshua Ballew, Gautam Iyer, C. David Levermore, Hailiang Liu, and Robert L. Pego. Global dynamics and photon loss in the Kompaneets equation. SIAM Journal on Mathematical Analysis, 55(5):5715–5750, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1533207.

Baumstark:2023:AHF

[BJ23] Julian Baumstark and Tobias Jahnke. Approximation of high-frequency wave propagation in dispersive media. SIAM Journal on Mathematical Analysis, 55(2):1214–1245, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1474035.

Bowman:2025:GEI

[BJ25] David Bowman and Sehyun Ji. Global existence for an isotropic Landau model. SIAM Journal on Mathematical Analysis, 57(2):1406–1451, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1657262.

Baumstark:2024:PHF

[BJL24] Julian Baumstark, Tobias Jahnke, and Christian Lubich. Polarized high-frequency wave propagation beyond the nonlinear Schrödinger approximation. SIAM Journal on Mathematical Analysis, 56(1):454– 473, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Bhavna:2024:MIO

[BJP24] Bhavna, Mathew A. Johnson, and Ashish Kumar Pandey. Modulational instability in the Ostrovsky equation and related models. SIAM Journal on Mathematical Analysis, 56(6):7390–7416, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1608665.

Balanzario:2020:RAS

[BK20a] E. P. Balanzario and E. I.

Kaikina. Regularity analysis for stochastic complex Landau–Ginzburg equation with Dirichlet white-noise boundary conditions. SIAM Journal on Mathematical Analysis, 52(4):3376–3396, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Balazs:2020:DES

[BK20b] István Balázs and Tibor Krisztin. A differential equation with a state-dependent queueing delay. SIAM Journal on Mathematical Analysis, 52 (4):3697–3737, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Belishev:2021:DNS

[BK21] Mikhail I. Belishev and Dmitrii V.

Korikov. On determination of nonorientable surface via its Dirichlet-to-Neumann operator. SIAM Journal on Mathematical Analysis, 53(5):5278– 5287, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buoso:2022:BRB

[BK22] Davide Buoso and James B. Kennedy. The bilaplacian with Robin boundary conditions. SIAM Journal on Mathematical Analysis, 54(1):36–78, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1363984.

Biswas:2023:EUN

[BK23a] Anup Biswas and Saibal Khan. Existence-uniqueness for nonlinear integro-differential equations with drift in R^{d}. SIAM Journal on Mathematical Analysis, 55(5):4378–4409, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1505931.

Bresciani:2023:RMP

[BK23b] Marco Bresciani and Martin Kruzík. A reduced model for plates arising as low-energy Γ-limit in nonlinear magnetoelasticity. SIAM Journal on Mathematical Analysis, 55 (4):3108–3168, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1446836.

Bardi:2024:LTB

[BK24] Martino Bardi and Hicham Kouhkouh. Long-time behavior of deterministic mean field games with nonmonotone interactions. SIAM Journal on Mathematical Analysis, 56(4):5079–5098, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1608100.

Bae:2023:BMT

[BKPY23] Gi-Chan Bae, Christian Klingenberg, Marlies Pirner, and Seok-Bae Yun. BGK model for two-component gases near a global Maxwellian. SIAM Journal on Mathematical Analysis, 55(2):1007–1047, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469535.

Bhattacharyya:2023:MRT

[BKS23] Sombuddha Bhattacharyya, Venkateswaran P. Krishnan, and Suman K. Sahoo. Momentum ray transforms and a partial data inverse problem for a polyharmonic operator. SIAM Journal on Mathematical Analysis, 55(4):4000–4038, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1500617.

Bona:2020:KEH

[BL20] Jerry L. Bona and Jonatan Lenells. The KdV equation on the half-line: Timeperiodicity and mass transport. SIAM Journal on Mathematical Analysis, 52(2):1009– 1039, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blaasten:2021:SCR

[BL21] Emilia L. K. Blåsten and Hongyu Liu. Scattering by curvatures, radiationless sources, transmission eigenfunctions, and inverse scattering problems. SIAM Journal on Mathematical Analysis, 53(4):3801– 3837, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bies:2023:GTR

[Bl23] Piotr Micha Bies and Tomasz Cie lak. Global-in-time regular unique solutions with positive temperature to onedimensional thermoelasticity. SIAM Journal on Mathematical Analysis, 55(6):7024–7038, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Blaustein:2023:DLV

[Bla23] Alain Blaustein. Diffusive limit of the Vlasov–Poisson–Fokker– Planck model: Quantitative and strong convergence results. SIAM Journal on Mathematical Analysis, 55(5):5464–5482, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1530197.

Borthagaray:2024:QLF

[BLN24a] Juan Pablo Borthagaray, Wenbo Li, and Ricardo H. Nochetto. Quasi-linear fractional-order operators in Lipschitz domains. SIAM Journal on Mathematical Analysis, 56 (3):4006–4039, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1575871.

Brennecke:2024:SOE

[BLN24b] Christian Brennecke, Jinyeop Lee, and Phan Thành Nam. Second order expansion of Gibbs state reduced density matrices in the Gross– Pitaevskii regime. SIAM Journal on Mathematical Analysis, 56(4):5262–5284, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1608215.

Balehowsky:2022:ICT

[BLPS22] Tracey Balehowsky, Matti Lassas, Pekka Pankka, and Ville Sirviö. Invisibility cloaking and transformation optics for three dimensional manifolds and applications in cosmology. SIAM Journal on Mathematical Analysis, 54(3):3420–3456, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1431096.

Bugiera:2022:SUG

[BLS22] Lars Bugiera, Enno Lenzmann, and Jérémy Sok. On symmetry and uniqueness of ground states for linear and nonlinear elliptic PDEs. SIAM Journal on Mathematical Analysis, 54(6):6119–6135, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1487576.

Boardman:2020:SBM

[BLW20] Nicki Boardman, Hongxia Lin, and Jiahong Wu. Stabilization of a background magnetic field on a 2 dimensional magnetohydrodynamic flow. SIAM Journal on Mathematical Analysis, 52(5):5001–5035, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bao:2024:CRN

[BLZ24] Weizhu Bao, Yong Lu, and Zhifei Zhang. Convergence rates in the nonrelativistic limit of the cubic Klein– Gordon equation. SIAM Journal on Mathematical Analysis, 56(5):6822–6860, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1572933.

Bouchitte:2024:ATG

[BM24] Guy Bouchitté and Rajesh Mahadevan. Asymptotic theory for a general class of shortrange interaction functionals. SIAM Journal on Mathematical Analysis, 56(6):7170–7193, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1623306.

Bedrossian:2022:LWD

[BMM22] Jacob Bedrossian, Nader Masmoudi, and Clément Mouhot. Linearized wave-damping structure of Vlasov–Poisson in R^{3}. SIAM Journal on Mathematical Analysis, 54(4):4379–4406, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1386141.

Bulicek:2021:IHC

[BMPS21] Miroslav Bulícek, Josef Málek, Vít Prusa, and Endre Süli. On incompressible heat-conducting viscoelastic rate-type fluids with stress-diffusion and purely spherical elastic response. SIAM Journal on Mathematical Analysis, 53(4): 3985–4030, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blechta:2020:CIF

[BMR20] Jan Blechta, Josef Málek, and K. R. Rajagopal. On the classification of incompressible fluids and a mathematical analysis of the equations that govern their motion. SIAM Journal on Mathematical Analysis, 52(2): 1232–1289, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beck:2023:ECE

[BMZ23] Lisa Beck, Daniel Matthes, and Martina Zizza. Exponential convergence to equilibrium for coupled systems of nonlinear degenerate drift diffusion equations. SIAM Journal on Mathematical Analysis, 55(3):1766–1809, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466980.

Baxendale:2024:ASS

[BN24] Peter H. Baxendale and N. Sri Namachchivaya. Almost-sure stability of the single mode solution of a noisy nonlinear autoparametric system. SIAM Journal on Mathematical Analysis, 56(6):7074–7107, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1556502.

Beeson:2022:AFE

[BNP22] Ryne Beeson, Navaratnam S. Namachchivaya, and Nicolas Perkowski. Approximation of the filter equation for multiple timescale, correlated, nonlinear systems. SIAM Journal on Mathematical Analysis, 54(3):3054–3090, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1379265.

Benesova:2023:FRB

[BNSS23] Barbora Benesová, Sárka Necasová, Jan Scherz, and Anja Schlömerkemper. Fluidrigid body interaction in an incompressible electrically conducting fluid. SIAM Journal on Mathematical Analysis, 55(2):929–965, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M148255X.

Boccardo:2021:ESR

[BO21] Lucio Boccardo and Luigi Orsina. An elliptic system related to the stationary thermistor problem. SIAM Journal on Mathematical Analysis, 53(6): 6910–6931, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brandolese:2024:FRD

[BO24] Lorenzo Brandolese and Takahiro Okabe. Forced rapidly dissipative Navier–Stokes flows. SIAM Journal on Mathematical Analysis, 56(1):412–432, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Bobrowski:2020:MDT

[Bob20] Adam Bobrowski. Modeling diffusion in thin 2D layers separated by a semipermeable membrane. SIAM Journal on Mathematical Analysis, 52(4): 3222–3251, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bobrowski:2023:CNF

[Bob23] Adam Bobrowski. Concatenation of nonhonest Feller processes, exit laws, and limit theorems on graphs. SIAM Journal on Mathematical Analysis, 55(4):3457–3508, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1487552.

Bocchi:2020:FSS

[Boc20]  Edoardo Bocchi.  Floating

structures in shallow water: Local well-posedness in the axisymmetric case. SIAM Journal on Mathematical Analysis, 52(1):306–339, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:2020:PBA

[BP20] Guillaume Bal and Benjamin Palacios. Pencil-beam approximation of stationary Fokker– Planck. SIAM Journal on Mathematical Analysis, 52(4): 3487–3519, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biswas:2021:CDA

[BP21] Animikh Biswas and Randy Price. Continuous data assimilation for the threedimensional Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 53(6): 6697–6723, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bruell:2023:SPT

[BP23] Gabriele Bruell and Long Pei.

Symmetry of periodic traveling waves for nonlocal dispersive equations. SIAM Journal on Mathematical Analysis, 55(1):486–507, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1433162.

Ben-Porat:2025:DYS

[BPIR25] Immanuel Ben-Porat, Mikaela Iacobelli, and Alexandre Rege. Derivation of Yudovich solutions of incompressible Euler from the Vlasov–Poisson system. SIAM Journal on Mathematical Analysis, 57 (1):886–906, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1648417.

Bengel:2024:PEL

[BPR24] Lukas Bengel, Dmitry Pelinovsky, and Wolfgang Reichel. Pinning in the extended Lugiato–Lefever equation. SIAM Journal on Mathematical Analysis, 56 (3):3679–3702, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1550700.

Blaasten:2023:FHM

[BPS23] Emilia L. K. Blåsten, Lassi Päivärinta, and Sadia Sadique. The Fourier, Hilbert, and Mellin transforms on a halfline. SIAM Journal on Mathematical Analysis, 55(6):7529– 7548, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Bulicek:2022:EUG

[BPSS22] Miroslav Bulícek, Victoria Patel, Endre Süli, and Yasemin Sengül. Existence and uniqueness of global weak solutions to strain-limiting viscoelasticity with Dirichlet boundary data. SIAM Journal on Mathematical Analysis, 54(6):6186–6222, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1455322.

Barbu:2023:NFP

[BR23] Viorel Barbu and Michael Röckner. Nonlinear Fokker– Planck equations with timedependent coefficients. SIAM Journal on Mathematical Analysis, 55(1):1–18, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145481X.

Bigorgne:2025:MSS

[BRR25] Léo Bigorgne, Anibal Velozo Ruiz, and Renato Velozo Ruiz. Modified scattering of small data solutions for the Vlasov– Poisson system with a repulsive potential. SIAM Journal on Mathematical Analysis, 57(1):714–752, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1631018.

Banerjee:2024:CPS

[BS24a] Agnid Banerjee and Soumen Senapati. The Calderón problem for space–time fractional parabolic operators with variable coefficients. SIAM Journal on Mathematical Analysis, 56(4):4759–4810, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1584137.

Buccheri:2024:VSN

[BS24b] Stefano Buccheri and Ulisse Stefanelli. Viscosity solutions for nonlocal equations with space-dependent operators. SIAM Journal on Mathematical Analysis, 56(1):336– 373, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Benabbas:2024:WPC

[BSH24] Imen Benabbas and Belkacem Said-Houari. The Westervelt– Pennes–Cattaneo model of nonlinear thermo-acoustics: Local well–posedness and singular limit for vanishing relaxation time. SIAM Journal on Mathematical Analysis, 56(3): 2968–3003, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Boudin:2022:ECT

[BST22] Laurent Boudin, Francesco Salvarani, and Emmanuel Trélat. Exponential convergence towards consensus for non-symmetric linear firstorder systems in finite and infinite dimensions. SIAM Journal on Mathematical Analysis, 54(3):2727–2752, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1416102.

Bai:2023:LGS

[BSW23] Ruobing Bai, Jia Shen, and Yifei Wu. Large global solutions for the energy-critical nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 55 (5):4193–4218, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1495123.

Bradshaw:2021:LES

[BT21] Zachary Bradshaw and Tai-Peng Tsai. Local energy solutions to the Navier–Stokes equations in Wiener amalgam spaces. SIAM Journal on Mathematical Analysis, 53(2): 1993–2026, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Basaric:2023:WPQ

[BT23] Danica Basarić and Tong Tang. On well-posedness of quantum fluid systems in the class of dissipative solutions. SIAM Journal on Mathematical Analysis, 55(3):2434–2466, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M148985X.

Buze:2025:CHK

[Buz25] Maciej Buze. Constrained Hellinger–Kantorovich barycenters: Least-cost soft and conic multimarginal formulations. SIAM Journal on Mathematical Analysis, 57(1):495–519, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1639804.

Brenier:2020:OTM

[BV20a] Yann Brenier and Dmitry Vorotnikov. On optimal transport of matrix-valued measures. SIAM Journal on Mathematical Analysis, 52(3):2849– 2873, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Budd:2020:GMB

[BV20b] Jeremy Budd and Yves Van Gennip. Graph Merriman– Bence–Osher as a SemiDiscrete implicit Euler scheme for graph Allen–Cahn flow. SIAM Journal on Mathematical Analysis, 52(5):4101–4139, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blochas:2024:MZM

[BW24] Paul Blochas and Aric Wheeler. Majda and ZND models for detonation: Nonlinear stability vs. formation of singularities. SIAM Journal on Mathematical Analysis, 56(5):6137–6191, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1544945.

Brkic:2025:WSA

[BW25] Patrick Brkic and Emil Wiedemann. Wild solutions of the 3D axisymmetric Euler equations. SIAM Journal on Mathematical Analysis, 57(1):996–1020, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1653914.

Bonnet-Weill:2024:VIP

[BWF24] Benoît Bonnet-Weill and Hélène Frankowska. On the viability and invariance of proper sets under continuity inclusions in Wasserstein spaces. SIAM Journal on Mathematical Analysis, 56(3):2863–2914, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Bae:2020:QBM

[BY20] Gi-Chan Bae and Seok-Bae Yun. Quantum BGK model near a global Fermi–Dirac distribution. SIAM Journal on Mathematical Analysis, 52(3): 2313–2352, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Byun:2023:DCG

[BY23] Sung-Soo Byun and Meng Yang. Determinantal Coulomb gas ensembles with a class of discrete rotational symmetric potentials. SIAM Journal on Mathematical Analysis, 55(6):6867–6897, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Brull:2024:SFB

[BY24] Stephane Brull and Seok-Bae Yun. Stationary flows of the ES-BGK model with the correct Prandtl number. SIAM Journal on Mathematical Analysis, 56(5):6361–6397, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1599628.

Becker:2025:STT

[BZ25] Simon Becker and Xiaowen Zhu. Spectral theory of twisted bilayer graphene in a magnetic field. SIAM Journal on Mathematical Analysis, 57 (2):1621–1651, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M163476X.

Bian:2024:GSI

[BZGZ24] Xingzhi Bian, Peicheng Zhu, Boling Guo, and Ying Zhang. Global solutions to an initialboundary value problem of a phase-field model for motion of grain boundaries. SIAM Journal on Mathematical Analysis, 56(4):4296–4323, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1477775.

Carrillo:2024:SVC

[CaDZ24] José A. Carrillo, Xu an Dou, and Zhennan Zhou. A simplified voltage-conductance kinetic model for interacting neurons and its asymptotic limit. SIAM Journal on Mathematical Analysis, 56(1):676– 726, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Cai:2023:UBH

[Cai23] Yuan Cai. Uniform bound of the highest-order energy of the 2D incompressible elastodynamics. SIAM Journal on Mathematical Analysis, 55 (5):5893–5918, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1519894.

Carrillo:2024:DMV

[CbGwG24] José A. Carrillo, Tomasz D.

biec, Piotr Gwiazda, and Agnieszka wierczewska Gwiazda. Dissipative measure-valued solutions to the Euler–Poisson equation. SIAM Journal on Mathematical Analysis, 56(1): 304–335, January 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Crin-Barat:2023:HPC

[CBHS23] Timothée Crin-Barat, Qingyou He, and Ling-Yun Shou. The hyperbolic-parabolic chemotaxis system for vasculogenesis: Global dynamics and relaxation limit toward a Keller– Segel model. SIAM Journal on Mathematical Analysis, 55(5):4445–4492, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1477295.

Cook:2022:RCC

[CC22] Brendan Cook and Jeff Calder.

Rates of convergence for the continuum limit of nondominated sorting. SIAM Journal on Mathematical Analysis, 54(1):872–911, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1344901.

Chicaud:2023:ATH

[CC23] Damien Chicaud and Patrick Ciarlet. Analysis of timeharmonic Maxwell impedance problems in anisotropic media. SIAM Journal on Mathematical Analysis, 55(3):1969–2000, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1485413.

Chemetov:2025:OCN

[CC25] Nikolai V. Chemetov and Fernanda Cipriano. Optimal control of Newtonian fluids in a stochastic environment. SIAM Journal on Mathematical Analysis, 57(1):364–403, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1623069.

Carlier:2024:WMR

[CCE24] G. Carlier, E. Chenchene, and K. Eichinger. Wasserstein medians: Robustness, PDE characterization, and numerics. SIAM Journal on Mathematical Analysis, 56(5):6483–6520, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1624786.

Carlone:2021:ETD

[CCFO21] Raffaele Carlone, Michele Correggi, Marco Falconi, and Marco Olivieri. Emergence of time-dependent point interactions in polaron models. SIAM Journal on Mathematical Analysis, 53(4):4657–4691, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Capoferri:2023:ELD

[CCi23] Matteo Capoferri, Mikhail Cherdantsev, and Igor Vel i. Eigenfunctions localised on a defect in high-contrast random media. SIAM Journal on Mathematical Analysis, 55(6):7449–7489, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Chen:2023:QBE

[CCJT23] Hebai Chen, Xingwu Chen, Man Jia, and Yilei Tang. A quintic z_{2}-equivariant Liénard system arising from the complex Ginzburg–Landau equation. SIAM Journal on Mathematical Analysis, 55(6):5993– 6038, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Chen:2021:HSS

[CCKW21] Xin Chen, Zhen-Qing Chen, Takashi Kumagai, and Jian Wang. Homogenization of symmetric stable-like processes in stationary ergodic media. SIAM Journal on Mathematical Analysis, 53(3):2957–3001, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cox:2023:HSF

[CCLM23] Graham Cox, Mitchell Curran, Yuri Latushkin, and Robert Marangell. Hamiltonian spectral flows, the Maslov index, and the stability of standing waves in the nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 55(5):4998–5050, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1533797.

Chicaud:2021:AVF

[CCM21] Damien Chicaud, Patrick Ciarlet, Jr., and Axel Modave. Analysis of variational formulations and low-regularity solutions for time-harmonic electromagnetic problems in complex anisotropic media. SIAM Journal on Mathematical Analysis, 53(3):2691–2717, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2022:HOR

[CCS22] Xinfu Chen, John Chadam, and David Saunders. Higherorder regularity of the free boundary in the inverse firstpassage problem. SIAM Journal on Mathematical Analysis, 54(4):4695–4720, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466797.

Carrillo:2023:MFL

[CCS23] José A. Carrillo, Andrea Clini, and Susanne Solem. The mean field limit of stochastic differential equation systems modeling grid cells. SIAM Journal on Mathematical Analysis, 55(4):3602–3634, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465640.

Cai:2024:LOT

[CCS24] Hong Cai, Geng Chen, and Yannan Shen. Lipschitz optimal transport metric for a wave system modeling nematic liquid crystals. SIAM Journal on Mathematical Analysis, 56(4):5144–5174, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1629547.

Cai:2023:SFR

[CCW23] Hong Cai, Geng Chen, and Tian-Yi Wang. Singularity formation for radially symmetric expanding wave of compressible Euler equations.

SIAM Journal on Mathematical Analysis, 55(4):2917–2947, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1487692.

Chen:2021:FSE

[CCZ21] Geng Chen, Gui-Qiang G. Chen, and Shengguo Zhu. Formation of singularities and existence of global continuous solutions for the compressible Euler equations. SIAM Journal on Mathematical Analysis, 53(6):6280–6325, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2022:GSV

[CCZ22] Zili Chen, Jing Chen, and Xianwen Zhang. Global solutions of the Vlasov–Poisson system with a radiation damping term for general initial data. SIAM Journal on Mathematical Analysis, 54(1):693–722, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1366101.

Colombo:2020:RTH

[CD20] Maria Colombo and Luigi De Rosa. Regularity in time of Hölder solutions of Euler and hypodissipative Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 52(1): 221–238, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2021:OII

[CD21] Jianfeng Cheng and Lili Du.

Oblique injection of incompressible ideal fluid from a slot into a free stream. SIAM Journal on Mathematical Analysis, 53(1):754–794, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coghi:2022:MVS

[CDF^{+}22] Michele Coghi, Wolfgang Dreyer, Peter K. Friz, Paul Gajewski, Clemens Guhlke, and Mario Maurelli. A McKean–Vlasov SDE and particle system with interaction from reflecting boundaries. SIAM Journal on Mathematical Analysis, 54(2):2251–2294, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1409421.

Corona:2023:NAV

[CDG23] D. Corona, A. Della Corte, and F. Giannoni. A nonautonomous variational problem describing a nonlinear Timoshenko beam. SIAM Journal on Mathematical Analysis, 55(5):4664–4694, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1493653.

Candiani:2020:MBR

[CDGH20] Valentina Candiani, Jérémi Dardé, Henrik Garde, and Nuutti Hyvönen. Monotonicitybased reconstruction of extreme inclusions in electrical impedance tomography. SIAM Journal on Mathematical Analysis, 52(6):6234–6259, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2022:ORM

[CDL22] Jongkeun Choi, Hongjie Dong, and Zongyuan Li. Optimal regularity of mixed Dirichlet-conormal boundary value problems for parabolic operators. SIAM Journal on Mathematical Analysis, 54 (2):1393–1427, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1461344.

Cao:2024:VPB

[CDL24] Chuqi Cao, Dingqun Deng, and Xingyu Li. The Vlasov– Poisson–Boltzmann/Landau system with polynomial perturbation near Maxwellian. SIAM Journal on Mathematical Analysis, 56(1):820–876, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Casado-Diaz:2020:DRP

[CDLLSG20] Juan Casado-Díaz, Manuel Luna-Laynez, and Francisco J. Suárez-Grau. A decomposition result for the pressure of a fluid in a thin domain and extensions to elasticity problems. SIAM Journal on Mathematical Analysis, 52(3):2201–2236, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2021:NGS

[CDLZ21] Xinlin Cao, Huaian Diao, Hongyu Liu, and Jun Zou. On novel geometric structures of Laplacian eigenfunctions in R^{3} and applications to inverse problems. SIAM Journal on Mathematical Analysis, 53(2): 1263–1294, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cesana:2024:SMS

[CDM24a] Pierluigi Cesana, Lucia De Luca, and Marco Morandotti. Semidiscrete modeling of systems of wedge disclinations and edge dislocations via the Airy stress function method. SIAM Journal on Mathematical Analysis, 56(1):79–136, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Conti:2024:ORE

[CDM24b] Sergio Conti, Georg Dolzmann, and Stefan Müller. Optimal rigidity estimates for maps of a compact Riemannian manifold to itself. SIAM Journal on Mathematical Analysis, 56(6):8070–8095, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1650168.

Capuani:2021:MEF

[CDN21] Rossana Capuani, Prerona Dutta, and Khai T. Nguyen. Metric entropy for functions of bounded total generalized variation. SIAM Journal on Mathematical Analysis, 53(1): 1168–1190, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Constantin:2020:EHE

[CDS20] Peter Constantin, Theodore D. Drivas, and Roman Shvydkoy. Entropy hierarchies for equations of compressible fluids and self-organized dynamics. SIAM Journal on Mathematical Analysis, 52(3):3073– 3092, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2021:FKT

[CDX21] Zhen-Qing Chen, Weihua Deng, and Pengbo Xu. Feynman– Kac transform for anomalous processes. SIAM Journal on Mathematical Analysis, 53(5): 6017–6047, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2022:GES

[CDX22] Jongkeun Choi, Hongjie Dong, and Longjuan Xu. Gradient estimates for Stokes and Navier–Stokes systems with piecewise DMO coefficients. SIAM Journal on Mathematical Analysis, 54(3):3609–3635, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1423518.

Chen:2024:SNS

[CDZ24] Weiquan Chen, Zhao Dong, and Xiangchan Zhu. Sharp nonuniqueness of solutions to stochastic Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 56(2):2248– 2285, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Caetano:2023:RSE

[CEGP23] Diogo Caetano, Charles M. Elliott, Maurizio Grasselli, and Andrea Poiatti. Regularization and separation for evolving surface Cahn–Hilliard equations. SIAM Journal on Mathematical Analysis, 55(6):6625– 6675, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Carlier:2021:EWB

[CEK21] Guillaume Carlier, Katharina Eichinger, and Alexey Kroshnin. Entropic-Wasserstein barycenters: PDE characterization, regularity, and CLT.

SIAM Journal on Mathematical Analysis, 53(5):5880–5914, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2021:CSS

[CEW21] Huyuan Chen, Gilles Evéquoz, and Tobias Weth. Complex solutions and stationary scattering for the nonlinear Helmholtz equation. SIAM Journal on Mathematical Analysis, 53(2): 2349–2372, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crasta:2020:BFB

[CF20] Graziano Crasta and Ilaria Fragalà. Bernoulli free boundary problem for the infinity Laplacian. SIAM Journal on Mathematical Analysis, 52(1): 821–844, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cardaliaguet:2021:HMT

[CF21] Pierre Cardaliaguet and Nicolas Forcadel. From heterogeneous microscopic traffic flow models to macroscopic models. SIAM Journal on Mathematical Analysis, 53(1):309– 322, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chauleur:2022:APW

[CF22] Quentin Chauleur and Erwan Faou. Around plane waves solutions of the Schrödinger– Langevin equation. SIAM Journal on Mathematical Analysis, 54(5):5103–5125, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458387.

Cuba:2025:EAV

[CF25] Edison Cuba and Lucas C. F.

Ferreira. Existence of asymmetric vortex patch for the generalized SQG equations. SIAM Journal on Mathematical Analysis, 57(1):118–152, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1543355.

Cristoferi:2025:HPS

[CFG25] Riccardo Cristoferi, Irene Fonseca, and Likhit Ganedi. Homogenization and phase separation with fixed wells — the supercritical case. SIAM Journal on Mathematical Analysis, 57(2):2138–2173, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1571861.

Chen:2020:LRS

[CFHX20] Gui-Qiang Chen, Mikhail Feldman, Jingchen Hu, and Wei Xiang. Loss of regularity of solutions of the Lighthill problem for shock diffraction for potential flow. SIAM Journal on Mathematical Analysis, 52(2): 1096–1114, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2025:ANV

[CFP25] Young-Pil Choi, Michał Fabisiak, and Jan Peszek. Alignment with nonlinear velocity couplings: Collision avoidance and micro-to-macro meanfield limits. SIAM Journal on Mathematical Analysis, 57 (5):5791–5820, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1697360.

Chafai:2021:CGU

[CFS21] Djalil Chafaï, Grégoire Ferré, and Gabriel Stoltz. Coulomb gases under constraint: Some theoretical and numerical results. SIAM Journal on Mathematical Analysis, 53(1):181– 220, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colli:2022:CHE

[CFS22] Pierluigi Colli, Takeshi Fukao, and Luca Scarpa. The Cahn– Hilliard equation with forwardbackward dynamic boundary condition via vanishing viscosity. SIAM Journal on Mathematical Analysis, 54 (3):3292–3315, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M142441X.

Chaumont-Frelet:2023:SFL

[CFS23] T. Chaumont-Frelet and E. A.

Spence. Scattering by finely layered obstacles: Frequencyexplicit bounds and homogenization. SIAM Journal on Mathematical Analysis, 55 (2):1319–1363, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450136.

Caroccia:2020:IRV

[CFV20] M. Caroccia, M. Focardi, and N. Van Goethem. On the integral representation of variational functionals on BD. SIAM Journal on Mathematical Analysis, 52(4):4022–4067, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Calsavara:2020:EGT

[CGG20] Bianca M. Calsavara and Francisco Guillen-Gonzalez. Existence of global-in-time weak solutions for a solidification model with convection in the liquid and rigid motion in the solid. SIAM Journal on Mathematical Analysis, 52(6):6260– 6280, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cances:2023:SOH

[CGG23] Éric Cancès, Louis Garrigue, and David Gontier. Secondorder homogenization of periodic Schrödinger operators with highly oscillating potentials. SIAM Journal on Mathematical Analysis, 55 (3):2288–2323, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1478446.

Chen:2024:STD

[CGJL24] Bochao Chen, Yixian Gao, Shuguan Ji, and Yang Liu. Stability for time-domain elastic wave equations. SIAM Journal on Mathematical Analysis, 56(1):433–453, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Chone:2023:WOT

[CGK23] Philippe Choné, Nathael Gozlan, and Francis Kramarz. Weak optimal transport with unnormalized kernels. SIAM Journal on Mathematical Analysis, 55(6):6039– 6092, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Chen:2023:WPS

[CGS23] Jie Chen, Boling Guo, and Jie Shao. Well-posedness and scattering for the generalized Boussinesq equation. SIAM Journal on Mathematical Analysis, 55(1):133–161, ???? 2023. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1464993.

Colli:2024:CEP

[CGSS24] Pierluigi Colli, Gianni Gilardi, Andrea Signori, and Jürgen Sprekels. Curvature effects in pattern formation: Wellposedness and optimal control of a sixth-order Cahn– Hilliard equation. SIAM Journal on Mathematical Analysis, 56(4):4928–4969, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1630372.

Carstea:2023:IPP

[CGU23] Catalin I. Cârstea, Tuhin Ghosh, and Gunther Uhlmann. An inverse problem for the porous medium equation with partial data and a possibly singular absorption term. SIAM Journal on Mathematical Analysis, 55(1):162–185, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465573.

Canizo:2021:SGG

[CGY21] José A. Cañizo, Pierre Gabriel, and Havva Yoldas. Spectral gap for the growthfragmentation equation via Harris’s theorem. SIAM Journal on Mathematical Analysis, 53(5):5185–5214, ???? 2021.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2020:SDQ

[CGZ20] Xing Cheng, Zihua Guo, and Zehua Zhao. On scattering for the defocusing quintic nonlinear Schrödinger equation on the two-dimensional cylinder. SIAM Journal on Mathematical Analysis, 52(5):4185–4237, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chamorro:2021:RTD

[CH21a] Diego Chamorro and Jiao He. Regularity theory for the dissipative solutions of the MagnetoHydroDynamic equations. SIAM Journal on Mathematical Analysis, 53(5):5288–5321, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Corcho:2021:CPA

[CH21b] Adán J. Corcho and Hichem Hajaiej. On the Cauchy problem associated with a nonequilibrium Bose–Einstein condensate of exciton polaritons. SIAM Journal on Mathematical Analysis, 53(6):6191–6215, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2023:IJF

[CH23] Jianfeng Cheng and Jinli Huang. Incompressible jet flow past an obstacle. SIAM Journal on Mathematical Analysis, 55(6):7077–7112, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Chaturvedi:2023:LEL

[Cha23] Sanchit Chaturvedi. Local existence for the Landau equation with hard potentials. SIAM Journal on Mathematical Analysis, 55(5):5345–5385, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1490107.

Carrillo:2024:CBL

[CHaW24] José A. Carrillo, Guangyi Hong, and Zhi an Wang. Convergence of boundary layers of chemotaxis models with physical boundary conditions I: Degenerate initial data. SIAM Journal on Mathematical Analysis, 56(6):7576–7643, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1628426.

Chen:2022:RBE

[Che22]  Hongxu Chen.  Regularity

of Boltzmann equation with Cercignani–Lampis boundary in convex domain. SIAM Journal on Mathematical Analysis, 54(3):3316–3378, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1421635.

Chihara:2022:MAP

[Chi22] Hiroyuki Chihara. Microlocal analysis of d-plane transform on the Euclidean space. SIAM Journal on Mathematical Analysis, 54(6):6254–6287, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1442322.

Cao:2024:PMS

[CHJ24] Chuqi Cao, Ling-Bing He, and Jie Ji. Propagation of moments and sharp convergence rate for inhomogeneous noncutoff Boltzmann equation with soft potentials. SIAM Journal on Mathematical Analysis, 56 (1):1321–1426, February 2024. CODEN SJMAAH. ISSN 1095-7154.

Cho:2023:SNH

[CHL23] Yonggeun Cho, Seokchang Hong, and Kiyeon Lee. Scattering and nonscattering of the Hartree-type nonlinear Dirac system at critical regularity. SIAM Journal on Mathematical Analysis, 55(4):3395–3419, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1437135.

Chou:2022:GES

[CHLL22] Shih-Wei Chou, John M. Hong, Hsin-Yi Lee, and Ying-Chieh Lin. Global entropy solutions and zero relaxation limit for Greenberg–Klar–Rascle multilane traffic flow model. SIAM Journal on Mathematical Analysis, 54(6):5949–5980, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1473716.

Cakoni:2025:LER

[CHLV25] Fioralba Cakoni, Narek Hovsepyan, Matti Lassas, and Michael Vogelius. On the lack of external response of a nonlinear medium in the secondharmonic generation process. SIAM Journal on Mathematical Analysis, 57(2):1370–1405, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1669700.

Chou:2025:GCS

[CHLW25] Shih-Wei Chou, John M. Hong, Hsin-Yi Lee, and Shih-Ming Wang. Global classical solutions for nonisentropic gas in nozzle flows. SIAM Journal on Mathematical Analysis, 57(2):1990–2012, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1691764.

Cipriani:2022:ZIL

[CHQ22] Cristina Cipriani, Hui Huang, and Jinniao Qiu. Zero-inertia limit: From particle swarm optimization to consensus-based optimization. SIAM Journal on Mathematical Analysis, 54(3):3091–3121, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1412323.

Chen:2022:GSW

[CHS22] Yazhou Chen, Bin Huang, and Xiaoding Shi. Global strong and weak solutions to the initial-boundary-value problem of two-dimensional compressible MHD system with large initial data and vacuum. SIAM Journal on Mathematical Analysis, 54(3):3817–3847, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1453438.

Chu:2023:HSL

[Chu23] Raymond Chu. A Hele–Shaw limit with a variable upper bound and drift. SIAM Journal on Mathematical Analysis, 55(5):4938–4976, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1482743.

Chen:2024:ILE

[CHW24] Gui-Qiang G. Chen, Fei-Min Huang, and Tian-Yi Wang. Isothermal limit of entropy solutions of the Euler equations for isentropic gas dynamics. SIAM Journal on Mathematical Analysis, 56(1):1300–1320, February 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Chen:2025:GSC

[CHWY25] Gui-Qiang G. Chen, Lin He, Yong Wang, and Difan Yuan. Global solutions of the compressible Euler–Poisson equations for plasma with doping profile for large initial data of spherical symmetry. SIAM Journal on Mathematical Analysis, 57(2):1227–1279, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1605806.

Chen:2024:SFF

[CHX24a] Geng Chen, Tao Huang, and Xiang Xu. Singularity formation for full Ericksen–Leslie system of nematic liquid crystal flows in dimension two. SIAM Journal on Mathematical Analysis, 56(3):3968–4005, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1571046.

Cheng:2024:UNS

[CHX24b] Tingzhi Cheng, Genggeng Huang, and Xianghui Xu. Uniqueness of nontrivial solutions for degenerate Monge– Ampere equations. SIAM Journal on Mathematical Analysis, 56(1):234–253, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Choi:2021:ASP

[CHXZ21] Young-Pil Choi, Seung-Yeal Ha, Qinghua Xiao, and Yinglong Zhang. Asymptotic stability of the phase-homogeneous solution to the Kuramoto– Sakaguchi equation with inertia. SIAM Journal on Mathematical Analysis, 53(3):3188– 3235, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2024:IBV

[CHZ24] Geng Chen, Yanbo Hu, and Qingtian Zhang. Initialboundary value problems for Poiseuille flow of nematic liquid crystal via full Ericksen– Leslie model. SIAM Journal on Mathematical Analysis, 56(2): 1809–1850, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Ciarlet:2020:AEF

[Cia20] Patrick Ciarlet, Jr. On the approximation of electromagnetic fields by edge finite elements. Part 3: Sensitivity to coefficients. SIAM Journal on Mathematical Analysis, 52(3): 3004–3038, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2023:NSS

[CiDRdTY23] Sitong Chen, Vicen iu D. R. dulescu, Xianhua Tang, and Shuai Yuan. Normalized solutions for Schrödinger equations with critical exponential growth in r^{2}. SIAM Journal on Mathematical Analysis, 55(6):7704– 7740, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Ciosmak:2021:MHI

[Cio21] Krzysztof J. Ciosmak. Matrix Hölder’s inequality and divergence formulation of optimal transport of vector measures. SIAM Journal on Mathematical Analysis, 53(6):6932–6958, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2021:GBW

[CJ21] Wentao Cao and Peng Jiang.

Global bounded weak entropy solutions to the Euler–Vlasov equations in fluid-particle system. SIAM Journal on Mathematical Analysis, 53(4):3958– 3984, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2024:MEE

[CJ24] Young-Pil Choi and Jinwook Jung. Modulated energy estimates for singular kernels and their applications to asymptotic analyses for kinetic equations. SIAM Journal on Mathematical Analysis, 56(2):1525– 1559, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Chowdhury:2024:FNP

[CJK24] Indranil Chowdhury, Espen R.

Jakobsen, and Miłosz Krupski.

On fully nonlinear parabolic mean field games with nonlocal and local diffusions. SIAM Journal on Mathematical Analysis, 56(5):6302–6336, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1615528.

Carlen:2021:ASE

[CJL21] Eric A. Carlen, Ian Jauslin, and Elliott H. Lieb. Analysis of a simple equation for the ground state of the Bose gas II: Monotonicity, convexity, and condensate fraction. SIAM Journal on Mathematical Analysis, 53(5):5322–5360, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2023:UMV

[CJL23] Mi-Ran Choi, Sangdon Jin, and Young-Ran Lee. Uniqueness of a minimizer for a variational problem related to the dispersion managed nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 55(4):2595–2614, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1499972.

Ciarlet:2022:MSH

[CK22] Patrick Ciarlet and Maryna Kachanovska. A mathematical study of a hyperbolic metamaterial in free space. SIAM Journal on Mathematical Analysis, 54(2):2216–2250, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1404223.

Cardone:2023:SLD

[CK23a] Giuseppe Cardone and Andrii Khrabustovskyi. Spectrum of the Laplacian on a domain perturbed by small resonators. SIAM Journal on Mathematical Analysis, 55(4):3677–3712, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M148207X.

Chang:2023:LRN

[CK23b] Tongkeun Chang and Kyungkeun Kang. Local regularity near boundary for the Stokes and Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 55(5):5051–5085, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M150397X.

Chung:2020:DNF

[CKKP20] Jaywan Chung, Yong-Jung Kim, Ohsang Kwon, and Xingbin Pan. Discontinuous nonlinearity and finite time extinction. SIAM Journal on Mathematical Analysis, 52(1): 894–926, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coron:2020:NTE

[CKP20] Jean-Michel Coron, Alexander Keimer, and Lukas Pflug. Nonlocal transport equations — existence and uniqueness of solutions and relation to the corresponding conservation laws. SIAM Journal on Mathematical Analysis, 52(6):5500–5532, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cornean:2022:TDQ

[CKP^{+}22] Horia Cornean, David Krejcirík, Thomas G. Pedersen, Nicolas Raymond, and Edgardo Stockmeyer. On the two-dimensional quantum confined Stark effect in strong electric fields. SIAM Journal on Mathematical Analysis, 54(2):2114–2127, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1386712.

Choi:2024:RTP

[CKP24] Jae-Hwan Choi, Jaehoon Kang, and Daehan Park. A regularity theory for parabolic equations with anisotropic nonlocal operators in l_{q}(l_{p}) spaces. SIAM Journal on Mathematical Analysis, 56(1): 1264–1299, February 2024. CODEN SJMAAH. ISSN 1095-7154.

Chen:2020:SRC

[CKW20] Lin Chen, Fanze Kong, and Qi Wang. Stationary ring and concentric-ring solutions of the Keller–Segel model with quadratic diffusion. SIAM Journal on Mathematical Analysis, 52(5):4565–4615, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2021:SCS

[CKZ21] Gui-Qiang G. Chen, Jie Kuang, and Yongqian Zhang. Stability of conical shocks in the three-dimensional steady supersonic isothermal flows past Lipschitz perturbed cones. SIAM Journal on Mathematical Analysis, 53(3):2811–2862, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlier:2020:DAM

[CL20] Guillaume Carlier and Maxime Laborde. A differential approach to the multi-marginal Schrödinger system. SIAM Journal on Mathematical Analysis, 52(1):709–717, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cantrell:2021:ESD

[CL21a] Robert Stephen Cantrell and King-Yeung Lam. On the evolution of slow dispersal in MultiSpecies communities. SIAM Journal on Mathematical Analysis, 53(4):4933–4964, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2021:ABS

[CL21b] Zhengzheng Chen and Yeping Li. Asymptotic behavior of solutions to an impermeable wall problem of the compressible fluid models of Korteweg type with density-dependent viscosity and capillarity. SIAM Journal on Mathematical Analysis, 53(2):1434–1473, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheskidov:2021:ADA

[CL21c] Alexey Cheskidov and Xiaoyutao Luo. Anomalous dissipation, anomalous work, and energy balance for the Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 53(4):3856–3887, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ciotti:2021:PDC

[CL21d] Benjamin Ciotti and Bo Li. Passing from discrete to continuum models of electrostatic energy. SIAM Journal on Mathematical Analysis, 53(4): 4568–4604, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2022:LSS

[CL22a] Yue Cao and Yachun Li. Local strong solutions to the full compressible Navier–Stokes system with temperaturedependent viscosity and heat conductivity. SIAM Journal on Mathematical Analysis, 54 (5):5588–5628, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419544.

Chiarini:2022:SEF

[CL22b] Leandro Chiarini and Claudio Landim. Schauder estimates for fractional Laplacians and NonLocal, onedimensional singular SPDEs. SIAM Journal on Mathematical Analysis, 54(2):2163–2215, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1382829.

Cheng:2023:SBT

[CL23] Mengyu Cheng and Zhenxin Liu. The Second Bogolyubov Theorem and global averaging principle for SPDEs with monotone coefficients. SIAM Journal on Mathematical Analysis, 55(2):1100–1144, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443698.

Charlier:2024:BEW

[CL24] C. Charlier and J. Lenells. Boussinesq’s equation for water waves: Asymptotics in sector V. SIAM Journal on Mathematical Analysis, 56 (3):4104–4142, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1587671.

Clark:2025:MRC

[CL25] Carrie Clark and Richard S. Laugesen. Maximizing Riesz capacity ratios: Conjectures and theorems. SIAM Journal on Mathematical Analysis, 57(5):5683–5717, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M171992X.

Claeys:2023:NOS

[Cla23] Xavier Claeys. NonLocal optimized Schwarz method for the Helmholtz equation with physical boundaries. SIAM Journal on Mathematical Analysis, 55(6):7490–7512, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Chen:2023:RTS

[CLLZ23] Ziang Chen, Jianfeng Lu, Yulong Lu, and Shengxuan Zhou. A regularity theory for static Schrödinger equations on R^{d} in spectral Barron spaces.

SIAM Journal on Mathematical Analysis, 55(1):557–570, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1478719.

Chen:2023:QNL

[CLMZ23] Liang Chen, Dongfang Li, Ming Mei, and Guojing Zhang. Quasi-neutral limit to steadystate hydrodynamic model of semiconductors with degenerate boundary. SIAM Journal on Mathematical Analysis, 55(4):2813–2837, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1504135.

Crismale:2024:VAR

[CLR24] Vito Crismale, Giuliano Lazzaroni, and Riccarda Rossi. On the viscoplastic approximation of a rate-independent coupled elastoplastic damage model. SIAM Journal on Mathematical Analysis, 56(6):7940–7988, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1620429.

Chen:2020:EEC

[CLWX20] Ming Chen, Zhilei Liang, Dehua Wang, and Runzhang Xu. Energy equality in compressible fluids with physical boundaries. SIAM Journal on Mathematical Analysis, 52(2):1363– 1385, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2023:SRD

[CLYaY23] Yuhui Chen, Minling Li, Qinghe Yao, and Zheng an Yao. Sharp rates of decay and global-in-time stability of large solutions to the three-dimensional incompressible Phan-Thien–Tanner system of polymeric flows. SIAM Journal on Mathematical Analysis, 55(5):4537–4569, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1435161.

Cao:2022:GRS

[CLZ22] Yue Cao, Hao Li, and Shengguo Zhu. Global regular solutions for one-dimensional degenerate compressible Navier– Stokes equations with large data and far field vacuum. SIAM Journal on Mathematical Analysis, 54(4):4658–4694, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1464609.

Caro:2020:OLW

[CM20a] Pedro Caro and Cristóbal J. Meroño. The observational limit of wave packets with noisy measurements. SIAM Journal on Mathematical Analysis, 52(5):5196–5212, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cogar:2020:MET

[CM20b] Samuel Cogar and Peter B.

Monk. Modified electromagnetic transmission eigenvalues in inverse scattering theory. SIAM Journal on Mathematical Analysis, 52(6):6412–6441, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chiba:2022:SBM

[CM22] Hayato Chiba and Georgi S.

Medvedev. Stability and bifurcation of mixing in the Kuramoto model with inertia. SIAM Journal on Mathematical Analysis, 54(2):1797–1819, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1427000.

Campos:2023:TSI

[CM23a] Luccas Campos and Jason Murphy. Threshold solutions for the intercritical inhomogeneous NLS. SIAM Journal on Mathematical Analysis, 55(4):3807–3843, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1497663.

Casteras:2023:HDC

[CM23b] Jean-Baptiste Casteras and Léonard Monsaingeon. Hidden dissipation and convexity for Kimura equations. SIAM Journal on Mathematical Analysis, 55(6):7361–7398, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Cavalcante:2023:ASK

[CM23c] Márcio Cavalcante and Claudio Muñoz. Asymptotic stability of KdV solitons on the half-line: a study in the energy space. SIAM Journal on Mathematical Analysis, 55 (5):4337–4377, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M146377X.

Cucchi:2020:CHM

[CMM20] Alessandro Cucchi, Antoine Mellet, and Nicolas Meunier. A Cahn–Hilliard model for cell motility. SIAM Journal on Mathematical Analysis, 52(4): 3843–3880, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuccagna:2024:ASC

[CMMS24] Scipio Cuccagna, Masaya Maeda, Federico Murgante, and Stefano Scrobogna. On asymptotic stability on a center hypersurface at the soliton for even solutions of the nonlinear Klein–Gordon equation when 2 ≥ p\frac{5}{3}. SIAM Journal on Mathematical Analysis, 56(4):5445–5473, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1590871.

Caravenna:2025:HRC

[CMP25] L. Caravenna, E. Marconi, and A. Pinamonti. Hölder regularity of continuous solutions to balance laws and applications in the Heisenberg group. SIAM Journal on Mathematical Analysis, 57(1):979–995, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M165291X.

Cupps:2021:UBR

[CMT21] Brian P. Cupps, Jeff Morgan, and Bao Quoc Tang. Uniform boundedness for reaction– diffusion systems with mass dissipation. SIAM Journal on Mathematical Analysis, 53(1): 323–350, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannarsa:2023:BCD

[CMU23] Piermarco Cannarsa, Patrick Martinez, and Cristina Urbani. Bilinear control of a degenerate hyperbolic equation. SIAM Journal on Mathematical Analysis, 55(6):6517–6553, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Chickering:2023:ASS

[CMVP23] Kyle R. Chickering, Ryan C.

Moreno-Vasquez, and Gavin Pandya. Asymptotically selfsimilar shock formation for 1D fractal Burgers’ equation. SIAM Journal on Mathematical Analysis, 55(6):7328–7360, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Chae:2022:GRN

[CMX22] Dongho Chae, Qianyun Miao, and Liutang Xue. Global regularity of NonDiffusive temperature fronts for the twodimensional viscous Boussinesq system. SIAM Journal on Mathematical Analysis, 54(4):4043–4103, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1457345. See erra-

tum [CMX24].

Chae:2024:EGR

[CMX24] Dongho Chae, Qianyun Miao, and Liutang Xue. Erratum: Global regularity of nondiffusive temperature fronts for the 2D viscous Boussinesq system. SIAM Journal on Mathematical Analysis, 56(6):8125–8130, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1637453. See [CMX22].

Chen:2022:TSS

[CMZZ22] Liang Chen, Ming Mei, Guojing Zhang, and Kaijun Zhang. Transonic steady-states of Euler–Poisson equations for semiconductor models with sonic boundary. SIAM Journal on Mathematical Analysis, 54(1):363–388, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1416369.

Cakoni:2021:DTE

[CN21] Fioralba Cakoni and Hoai-Minh Nguyen. On the discreteness of transmission eigenvalues for the Maxwell equations. SIAM Journal on Mathematical Analysis, 53(1):888– 913, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2023:PPI

[CN23] Ke Chen and Quoc-Hung Nguyen. The Peskin problem with {⌢ \atop b_{∞,∞}^{1}} initial data. SIAM Journal on Mathematical Analysis, 55(6):6262–6304, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Chambolle:2024:GFC

[CN24] Antonin Chambolle and Matteo Novaga. L^{1}-gradient flow of convex functionals. SIAM Journal on Mathematical Analysis, 56(5):5747–5781, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1527556.

Cardone:2023:AES

[CNT23] Giuseppe Cardone, Sergey A. Nazarov, and Jari Taskinen. Asymptotic expansions of solutions to the Poisson equation with alternating boundary conditions on an open arc. SIAM Journal on Mathematical Analysis, 55(6):6940–6964, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Carstea:2020:UEG

[CNW20] Catalin I. Cârstea, Tu Nguyen, and Jenn-Nan Wang. Uniqueness estimates for the general complex conductivity equation and their applications to inverse problems. SIAM Journal on Mathematical Analysis, 52 (1):570–580, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2021:GDO

[CNW21] Yan-Yu Chen, Hirokazu Ninomiya, and Chang-Hong Wu. Global dynamics on onedimensional excitable media. SIAM Journal on Mathematical Analysis, 53(6):7081–7112, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Comech:2022:SCS

[Com22] Andrew Comech. On solutions with compact spectrum to nonlinear Klein–Gordon and Schrödinger equations. SIAM Journal on Mathematical Analysis, 54(2):2128–2141, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1411330.

Cooperman:2023:EMS

[Coo23] William Cooperman. Exponential mixing by shear flows. SIAM Journal on Mathematical Analysis, 55(6):7513–7528, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Correia:2024:SLW

[COS24] Simão Correia, Filipe Oliveira, and Jorge Drumond Silva. Sharp local well-posedness and nonlinear smoothing for dispersive equations through frequency-restricted estimates. SIAM Journal on Mathematical Analysis, 56(4):5604–5633, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M156923X.

Covi:2022:UFC

[Cov22] Giovanni Covi. Uniqueness for the fractional Calderón problem with quasilocal perturbations. SIAM Journal on Mathematical Analysis, 54 (6):6136–6163, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1478641.

Chata:2024:END

[CP24] Juan Carlos Ortiz Chata and Francesco Petitta. Existence, nonexistence, and degeneracy of limit solutions to p-Laplace problems involving Hardy potentials as p → 1^{+}. SIAM Journal on Mathematical Analysis, 56(6):7014–7042, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1634850.

Cardenas:2024:DBE

[CPW24] Esteban Cárdenas, Natasa Pavlović, and William Warner. Derivation of a Boltzmann equation with higher-order collisions from a generalized Kac model. SIAM Journal on Mathematical Analysis, 56 (4):5409–5444, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1606150.

Chen:2025:MAA

[CQ25] Xinfu Chen and Yuanwei Qi. Mathematical analysis of an acid induced cancer invasion model-the existence of traveling wave. SIAM Journal on Mathematical Analysis, 57 (5):5342–5357, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1660358.

Crismale:2021:BVS

[CR21] Vito Crismale and Riccarda Rossi. Balanced viscosity solutions to a rate-independent coupled elasto-plastic damage system. SIAM Journal on Mathematical Analysis, 53(3): 3420–3492, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carter:2021:PRA

[CRS21] Paul Carter, Jens D. M.

Rademacher, and Björn Sandstede. Pulse replication and accumulation of eigenvalues. SIAM Journal on Mathematical Analysis, 53(3):3520–3576, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Covi:2024:SEI

[CRTZ24] Giovanni Covi, Jesse Railo, Teemu Tyni, and Philipp Zimmermann. Stability estimates for the inverse fractional conductivity problem. SIAM Journal on Mathematical Analysis, 56(2):2456–2487, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Cardaliaguet:2022:MSM

[CS22] Pierre Cardaliaguet and Panagiotis Souganidis. Monotone solutions of the master equation for mean field games with idiosyncratic noise. SIAM Journal on Mathematical Analysis, 54(4):4198–4237, ???? 2022. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1450008.

Cui:2023:SLS

[CS23] Jianbo Cui and Liying Sun. Stochastic logarithmic Schrödinger equations: Energy regularized approach. SIAM Journal on Mathematical Analysis, 55(4):3044–3080, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1442425.

Caillet:2024:DND

[CS24] Thibault Caillet and Filippo Santambrogio. Doubly nonlinear diffusive PDEs: New existence results via generalized Wasserstein gradient flows. SIAM Journal on Mathematical Analysis, 56(6):7043–7073, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1639427.

Chen:2021:SDP

[CT21] Hao Chen and Martin Traizet. Stacking disorder in periodic minimal surfaces. SIAM Journal on Mathematical Analysis, 53(1):855–887, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Claudianor:2023:EMN

[CT23] O. Alves Claudianor and Nguyen V. Thin. On existence of multiple normalized solutions to a class of elliptic problems in whole RN via Lusternik–Schnirelmann category. SIAM Journal on Mathematical Analysis, 55 (2):1264–1283, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1470694.

Calder:2022:LRG

[CTL22] Jeff Calder, Nicolás García Trillos, and Marta Lewicka. Lipschitz regularity of graph Laplacians on random data clouds. SIAM Journal on Mathematical Analysis, 54 (1):1169–1222, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1356610.

Cai:2021:GWP

[CW21] Yuan Cai and Fan Wang.

Global well-posedness of incompressible elastodynamics in three-dimensional thin domain. SIAM Journal on Mathematical Analysis, 53(6):6654– 6696, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2022:MSV

[CW22] Daomin Cao and Jie Wan.

Multiscale steady vortex patches for 2D incompressible Euler equations. SIAM Journal on Mathematical Analysis, 54(2):1488–1514, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1390529.

Cui:2024:SBM

[CW24] Shikun Cui and Wendong Wang. Suppression of blow-up in multispecies Patlak–Keller– Segel–Navier–Stokes system via the Poiseuille flow in a finite channel. SIAM Journal on Mathematical Analysis, 56(6):7683–7712, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1629122.

Chaban:2025:IQB

[CW25] J. Chaban and M. I. Weinstein. Instability of quadratic band degeneracies and the emergence of Dirac points. SIAM Journal on Mathematical Analysis, 57(2):1892–1938, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1654889.

Chandler-Wilde:2020:HFB

[CWSGS20] S. N. Chandler-Wilde, E. A. Spence, A. Gibbs, and V. P. Smyshlyaev. High-frequency bounds for the Helmholtz equation under parabolic trapping and applications in numerical analysis. SIAM Journal on Mathematical Analysis, 52(1):845–893, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2020:DVT

[CWZ20] Daomin Cao, Guodong Wang, and Weicheng Zhan. Desingularization of vortices for two-dimensional steady Euler flows via the vorticity method. SIAM Journal on Mathematical Analysis, 52(6):5363–5388, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2024:AAR

[CWZ24] Pengyu Chen, Renhai Wang, and Xuping Zhang. Asymptotically autonomous robustness of random attractors for 3D BBM equations driven by nonlinear colored noise. SIAM Journal on Mathematical Analysis, 56(1):254–274, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Chen:2022:SFP

[CXZ22] Jun Chen, Zhouping Xin, and Aibin Zang. Subsonic flows past a profile with a vortex line at the trailing edge. SIAM Journal on Mathematical Analysis, 54(1):912–939, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1384099.

Chen:2021:VSC

[CY21a] De-Han Chen and Irwin Yousept. Variational source conditions in L^{p}-spaces. SIAM Journal on Mathematical Analysis, 53(3):2863–2889, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2021:EHL

[CY21b] Young-Pil Choi and Seok-Bae Yun. Existence and hydrodynamic limit for a Paveri– Fontana type kinetic traffic model. SIAM Journal on Mathematical Analysis, 53(2): 2631–2659, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2022:STM

[CY22] Liang-Biao Chen and Peng-Fei Yao. Strain tensors and matching property on degenerated hyperbolic surfaces. SIAM Journal on Mathematical Analysis, 54(4):4930–4953, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1440906.

Cao:2023:EMB

[CYY23] Daomin Cao, Shusen Yan, and Weilin Yu. Existence of multiple boundary vortexes in planar incompressible Euler flows.

SIAM Journal on Mathematical Analysis, 55(5):5189–5220, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1500186.

Dai:2021:GLS

[Dai21a]  Jia-Yuan Dai.  Ginzburg–

Landau spiral waves in circular and spherical geometries. SIAM Journal on Mathematical Analysis, 53(1):1004–1028, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2021:NWS

[Dai21b] Mimi Dai. Nonunique weak solutions in Leray–Hopf class for the three-dimensional Hall– MHD system. SIAM Journal on Mathematical Analysis, 53 (5):5979–6016, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

David:2024:DCD

[DbMS24] Noemi David, Tomasz D. biec, Mainak Mandal, and Markus Schmidtchen. A degenerate cross-diffusion system as the inviscid limit of a nonlocal tissue growth model. SIAM Journal on Mathematical Analysis, 56(2):2090–2114, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Dhia:2022:CSH

[DCWF^{+}22] Anne-Sophie Bonnet-Ben Dhia, Simon N. Chandler-Wilde, Sonia Fliss, Christophe Hazard, Karl-Mikael Perfekt, and Yohanes Tjandrawidjaja. The complex-scaled half-space matching method. SIAM Journal on Mathematical Analysis, 54(1):512–557, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1387122.

Davoli:2024:BBM

[DDG24] Elisa Davoli, Giovanni Di Fratta, and Rossella Giorgio. A Bourgain–Brezis–Mironescu formula accounting for nonlocal antisymmetric exchange interactions. SIAM Journal on Mathematical Analysis, 56 (6):6995–7013, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1632577.

Unanue:2025:DL

[dDUFJK25] Alvaro de Diego Unanue, Gary Froyland, Oliver Junge, and Péter Koltai. A dynamic p-Laplacian. SIAM Journal on Mathematical Analysis, 57 (2):1725–1752, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1632528.

Dechicha:2024:GRA

[Dec24] Dahmane Dechicha. Gevrey regularity and analyticity for the solutions of the Vlasov–Navier–Stokes system. SIAM Journal on Mathematical Analysis, 56(6):7903–7939, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1621617.

DelZotto:2023:EDT

[Del23]  Augusto Del Zotto.  En-

hanced dissipation and transition threshold for the Poiseuille flow in a periodic strip. SIAM Journal on Mathematical Analysis, 55(5):4410–4424, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1444011.

Deng:2020:DSN

[Den20] Dingqun Deng. Dissipation and semigroup on H_{n}^{k}: Noncutoff linearized Boltzmann operator with soft potential. SIAM Journal on Mathematical Analysis, 52(3):3093–3113, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:2021:RNC

[Den21]  Dingqun Deng.  Regular-

ity of non-cutoff Boltzmann equation with hard potential. SIAM Journal on Mathematical Analysis, 53(4):4513–4536, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dipierro:2025:DEN

[DFGV25] Serena Dipierro, Alberto Farina, Giovanni Giacomin, and Enrico Valdinoci. Density estimates for a nonlocal variational model with a degenerate double-well potential via the Sobolev inequality. SIAM Journal on Mathematical Analysis, 57(5):5628–5682, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/25M1726005.

DeLuca:2025:UCC

[DFV25] Alessandra De Luca, Veronica Felli, and Stefano Vita. Unique continuation from conical boundary points for fractional equations. SIAM Journal on Mathematical Analysis, 57(1):907–950, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1663193.

Ducrot:2020:SSB

[DG20] Arnaud Ducrot and Alexandre Genadot. Self-similar behavior of a nonlocal diffusion equation with time delay. SIAM Journal on Mathematical Analysis, 52 (3):2275–2312, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dweik:2023:OTA

[DG23] Samer Dweik and Wojciech Górny. Optimal transport approach to Sobolev regularity of solutions to the weighted least gradient problem. SIAM Journal on Mathematical Analysis, 55(3):1916–1948, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1468358.

Dunlap:2025:URL

[DG25] Alexander Dunlap and Cole Graham. Uniqueness and root-Lipschitz regularity for a degenerate heat equation. SIAM Journal on Mathematical Analysis, 57(2):1866–1891, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1596260.

Dang:2023:HNS

[DGB23] Thuyen Dang, Yuliya Gorb, and Silvia Jiménez Bolaños. Homogenization of a NonLinear strongly coupled model of magnetorheological fluids. SIAM Journal on Mathematical Analysis, 55(1):102–132, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1476939.

Dovetta:2020:PLA

[DGMP20] Simone Dovetta, Marco Ghimenti, Anna Maria Micheletti, and Angela Pistoia. Peaked and low action solutions of NLS equations on graphs with terminal edges. SIAM Journal on Mathematical Analysis, 52 (3):2874–2894, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2024:LWP

[DGOY24] Hongjie Dong, Yan Guo, Zhimeng Ouyang, and Timur Yastrzhembskiy. The local wellposedness of the relativistic Vlasov–Maxwell–Landau system with the specular reflection boundary condition. SIAM Journal on Mathematical Analysis, 56(5):6613–6688, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1608938.

Ding:2021:SSD

[DGR21] Yanheng Ding, Qi Guo, and Bernhard Ruf. Stationary states of Dirac–Klein–Gordon systems with nonlinear interacting terms. SIAM Journal on Mathematical Analysis, 53 (5):5731–5755, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dareiotis:2020:ESP

[DGT20] Konstantinos Dareiotis, Benjamin Gess, and Pavlos Tsatsoulis. Ergodicity for stochastic porous media equations with multiplicative noise. SIAM Journal on Mathematical Analysis, 52(5):4524–4564, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dietert:2022:IPT

[DGV22] Helge Dietert and David Gérard-Varet. On the illposedness of the triple deck model. SIAM Journal on Mathematical Analysis, 54 (2):2611–2633, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1427401.

Dong:2023:GWP

[DGWZ23] Zhao Dong, Boling Guo, Jiang-Lun Wu, and Guoli Zhou. Global well-posedness and regularity of stochastic 3D Burgers equation with multiplicative noise. SIAM Journal on Mathematical Analysis, 55 (3):1847–1882, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413377.

Dreier:2021:RID

[DH21] Florian Dreier and Markus Haltmeier. Recovering the initial data of the wave equation from Neumann traces. SIAM Journal on Mathematical Analysis, 53(2):2427–2451, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolz:2024:PSH

[DH24] Jürgen Dölz and Fernando Henríquez. Parametric shape holomorphy of boundary integral operators with applications. SIAM Journal on Mathematical Analysis, 56 (5):6731–6767, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1576451.

deHoop:2024:REV

[dHKLN24] Maarten V. de Hoop, Masato Kimura, Ching-Lung Lin, and Gen Nakamura. Resolvent estimates for viscoelastic systems of extended Maxwell type and their applications. SIAM Journal on Mathematical Analysis, 56(5):5782–5806, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1592456.

Ding:2023:NLG

[Din23] Min Ding. Newtonian limits and global stability of shock solutions for piston problem to the relativistic full Euler system. SIAM Journal on Mathematical Analysis, 55(2):628–673, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1463422.

DiFrancesco:2025:AQP

[DIS25] M. Di Francesco, V. Iorio, and M. Schmidtchen. The approximation of the quadratic porous medium equation via nonlocal interacting particles subject to repulsive Morse potential. SIAM Journal on Mathematical Analysis, 57(5):4631–4679, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1632760.

Druet:2020:ACD

[DJ20] Pierre-Etienne Druet and Ansgar Jüngel. Analysis of crossdiffusion systems for fluid mixtures driven by a pressure gradient. SIAM Journal on Mathematical Analysis, 52(2):2179– 2197, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DallAcqua:2024:CCC

[DJLR24] Anna Dall’Acqua, Gaspard Jankowiak, Leonie Langer, and Fabian Rupp. Conservation, convergence, and computation for evolving heterogeneous elastic wires. SIAM Journal on Mathematical Analysis, 56(4):4494–4529, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M159086X.

Dinh:2021:LTD

[DK21a] Van Duong Dinh and Sahbi Keraani. Long time dynamics of nonradial solutions to inhomogeneous nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 53(4):4765–4811, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2021:GFN

[DK21b] Hongjie Dong and Seick Kim. Green’s function for nondivergence elliptic operators in two dimensions. SIAM Journal on Mathematical Analysis, 53(4): 4637–4656, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunst:2021:DSS

[DK21c] Kamil Dunst and Piotr Kokocki. Double spiral singularities for a flow of regular planar curves. SIAM Journal on Mathematical Analysis, 53(4):4727–4743, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davies:2023:NSB

[DK23] Joseph P. Davies and Gabriel S. Koch. Navier–Stokes blowup rates in certain Besov spaces whose regularity exceeds the critical value by ϵ ∈ [1, 2]. SIAM Journal on Mathematical Analysis, 55(5):5432–5444, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1520268.

Doppler:2025:LWP

[DKLT25] Sabine Doppler, Rupert Klein, Xin Liu, and Edriss S. Titi. Local well-posedness of a system coupling compressible atmospheric dynamics and a microphysics model of moisture in air. SIAM Journal on Mathematical Analysis, 57 (5):5161–5188, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1699838.

Ding:2021:EKS

[DL21] Zhiyan Ding and Qin Li.

Ensemble Kalman sampler: Mean-field limit and convergence analysis. SIAM Journal on Mathematical Analysis, 53 (2):1546–1578, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2022:NRG

[DL22a] Feng Dai and Bin Liu. A new result for global solvability of a two species cancer invasion haptotaxis model with tissue remodeling. SIAM Journal on Mathematical Analysis, 54(1):1–35, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1309870.

Ding:2022:GSC

[DL22b] Mengyao Ding and Johannes Lankeit. Generalized solutions to a Chemotaxis-Navier– Stokes system with arbitrary superlinear degradation. SIAM Journal on Mathematical Analysis, 54(1):1022–1052, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M140907X.

Dai:2024:ATI

[DL24a] Dan Dai and Wen-Gao Long. Asymptotics and total integrals of the P_{I}^{2} Tritronquée solution and its Hamiltonian. SIAM Journal on Mathematical Analysis, 56(4):5350–5371, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1592304.

Ding:2024:SSF

[DL24b] Min Ding and Yachun Li. Steady supersonic flows past Lipschitz wedges for twodimensional relativistic Euler equations. SIAM Journal on Mathematical Analysis, 56(4):5474–5520, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1600530.

Deng:2020:ASP

[DLL20] Youjun Deng, Hongjie Li, and Hongyu Liu. Analysis of surface polariton resonance for nanoparticles in elastic system. SIAM Journal on Mathematical Analysis, 52(2):1786–1805, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2021:CNS

[DLQ21] Wei Dai, Zhao Liu, and Guolin Qin. Classification of nonnegative solutions to static Schrödinger–Hartree–Maxwell type equations. SIAM Journal on Mathematical Analysis, 53 (2):1379–1410, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2022:ADP

[DLW22] Qiang Du, Xin Yang Lu, and Chong Wang. The average distance problem with perimeterto-area ratio penalization. SIAM Journal on Mathematical Analysis, 54(3):3122–3138, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422987.

Du:2024:ESD

[DLW24] Kai Du, Ruoyang Liu, and Yuxing Wang. Entropy solutions to the Dirichlet problem for nonlinear diffusion equations with conservative noise. SIAM Journal on Mathematical Analysis, 56(1):637–675, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Ding:2021:VPB

[DLX21] Shijin Ding, Zhilin Lin, and Feng Xie. Verification of Prandtl boundary layer ansatz for steady electrically conducting fluids with a moving physical boundary. SIAM Journal on Mathematical Analysis, 53 (5):4997–5059, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duan:2021:GCS

[DLY21] Renjun Duan, Shuangqian Liu, and Tong Yang. Global classical solutions for the Vlasov– Nordström–Fokker–Planck system. SIAM Journal on Mathematical Analysis, 53(5):6164– 6190, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:2025:DSM

[DLZ25] Ming-Hui Ding, Hongyu Liu, and Guang-Hui Zheng. Determining a stationary mean field game system from full/ partial boundary measurement. SIAM Journal on Mathematical Analysis, 57 (1):661–681, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1594327.

Ding:2020:DTP

[DM20a] Weiwei Ding and Hiroshi Matano. Dynamics of timeperiodic reaction–diffusion equations with front-like initial data on R. SIAM Journal on Mathematical Analysis, 52(3):2411– 2462, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Donatelli:2020:CIQ

[DM20b] Donatella Donatelli and Prince Romeo Mensah. The combined incompressible quasineutral limit of the stochastic Navier–Stokes– Poisson system. SIAM Journal on Mathematical Analysis, 52 (5):5090–5120, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2021:EMR

[DM21] Xiaojing Dong and Anmin Mao. Existence and multiplicity results for general quasilinear elliptic equations. SIAM Journal on Mathematical Analysis, 53(4):4965–4984, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duan:2022:GSW

[DM22] Senhao Duan and Yue Ma.

Global solutions of wave-Klein–Gordon systems in 2 + 1 dimensional space-time with strong couplings in divergence form. SIAM Journal on Mathematical Analysis, 54 (3):2691–2726, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1377229.

Dominguez:2024:IDH

[DM24] Tomas Dominguez and Jean-Christophe Mourrat. Infinitedimensional Hamilton–Jacobi equations for statistical inference on sparse graphs. SIAM Journal on Mathematical Analysis, 56(4):4530–4593, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1527696.

DiFratta:2020:VPM

[DMRS20] Giovanni Di Fratta, Cyrill B. Muratov, Filipp N. Rybakov, and Valeriy V. Slastikov. Variational principles of micromagnetics revisited. SIAM Journal on Mathematical Analysis, 52 (4):3580–3599, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davoli:2022:ECH

[DMS22a] Elisa Davoli, Anastasia Molchanova, and Ulisse Stefanelli. Equilibria of charged hyperelastic solids. SIAM Journal on Mathematical Analysis, 54 (2):1470–1487, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413286.

DiFratta:2022:SPM

[DMS22b] Giovanni Di Fratta, Antonin Monteil, and Valeriy Slastikov. Symmetry properties of minimizers of a perturbed Dirichlet energy with a boundary penalization. SIAM Journal on Mathematical Analysis, 54 (3):3636–3653, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M143011X.

Dovetta:2022:NRR

[DMS22c] Simone Dovetta, Elio Marconi, and Laura V. Spinolo. New regularity results for scalar conservation laws, and applications to a source-destination model for traffic flows on networks. SIAM Journal on Mathematical Analysis, 54 (3):3019–3053, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1434283.

Didier:2020:AAM

[DN20] Gustavo Didier and Hung Nguyen. Asymptotic analysis of the mean squared displacement under fractional memory kernels. SIAM Journal on Mathematical Analysis, 52(4): 3818–3842, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2022:HDF

[DN22] Yihong Du and Wenjie Ni. The high dimensional Fisher–KPP nonlocal diffusion equation with free boundary and radial symmetry, Part 1. SIAM Journal on Mathematical Analysis, 54(3):3930–3973, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1451920.

doNascimento:2023:DEU

[dN23] Arnaldo Simal do Nascimento. The destabilizing effect of unbalanced reaction as diffusivity decreases on reaction– diffusion equations in boundaryless surfaces. SIAM Journal on Mathematical Analysis, 55(1):68–101, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1448082.

Du:2025:LTD

[DNS25] Yihong Du, Wenjie Ni, and Linfei Shi. Long-time dynamics of a competition model with nonlocal diffusion and free boundaries: Vanishing and spreading of the invader. SIAM Journal on Mathematical Analysis, 57(2):1195–1226, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1647473.

Dodson:2021:SCS

[Dod21] Benjamin Dodson. The L^{2} sequential convergence of a solution to the one-dimensional, mass-critical NLS above the ground state. SIAM Journal on Mathematical Analysis, 53 (4):4744–4764, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2022:GSK

[Don22] Shijie Dong. Global solution to the Klein–Gordon–Zakharov equations with uniform energy bounds. SIAM Journal on Mathematical Analysis, 54(1):595–615, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1395235.

Dong:2025:NSS

[Don25] Wenchao Dong. Nonlinear stability of solution on the pressure problem of Navier– Stokes–Fourier system with temperature-dependent transport coefficients. SIAM Journal on Mathematical Analysis, 57(1):520–562, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M162243X.

DeRosa:2025:NAD

[DP25] Luigi De Rosa and Jaemin Park. No anomalous dissipation in two-dimensional incompressible fluids. SIAM Journal on Mathematical Analysis, 57(5):5771–5790, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/25M1773167.

DeLuca:2024:TSF

[DPS24] Lucia De Luca, Marcello Ponsiglione, and Emanuele Spadaro. Two slope functions minimizing fractional seminorms and applications to misfit dislocations. SIAM Journal on Mathematical Analysis, 56(1):1179–1196, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Dai:2023:CSC

[DQ23] Wei Dai and Guolin Qin. Classification of solutions to conformally invariant systems with mixed order and exponentially increasing or nonlocal nonlinearity. SIAM Journal on Mathematical Analysis, 55(3):2111–2149, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1499650.

Daneri:2020:PFL

[DR20a] Sara Daneri and Eris Runa. Pattern formation for a Local/Nonlocal interaction functional arising in colloidal systems. SIAM Journal on Mathematical Analysis, 52(3):2531– 2560, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duchene:2020:LTA

[DR20b] V. Duchêne and L. M. Rodrigues. Large-time asymptotic stability of Riemann shocks of scalar balance laws. SIAM Journal on Mathematical Analysis, 52(1):792–820, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2024:NEP

[DR24] Hongjie Dong and Junhee Ryu.

Nonlocal elliptic and parabolic equations with general stable operators in weighted Sobolev spaces. SIAM Journal on Mathematical Analysis, 56(4):4623–4661, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M160061X.

DeSilva:2021:CDO

[DS21a] Daniela De Silva and Ovidiu Savin. On certain degenerate one-phase free boundary problems. SIAM Journal on Mathematical Analysis, 53(1): 649–680, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:2021:ETD

[DS21b] Shengfu Deng and Shu-Ming Sun. The existence of three-dimensional multi-Hump gravity-capillary surface waves on water of finite depth. SIAM Journal on Mathematical Analysis, 53(2):2134–2205, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dodson:2022:BSW

[DS22] Stephanie Dodson and Björn Sandstede. Behavior of spiral wave spectra with a rank-deficient diffusion matrix. SIAM Journal on Mathematical Analysis, 54(3):3789–3816, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1455917.

Debiec:2023:CHM

[DS23] Tomasz Dȩbiec and Endre Süli. Corotational Hookean models of dilute polymeric fluids: Existence of global weak solutions, weak-strong uniqueness, equilibration, and macroscopic closure. SIAM Journal on Mathematical Analysis, 55(1):310–346, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M149867X.

Dinvay:2020:WPD

[DST20] E. Dinvay, S. Selberg, and A. Tesfahun. Well-posedness for a dispersive system of the Whitham–Boussinesq type. SIAM Journal on Mathematical Analysis, 52(3):2353–2382, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dohnal:2023:REA

[DST23] Tomá Dohnal, Roland Schnaubelt, and Daniel P. Tietz. Rigorous envelope approximation for interface wave packets in Maxwell’s equations with two dimensional localization. SIAM Journal on Mathematical Analysis, 55(6):6898–6939, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Duan:2024:BAN

[DSU24] Renjun Duan, Shota Sakamoto, and Yoshihiro Ueda. An l_{k}^{1} ∩ l_{k}^{p} approach for the noncutoff Boltzmann equation in r^{3}. SIAM Journal on Mathematical Analysis, 56(1):762– 800, January 2024. CODEN SJMAAH. ISSN 1095-7154.

DeNitti:2024:IPP

[DT24] Nicola De Nitti and Roman M. Taranets. Interface propagation properties for a nonlocal thin-film equation. SIAM Journal on Mathematical Analysis, 56(1):173–196, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Deconinck:2020:OSE

[DU20] Bernard Deconinck and Jeremy Upsal. The orbital stability of elliptic solutions of the focusing nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 52(1): 1–41, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duc:2023:PBO

[Duc23] Tho Nguyen Duc. Pseudomodes for biharmonic operators with complex potentials. SIAM Journal on Mathematical Analysis, 55(6):6580–6624, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Dunster:2021:NKF

[Dun21]  T. M. Dunster.  Nield–

Kuznetsov functions and Laplace transforms of parabolic cylinder functions. SIAM Journal on Mathematical Analysis, 53 (5):5915–5947, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:2024:UAE

[Dun24]  T. M. Dunster.  Uniform

asymptotic expansions for the zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 56(5):6521–6550, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1611269.

Dong:2020:BRC

[DW20] Hongjie Dong and Kunrui Wang. Boundary εregularity criteria for the threedimensional Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 52(2): 1290–1309, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2021:SCT

[DW21] Lili Du and Yongfu Wang.

Steady collision of two jets issuing from two axially symmetric channels. SIAM Journal on Mathematical Analysis, 53(2): 2523–2566, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deWolff:2023:EPC

[dW23] Babette A. J. de Wolff. Equivariant pyragas control of discrete waves. SIAM Journal on Mathematical Analysis, 55(6):6707–6739, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Dong:2024:GBE

[DY24] Hongjie Dong and Timur Yastrzhembskiy. Global L_{p} estimates for kinetic Kolmogorov– Fokker–Planck equations in divergence form. SIAM Journal on Mathematical Analysis, 56 (1):1223–1263, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Duan:2021:HSP

[DYZ21] Renjun Duan, Haiyan Yin, and Changjiang Zhu. A half-space problem on the full Euler– Poisson system. SIAM Journal on Mathematical Analysis, 53 (5):6094–6121, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2023:TAN

[DZZ23] Hongjie Dong, Chulan Zeng, and Qi S. Zhang. Time analyticity for nonlocal parabolic equations. SIAM Journal on Mathematical Analysis, 55 (3):1883–1915, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1490740.

Evgrafov:2023:NBP

[EB23] Anton Evgrafov and José C. Bellido. Nonlocal basis pursuit: Nonlocal optimal design of conductive domains in the vanishing material limit. SIAM Journal on Mathematical Analysis, 55(4):2740–2773, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1479166.

Eceizabarrena:2022:ASF

[ED22] Daniel Eceizabarrena and Victor Vilaça Da Rocha. An analytical study of flatness and intermittency through Riemann’s nondifferentiable functions. SIAM Journal on Mathematical Analysis, 54 (3):3575–3608, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1411512.

Eberle:2021:SBI

[EFHS21] S. Eberle, F. Florian, R. Hiptmair, and S. A. Sauter. A stable boundary integral formulation of an acoustic wave transmission problem with mixed boundary conditions. SIAM Journal on Mathematical Analysis, 53(2):1492–1508, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Emmert:2020:LDM

[EFK20] Lukas Emmert, Rupert L.

Frank, and Tobias König. Liquid drop model for nuclear matter in the dilute limit. SIAM Journal on Mathematical Analysis, 52(2):1980–1999, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esposito:2024:PLS

[EFP^{+}24] Antonio Corbo Esposito, Luisa Faella, Gianpaolo Piscitelli, Ravi Prakash, and Antonello Tamburrino. The p-Laplace “Signature” for quasilinear inverse problems with large boundary data. SIAM Journal on Mathematical Analysis, 56(1):275–303, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Ehrnstrom:2022:EDS

[EGN22] Mats Ehrnström, Mark D. Groves, and Dag Nilsson. Existence of Davey–Stewartson type solitary waves for the fully dispersive Kadomtsev– Petviashvilii equation. SIAM Journal on Mathematical Analysis, 54(4):4954–4986, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1451518.

Entekhabi:2020:ISA

[EI20] Mozhgan Entekhabi and Victor Isakov. Increasing stability in acoustic and elastic inverse source problems. SIAM Journal on Mathematical Analysis, 52(5):5232–5256, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ennaji:2022:QCH

[EIN22] Hamza Ennaji, Noureddine Igbida, and Van Thanh Nguyen. Quasi-convex Hamilton–Jacobi equations via Finsler p-Laplace-type operators. SIAM Journal on Mathematical Analysis, 54(4):5041–5063, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M143306X.

Eiter:2022:OTR

[Eit22] Thomas Eiter. On the Oseentype resolvent problem associated with time-periodic flow past a rotating body. SIAM Journal on Mathematical Analysis, 54(4):4987–5012, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1456728.

Eleuteri:2020:ASS

[EK20] Michela Eleuteri and Pavel Krejcí. Asymptotic stability of solutions to the porous media system with hysteresis. SIAM Journal on Mathematical Analysis, 52(4):3962–3989, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Epshteyn:2021:MGB

[ELM21] Yekaterina Epshteyn, Chun Liu, and Masashi Mizuno. Motion of grain boundaries with dynamic lattice misorientations and with triple junctions drag. SIAM Journal on Mathematical Analysis, 53(3): 3072–3097, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erceg:2022:STD

[EM22] Marko Erceg and Darko Mitrović. Strong traces to degenerate parabolic equations. SIAM Journal on Mathematical Analysis, 54(2):1775–1796, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1425530.

Emerald:2021:RDW

[Eme21]  Louis Emerald.  Rigorous

derivation from the water waves equations of some full dispersion shallow water models. SIAM Journal on Mathematical Analysis, 53(4):3772– 3800, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Einav:2020:IDE

[EMT20] Amit Einav, Jeffrey J. Morgan, and Bao Q. Tang. Indirect diffusion effect in degenerate reaction–diffusion systems. SIAM Journal on Mathematical Analysis, 52(5):4314–4361, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eckstein:2022:QSR

[EN22] Stephan Eckstein and Marcel Nutz. Quantitative stability of regularized optimal transport and convergence of Sinkhorn’s algorithm. SIAM Journal on Mathematical Analysis, 54 (6):5922–5948, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M145505X.

Efendiev:2020:MAP

[EOE20] Messoud A. Efendiev, Mitsuharu Otani, and Hermann J. Eberl. Mathematical analysis of a PDE–ODE coupled model of mitochondrial swelling with degenerate calcium ion diffusion. SIAM Journal on Mathematical Analysis, 52(1):543– 569, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eptaminitakis:2024:WNG

[ES24] Nikolas Eptaminitakis and Plamen Stefanov. Weakly nonlinear geometric optics for the Westervelt equation and recovery of the nonlinearity. SIAM Journal on Mathematical Analysis, 56(1):801–819, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Evans:2021:HPE

[Eva21] Josephine Evans. Hypocoercivity in phi-entropy for the linear relaxation Boltzmann equation on the torus. SIAM Journal on Mathematical Analysis, 53 (2):1357–1378, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evans:2023:ABR

[EY23] Josephine A. Evans and Havva Yolda. On the asymptotic behavior of a run and tumble equation for bacterial chemotaxis. SIAM Journal on Mathematical Analysis, 55(6):7635– 7664, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Esteve-Yague:2022:DRI

[EYZ22] Carlos Esteve-Yagüe and Enrique Zuazua. Differentiability with respect to the initial condition for Hamilton– Jacobi equations. SIAM Journal on Mathematical Analysis, 54(5):5388–5423, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469353.

Esteve:2020:IPH

[EZ20] Carlos Esteve and Enrique Zuazua. The inverse problem for Hamilton–Jacobi equations and semiconcave envelopes. SIAM Journal on Mathematical Analysis, 52(6):5627–5657, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feizmohammadi:2025:REE

[Fei25] Ali Feizmohammadi. Reconstruction of 1D evolution equations and their initial data from one passive measurement. SIAM Journal on Mathematical Analysis, 57(5):5089–5106, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/25M1728533.

Feng:2025:EAB

[Fen25]  Baowei Feng.  Existence

and asymptotic behavior for a Lamé system with timedependent delay and analytic nonlinearity. SIAM Journal on Mathematical Analysis, 57 (2):1753–1780, ???? 2025.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1590159.

Feppon:2021:HOH

[Fep21] Florian Feppon. High order homogenization of the Stokes system in a periodic porous medium. SIAM Journal on Mathematical Analysis, 53(3): 2890–2924, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fantini:2025:RSV

[FF25] Veronica Fantini and Aaron Fenyes. Regular singular Volterra equations on complex domains. SIAM Journal on Mathematical Analysis, 57(1):336–363, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1613876.

Francini:2020:ERE

[FFV20] Elisa Francini, Giovanni Franzina, and Sergio Vessella. Existence and regularity for eddy current system with nonsmooth conductivity. SIAM Journal on Mathematical Analysis, 52(2):2134–2157, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ferreira:2021:HQC

[FFV21] Rita Ferreira, Irene Fonseca, and Raghavendra Venkatraman. Homogenization of quasicrystalline functionals via twoscale-cut-and-project convergence. SIAM Journal on Mathematical Analysis, 53(2):1785– 1817, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2022:APT

[FG22] Beixiang Fang and Xin Gao. On admissible positions of transonic shocks for steady isothermal Euler flows in a horizontal flat nozzle under vertical gravity. SIAM Journal on Mathematical Analysis, 54(5):5223–5267, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M146733X.

Fonseca:2021:MCD

[FGW21] Irene Fonseca, Janusz Ginster, and Stephan Wojtowytsch. On the motion of curved dislocations in three dimensions: Simplified linearized elasticity. SIAM Journal on Mathematical Analysis, 53(2):2373–2426, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fischer:2022:GEA

[FHKM22] Julian Fischer, Katharina Hopf, Michael Kniely, and Alexander Mielke. Global existence analysis of energy– reaction–diffusion systems. SIAM Journal on Mathematical Analysis, 54(1):220–267, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1387237.

Feng:2020:CGL

[FHM20] Zhewen Feng, Min-Chun Hong, and Yu Mei. Convergence of the Ginzburg–Landau approximation for the Ericksen– Leslie system. SIAM Journal on Mathematical Analysis, 52 (1):481–523, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feng:2024:RTLa

[FHM^{+}24a] Yue-Hong Feng, Haifeng Hu, Ming Mei, Yue-Jun Peng, and Guo-Jing Zhang. Relaxation time limits of subsonic steady states for hydrodynamic model of semiconductors with sonic or nonsonic boundary. SIAM Journal on Mathematical Analysis, 56(3):3452–3477, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Feng:2024:RTLb

[FHM^{+}24b] Yue-Hong Feng, Haifeng Hu, Ming Mei, Gantumur Tsogtgerel, and Guojing Zhang. Relaxation time limits of subsonic steady states for multidimensional hydrodynamic model of semiconductors. SIAM Journal on Mathematical Analysis, 56(5):6933–6962, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1648879.

Flandoli:2024:SCE

[FHP24] Franco Flandoli, Ruojun Huang, and Andrea Papini. Smoluchowski coagulation equation with velocity dependence. SIAM Journal on Mathematical Analysis, 56(4):5634–5677, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1540594.

Florian:2024:SIE

[FHS24] F. Florian, R. Hiptmair, and S. A. Sauter. Skeleton integral equations for acoustic transmission problems with varying coefficients. SIAM Journal on Mathematical Analysis, 56(5):6232–6267, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1572106.

Feppon:2022:HOH

[FJ22] Florian Feppon and Wenjia Jing. High order homogenized Stokes models capture all three regimes. SIAM Journal on Mathematical Analysis, 54(4):5013–5040, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1390232.

Feireisl:2025:CRC

[FK25] Eduard Feireisl and Young-Sam Kwon. Conditional regularity for the compressible MHD system with inhomogeneous boundary data. SIAM Journal on Mathematical Analysis, 57(1):14–29, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1580760.

Ftouhi:2021:BSD

[FL21] Ilias Ftouhi and Jimmy Lamboley. Blaschke–Santaló diagram for volume, perimeter, and first Dirichlet eigenvalue. SIAM Journal on Mathematical Analysis, 53(2):1670–1710, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Flandoli:2024:BHS

[FL24] Franco Flandoli and Dejun Luo. On the Boussinesq hypothesis for a stochastic Proudman–Taylor model. SIAM Journal on Mathematical Analysis, 56(3):3886–3923, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1587944.

Ferreira:2024:NLC

[FLNV24] Marina A. Ferreira, Jani Lukkarinen, Alessia Nota, and Juan J. L. Velázquez. Nonpower law constant flux solutions for the Smoluchowski coagulation equation. SIAM Journal on Mathematical Analysis, 56(3):2783–2817, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Filonov:2024:UEP

[FLPS24] Nikolay Filonov, Michael Levitin, Iosif Polterovich, and David A. Sher. Uniform enclosures for the phase and zeros of Bessel functions and their derivatives. SIAM Journal on Mathematical Analysis, 56(6):7644–7682, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1642032.

Feng:2024:PST

[FLRX24] Yan-Xia Feng, Wan-Tong Li, Shigui Ruan, and Ming-Zhen Xin. Principal spectral theory of time-periodic nonlocal dispersal cooperative systems and applications. SIAM Journal on Mathematical Analysis, 56(3):4040–4083, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1543902.

Fischer:2020:CRA

[FLS20] Julian Fischer, Tim Laux, and Theresa M. Simon. Convergence rates of the Allen–Cahn equation to mean curvature flow: a short proof based on relative entropies. SIAM Journal on Mathematical Analysis, 52(6):6222–6233, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedrich:2025:SEF

[FLS25] Manuel Friedrich, Camille Labourie, and Kerrek Stinson. Strong existence for free discontinuity problems in linear elasticity. SIAM Journal on Mathematical Analysis, 57(2):1652–1679, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1640847.

Faustmann:2022:WAR

[FMMS22] Markus Faustmann, Carlo Marcati, Jens Markus Melenk, and Christoph Schwab. Weighted analytic regularity for the integral fractional Laplacian in polygons. SIAM Journal on Mathematical Analysis, 54(6):6323–6357, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M146569X.

Frid:2021:GSS

[FMN21] Hermano Frid, Daniel Marroquin, and João F. C. Nariyoshi. Global smooth solutions with large data for a system modeling Aurora type phenomena in the 2-torus. SIAM Journal on Mathematical Analysis, 53 (1):1122–1167, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Forkert:2022:ECE

[FMP22] Dominik Forkert, Jan Maas, and Lorenzo Portinale. Evolutionary Γ-convergence of entropic gradient flow structures for Fokker–Planck equations in multiple dimensions. SIAM Journal on Mathematical Analysis, 54(4):4297–4333, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410968.

Friesecke:2021:BHK

[FMS21] Gero Friesecke, Daniel Matthes, and Bernhard Schmitzer. Barycenters for the Hellinger– Kantorovich distance over R^{d}. SIAM Journal on Mathematical Analysis, 53(1):62–110, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fitzgibbon:2021:RDA

[FMTY21] William E. Fitzgibbon, Jeffrey J. Morgan, Bao Q. Tang, and Hong-Ming Yin. Reaction– diffusion–advection systems with discontinuous diffusion and mass control. SIAM Journal on Mathematical Analysis, 53(6):6771–6803, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fall:2023:EPN

[FMW23] Mouhamed Moustapha Fall, Ignace Aristide Minlend, and Tobias Weth. On an electrostatic problem and a new class of exceptional subdomains of R3. SIAM Journal on Mathematical Analysis, 55(3):2347–2376, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487540.

Fjordholm:2024:VVL

[FMW24] Ulrik Skre Fjordholm, Siddhartha Mishra, and Franziska Weber. On the vanishing viscosity limit of statistical solutions of the incompressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 56(4):5099–5143, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1566261.

Fornerod:2023:WLT

[FN23] Jean Fornerod and Hoai-Minh Nguyen. The Weyl law of transmission eigenvalues and the completeness of generalized transmission eigenfunctions without complementing conditions. SIAM Journal on Mathematical Analysis, 55 (4):3959–3999, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1486157.

Flandoli:2020:UAD

[FOS20] Franco Flandoli, Christian Olivera, and Marielle Simon. Uniform approximation of 2 dimensional Navier–Stokes equation by stochastic interacting particle systems. SIAM Journal on Mathematical Analysis, 52(6):5339–5362, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Foss:2021:TGS

[Fos21] Mikil Foss. Traces on general sets in R^{n} for functions with no differentiability requirements. SIAM Journal on Mathematical Analysis, 53(4):4212–4251, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Freistuhler:2020:DBV

[FP20] Heinrich Freistühler and Valentin Pellhammer. Dependence on the background viscosity of solutions to a prototypical Non-Strictly hyperbolic system of conservation laws. SIAM Journal on Mathematical Analysis, 52(6):5658–5674, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fournier:2021:NTD

[FP21] Nicolas Fournier and Benoît Perthame. A nonexpanding transport distance for some structured equations. SIAM Journal on Mathematical Analysis, 53(6):6847–6872, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2022:NTW

[FP22] Jian Fang and Yingli Pan.

NonMonotonicity of traveling wave profiles for a unimodal recursive system. SIAM Journal on Mathematical Analysis, 54(2):1669–1694, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M139236X.

Faye:2023:EAS

[FR23] Grégory Faye and Luis Miguel Rodrigues. Exponential asymptotic stability of Riemann shocks of hyperbolic systems of balance laws. SIAM Journal on Mathematical Analysis, 55(6):6425–6456, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Feliciangeli:2020:UNM

[FS20] Dario Feliciangeli and Robert Seiringer. Uniqueness and NonDegeneracy of minimizers of the Pekar functional on a ball. SIAM Journal on Mathematical Analysis, 52(1):605– 622, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonte:2022:LTB

[FS22] Claudia Fonte and Valentin Schmutz. Long time behavior of an age- and leaky memorystructured neuronal population equation. SIAM Journal on Mathematical Analysis, 54(4):4721–4756, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1428571.

Fefferman:2022:LBQ

[FSW22] Charles Fefferman, Jacob Shapiro, and Michael I. Weinstein. Lower bound on quantum tunneling for strong magnetic fields. SIAM Journal on Mathematical Analysis, 54 (1):1105–1130, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1429412.

Fritz:2024:ADP

[FSW24] Marvin Fritz, Endre Süli, and Barbara Wohlmuth. Analysis of a dilute polymer model with a time-fractional derivative. SIAM Journal on Mathematical Analysis, 56(2):2063– 2089, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Fan:2024:DPN

[FSZ24] Chenjie Fan, Gigliola Staffilani, and Zehua Zhao. On decaying properties of nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 56(3):3082–3109, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Fuest:2020:GSN

[Fue20]  Mario Fuest.  Global so-

lutions near homogeneous steady states in a multidimensional population model with both predator- and prey-taxis. SIAM Journal on Mathematical Analysis, 52(6):5865–5891, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fanelli:2025:EVB

[FV25] Francesco Fanelli and Alexis F.

Vasseur. Effective velocity and L^{∞}-based well-posedness for incompressible fluids with odd viscosity. SIAM Journal on Mathematical Analysis, 57(1):153–189, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1635661.

Franz:2021:ICR

[FW21] Tino Franz and Holger Wendland. An improved convergence result for the smoothed particle hydrodynamics method.

SIAM Journal on Mathematical Analysis, 53(2):1239–1262, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fan:2021:WZT

[FX21] Chenjie Fan and Weijun Xu. A Wong–Zakai theorem for the stochastic mass-critical nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 53(3):3681–3715, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2020:PSN

[FXX20] Beixiang Fang, Wei Xiang, and Feng Xiao. Persistence of the steady normal shock structure for the unsteady potential flow. SIAM Journal on Mathematical Analysis, 52(6):6033–6104, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feng:2021:ESM

[FZ21] Chunrong Feng and Huaizhong Zhao. Ergodicity of sublinear Markovian semigroups. SIAM Journal on Mathematical Analysis, 53(5):5646–5681, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gahn:2025:DBP

[Gah25] Markus Gahn. Derivation of a Biot plate system for a thin poroelastic layer. SIAM Journal on Mathematical Analysis, 57(5):5303–5341, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1644535.

Gallenmuller:2023:MVL

[Gal23] Dennis Gallenmüller. Measurevalued low Mach number limits of ideal fluids. SIAM Journal on Mathematical Analysis, 55(2):1145–1169, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1467596.

Gao:2021:APC

[Gao21] Peng Gao. Averaging principle for complex Ginzburg–Landau equation perturbated by mixing random forces. SIAM Journal on Mathematical Analysis, 53(1):32–61, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garenaux:2024:NCS

[Gar24] Louis Garénaux. Nonlinear convective stability of a critical pulled front undergoing a Turing bifurcation at its back: A case study. SIAM Journal on Mathematical Analysis, 56(3): 3275–3325, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Gassot:2023:LEZ

[Gas23] Louise Gassot. Lax eigenvalues in the zero-dispersion limit for the Benjamin–Ono equation on the torus. SIAM Journal on Mathematical Analysis, 55(5):5782–5822, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M154635X.

Granero-Belinchon:2020:AMF

[GBS20] Rafael Granero-Belinchón and Stefano Scrobogna. On an asymptotic model for free boundary Darcy flow in porous media. SIAM Journal on Mathematical Analysis, 52(5): 4937–4970, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gan:2024:CND

[GDS24] Weiye Gan, Qiang Du, and Zuoqiang Shi. -convergence of nonlocal Dirichlet energies with penalty formulations of Dirichlet boundary data. SIAM Journal on Mathematical Analysis, 56(6):7439–7462, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1604746.

Geng:2025:RRT

[Gen25] Zhiyuan Geng. Rigidity results for a triple junction solution of the Allen–Cahn system. SIAM Journal on Mathematical Analysis, 57(5):5570–5597, ???? 2025. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/25M1736815.

Geredeli:2021:BSW

[Ger21] Pelin G. Geredeli. Bounded semigroup wellposedness for a linearized compressible flow structure PDE interaction with material derivative. SIAM Journal on Mathematical Analysis, 53(2):1711–1744, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gervais:2023:CBN

[Ger23] Pierre Gervais. On the convergence from Boltzmann to Navier–Stokes–Fourier for general initial data. SIAM Journal on Mathematical Analysis, 55(2):805–848, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1471687.

Gerner:2024:PBS

[Ger24] Wadim Gerner. Properties of the Biot–Savart operator acting on surface currents. SIAM Journal on Mathematical Analysis, 56(5):6446–6482, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1615693.

Gerard:2020:DSE

[GG20] Patrick Gérard and Sandrine Grellier. On a damped Szegő equation (with an appendix in collaboration with Christian Klein). SIAM Journal on Mathematical Analysis, 52(5): 4391–4420, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galeati:2023:MGR

[GG23] Lucio Galeati and Massimiliano Gubinelli. Mixing for generic rough shear flows. SIAM Journal on Mathematical Analysis, 55(6):7240–7272, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Ghisi:2024:AGE

[GG24] Marina Ghisi and Massimo Gobbino. Almost global existence for Kirchhoff equations around global solutions. SIAM Journal on Mathematical Analysis, 56(2):1936–1958, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Gancedo:2025:LTI

[GGBS25] Francisco Gancedo, Rafael Granero-Belinchón, and Elena Salguero. Long time interface dynamics for gravity Stokes flow. SIAM Journal on Mathematical Analysis, 57 (2):1680–1724, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1667488.

Golding:2024:NRE

[GGL24] William Golding, Maria Pia Gualdani, and Amélie Loher. Nonlinear regularization estimates and global wellposedness for the Landau– Coulomb equation near equilibrium. SIAM Journal on Mathematical Analysis, 56 (6):8037–8069, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1573951.

Gong:2021:CMT

[GH21a] Yishu Gong and Siming He. On the 8π-critical-mass threshold of a Patlak–Keller– Segel–Navier–Stokes system. SIAM Journal on Mathematical Analysis, 53(3):2925–2956, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grabovsky:2021:FEC

[GH21b] Yury Grabovsky and Narek Hovsepyan. On the feasibility of extrapolation of the complex electromagnetic permittivity function using Kramers– Kronig relations. SIAM Journal on Mathematical Analysis, 53(6):6993–7023, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2024:ESP

[GH24a] Yuetian Gao and Fangyu Han. Existence of solutions to the Patlak–Keller–Segel–Navier– Stokes system. SIAM Journal on Mathematical Analysis, 56(5):6798–6821, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1569307.

Garde:2024:LCP

[GH24b] Henrik Garde and Nuutti Hyvönen. Linearized Calderón problem: Reconstruction and Lipschitz stability for infinitedimensional spaces of unbounded perturbations. SIAM Journal on Mathematical Analysis, 56(3):3588–3604, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Garde:2020:RLF

[GHK20] Henrik Garde, Nuutti Hyvönen, and Topi Kuutela. On regularity of the logarithmic forward map of electrical impedance tomography. SIAM Journal on Mathematical Analysis, 52 (1):197–220, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gan:2022:DLC

[GHL22] Zaihui Gan, Xianpeng Hu, and Fanghua Lin. Defects in liquid crystal flows. SIAM Journal on Mathematical Analysis, 54(2):1695–1717, ???? 2022.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1396010.

Ghattassi:2022:DLR

[GHM22] Mohamed Ghattassi, Xiaokai Huo, and Nader Masmoudi. On the diffusive limits of radiative heat transfer system I: Well-prepared initial and boundary conditions. SIAM Journal on Mathematical Analysis, 54(5):5335–5387, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1455267.

Garcia:2023:TPD

[GHM23] Claudia García, Taoufik Hmidi, and Joan Mateu. Time periodic doubly connected solutions for the 3D quasigeostrophic model. SIAM Journal on Mathematical Analysis, 55(6):6133–6193, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Ginster:2024:SIL

[GHPZ24] Janusz Ginster, Gurgen Hayrapetyan, Anastasija Pešić, and Barbara Zwicknagl. A sharp interface limit of a nonlocal variational model for pattern formation in biomembranes. SIAM Journal on Mathematical Analysis, 56 (3):2818–2862, May 2024. CO-DEN SJMAAH. ISSN 1095-

Geng:2021:CGB

[GHW21] Shifeng Geng, Feimin Huang, and Xiaochun Wu. L^{1}- convergence to generalized Barenblatt solution for compressible Euler equations with time-dependent damping. SIAM Journal on Mathematical Analysis, 53(5):6048–6072, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gibson:2024:ESS

[Gib24] Peter C. Gibson. Explicit solution of the 1D Schrödinger equation. SIAM Journal on Mathematical Analysis, 56 (4):4466–4493, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1514441.

Ginsberg:2021:BSI

[Gin21] Daniel Ginsberg. On the breakdown of solutions to the incompressible Euler equations with free surface boundary. SIAM Journal on Mathematical Analysis, 53(3):3366–3384, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Girardin:2025:PES

[Gir25] Léo Girardin. Persistence, extinction, and spreading properties of noncooperative Fisher– KPP systems in space-time periodic media. SIAM Journal on Mathematical Analysis, 57(1):233–261, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1576086.

Gorbachev:2020:UPE

[GIT20] Dmitry Gorbachev, Valerii Ivanov, and Sergey Tikhonov. Uncertainty principles for eventually constant sign bandlimited functions. SIAM Journal on Mathematical Analysis, 52(5):4751–4782, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghoshal:2021:NRE

[GJ21] Shyam Sundar Ghoshal and Animesh Jana. Nonexistence of the BV regularizing effect for scalar conservation laws in several space dimensions for C^{2} fluxes. SIAM Journal on Mathematical Analysis, 53(2): 1908–1943, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghoshal:2024:HRE

[GJP24] S. S. Ghoshal, S. Junca, and A. Parmar. Higher regularity for entropy solutions of conservation laws with geometrically constrained discontinuous flux. SIAM Journal on Mathematical Analysis, 56(5):6121–6136, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1604199.

Garcke:2020:WSC

[GK20] Harald Garcke and Patrik Knopf. Weak solutions of the Cahn–Hilliard system with dynamic boundary conditions: a gradient flow approach. SIAM Journal on Mathematical Analysis, 52(1):340–369, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gavioli:2021:VPM

[GK21a] Chiara Gavioli and Pavel Krejcí. On a viscoelastoplastic porous medium problem with nonlinear interaction. SIAM Journal on Mathematical Analysis, 53(1):1191–1213, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gong:2021:WMV

[GK21b] Xiaoqian Gong and Matthias Kawski. Weak measurevalued solutions of a nonlinear hyperbolic conservation law. SIAM Journal on Mathematical Analysis, 53(4):4417–4444, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gebhard:2022:BTD

[GK22] Björn Gebhard and József J. Kolumbán. On bounded twodimensional globally dissipative Euler flows. SIAM Journal on Mathematical Analysis, 54(3):3457–3479, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1454675.

Gu:2023:FWN

[GK23] Yu Gu and Tomasz Komorowski. Fluctuations of the winding number of a directed polymer on a cylinder. SIAM Journal on Mathematical Analysis, 55(4):3262–3286, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1512508.

Grun:2025:FSP

[GK25] Günther Grün and Lorenz Klein. Finite speed of propagation for a class of stochastic thin-film equations. SIAM Journal on Mathematical Analysis, 57(5):4822–4866, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1705330.

Giletti:2022:TSN

[GKK22] Thomas Giletti, Ho-Youn Kim, and Yong-Jung Kim. Terrace solutions for non-Lipschitz multistable nonlinearities. SIAM Journal on Mathematical Analysis, 54 (4):4785–4805, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1445430.

Ghoussoub:2021:HCP

[GKLP21] Nassif Ghoussoub, Young-Heon Kim, Hugo Lavenant, and Aaron Zeff Palmer. Hidden convexity in a problem of nonlinear elasticity. SIAM Journal on Mathematical Analysis, 53 (1):1070–1087, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gladbach:2020:SLD

[GKM20] Peter Gladbach, Eva Kopfer, and Jan Maas. Scaling limits of discrete optimal transport. SIAM Journal on Mathematical Analysis, 52(3):2759–2802, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gu:2020:FNS

[GL20] Yu Gu and Jiawei Li. Fluctuations of a nonlinear stochastic heat equation in dimensions three and higher. SIAM Journal on Mathematical Analysis, 52(6):5422–5440, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giri:2021:ENC

[GL21] Ankik Kumar Giri and Philippe Laurençot. Existence and NonExistence for the collisioninduced breakage equation.

SIAM Journal on Mathematical Analysis, 53(4):4605–4636, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2023:SPW

[GL23a] Yuan Gao and Jian-Guo Liu.

A selection principle for weak KAM solutions via Freidlin– Wentzell large deviation principle of invariant measures. SIAM Journal on Mathematical Analysis, 55(6):6457–6495, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Girodroux-Lavigne:2023:DER

[GL23b] Alexandre Girodroux-Lavigne.

Derivation of an effective rheology for dilute suspensions of microswimmers. SIAM Journal on Mathematical Analysis, 55(6):6305–6341, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Grohs:2023:NUT

[GL23c] Philipp Grohs and Lukas Liehr. Non-uniqueness theory in sampled STFT phase retrieval. SIAM Journal on Mathematical Analysis, 55 (5):4695–4726, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1510224.

Gaidashev:2024:REF

[GL24a] Denis Gaidashev and Alejandro Luque. Renormalization and existence of finitetime blow-up solutions for a one-dimensional analogue of the Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 56(4):4356–4374, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1551481.

Goudey:2024:LEH

[GL24b] Rémi Goudey and Claude Le Bris. Linear elliptic homogenization for a class of highly oscillating nonperiodic potentials. SIAM Journal on Mathematical Analysis, 56(2):2738– 2782, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Geng:2020:ABS

[GLM20] Shifeng Geng, Yanping Lin, and Ming Mei. Asymptotic behavior of solutions to Euler equations with timedependent damping in critical case. SIAM Journal on Mathematical Analysis, 52(2):1463– 1488, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2023:EAB

[GLP23] Yujin Guo, Yong Luo, and Shuangjie Peng. Existence and asymptotic behavior of ground states for rotating Bose–Einstein condensates. SIAM Journal on Mathematical Analysis, 55(2):773–804, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1474680.

Giorgini:2022:ESS

[GLRS22] Andrea Giorgini, Kei Fong Lam, Elisabetta Rocca, and Giulio Schimperna. On the existence of strong solutions to the Cahn–Hilliard–Darcy system with mass source. SIAM Journal on Mathematical Analysis, 54(1):737–767, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1376443.

Galkowski:2023:PML

[GLS23] Jeffrey Galkowski, David Lafontaine, and Euan Spence. Perfectly-matched-layer truncation is exponentially accurate at high frequency. SIAM Journal on Mathematical Analysis, 55(4):3344–3394, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443716.

Galkowski:2023:DHF

[GLSW23] J. Galkowski, D. Lafontaine, E. A. Spence, and J. Wunsch. Decompositions of highfrequency Helmholtz solutions via functional calculus, and application to the finite element method. SIAM Journal on Mathematical Analysis, 55(4):3903–3958, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1409160.

Gao:2022:RSC

[GLW22] Yu Gao, Hao Liu, and Tak Kwong Wong. Regularity structure of conservative solutions to the Hunter– Saxton equation. SIAM Journal on Mathematical Analysis, 54(1):423–452, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1427590.

Gao:2023:ABC

[GLW23] Yu Gao, Hao Liu, and Tak Kwong Wong. Asymptotic behavior of conservative solutions to the Hunter– Saxton equation. SIAM Journal on Mathematical Analysis, 55(5):5483–5525, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1540491.

Gao:2020:STS

[GLY20] Junlei Gao, Li Liu, and Hairong Yuan. On stability of transonic shocks for stationary Rayleigh flows in two-dimensional ducts. SIAM Journal on Mathematical Analysis, 52(5):5287–5337, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

gangLi:2025:SCB

[gLZ25] Hai gang Li and Peihao Zhang. Stress concentration between two adjacent rigid particles in Navier–Stokes flow. SIAM Journal on Mathematical Analysis, 57(5):4756–4794, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1695427.

Garavello:2020:GWS

[GM20a] Mauro Garavello and Francesca Marcellini. Global weak solutions to the Cauchy problem for a two-phase model at a node. SIAM Journal on Mathematical Analysis, 52(2): 1567–1590, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grava:2020:LTA

[GM20b] Tamara Grava and Alexander Minakov. On the long-time asymptotic behavior of the modified Korteweg–de Vries equation with step-like initial data. SIAM Journal on Mathematical Analysis, 52(6):5892– 5993, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gallouet:2022:CLD

[GMN22] Thomas O. Gallouët, Quentin Mérigot, and Andrea Natale. Convergence of a Lagrangian discretization for barotropic fluids and porous media flow. SIAM Journal on Mathematical Analysis, 54(3):2990–3018, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422756.

Galkowski:2021:ETH

[GMS21a] Jeffrey Galkowski, Pierre Marchand, and Euan A. Spence. Eigenvalues of the truncated Helmholtz solution operator under strong trapping. SIAM Journal on Mathematical Analysis, 53(6):6724– 6770, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gangbo:2021:FDA

[GMŚ21b] Wilfrid Gangbo, Sergio Mayorga, and Andrzej Świȩch. Finite dimensional approximations of Hamilton–Jacobi– Bellman equations in spaces of probability measures. SIAM Journal on Mathematical Analysis, 53(2):1320–1356, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garroni:2021:DLT

[GMS21c] Adriana Garroni, Roberta Marziani, and Riccardo Scala. Derivation of a line-tension model for dislocations from a nonlinear three-dimensional energy: The case of quadratic growth. SIAM Journal on Mathematical Analysis, 53(4):

4252–4302, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gaczkowski:2023:MSC

[GMS23] Michał Gaczkowski, Jarosław Mederski, and Jacopo Schino. Multiple solutions to cylindrically symmetric curl–curl problems and related Schrödinger equations with singular potentials. SIAM Journal on Mathematical Analysis, 55 (5):4425–4444, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1494786.

Goodwill:2023:IFL

[GN23] Tristan Goodwill and Michael O.

Neil. An interface formulation of the Laplace–Beltrami problem on piecewise smooth surfaces. SIAM Journal on Mathematical Analysis, 55(6):7575– 7615, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Gaal:2020:MEP

[GNNCR20] Marcell Gaál, Béla Nagy, Zsuzsanna Nagy-Csiha, and Szilárd Gy. Révész. Minimal energy point systems on the unit circle and the real line. SIAM Journal on Mathematical Analysis, 52(6):6281–6296, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grenier:2022:PEN

[GNR22] Emmanuel Grenier, Toan T. Nguyen, and Igor Rodnianski. Plasma echoes near stable Penrose data. SIAM Journal on Mathematical Analysis, 54(1):940–953, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1392553.

Golding:2023:URT

[Gol23] William Golding. Unconditional regularity and trace results for the isentropic Euler equations with γ = 3. SIAM Journal on Mathematical Analysis, 55(5):5751–5781, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1554667.

Griso:2020:ABT

[GOW20] Georges Griso, Julia Orlik, and Stephan Wackerle. Asymptotic behavior for textiles. SIAM Journal on Mathematical Analysis, 52(2):1639–1689, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gobbino:2024:MPL

[GP24] Massimo Gobbino and Nicola Picenni. Monotonicity properties of limits of solutions to the semidiscrete scheme for a class of Perona–Malik type equations. SIAM Journal on Mathematical Analysis, 56(2):2034– 2062, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Gao:2023:ASD

[GR23] Yuan Gao and Jean-Michel Roquejoffre. Asymptotic stability for diffusion with dynamic boundary reaction from Ginzburg–Landau energy. SIAM Journal on Mathematical Analysis, 55(2):1246–1263, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469791.

Gell-Redman:2020:EAN

[GRHSZ20] Jesse Gell-Redman, Andrew Hassell, Jacob Shapiro, and Junyong Zhang. Existence and asymptotics of nonlinear Helmholtz eigenfunctions. SIAM Journal on Mathematical Analysis, 52(6):6180–6221, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golab:2023:TDH

[GRR23] Adam J. Golab, James C.

Robinson, and Jose L. Rodrigo. On two-dimensional harmonic extensions of vector fields and stellarator coils. SIAM Journal on Mathematical Analysis, 55(5):4880–4911, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458661.

Geng:2024:GEW

[GRZ24] Zhiyuan Geng, Arnab Roy, and Arghir Zarnescu. Global existence of weak solutions for a model of nematic liquid crystal-colloidal interactions. SIAM Journal on Mathematical Analysis, 56(4):4324–4355, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M161149X.

Gonzalez-Sanz:2024:NIO

[GSGM24] Alberto González-Sanz, Michel Groppe, and Axel Munk. Nonlinear inverse optimal transport: Identifiability of the transport cost from its marginals and optimal values. SIAM Journal on Mathematical Analysis, 56(6):7808–7829, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1624403.

Grasselli:2023:PFM

[GSS23] Maurizio Grasselli, Luca Scarpa, and Andrea Signori. On a phase field model for RNA-Protein dynamics. SIAM Journal on Mathematical Analysis, 55(1):405–457, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1483086.

Giunti:2021:ESM

[GV21] Arianna Giunti and Juan J. López Velázquez. Edge states for the magnetic Laplacian in domains with smooth boundary. SIAM Journal on Mathematical Analysis, 53(3): 3602–3643, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garde:2024:RCC

[GV24] Henrik Garde and Michael Vogelius. Reconstruction of cracks in Calderón’s inverse conductivity problem using energy comparisons. SIAM Journal on Mathematical Analysis, 56(1):727–745, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Guo:2021:GSS

[GWX21] Zhengguang Guo, Yun Wang, and Chunjing Xie. Global strong solutions to the inhomogeneous incompressible Navier–Stokes system in the exterior of a cylinder. SIAM Journal on Mathematical Analysis, 53(6):6804–6821, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gong:2021:PRS

[GWZ21] Huajun Gong, Changyou Wang, and Xiaotao Zhang. Partial regularity of suitable weak solutions of the Navier– Stokes–Planck–Nernst–Poisson equation. SIAM Journal on Mathematical Analysis, 53(3): 3306–3337, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2024:MNS

[GY24] Yuxia Guo and Yuanyang Yu. Multiple normalized solutions for first order Hamiltonian systems. SIAM Journal on Mathematical Analysis, 56(3):3861–3885, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1584575.

Guo:2025:SSI

[GY25] Yan Guo and Zhuolun Yang. Small scales in inviscid limits of steady fluids. SIAM Journal on Mathematical Analysis, 57(5):5216–5242, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1693726.

Geng:2020:OIP

[GZ20a] Jun Geng and Jinping Zhuge. Oscillatory integrals and periodic homogenization of Robin boundary value problems. SIAM Journal on Mathematical Analysis, 52(1):104–134, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gladkov:2020:ESM

[GZ20b] Nikita A. Gladkov and Alexander P. Zimin. An explicit solution for a multimarginal mass transportation problem. SIAM Journal on Mathematical Analysis, 52(4):3666–3696, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gui:2022:SLG

[GZ22] Guilong Gui and Ping Zhang.

Semiclassical limit of Gross– Pitaevskii equation with Dirichlet boundary condition. SIAM Journal on Mathematical Analysis, 54(1):1053–1104, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1369749.

Gong:2021:GSS

[GZW21] Weihua Gong, Fujun Zhou, and Weijun Wu. Global strong solution and incompressible Navier–Stokes–Fourier–Poisson limit of the Vlasov–Poisson– Boltzmann system. SIAM Journal on Mathematical Analysis, 53(6):6424–6470, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2022:TWP

[GZZ22] Haimo Guo, Meirong Zhang, and Yi Zhu. ThreeFold Weyl points for the Schrödinger operator with periodic potentials.

SIAM Journal on Mathematical Analysis, 54(3):3654–3695, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1410464.

Halla:2022:TED

[Hal22] Martin Halla. On the treatment of exterior domains for the time-harmonic equations of stellar oscillations. SIAM Journal on Mathematical Analysis, 54(5):5268–5290, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1418812.

Halla:2023:ESM

[Hal23] Martin Halla. On the existence and stability of modified Maxwell Steklov eigenvalues. SIAM Journal on Mathematical Analysis, 55(5):5445–5463, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1509266.

Hamfeldt:2025:SCP

[Ham25] Brittany Froese Hamfeldt. A strong comparison principle for the generalized Dirichlet problem for Monge-Ampère. SIAM Journal on Mathematical Analysis, 57(1):1–13, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1577006.

Han:2020:SPV

[Han20]  Weimin Han.  Singular

perturbations of variationalhemivariational inequalities. SIAM Journal on Mathematical Analysis, 52(2):1549–1566, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harutyunyan:2020:SKI

[Har20] Davit Harutyunyan. The sharp L^{p} Korn interpolation and second inequalities in thin domains. SIAM Journal on Mathematical Analysis, 52(6): 5775–5791, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrach:2023:CPF

[Har23] Bastian Harrach. The Calderón problem with finitely many unknowns is equivalent to convex semidefinite optimization. SIAM Journal on Mathematical Analysis, 55(5):5666–5684, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1544854.

Hartman:2024:SRN

[HBK24] Emmanuel Hartman, Martin Bauer, and Eric Klassen. Square root normal fields for Lipschitz surfaces and the Wasserstein Fisher Rao metric. SIAM Journal on Mathematical Analysis, 56(2):2171– 2190, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Hu:2021:SSS

[HC21] Yanbo Hu and Jianjun Chen. Sonic-supersonic solutions to a mixed-type boundary value problem for the two-dimensional full Euler equations. SIAM Journal on Mathematical Analysis, 53(2):1579–1629, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haas:2020:MEN

[HdRS20] Tobias Haas, Björn de Rijk, and Guido Schneider. Modulation equations near the Eckhaus boundary: The KdV equation. SIAM Journal on Mathematical Analysis, 52(6): 5389–5421, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2023:EDB

[He23] Siming He. Enhanced dissipation and blow-up suppression in a chemotaxis-fluid system. SIAM Journal on Mathematical Analysis, 55(4):2615–2643, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1517159.

Heihoff:2023:TNF

[Hei23] Frederic Heihoff. Two new functional inequalities and their application to the eventual smoothness of solutions to a chemotaxis-Navier–Stokes system with rotational flux. SIAM Journal on Mathematical Analysis, 55(6):7113–7154, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Hamster:2020:STW

[HH20] C. H. S. Hamster and H. J.

Hupkes. Stability of traveling waves for systems of reaction– diffusion equations with multiplicative noise. SIAM Journal on Mathematical Analysis, 52 (2):1386–1426, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Halla:2021:WPD

[HH21] Martin Halla and Thorsten Hohage. On the well-posedness of the damped time-harmonic Galbrun equation and the equations of stellar oscillations. SIAM Journal on Mathematical Analysis, 53(4):4068–4095, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hamamuki:2023:DAL

[HH23] Nao Hamamuki and Kazuya Hirose. A dynamical approach to lower gradient estimates for viscosity solutions of Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 55(4):3169–3204, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480483.

Huang:2024:ICL

[HH24] Yichen Huang and Aram W. Harrow. Improved concentration of Laguerre and Jacobi ensembles. SIAM Journal on Mathematical Analysis, 56(1): 554–567, January 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Hao:2024:HHL

[HHJ24] Xiaonan Hao, Jiaxi Huang, and Ning Jiang. On the hydrodynamics of hyperbolic liquid crystal elastomers. SIAM Journal on Mathematical Analysis, 56(6):7463–7507, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1587166.

Hieber:2023:GSW

[HHS23] Matthias Hieber, Amru Hussein, and Martin Saal. Global strong well-posedness of the stochastic bidomain equations with FitzHugh–Nagumo transport. SIAM Journal on Mathematical Analysis, 55 (4):4140–4161, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1478008.

Hittmeir:2011:CDP

[HJ11] Sabine Hittmeir and Ansgar Jüngel. Cross diffusion preventing blow-up in the two-dimensional Keller–Segel model. SIAM Journal on Mathematical Analysis, 43(2): 997–1022, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/1/
sjmaah/v43/i2/p997_s1. See

corrigendum [BCJ20].

Hong:2024:UOS

[HJ24] Younghun Hong and Sangdon Jin. Uniqueness and orbital stability of standing waves for the nonlinear Schrödinger equation with a partial confinement. SIAM Journal on Mathematical Analysis, 56(2):2714– 2737, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Ha:2020:LSA

[HJJ20] Seung-Yeal Ha, Shi Jin, and Jinwook Jung. Local sensitivity analysis for the Kuramoto– Daido model with random inputs in a large coupling regime. SIAM Journal on Mathematical Analysis, 52(2):2000–2040, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hong:2025:ISD

[HJK25] Jiho Hong, Bangti Jin, and Yavar Kian. Identification of a spatially-dependent variable order in one-dimensional subdiffusion. SIAM Journal on Mathematical Analysis, 57 (2):1315–1341, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1673371.

Huang:2021:SDG

[HJLZ21] Jiaxi Huang, Ning Jiang, Yi-Long Luo, and Lifeng Zhao. Small data global regularity for the 3-D Ericksen–Leslie hyperbolic liquid crystal model without kinematic transport. SIAM Journal on Mathematical Analysis, 53(1):530–573, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herrmann:2025:FDF

[HJM25] Michael Herrmann, Guillaume James, and Karsten Matthies. Fronts in dissipative Fermi– Pasta–Ulam–Tsingou chains. SIAM Journal on Mathematical Analysis, 57(5):5718–5745, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/25M1744885.

Huo:2022:WSU

[HJT22] Xiaokai Huo, Ansgar Jüngel, and Athanasios E. Tzavaras. Weak-strong uniqueness for Maxwell–Stefan systems. SIAM Journal on Mathematical Analysis, 54(3):3215–3252, ???? 2022. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145210X.

Hu:2023:GWP

[HK23] Kai Hu and Jie Kuang. Global well-posedness of shock front solutions to one-dimensional piston problem for combustion Euler flows. SIAM Journal on Mathematical Analysis, 55(3):2042–2110, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1493586.

Han:2023:LTB

[HKK23] Sungho Han, Moon-Jin Kang, and Jeongho Kim. Largetime behavior of composite waves of viscous shocks for the barotropic Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 55 (5):5526–5574, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1520712.

Hasik:2024:BWD

[HKlT24a] Karel Hasík, Jana Kopfová, Petra Náb lková, and Sergei Trofimchuk. Bistable wavefronts in the delayed Belousov–Zhabotinsky reaction. SIAM Journal on Mathematical Analysis, 56(1):1197– 1222, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Hellmuth:2024:KCT

[HKLT24b] Kathrin Hellmuth, Christian Klingenberg, Qin Li, and Min Tang. Kinetic chemotaxis tumbling kernel determined from macroscopic quantities. SIAM Journal on Mathematical Analysis, 56(1):568–587, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Hiltunen:2024:NPD

[HKSW24] Erik Orvehed Hiltunen, Joseph Kraisler, John C. Schotland, and Michael I. Weinstein. Nonlocal partial differential equations and quantum optics: Bound states and resonances. SIAM Journal on Mathematical Analysis, 56(3):3802–3831, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M158142X.

Han:2023:CSP

[HKY23] Beom-Seok Han, Kunwoo Kim, and Jaeyun Yi. The compact support property for solutions to the stochastic partial differential equations with colored noise. SIAM Journal on Mathematical Analysis, 55(6):7665– 7703, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Harrach:2020:MBI

[HL20a] Bastian Harrach and Yi-Hsuan Lin. Monotonicitybased inversion of the fractional Schödinger equation II. General potentials and stability. SIAM Journal on Mathematical Analysis, 52(1):402– 436, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2020:ISP

[HL20b] Guanghui Hu and Jingzhi Li. Inverse source problems in an inhomogeneous medium with a single far-field pattern. SIAM Journal on Mathematical Analysis, 52(5):5213–5231, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hitruhin:2021:LCH

[HL21] Lauri Hitruhin and Sauli Lindberg. Lamination convex hull of stationary incompressible porous media equations. SIAM Journal on Mathematical Analysis, 53(1):491–508, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2022:GWP

[HL22a] Xiangdi Huang and Jing Li. Global well-posedness of classical solutions to the Cauchy problem of twodimensional barotropic compressible Navier–Stokes system with vacuum and large initial data. SIAM Journal on Mathematical Analysis, 54 (3):3192–3214, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1440943.

Huicheng:2022:FCM

[HL22b] Yin Huicheng and Zhu Lu. Formation and construction of a multidimensional shock wave for the first-order hyperbolic conservation law with smooth initial data. SIAM Journal on Mathematical Analysis, 54 (2):2587–2610, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1406581.

Hao:2023:IVC

[HL23a] Tiantian Hao and Yanlin Liu. The influence of viscous coefficients on the lifespan of 3-D anisotropic Navier–Stokes system. SIAM Journal on Mathematical Analysis, 55 (3):2186–2210, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1461459.

Hensel:2023:SIL

[HL23b] Sebastian Hensel and Yuning Liu. The sharp interface limit of a Navier-Stokes/Allen– Cahn system with constant mobility: Convergence rates by a relative energy approach. SIAM Journal on Mathematical Analysis, 55(5):4751–4787, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1500587.

Hong:2021:FWT

[HLL21] Wei Hong, Shihu Li, and Wei Liu. Freidlin–Wentzell type large deviation principle for multiscale locally monotone SPDEs. SIAM Journal on Mathematical Analysis, 53(6): 6517–6561, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2023:VPB

[HLlZ23] Ling-Bing He, Yuanjie Lei, and Yu long Zhou. From the Vlasov–Poisson–Boltzmann system to the Vlasov–Poisson– Landau system with Coulomb potential. SIAM Journal on Mathematical Analysis, 55(2):701–772, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1515963.

Hajaiej:2024:ESR

[HLS24] Hichem Hajaiej, Yongming Luo, and Lingjie Song. On existence and stability results for normalized ground states of mass-subcritical biharmonic nonlinear Schrödinger equation on R^{d} × T^{n}. SIAM Journal on Mathematical Analysis, 56(4):4415–4439, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1543707.

Hu:2024:IPH

[HLZ24] Dian Hu, Qianfeng Li, and Yongqian Zhang. An inverse problem for hypersonic flow past a curved cone. SIAM Journal on Mathematical Analysis, 56(5):6915–6932, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1614870.

Hao:2025:SAS

[HM25] Zeming Hao and Shuang Miao. On stability analysis for steady states of the free boundary hard phase model in general relativity. SIAM Journal on Mathematical Analysis, 57(5):4908–4952, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1690333.

Hilhorst:2020:GIS

[HMNW20] Danielle Hilhorst, Hiroshi Matano, Thanh Nam Nguyen, and Hendrik Weber. Generation of interface for solutions of the mass conserved Allen– Cahn equation. SIAM Journal on Mathematical Analysis, 52 (3):2624–2654, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2024:ARS

[HMS24] Yanchen He, Carlo Marcati, and Christoph Schwab. Analytic regularity of solutions to the Navier–Stokes equations with mixed boundary conditions in polygons. SIAM Journal on Mathematical Analysis, 56(2):2488–2520, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Hening:2023:RSE

[HNNW23] Alexandru Hening, Dang H.

Nguyen, Nhu Nguyen, and Harrison Watts. Random switching in an ecosystem with two prey and one predator. SIAM Journal on Mathematical Analysis, 55(1):347–366, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459836.

Hartmann:2020:CGN

[HNS20] Carsten Hartmann, Lara Neureither, and Upanshu Sharma. Coarse graining of nonreversible stochastic differential equations: Quantitative results and connections to averaging. SIAM Journal on Mathematical Analysis, 52(3): 2689–2733, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2025:NLD

[HNX25] Lingjia Huang, Quoc-Hung Nguyen, and Yiran Xu. Nonlinear Landau damping for the 2D Vlasov–Poisson system with massless electrons around Penrose-stable equilibrium. SIAM Journal on Mathematical Analysis, 57 (2):1939–1963, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1595382.

Hogancamp:2021:BGF

[Hog21] Thomas Hogancamp. Broadening global families of anti-plane shear equilibria. SIAM Journal on Mathematical Analysis, 53 (5):5853–5879, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holle:2022:EDJ

[Hol22]  Yannick Holle.  Entropy

dissipation at the junction for macroscopic traffic flow models. SIAM Journal on Mathematical Analysis, 54 (1):954–985, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1423920.

Hayashi:2025:GSG

[HOV25] Masayuki Hayashi, Tohru Ozawa, and Nicola Visciglia. Global H^{2}-solutions for the generalized derivative NLS on T. SIAM Journal on Mathematical Analysis, 57 (2):1483–1501, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1668317.

Heinze:2023:NCI

[HPS23] Georg Heinze, Jan-Frederik Pietschmann, and Markus Schmidtchen. Nonlocal crossinteraction systems on graphs: Nonquadratic Finslerian structure and nonlinear mobilities. SIAM Journal on Mathematical Analysis, 55(6):7039–7076, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Hoang:2020:SSN

[HR20] Vu Hoang and Maria Radosz.

Singular solutions for nonlocal systems of evolution equations with vorticity stretching. SIAM Journal on Mathematical Analysis, 52(2):2158–2178, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsu:2021:ROE

[HR21] Ting-Hao Hsu and Shigui Ruan. Relaxation oscillations and the entry-exit function in multidimensional slow-fast systems. SIAM Journal on Mathematical Analysis, 53(4):3717– 3758, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2023:GEV

[HR23] Yuxi Hu and Reinhard Racke.

Global existence versus blowup for multidimensional hyperbolized compressible Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 55(5):4788–4815, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1497468.

Hwang:2022:RBM

[HRY22] Byung-Hoon Hwang, Tommaso Ruggeri, and Seok-Bae Yun. On a relativistic BGK model for polyatomic gases near equilibrium. SIAM Journal on Mathematical Analysis, 54(3):2906–2947, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1404946.

Hao:2024:SCP

[HRZ24] Zimo Hao, Michael Röckner, and Xicheng Zhang. Strong convergence of propagation of chaos for McKean–Vlasov SDEs with singular interactions. SIAM Journal on Mathematical Analysis, 56(2):2661– 2713, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Hosseinkhan:2023:SDB

[HS23] Alireza Hosseinkhan and Ralph E. Showalter. Semilinear degenerate Biot–Signorini system. SIAM Journal on Mathematical Analysis, 55(5):5643–5665, ???? 2023. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1505335.

Hilder:2024:QCG

[HS24] Bastian Hilder and Upanshu Sharma. Quantitative coarsegraining of Markov chains. SIAM Journal on Mathematical Analysis, 56(1):913–954, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

He:2024:LSK

[HSSY24] Xiaolong He, Jia Shi, Yunfeng Shi, and Xiaoping Yuan. On linear stability of KAM tori via the Craig–Wayne– Bourgain method. SIAM Journal on Mathematical Analysis, 56(3):3605–3645, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Hsiao:2020:TDB

[HSV20] George C. Hsiao and Tonatiuh Sánchez-Vizuet. Timedomain boundary integral methods in linear thermoelasticity. SIAM Journal on Mathematical Analysis, 52(3):2463– 2490, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsiao:2021:BIF

[HSV21] George C. Hsiao and Tonatiuh Sánchez-Vizuet. Boundary integral formulations for transient linear thermoelasticity with combined-type boundary conditions. SIAM Journal on Mathematical Analysis, 53(4):

3888–3911, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsiao:2025:BFF

[HSVW25] George C. Hsiao, Tonatiuh Sánchez-Vizuet, and Wolfgang L. Wendland. Boundaryfield formulation for transient electromagnetic scattering by dielectric scatterers and coated conductors. SIAM Journal on Mathematical Analysis, 57 (2):1502–1525, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1667610.

Ha:2020:DLP

[HSZ20] Seung-Yeal Ha, Woojoo Shim, and Yinglong Zhang. A diffusion limit for the parabolic Kuramoto–Sakaguchi equation with inertia. SIAM Journal on Mathematical Analysis, 52(2): 1591–1638, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hraivoronska:2023:DLR

[HT23] Anastasiia Hraivoronska and Oliver Tse. Diffusive limit of random walks on tessellations via generalized gradient flows. SIAM Journal on Mathematical Analysis, 55(4):2948–2995, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1478306.

Han:2021:SWW

[HTY21] Fangyu Han, Zhong Tan, and Ganshan Yang. Stability of the Walker wall in two dimensions. SIAM Journal on Mathematical Analysis, 53(6):7024–7061, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2022:SSP

[Hu22] Yanbo Hu. On a supersonicsonic patch in the threedimensional steady axisymmetric transonic flows. SIAM Journal on Mathematical Analysis, 54(2):1515–1542, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1393108.

Hu:2024:GTW

[Hu24a] Yiran Hu. Global in time weak solutions to singular threedimensional quasi-geostrophic systems. SIAM Journal on Mathematical Analysis, 56 (5):5881–5914, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1552917.

Hu:2024:SCS

[Hu24b] Zhongtian Hu. Suppression of chemotactic singularity via viscous flow with large buoyancy.

SIAM Journal on Mathematical Analysis, 56(6):7866–7902, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1669001.

Huang:2024:ESD

[Hua24]  Bingkang Huang.  Exis-

tence and stability of dissipative measure-valued solutions to the full compressible magnetohydrodynamic flows. SIAM Journal on Mathematical Analysis, 56(2):2213–2247, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Hassainia:2022:MVP

[HW22a] Zineb Hassainia and Miles H. Wheeler. Multipole vortex patch equilibria for active scalar equations. SIAM Journal on Mathematical Analysis, 54(6):6054–6095, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1415339.

Henderson:2022:LWP

[HW22b] Christopher Henderson and Weinan Wang. Local wellposedness for the Boltzmann equation with very soft potential and polynomially decaying initial data. SIAM Journal on Mathematical Analysis, 54(3):2845–2875, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1427504.

Huang:2022:BLS

[HW22c] Feimin Huang and Yong Wang.

Boundary layer solution of the Boltzmann equation for diffusive reflection boundary conditions in half-space. SIAM Journal on Mathematical Analysis, 54(3):3480–3534, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M142945X.

Holler:2024:ERR

[HW24] Martin Holler and Benedikt Wirth. Exact reconstruction and reconstruction from noisy data with anisotropic total variation. SIAM Journal on Mathematical Analysis, 56(3): 2938–2967, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Huang:2024:UCW

[HWWW24] Feimin Huang, Weiqiang Wang, Yi Wang, and Yong Wang. Uniqueness of composite wave of shock and rarefaction in the inviscid limit of Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 56(3):3924–3967, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M156584X.

Hu:2025:THE

[HWXZ25] Guanghui Hu, Tianjiao Wang, Xiang Xu, and Yue Zhao. Time-harmonic elastic scattering by unbounded deterministic and random rough surfaces in three dimensions. SIAM Journal on Mathematical Analysis, 57(2):1964–1989, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1640306.

Huang:2022:PDT

[HWZ22] Mingdi Huang, Shi-Liang Wu, and Xiao-Qiang Zhao. Propagation dynamics for timeperiodic and partially degenerate reaction–diffusion systems. SIAM Journal on Mathematical Analysis, 54(2):1860–1897, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1397234.

He:2022:LSA

[HXH22] Wenhua He, Ruixiang Xing, and Bei Hu. Linear stability analysis for a free boundary problem modeling multilayer tumor growth with time delay. SIAM Journal on Mathematical Analysis, 54(4):4238–4276, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437494.

Hwang:2021:SSA

[HY21] Byung-Hoon Hwang and Seok-Bae Yun. Stationary solutions to the Anderson–Witting model of the relativistic Boltzmann equation in a bounded interval. SIAM Journal on Mathematical Analysis, 53(1): 730–753, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsieh:2022:ESP

[HY22a] Chia-Yu Hsieh and Yong Yu.

Existence of solutions to the Poisson–Nernst–Planck system with singular permanent charges in R^{2}. SIAM Journal on Mathematical Analysis, 54(1):1223–1240, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1399841.

Huang:2022:CGC

[HY22b] Feimin Huang and Yue Yao.

Continued gravitational collapse for gaseous star and pressureless Euler–Poisson system. SIAM Journal on Mathematical Analysis, 54(3):3139–3160, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450902.

Holmer:2022:BOS

[HZ22a] Justin Holmer and Katherine Zhiyuan Zhang. Benjamin– Ono soliton dynamics in a slowly varying potential revisited. SIAM Journal on Mathematical Analysis, 54 (2):2634–2690, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1425177.

Huang:2022:HOA

[HZ22b] Lin Huang and Lun Zhang. Higher order Airy and Painlevé asymptotics for the mKdV hierarchy. SIAM Journal on Mathematical Analysis, 54 (5):5291–5334, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448008.

He:2025:CBS

[HZ25] Yi He and Jianjun Zhang. Concentrating bound states for singularly perturbed nonlinear Dirichlet problems involving critical growth. SIAM Journal on Mathematical Analysis, 57(2):1832–1865, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1675898.

Ishii:2022:NNP

[IK22] Hitoshi Ishii and Taiga Kumagai. Nonlinear Neumann problems for fully nonlinear elliptic PDEs on a quadrant.

SIAM Journal on Mathematical Analysis, 54(6):5854–5887, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1444199.

Ignat:2020:GUE

[IKL20] Radu Ignat, Matthias Kurzke, and Xavier Lamy. Global uniform estimate for the modulus of two-dimensional Ginzburg– Landau vortexless solutions with asymptotically infinite boundary energy. SIAM Journal on Mathematical Analysis, 52(1):524–542, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iglesias:2021:CLS

[IM21a] José A. Iglesias and Gwenael Mercier. Convergence of level sets in total variation denoising through variational curvatures in unbounded domains. SIAM Journal on Mathematical Analysis, 53(2):1509–1545, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ilmavirta:2021:XRT

[IM21b] Joonas Ilmavirta and Keijo Mönkkönen. X-ray tomography of one-forms with partial data. SIAM Journal on Mathematical Analysis, 53(3):3002– 3015, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ishige:2020:BFO

[IMO20] Kazuhiro Ishige, Nobuhito Miyake, and Shinya Okabe. Blowup for a fourth-order parabolic equation with gradient nonlinearity. SIAM Journal on Mathematical Analysis, 52(1):927–953, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Imbert:2020:GLB

[IMS20] Cyril Imbert, Clément Mouhot, and Luis Silvestre. Gaussian lower bounds for the Boltzmann equation without cutoff. SIAM Journal on Mathematical Analysis, 52(3):2930–2944, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Isaacson:2022:MFL

[IMS22] Samuel A. Isaacson, Jingwei Ma, and Konstantinos Spiliopoulos. Mean field limits of particle-based stochastic reaction–diffusion models. SIAM Journal on Mathematical Analysis, 54(1):453–511, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1365600.

Iwaniec:2021:NMP

[IOPR21] Tadeusz Iwaniec, Jani Onninen, Pekka Pankka, and Teresa Radice. A neohookean model of plates. SIAM Journal on Mathematical Analysis, 53(1): 509–529, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ifrim:2022:NPC

[IPTT22] Mihaela Ifrim, Ben Pineau, Daniel Tataru, and Mitchell Taylor. No pure capillary solitary waves exist in 2D finite depth. SIAM Journal on Mathematical Analysis, 54 (4):4452–4464, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1414723.

Ignatova:2021:GSN

[IS21] Mihaela Ignatova and Jingyang Shu. Global solutions of the Nernst–Planck–Euler equations. SIAM Journal on Mathematical Analysis, 53(5):5507– 5547, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iyer:2022:BHT

[IV22] Gautam Iyer and Truong-Son Van. Bounds on the heat transfer rate via passive advection. SIAM Journal on Mathematical Analysis, 54(2):1927–1965, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1394497.

Imanuvilov:2020:CEL

[IY20] Oleg Yu. Imanuvilov and Masahiro Yamamoto. Carleman estimate for linear viscoelasticity equations and an inverse source problem. SIAM Journal on Mathematical Analysis, 52(1):718–791, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iyer:2023:QDE

[IZ23] Gautam Iyer and Hongyi Zhou. Quantifying the dissipation enhancement of cellular flows. SIAM Journal on Mathematical Analysis, 55(6):6496–6516, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Jacobs:2023:ESR

[Jac23] Matt Jacobs. Existence of solutions to reaction cross diffusion systems. SIAM Journal on Mathematical Analysis, 55(6):6991–7023, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Jia:2020:LID

[Jia20a] Hao Jia. Linear inviscid damping near monotone shear flows. SIAM Journal on Mathematical Analysis, 52(1):623–652, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2020:SRE

[Jia20b] Jin-Cheng Jiang. Sharp regularizing estimates for the gain term of the Boltzmann collision operator. SIAM Journal on Mathematical Analysis, 52 (4):4068–4100, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jing:2020:UHA

[Jin20] Wenjia Jing. A unified homogenization approach for the Dirichlet problem in perforated domains. SIAM Journal on Mathematical Analysis, 52(2): 1192–1220, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2021:NDL

[JJL21] Song Jiang, Qiangchang Ju, and Yongkai Liao. Nonequilibriumdiffusion limit of the compressible Euler–P1 approximation radiation model at low Mach number. SIAM Journal on Mathematical Analysis, 53(2): 2491–2522, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2022:GST

[JJZ22] Fei Jiang, Song Jiang, and Youyi Zhao. Global solutions of the three-dimensional incompressible ideal MHD equations with velocity damping in horizontally periodic domains. SIAM Journal on Mathematical Analysis, 54(4):4891–4929, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1437974.

Jean:2024:DOC

[JJZ24] Frédéric Jean, Othmane Jerhaoui, and Hasnaa Zidani. Deterministic optimal control on Riemannian manifolds under probability knowledge of the initial condition. SIAM Journal on Mathematical Analysis, 56(3):3326–3356, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Jin:2022:DKT

[JK22] Jiaxin Jin and Chanwoo Kim. Damping of kinetic transport equation with diffuse boundary condition. SIAM Journal on Mathematical Analysis, 54(5):5524–5550, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1455358.

Jo:2024:NRS

[JKL24] Min Jun Jo, Junha Kim, and Jihoon Lee. Nonconvergence of the rotating stratified flows toward the quasi-geostrophic dynamics. SIAM Journal on Mathematical Analysis, 56(3): 3357–3385, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Jin:2021:RST

[JKZ21] Bangti Jin, Yavar Kian, and Zhi Zhou. Reconstruction of a space-time-dependent source in subdiffusion models via a perturbation approach. SIAM Journal on Mathematical Analysis, 53(4):4445–4473, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jung:2021:SEL

[JL21] Younghoon Jung and Mikyoung Lim. Series expansions of the layer potential operators using the Faber polynomials and their applications to the transmission problem. SIAM Journal on Mathematical Analysis, 53(2):1630–1669, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiao:2023:DNN

[JLL^{+}23] Yuling Jiao, Yanming Lai, Xiliang Lu, Fengru Wang, Jerry Zhijian Yang, and Yuanyuan Yang. Deep neural networks with ReLU-sine-exponential activations break curse of dimensionality in approximation on Hölder class. SIAM Journal on Mathematical Analysis, 55(4):3635–3649, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144431X.

Jansen:2023:LTB

[JLN23] Jonas Jansen, Christina Lienstromberg, and Katerina Nik. Long-time behavior and stability for quasilinear doubly degenerate parabolic equations of higher order. SIAM Journal on Mathematical Analysis, 55(2):674–700, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1491137.

Jang:2025:CDS

[JLS25] Juhi Jang, Jiaqi Liu, and Matthew Schrecker. Converging/Diverging self-similar shock waves: From collapse to reflection. SIAM Journal on Mathematical Analysis, 57(1):190–232, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1653240.

Jin:2023:SCS

[JLX23] Kun-Peng Jin, Jin Liang, and Ti-Jun Xiao. Stability for a coupled system of second-order evolution equations with indirect memorydamping. SIAM Journal on Mathematical Analysis, 55(6): 7155–7188, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Jiang:2025:AQF

[JLZ25] Kai Jiang, Shifeng Li, and Pingwen Zhang. On the approximation of quasiperiodic functions with Diophantine frequencies by periodic functions. SIAM Journal on Mathematical Analysis, 57(1):951–978, ???? 2025. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M165925X.

Ji:2023:ODR

[JM23] Shanming Ji and Ming Mei.

Optimal decay rates of the compressible Euler equations with time-dependent damping in R^{n}: (II) overdamping case. SIAM Journal on Mathematical Analysis, 55(2):1048–1099, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144476X.

Jimenez:2023:DSH

[JMQ23] Chloé Jimenez, Antonio Marigonda, and Marc Quincampoix. Dynamical systems and Hamilton– Jacobi–Bellman equations on the Wasserstein space and their L^{2} representations. SIAM Journal on Mathematical Analysis, 55(5):5919–5966, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1503877.

Johnson:2020:MIV

[JP20] Mathew A. Johnson and Wesley R. Perkins. Modulational instability of viscous fluid conduit periodic waves. SIAM Journal on Mathematical Analysis, 52(1):277–305, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ji:2021:CPP

[JQSY21] Min Ji, Weiwei Qi, Zhongwei Shen, and Yingfei Yi. Convergence to periodic probability solutions in Fokker–Planck equations. SIAM Journal on Mathematical Analysis, 53(2): 1958–1992, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jenssen:2020:ODS

[JR20] Helge Kristian Jenssen and Johanna Ridder. Onedimensional scalar conservation laws with regulated data. SIAM Journal on Mathematical Analysis, 52(3):3114–3130, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Josien:2021:QHC

[JR21] Marc Josien and Claudia Raithel. Quantitative homogenization for the case of an interface between two heterogeneous media. SIAM Journal on Mathematical Analysis, 53 (1):813–854, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Josien:2024:SHG

[JRS24] Marc Josien, Claudia Raithel, and Mathias Schäffner. Stochastic homogenization and geometric singularities: A study on corners. SIAM Journal on Mathematical Analysis, 56(2):

2395–2455, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Jang:2023:ERS

[JSW23] Juhi Jang, Walter A. Strauss, and Yilun Wu. Existence of rotating stars with variable entropy. SIAM Journal on Mathematical Analysis, 55(4):2704–2739, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1498516.

Jing:2023:EFP

[JTY23] Wenjia Jing, Hung V. Tran, and Yifeng Yu. Effective fronts of polygon shapes in two dimensions. SIAM Journal on Mathematical Analysis, 55(6):6764–6777, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Juillet:2020:SMT

[Jui20] Nicolas Juillet. On a solution to the Monge transport problem on the real line arising from the strictly concave case. SIAM Journal on Mathematical Analysis, 52(5):4783–4805, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jung:2025:GDE

[Jun25] Junhwa Jung. Global diffusive expansion of Boltzmann equation in exterior domain. SIAM Journal on Mathematical Analysis, 57(2):1781–1831, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1669992.

Jang:2024:TDB

[JV24a] Jin Woo Jang and Juan J. L. Velázquez. On the temperature distribution of a body heated by radiation. SIAM Journal on Mathematical Analysis, 56(3): 3478–3508, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Jungel:2024:DDD

[JV24b] Ansgar Jüngel and Martin Vetter. Degenerate driftdiffusion systems for memristors. SIAM Journal on Mathematical Analysis, 56 (6):7780–7807, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1620235.

Ju:2023:ILI

[JW23] Qiangchang Ju and Jiawei Wang. Incompressible limit of ideal magnetohydrodynamic equations with a perfectly conducting wall condition. SIAM Journal on Mathematical Analysis, 55(6):7549–7574, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Jin:2023:GSS

[JWW23] Hai-Yang Jin, Zhi-An Wang, and Leyun Wu. Global solvability and stability of an alarm-taxis system. SIAM Journal on Mathematical Analysis, 55(4):2838–2876, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1477143.

Ji:2022:GWP

[JWX22] Ruihong Ji, Jiahong Wu, and Xiaojing Xu. Global wellposedness of the 2D MHD equations of damped wave type in Sobolev space. SIAM Journal on Mathematical Analysis, 54(6):6018–6053, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465342.

Ji:2022:ASS

[JZZ22] Shanming Ji, Minyi Zhang, and Changjiang Zhu. Asymptotic stability of solutions to a hyperbolic-elliptic coupled system of radiating gas on the half line. SIAM Journal on Mathematical Analysis, 54(3):3884–3929, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459848.

Klibanov:2024:LSE

[KA24] Michael V. Klibanov and Yurii Averboukh. Lipschitz stability estimate and uniqueness in the retrospective analysis for the mean field games system via two Carleman estimates.

SIAM Journal on Mathematical Analysis, 56(1):616–636, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Katsevich:2020:RAI

[Kat20] Alexander Katsevich. Resolution analysis of inverting the generalized Radon transform from discrete data in R^{3}. SIAM Journal on Mathematical Analysis, 52(4):3990–4021, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Katsevich:2023:NRA

[Kat23] Alexander Katsevich. Novel resolution analysis for the Radon transform in R2 for functions with rough edges. SIAM Journal on Mathematical Analysis, 55(5):4255–4296, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1502252.

Korikov:2023:SDR

[KB23] Dmitrii V. Korikov and Mikhail I. Belishev. Stability of determination of Riemann surface from its Dirichlet-to-Neumann map in terms of Teichmüller distance. SIAM Journal on Mathematical Analysis, 55(6):7426–7448, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Kang:2021:QPC

[KbP21] Kyungkeun Kang and Xing bin Pan. On a quasilinear parabolic curl system motivated by time evolution of Meissner states of superconductors. SIAM Journal on Mathematical Analysis, 53(6): 6471–6516, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kilque:2022:WNM

[Kil22] Corentin Kilque. Weakly nonlinear multiphase geometric optics for hyperbolic quasilinear boundary value problems: Construction of a leading profile. SIAM Journal on Mathematical Analysis, 54 (2):2413–2507, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413596.

Kim:2021:HAL

[Kim21] Soojung Kim. Hydrodynamics of anisotropic liquid crystals in an applied magnetic field. SIAM Journal on Mathematical Analysis, 53(3):3123–3157, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Karch:2020:FNN

[KKK20] Grzegorz Karch, Moritz Kassmann, and Miłosz Krupski. A framework for nonlocal, nonlinear initial value problems. SIAM Journal on Mathematical Analysis, 52(3):2383–2410, ???? 2020. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Kracmar:2022:WSB

[KKNN22] Stanislav Kracmar, Young-Sam Kwon, Sárka Necasová, and Antonín Novotný. Weak solutions for a bifluid model for a mixture of two compressible noninteracting fluids with general boundary data. SIAM Journal on Mathematical Analysis, 54(1):818–871, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1419246.

Kagaya:2021:SNB

[KL21] Takashi Kagaya and Qing Liu. Singular Neumann boundary problems for a class of fully nonlinear parabolic equations in one dimension. SIAM Journal on Mathematical Analysis, 53(4):4350–4385, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kupferman:2022:SVC

[KL22] Raz Kupferman and Roee Leder. On Saint-Venant compatibility and stress potentials in manifolds with boundary and constant sectional curvature. SIAM Journal on Mathematical Analysis, 54 (4):4625–4657, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466736.

Kim:2023:TGN

[KL23] Inwon Kim and Jona Lelmi.

Tumor growth with nutrients: Stability of the tumor patches. SIAM Journal on Mathematical Analysis, 55(5):5862–5892, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1542222.

Kachmar:2024:GBL

[KL24a] Ayman Kachmar and Vladimir Lotoreichik. A geometric bound on the lowest magnetic Neumann eigenvalue via the torsion function. SIAM Journal on Mathematical Analysis, 56(4):5723–5745, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1624658.

Kim:2024:VCB

[KL24b] Chanwoo Kim and Joonhyun La. Vorticity convergence from Boltzmann to 2D incompressible Euler equations below Yudovich class. SIAM Journal on Mathematical Analysis, 56(3): 3144–3202, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Kang:2025:AGT

[KLLT25] Kyungkeun Kang, Baishun Lai, Chen-Chih Lai, and Tai-Peng Tsai. Applications of the Green tensor estimates of the nonstationary Stokes system in the half space. SIAM Journal on Mathematical Analysis, 57(1):1137–1194, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1675801.

Kagaya:2024:RFV

[KLM24] Takashi Kagaya, Qing Liu, and Hiroyoshi Mitake. A representation formula for viscosity solutions of nonlocal Hamilton– Jacobi equations and applications. SIAM Journal on Mathematical Analysis, 56 (5):5807–5839, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1608136.

Kavallaris:2023:DDB

[KLS23] Nikos I. Kavallaris, Evangelos Latos, and Takashi Suzuki. Diffusion-driven blow-up for a nonlocal Fisher–KPP type model. SIAM Journal on Mathematical Analysis, 55 (3):2411–2433, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145519X.

Khimin:2024:AST

[KLSW24] Denis Khimin, Johannes Lankeit, Marc C. Steinbach, and Thomas Wick. Analysis of a space-time phase-field fracture complementarity model and its optimal control formulation. SIAM Journal on Mathematical Analysis, 56(5):6192–6212, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1605314.

Kostianko:2022:IMS

[KLSZ22] Anna Kostianko, Xinhua Li, Chunyou Sun, and Sergey Zelik. Inertial manifolds via spatial averaging revisited. SIAM Journal on Mathematical Analysis, 54(1):268–305, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1375437.

Korobkov:2022:EHI

[KLW22a] Mikhail Korobkov, Wenqi Lyu, and Shangkun Weng. On the existence of helical invariant solutions to steady Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 54(3):3974–3996, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1343324.

Kow:2022:CPF

[KLW22b] Pu-Zhao Kow, Yi-Hsuan Lin, and Jenn-Nan Wang. The Calderón problem for the fractional wave equation: Uniqueness and optimal stability.

SIAM Journal on Mathematical Analysis, 54(3):3379–3419, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1444941.

Kawamoto:2020:LSI

[KM20] Atsushi Kawamoto and Manabu Machida. Lipschitz stability in inverse source and inverse coefficient problems for a first- and half-order timefractional diffusion equation. SIAM Journal on Mathematical Analysis, 52(1):967–1005, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2023:ILP

[KM23] Inwon Kim and Antoine Mellet. Incompressible limit of a porous media equation with bistable and monostable reaction term. SIAM Journal on Mathematical Analysis, 55(5):5318–5344, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1517147.

Kinra:2024:TWA

[KM24] Kush Kinra and Manil T. Mohan. Theory of weak asymptotic autonomy of pullback stochastic weak attractors and its applications to 2D stochastic Euler equations driven by multiplicative noise.

SIAM Journal on Mathematical Analysis, 56(5):6268–6301, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1637878.

Kupferman:2025:LIE

[KM25] Raz Kupferman and Cy Maor.

Linearization in incompatible elasticity for general ambient spaces. SIAM Journal on Mathematical Analysis, 57 (5):5598–5627, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1701198.

Korotyaev:2024:IRP

[KMM24] Evgeny Korotyaev, Andrea Mantile, and Dmitrii Mokeev. Inverse resonance problems for energy-dependent potentials on the half-line. SIAM Journal on Mathematical Analysis, 56(2):2115–2148, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Killip:2021:SCQ

[KMV21] Rowan Killip, Jason Murphy, and Monica Visan. Scattering for the cubic-quintic NLS: Crossing the virial threshold. SIAM Journal on Mathematical Analysis, 53(5):5803–5812, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kamenskii:2020:CPP

[KMW20] Mikhail Kamenskii, Oleg Makarenkov, and Lakmi N. Wadippuli. A continuation principle for periodic BV-Continuous state-dependent sweeping processes. SIAM Journal on Mathematical Analysis, 52(6):5598–5626, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalinin:2024:SIB

[KMW24] Konstantin Kalinin, Govind Menon, and Bian Wu. Scale invariant bounds for mixing in the Rayleigh–Taylor instability. SIAM Journal on Mathematical Analysis, 56 (6):7846–7865, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1630402.

Kaneko:2020:APS

[KMY20] Yuki Kaneko, Hiroshi Matsuzawa, and Yoshio Yamada. Asymptotic profiles of solutions and propagating terrace for a free boundary problem of nonlinear diffusion equation with positive bistable nonlinearity. SIAM Journal on Mathematical Analysis, 52(1): 65–103, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaltenbacher:2022:PAQ

[KN22] Barbara Kaltenbacher and Vanja Nikolić. Parabolic approximation of quasilinear wave equations with applications in nonlinear acoustics. SIAM Journal on Mathematical Analysis, 54(2):1593–1622, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1380430.

Kehoe:2023:JGF

[KN23] Matthew Kehoe and David P.

Nicholls. Joint geometry/ frequency analyticity of fields scattered by periodic layered media. SIAM Journal on Mathematical Analysis, 55 (3):1737–1765, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1477568.

Khoshnevisan:2021:SSE

[KNP21] Davar Khoshnevisan, David Nualart, and Fei Pu. Spatial stationarity, ergodicity, and CLT for parabolic Anderson model with delta initial condition in dimension d ≥ 1. SIAM Journal on Mathematical Analysis, 53(2):2084–2133, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kukavica:2024:MEB

[KO24] Igor Kukavica and Wojciech S.

Ożański. On a model of an elastic body fully immersed in a viscous incompressible fluid with small data. SIAM Journal on Mathematical Analysis, 56(1):746–761, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Kofroth:2023:ILE

[Kof23] Collin Kofroth. Integrated local energy decay for the damped wave equation on stationary space-times. SIAM Journal on Mathematical Analysis, 55(5):5086–5126, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1490545.

Korolev:2022:TLN

[Kor22] Yury Korolev. Two-layer neural networks with values in a Banach space. SIAM Journal on Mathematical Analysis, 54(6):6358–6389, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1458144.

Kreutz:2021:MVV

[KP21] Leonard C. Kreutz and Paolo Piovano. Microscopic validation of a variational model of epitaxially strained crystalline films. SIAM Journal on Mathematical Analysis, 53(1):453– 490, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kapitula:2020:RKM

[KPS20] Todd Kapitula, Ross Parker, and Björn Sandstede. A reformulated Krein matrix for star-even polynomial operators with applications. SIAM Journal on Mathematical Analysis, 52(5):4705–4750, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Korner:2021:SLS

[KR21] Jacob Körner and Gerhard Rein. Strong Lagrangian solutions of the (relativistic) Vlasov–Poisson system for NonSmooth, spherically symmetric data. SIAM Journal on Mathematical Analysis, 53(4): 4985–4996, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaltenbach:2023:ESS

[KR23] Alex Kaltenbach and Michael Ruzicka. Existence of steady solutions for a general model for micropolar electrorheological fluid flows. SIAM Journal on Mathematical Analysis, 55(3):2238–2260, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1500599.

Kress:2021:DBC

[Kre21]  Rainer Kress.  The DB

boundary condition in electromagnetic scattering revisited.

SIAM Journal on Mathematical Analysis, 53(3):3661–3680, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krishnan:2021:SFD

[KRS21] Venkateswaran P. Krishnan, Rakesh Rakesh, and Soumen Senapati. Stability for a formally determined inverse problem for a hyperbolic PDE with space and time dependent coefficients. SIAM Journal on Mathematical Analysis, 53(6): 6822–6846, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurt:2021:FTB

[KS21] Halil Ibrahim Kurt and Wenxian Shen. Finite-time blow-up prevention by logistic source in parabolic–elliptic chemotaxis models with singular sensitivity in any dimensional setting. SIAM Journal on Mathematical Analysis, 53(1):973–1003, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [KS24a].

Kurt:2024:EFT

[KS24a] Halil Ibrahim Kurt and Wenxian Shen. Erratum: Finitetime blow-up prevention by logistic source in parabolic– elliptic chemotaxis models with singular sensitivity in any dimensional setting. SIAM Journal on Mathematical Analysis, 56(6):8136–8148, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M163757X. See [KS21].

Kuto:2024:MIS

[KS24b] Kousuke Kuto and Homare Sato. Morse index of steadystates to the SKT model with Dirichlet boundary conditions. SIAM Journal on Mathematical Analysis, 56(4):5386–5408, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1627705.

Klein:2023:LBR

[KSS23] Christian Klein, Johannes Sjöstrand, and Nikola Stoilov. Large |k| behavior for the reflection coefficient for Davey– Stewartson II equations. SIAM Journal on Mathematical Analysis, 55(5):4912–4937, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487916.

Kow:2024:SBC

[KSS24] Pu-Zhao Kow, Mikko Salo, and Henrik Shahgholian. On scattering behavior of corner domains with anisotropic inhomogeneities. SIAM Journal on Mathematical Analysis, 56(4):4834–4853, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1603029.

Kian:2023:LSR

[KST23] Yavar Kian, Éric Soccorsi, and Faouzi Triki. Logarithmic stable recovery of the source and the initial state of time fractional diffusion equations. SIAM Journal on Mathematical Analysis, 55(4):3888–3902, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1504743.

Kraisler:2024:DDE

[KSW24] Joseph Kraisler, Amir Sagiv, and Michael I. Weinstein. Dispersive decay estimates for Dirac equations with a domain wall. SIAM Journal on Mathematical Analysis, 56 (6):7194–7227, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1625378.

Kim:2020:EUR

[KT20] Hyunseok Kim and Tai-Peng Tsai. Existence, uniqueness, and regularity results for elliptic equations with drift terms in critical weak spaces. SIAM Journal on Mathematical Analysis, 52(2):1146–1191, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kappeler:2022:ALT

[KT22a] T. Kappeler and P. Topalov.

Arnold–Liouville theorem for integrable PDEs: a case study of the focusing NLS equation. SIAM Journal on Mathematical Analysis, 54(4):4334–4378, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443169.

Kiselev:2022:GRN

[KT22b] Alexander Kiselev and Changhui Tan. Global regularity for a nonlocal PDE describing evolution of polynomial roots under differentiation. SIAM Journal on Mathematical Analysis, 54(3):3161–3191, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422859.

Korn:2024:GWP

[KT24] Peter Korn and E. S. Titi.

Global well-posedness of the primitive equations of largescale ocean dynamics with the Gent–McWilliams–Redi eddy parametrization model. SIAM Journal on Mathematical Analysis, 56(6):8011–8036, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1564316.

Kim:2020:SCS

[KTT20] Yeoneung Kim, Hung V. Tran, and Son N. Tu. Stateconstraint static Hamilton– Jacobi equations in nested domains. SIAM Journal on Mathematical Analysis, 52(5): 4161–4184, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kow:2021:OIS

[KUW21] Pu-Zhao Kow, Gunther Uhlmann, and Jenn-Nan Wang. Optimality of increasing stability for an inverse boundary value problem. SIAM Journal on Mathematical Analysis, 53(6): 7062–7080, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krupa:2020:SUP

[KV20] Sam G. Krupa and Alexis F. Vasseur. Stability and uniqueness for piecewise smooth solutions to a nonlocal scalar conservation law with applications to Burgers–Hilbert equation. SIAM Journal on Mathematical Analysis, 52(3):2491–2530, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koumatos:2021:QGI

[KV21] Konstantinos Koumatos and Andreas P. Vikelis. Aquasiconvexity, Gårding inequalities, and applications in PDE constrained problems in dynamics and statics. SIAM Journal on Mathematical Analysis, 53(4):4178–4211, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koralov:2022:GLT

[KV22] L. Koralov and B. Vainberg. Global limit theorem for parabolic equations with a potential. SIAM Journal on Mathematical Analysis, 54 (2):2097–2113, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1429370.

Kimura:2021:QEC

[KvM21] Masato Kimura and Patrick van Meurs. Quantitative estimate of the continuum approximations of interacting particle systems in one dimension. SIAM Journal on Mathematical Analysis, 53(1):681–709, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kong:2023:ESL

[KW23] Fanze Kong and Juncheng Wei.

Existence and stability of localized patterns in the population models with large advection and strong Allee effect. SIAM Journal on Mathematical Analysis, 55(4):2505–2552, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M148625X.

Kweon:2023:CDR

[Kwe23]  Jae Ryong Kweon.  Cav-

ity dynamics and regularity for compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 55 (5):4599–4639, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M152270X.

Kenig:2021:PSE

[KZ21a] Carlos Kenig and Jiuyi Zhu. Propagation of smallness in elliptic periodic homogenization. SIAM Journal on Mathematical Analysis, 53(1):111–132, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kreisbeck:2021:LDI

[KZ21b] Carolin Kreisbeck and Elvira Zappale. Loss of doubleintegral character during relaxation. SIAM Journal on Mathematical Analysis, 53(1): 351–385, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lafleche:2020:FFP

[Laf20] Laurent Lafleche. Fractional Fokker–Planck equation with general confinement force. SIAM Journal on Mathematical Analysis, 52(1):164–196, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2020:GNR

[Lai20] Geng Lai. Global NonIsentropic rotational supersonic flows in a semi-infinite divergent duct. SIAM Journal on Mathematical Analysis, 52(5): 5121–5154, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lear:2023:GEL

[Lea23] Daniel Lear. Global existence and limiting behavior of unidirectional flocks for the fractional Euler alignment system. SIAM Journal on Mathematical Analysis, 55(4):3731–3754, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1514039.

Lewicka:2025:MAS

[Lew25] Marta Lewicka. The Monge– Ampère system: Convex integration in arbitrary dimension and codimension. SIAM Journal on Mathematical Analysis, 57(1):601–636, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1634126.

Liu:2021:NSL

[LG21] Jinjing Liu and Zhenhua Guo.

Nonlinear stability of large amplitude viscous shock waves to the one-dimensional system of viscoelasticity. SIAM Journal on Mathematical Analysis, 53 (2):1818–1830, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2020:AEF

[Li20] Haigang Li. Asymptotics for the electric field concentration in the perfect conductivity problem. SIAM Journal on Mathematical Analysis, 52(4): 3350–3375, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2021:FPI

[Li21] Li Li. A fractional parabolic inverse problem involving a timedependent magnetic potential. SIAM Journal on Mathematical Analysis, 53(1):435–452, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2022:POO

[Li22a] Liangpan Li. On the placement of an obstacle so as to optimize the Dirichlet heat content. SIAM Journal on Mathematical Analysis, 54(3):3275–3291, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1433411.

Li:2022:LAS

[Li22b] Zongguang Li. Large amplitude solutions in L_{v}^{p}L_{T}^{∞}L_{x}^{∞} to the Boltzmann equation for soft potentials. SIAM Journal on Mathematical Analysis, 54(4):4163–4197, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443790.

Li:2023:PSS

[Li23] Shanbing Li. Positive steadystate solutions for a class of prey-predator systems with indirect prey-taxis. SIAM Journal on Mathematical Analysis, 55(6):6342–6374, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Liao:2022:LCW

[Lia22] Naian Liao. Local continuity of weak solutions to the Stefan problem involving the singular p-Laplacian. SIAM Journal on Mathematical Analysis, 54(2):2570–2586, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1402443.

Liao:2024:HES

[Lia24]  Naian Liao.  Hölder esti-

mates for the Stefan problem. SIAM Journal on Mathematical Analysis, 56(6):7989–8010, ???? 2024. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1600013.

Liu:2023:BLD

[Liu23]  Chao Liu.  Blowups and

longtime developments with near-boundary mass accretions of irregularly shaped Euler– Poisson dominated molecular clouds in astrophysics. SIAM Journal on Mathematical Analysis, 55(4):2553–2594, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469420.

Liu:2024:NPE

[Liu24] Shuang Liu. Nonmonotonicity of principal eigenvalues in diffusion rate for some non-self-adjoint operators with large advection. SIAM Journal on Mathematical Analysis, 56 (5):6025–6056, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1603947.

List:2020:RUU

[LKPR20] Florian List, Kundan Kumar, Iuliu Sorin Pop, and Florin Adrian Radu. Rigorous upscaling of unsaturated flow in fractured porous media. SIAM Journal on Mathematical Analysis, 52(1):239–276, ???? 2020. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2020:PRG

[LL20] Ru-Yu Lai and Qin Li. Parameter reconstruction for general transport equation. SIAM Journal on Mathematical Analysis, 52(3):2734–2758, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2021:GLO

[LL21] Tiejun Li and Xiaoguang Li.

Gamma-limit of the Onsager– Machlup functional on the space of curves. SIAM Journal on Mathematical Analysis, 53(1):1–31, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2022:DES

[LL22] Jing Li and Kuijie Li. The defocusing energy-supercritical nonlinear Schrödinger equation in high dimensions. SIAM Journal on Mathematical Analysis, 54(3):3253–3274, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1456443.

Long:2023:CPF

[LLhW23] Wen-Gao Long, Yu-Tian Li, and Qing hai Wang. Connection problem of the first Painlevé transcendent between poles and negative infinity.

SIAM Journal on Mathematical Analysis, 55(6):6676–6706, November 2023. CODEN SJ-MAAH. ISSN 1095-7154.

Lim:2020:QPC

[LLN20] Tau Shean Lim, Yulong Lu, and James H. Nolen. Quantitative propagation of chaos in a bimolecular chemical reaction– diffusion model. SIAM Journal on Mathematical Analysis, 52 (2):2098–2133, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2021:APE

[LLPZ21] Shuang Liu, Yuan Lou, Rui Peng, and Maolin Zhou. Asymptotics of the principal eigenvalue for a linear timeperiodic parabolic operator i: Large advection. SIAM Journal on Mathematical Analysis, 53(5):5243–5277, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2020:CPS

[LLR20] Ru-Yu Lai, Yi-Hsuan Lin, and Angkana Rüland. The Calderón problem for a spacetime fractional parabolic equation. SIAM Journal on Mathematical Analysis, 52(3):2655– 2688, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lindblad:2020:DAO

[LLS20] Hans Lindblad, Jonas Lührmann, and Avy Soffer. Decay and asymptotics for the one-dimensional Klein–Gordon equation with variable coefficient cubic nonlinearities. SIAM Journal on Mathematical Analysis, 52(6):6379–6411, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2025:UAP

[LLS25] Qianxiao Li, Ting Lin, and Zuowei Shen. On the universal approximation property of deep fully convolutional neural networks. SIAM Journal on Mathematical Analysis, 57(5):5275–5302, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1570119.

Lin:2020:RBE

[LLW20] Yu-Chu Lin, Ming-Jiea Lyu, and Kung-Chien Wu. Relativistic Boltzmann equation: Large time behavior and finite speed of propagation. SIAM Journal on Mathematical Analysis, 52(6):5994–6032, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2021:CPC

[LLW21] Sili Liu, Yong Lu, and Huanyao Wen. On the Cauchy problem for a compressible Oldroyd-B model without stress diffusion. SIAM Journal on Mathematical Analysis, 53(6):6216–6242, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2022:IES

[LLW22] Jianliang Li, Peijun Li, and Xu Wang. Inverse elastic scattering for a random potential. SIAM Journal on Mathematical Analysis, 54(5):5126–5159, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1430200.

Lin:2024:GES

[LLW24] Quyuan Lin, Rongchang Liu, and Weinan Wang. Global existence for the stochastic Boussinesq equations with transport noise and small rough data. SIAM Journal on Mathematical Analysis, 56(1): 501–528, January 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Li:2024:FHT

[LLWZ24] Menghui Li, Xiao Luo, Juncheng Wei, and Maoding Zhen. Free and harmonic trapped spin-1 Bose–Einstein condensates in R^{3}. SIAM Journal on Mathematical Analysis, 56 (4):4375–4414, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1572222.

Li:2025:GSS

[LLWZ25] Menghui Li, Xiao Luo, Juncheng Wei, and Maoding Zhen. On ground states of spin-1 Bose– Einstein condensates with Ioffe–Pritchard magnetic field in two and three dimensions. SIAM Journal on Mathematical Analysis, 57(5):5039–5079, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1674091.

Liu:2021:RGF

[LLX21] Yuning Liu, Xin Yang Lu, and Xiang Xu. Regularity of a gradient flow generated by the anisotropic Landau–de Gennes energy with a singular potential. SIAM Journal on Mathematical Analysis, 53(3):3338– 3365, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2024:PDI

[LLZ24] Ru-Yu Lai, Xuezhu Lu, and Ting Zhou. Partial data inverse problems for the nonlinear time-dependent Schrödinger equation. SIAM Journal on Mathematical Analysis, 56 (4):4712–4741, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1587993.

Lopez:2020:SCD

[LM20] Robert Lopez and Amir Moradifam. Stability of current density impedance imaging. SIAM Journal on Mathematical Analysis, 52(5):4506– 4523, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Linares:2021:LTB

[LM21] Felipe Linares and Argenis J. Mendez. On long time behavior of solutions of the Schrödinger–Korteweg–de Vries system. SIAM Journal on Mathematical Analysis, 53 (4):3838–3855, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

LePeutrec:2020:RQS

[LN20] Dorian Le Peutrec and Boris Nectoux. Repartition of the quasi-stationary distribution and first exit point density for a double-well potential. SIAM Journal on Mathematical Analysis, 52(1):581–604, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lassas:2024:DDI

[LNOY24] Matti Lassas, Medet Nursultanov, Lauri Oksanen, and Lauri Ylinen. Disjoint data inverse problem on manifolds with quantum chaos bounds. SIAM Journal on Mathematical Analysis, 56(6):7748–7779, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1606897.

Logioti:2021:PFB

[LNRV21] A. Logioti, B. Niethammer, M. Röger, and J. J. L. Velázquez. A parabolic free boundary problem arising in a model of cell polarization. SIAM Journal on Mathematical Analysis, 53(1):1214–1238, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2020:ISW

[LOWX20] Bing Li, Masahito Ohta, Yifei Wu, and Jun Xue. Instability of the solitary waves for the generalized Boussinesq equations. SIAM Journal on Mathematical Analysis, 52(4):3192– 3221, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leopold:2020:DMS

[LP20] Nikolai Leopold and Peter Pickl. Derivation of the Maxwell–Schrödinger equations from the Pauli–Fierz Hamiltonian. SIAM Journal on Mathematical Analysis, 52 (5):4900–4936, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lafortune:2022:SIP

[LP22] Stéphane Lafortune and Dmitry E.

Pelinovsky. Spectral instability of peakons in the b-family of the Camassa–Holm equations.

SIAM Journal on Mathematical Analysis, 54(4):4572–4590, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458776.

Liao:2023:EAL

[LP23] Xian Liao and Michael Plum. Eigenvalue analysis of the Lax operator for the onedimensional cubic nonlinear defocusing Schrödinger equation. SIAM Journal on Mathematical Analysis, 55(6):7399– 7425, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Li:2022:IME

[LPS22] Wei Li, Karl-Mikael Perfekt, and Stephen P. Shipman. Infinitely many embedded eigenvalues for the Neumann– Poincaré operator in 3D. SIAM Journal on Mathematical Analysis, 54(1):343–362, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1400365.

Li:2021:SLT

[LPW21] Hai-Liang Li, Benoít Perthame, and Xinmei Wen. Sharp longtime asymptotics for chemotaxis with free boundary. SIAM Journal on Mathematical Analysis, 53(2):2027–2083, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2022:ASD

[LPW22] Qingqing Liu, Hongyun Peng, and Zhi-An Wang. Asymptotic stability of diffusion waves of a quasi-linear hyperbolicparabolic model for vasculogenesis. SIAM Journal on Mathematical Analysis, 54 (1):1313–1346, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1418150.

Li:2023:GSQ

[LPZ23] Wei-Xi Li, Marius Paicu, and Ping Zhang. Gevrey solutions of quasi-linear hyperbolic hydrostatic Navier–Stokes system. SIAM Journal on Mathematical Analysis, 55(6):6194– 6228, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Linares:2021:MFE

[LR21] Felipe Linares and João P. G.

Ramos. Maximal function estimates and local wellposedness for the generalized Zakharov–Kuznetsov equation. SIAM Journal on Mathematical Analysis, 53(1):914–936, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Luo:2020:NLT

[LRW20] Senping Luo, Xiaofeng Ren, and Juncheng Wei. Nonhexagonal lattices from a two species interacting system.

SIAM Journal on Mathematical Analysis, 52(2):1903–1942, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lamberti:2020:ICO

[LS20] Pier Domenico Lamberti and Ioannis G. Stratis. On an interior Calderón operator and a related Steklov eigenproblem for Maxwell’s equations. SIAM Journal on Mathematical Analysis, 52(5):4140–4160, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurencot:2021:MTI

[LS21a] Philippe Laurençot and Christian Stinner. Mass threshold for infinite-time blowup in a chemotaxis model with split population. SIAM Journal on Mathematical Analysis, 53(3): 3385–3419, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2021:GTW

[LS21b] Yang Li and Yongzhong Sun. On global-in-time weak solutions to a two-dimensional full compressible NonResistive MHD system. SIAM Journal on Mathematical Analysis, 53 (4):4142–4177, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2023:SRJ

[LS23a] Geng Lai and Wancheng Sheng. Supersonic reacting jet flows from a threedimensional divergent conical nozzle. SIAM Journal on Mathematical Analysis, 55 (5):5260–5317, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1529099.

Leopold:2023:PML

[LS23b] Nikolai Leopold and Chiara Saffirio. Propagation of moments for large data and semiclassical limit to the relativistic Vlasov equation. SIAM Journal on Mathematical Analysis, 55(3):1676–1706, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1493616.

Li:2023:GEO

[LS23c] Hai-Liang Li and Ling-Yun Shou. Global existence and optimal time-decay rates of the compressible Navier–Stokes– Euler system. SIAM Journal on Mathematical Analysis, 55(3):1810–1846, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1497857.

Liu:2023:CNL

[LSW23] Gi-Ren Liu, Yuan-Chung Sheu, and Hau-Tieng Wu. Central and noncentral limit theorems arising from the scattering transform and its neural activation generalization. SIAM Journal on Mathematical Analysis, 55(2):1170–1213, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1454511.

Liu:2024:PDI

[LSY24] Boya Liu, Teemu Saksala, and Lili Yan. Partial data inverse problem for hyperbolic equation with time-dependent damping coefficient and potential. SIAM Journal on Mathematical Analysis, 56 (4):5678–5722, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1588676.

Lu:2021:DNA

[LSYZ21] Jianfeng Lu, Zuowei Shen, Haizhao Yang, and Shijun Zhang. Deep network approximation for smooth functions. SIAM Journal on Mathematical Analysis, 53(5):5465–5506, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lauriere:2022:CLP

[LT22] Mathieu Laurière and Ludovic Tangpi. Convergence of large population games to mean field games with interaction through the controls. SIAM Journal on Mathematical Analysis, 54(3):3535–3574, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469328.

Laurencot:2024:NPF

[LT24] Philippe Laurençot and Ariane Trescases. No pattern formation in a quasilinear chemotaxis model with local sensing. SIAM Journal on Mathematical Analysis, 56(5):6861–6884, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1654907.

Lai:2021:RCK

[LUY21] Ru-Yu Lai, Gunther Uhlmann, and Yang Yang. Reconstruction of the collision kernel in the nonlinear Boltzmann equation. SIAM Journal on Mathematical Analysis, 53(1):1049– 1069, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2020:RJU

[LW20a] Sirui Li and Wei Wang. Rigorous justification of the uniaxial limit from the Qian– Sheng inertial Q-tensor theory to the Ericksen–Leslie theory. SIAM Journal on Mathematical Analysis, 52(5):4421–4468, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2020:MTW

[LW20b] Yong Liu and Juncheng Wei. Multivortex traveling waves for the Gross–Pitaevskii equation and the Adler–Moser polynomials. SIAM Journal on Mathematical Analysis, 52(4):3546– 3579, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2021:OIT

[LW21] Ze Li and Ming Wang. Observability inequality at two time points for KdV equations. SIAM Journal on Mathematical Analysis, 53(2):1944–1957, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liang:2022:GMF

[LW22] Rui Liang and Yuzhao Wang. Gibbs measure for the focusing fractional NLS on the torus. SIAM Journal on Mathematical Analysis, 54(6):6096–6118, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1445946.

Li:2023:GEP

[LW23a] Fucai Li and Yichun Wang. Global Euler–Poisson limit to the Vlasov–Poisson–Boltzmann system with soft potential.

SIAM Journal on Mathematical Analysis, 55(4):2877–2916, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1474254.

Luo:2023:WPL

[LW23b] Dejun Luo and Danli Wang.

Well posedness and limit theorems for a class of stochastic dyadic models. SIAM Journal on Mathematical Analysis, 55(2):1464–1498, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1511497.

Li:2024:ISB

[LW24] Peijun Li and Xu Wang. Inverse scattering for the biharmonic wave equation with a random potential. SIAM Journal on Mathematical Analysis, 56(2):1959–1995, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Liu:2022:NES

[LWC22] Yarong Liu, Yejuan Wang, and Tomas Caraballo. Nontrivial equilibrium solutions and general stability for stochastic evolution equations with pantograph delay and tempered fractional noise. SIAM Journal on Mathematical Analysis, 54(5):5629–5661, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1496876.

Lu:2020:DOD

[LWW20] Jianfeng Lu, Alexander B. Watson, and Michael I. Weinstein. Dirac operators and domain walls. SIAM Journal on Mathematical Analysis, 52(2): 1115–1145, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2024:PEW

[LWX24] Yun Li, Fuke Wu, and Longjie Xie. Poisson equation on Wasserstein space and diffusion approximations for multiscale McKean–Vlasov equation. SIAM Journal on Mathematical Analysis, 56(2):1495– 1524, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Lai:2021:SPA

[LWZ21a] Suhua Lai, Jiahong Wu, and Jianwen Zhang. Stabilizing phenomenon for 2D anisotropic magnetohydrodynamic system near a background magnetic field. SIAM Journal on Mathematical Analysis, 53(5):6073– 6093, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2021:WPF

[LWZ21b] Hui Li, Wei Wang, and Zhifei Zhang. Well-posedness of the free boundary problem in elastodynamics with mixed stability condition. SIAM Journal on Mathematical Analysis, 53 (5):5405–5435, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2023:GSI

[LWZ23] Hongxia Lin, Jiahong Wu, and Yi Zhu. Global solutions to 3D incompressible MHD system with dissipation in only one direction. SIAM Journal on Mathematical Analysis, 55(5):4570–4598, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1471274.

Liu:2022:RJF

[LWZZ22] Jian-Guo Liu, Ziheng Wang, Yuan Zhang, and Zhennan Zhou. Rigorous justification of the Fokker–Planck equations of neural networks based on an iteration perspective. SIAM Journal on Mathematical Analysis, 54(1):1270–1312, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1338368.

Liu:2021:EIL

[LX21] Jian-Guo Liu and Xiangsheng Xu. Existence and incompressible limit of a tissue growth model with autophagy. SIAM Journal on Mathematical Analysis, 53(5):5215–5242, ???? 2021. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2023:ESA

[LX23] Yanlin Liu and Li Xu. On the existence and structures of almost axisymmetric solutions to 3-D Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 55(1):458–485, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1469912.

Li:2024:IUE

[LX24] Jinkai Li and Zhouping Xin. Instantaneous unboundedness of the entropy and uniform positivity of the temperature for the compressible Navier– Stokes equations with fast decay density. SIAM Journal on Mathematical Analysis, 56(3): 3004–3041, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Li:2025:SDE

[LX25] Haigang Li and Longjuan Xu. Second derivative estimates and gradient asymptotics for closely located inclusions in linear elasticity. SIAM Journal on Mathematical Analysis, 57(1):753–788, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1656888.

Li:2023:SBA

[LXZ23] Haigang Li, Longjuan Xu, and Peihao Zhang. Stress blowup analysis when a suspending rigid particle approaches the boundary in Stokes flow: 2dimensional case. SIAM Journal on Mathematical Analysis, 55(5):4493–4536, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1497407.

Luo:2025:QES

[LXZ25] Dejun Luo, Bin Xie, and Guohuan Zhao. Quantitative estimates for SPDEs on the full space with transport noise and L^{p}-initial data. SIAM Journal on Mathematical Analysis, 57(5):4795–4821, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1706815.

Li:2021:WPM

[LY21] Wei-Xi Li and Tong Yang.

Well-posedness of the MHD boundary layer system in Gevrey function space without structural assumption. SIAM Journal on Mathematical Analysis, 53(3):3236–3264, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2022:HSP

[LY22] Genghong Lin and Jianshe Yu.

Homoclinic solutions of periodic discrete Schrödinger equations with local superquadratic conditions. SIAM Journal on Mathematical Analysis, 54 (2):1966–2005, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413201.

Liu:2023:IBE

[LY23] Xin Liu and Yuan Yuan. Immediate blowup of entropybounded classical solutions to the vacuum free boundary problem of nonisentropic compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 55 (3):1524–1544, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1493732.

Li:2024:LMG

[LY24] Qinfeng Li and Ruofei Yao. On location of maximum of gradient of torsion function. SIAM Journal on Mathematical Analysis, 56(4):5372–5385, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1609452.

Liu:2023:VPE

[LYZ23] Cheng-Jie Liu, Tong Yang, and Zhu Zhang. Validity of Prandtl expansions for steady MHD in the Sobolev framework. SIAM Journal on Mathematical Analysis, 55(3):2377–2410, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1507139.

Li:2024:SRS

[LYZ24] Peijun Li, Xiaohua Yao, and Yue Zhao. The scattering resonances for schrödinger-type operators with unbounded potentials. SIAM Journal on Mathematical Analysis, 56(2):2149– 2170, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Li:2024:UFF

[LYZZ24] Long Li, Jiansheng Yang, Bo Zhang, and Haiwen Zhang. Uniform far-field asymptotics of the two-layered Green function in two dimensions and application to wave scattering in a two-layered medium. SIAM Journal on Mathematical Analysis, 56(3):4143–4184, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1525910.

Li:2020:BBA

[LZ20] Haigang Li and Zhiwen Zhao.

Boundary blow-up analysis of gradient estimates for Lamé systems in the presence of m-convex hard inclusions. SIAM Journal on Mathematical Analysis, 52(4):3777–3817, ???? 2020. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Luo:2021:PEI

[LZ21] Chenyun Luo and Junyan Zhang. A priori estimates for the incompressible freeboundary magnetohydrodynamics equations with surface tension. SIAM Journal on Mathematical Analysis, 53(2): 2595–2630, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lamy:2022:SRS

[LZ22] Xavier Lamy and Andres Zuniga. On the stability of radial solutions to an anisotropic Ginzburg–Landau equation. SIAM Journal on Mathematical Analysis, 54(1):723–736, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1433939.

Li:2023:SEC

[LZ23a] Fucai Li and Zhipeng Zhang. Stabilizing effect of capillarity in the Rayleigh–Taylor problem to the viscous incompressible capillary fluids. SIAM Journal on Mathematical Analysis, 55(4):3287–3315, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1456327.

Liard:2023:ANS

[LZ23b] Thibault Liard and Enrique Zuazua. Analysis and numerical solvability of backward– forward conservation laws. SIAM Journal on Mathematical Analysis, 55(3):1949–1968, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1478720.

Liu:2024:OTV

[LZ24a] Yanlin Liu and Ping Zhang.

On the one time-varying component regularity criteria for 3-D Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 56(5):6213–6231, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1623124.

Luo:2024:GBK

[LZ24b] Chenyun Luo and Kai Zhou.

A generalized Beale–Kato– Majda breakdown criterion for the free-boundary problem in Euler equations with surface tension. SIAM Journal on Mathematical Analysis, 56(1): 374–411, January 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Liao:2021:ODV

[LZZ21] Yongkai Liao, Huijiang Zhao, and Jiawei Zhou. Onedimensional viscous and heatconducting ionized gas with density-dependent viscosity. SIAM Journal on Mathematical Analysis, 53(5):5580–5612, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ma:2022:LSI

[Ma22] Xiao Ma. Linear stability of the 2D irrotational circulation flow around an elliptical cylinder. SIAM Journal on Mathematical Analysis, 54(5):5576–5587, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M139671X.

Martin:2022:TDP

[Mar22] Donati Martin. Two-dimensional point vortex dynamics in bounded domains: Global existence for almost every initial data. SIAM Journal on Mathematical Analysis, 54(1):79–113, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413213.

Minarcik:2020:LTB

[MB20] Jirí Minarcík and Michal Benes. Long-term behavior of curve shortening flow in R^{3}. SIAM Journal on Mathematical Analysis, 52(2):1221–1231, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meftahi:2021:ULS

[Mef21] Houcine Meftahi. Uniqueness, Lipschitz stability, and reconstruction for the inverse optical tomography problem. SIAM Journal on Mathematical Analysis, 53(6):6326–6354, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Mef24].

Meftahi:2024:EUL

[Mef24] Houcine Meftahi. Erratum:

Uniqueness, Lipschitz stability, and reconstruction for the inverse optical tomography problem. SIAM Journal on Mathematical Analysis, 56(6):8131–8135, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1605260. See [Mef21].

Menard:2024:MFL

[Mén24] Matthieu Ménard. Mean-field limit derivation of a monokinetic spray model with gyroscopic effects. SIAM Journal on Mathematical Analysis, 56 (1):1068–1113, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Mizoguchi:2020:CIE

[Miz20] Noriko Mizoguchi. Criterion on initial energy for finitetime blowup in parabolicparabolic Keller–Segel system. SIAM Journal on Mathematical Analysis, 52(6):5840–5864, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mariano:2023:CNS

[MM23] Paolo Maria Mariano and Domenico Mucci. Crack nucleation in shells with throughthe-thickness microstructure. SIAM Journal on Mathematical Analysis, 55(5):4977–4997, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1490004.

Meszaros:2024:MFG

[MM24] Alpár R. Mészáros and Chenchen Mou. Mean field games systems under displacement monotonicity. SIAM Journal on Mathematical Analysis, 56(1):529–553, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Matano:2023:SFS

[MMN23] Hiroshi Matano, Yoichiro Mori, and Mitsunori Nara. Stability of front solutions of the bidomain Allen–Cahn equation on an infinite strip. SIAM Journal on Mathematical Analysis, 55(3):1545–1595, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1418095.

Mateu:2023:SEE

[MMR^{+}23] Joan Mateu, Maria Giovanna Mora, Luca Rondi, Lucia Scardia, and Joan Verdera. Stability of ellipsoids as the energy minimizers of perturbed Coulomb energies. SIAM Journal on Mathematical Analysis, 55(4):3650–3676, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1479695.

Menovschikov:2021:EVT

[MMS21] Alexander Menovschikov, Anastasia Molchanova, and Luca Scarpa. An extended variational theory for nonlinear evolution equations via modular spaces. SIAM Journal on Mathematical Analysis, 53(4): 4865–4907, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Montoro:2025:CAW

[MMS25] Luigi Montoro, Luigi Muglia, and Berardino Sciunzi. The classification of all weak solutions to −Δu = u^{−γ} in the half space. SIAM Journal on Mathematical Analysis, 57(5):5080–5088, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M167189X.

Ma:2021:SDW

[MMSH21] To Fu Ma, Jaqueline G.

Mesquita, and Paulo N. Seminario-Huertas. Smooth dynamics of weakly damped Lamé systems with delay. SIAM Journal on Mathematical Analysis, 53(4):3759–3771, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Matsumura:2024:ABS

[MN24a] Akitaka Matsumura and Kenji Nishihara. Asymptotic behavior of solutions to the Cauchy problem for 1D psystem with space-dependent damping. SIAM Journal on Mathematical Analysis, 56(1): 993–1015, January 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Milne:2024:OTA

[MN24b] Tristan Milne and Adrian Nachman. An optimal transport analogue of the Rudin– Osher–Fatemi model and its corresponding multiscale theory. SIAM Journal on Mathematical Analysis, 56(1):1114– 1148, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Monard:2020:FRS

[Mon20]  François Monard.  Func-

tional relations, sharp mapping properties, and regularization of the X-ray transform on disks of constant curvature. SIAM Journal on Mathematical Analysis, 52(6):5675–5702, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moussa:2020:NCS

[Mou20] Ayman Moussa. From nonlocal to classical Shigesada– Kawasaki–Teramoto systems: Triangular case with bounded coefficients. SIAM Journal on Mathematical Analysis, 52(1): 42–64, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maheux:2020:KSS

[MP20a] Patrick Maheux and Vittoria Pierfelice. The Keller– Segel system on the twodimensional-hyperbolic space. SIAM Journal on Mathematical Analysis, 52(5):5036–5065, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Migorski:2020:AHD

[MP20b] Stanisław Migórski and Dariusz Pa̧czka. Almost historydependent variational-hemivariational inequality for frictional contact problems. SIAM Journal on Mathematical Analysis, 52(5): 4362–4390, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Madiyeva:2021:GPP

[MP21a] Aigerim Madiyeva and Dmitry E.

Pelinovsky. Growth of perturbations to the peaked periodic waves in the Camassa–Holm equation. SIAM Journal on Mathematical Analysis, 53(3): 3016–3039, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Montalto:2021:LSE

[MP21b] Riccardo Montalto and Michela Procesi. Linear Schrödinger equation with an almost periodic potential. SIAM Journal on Mathematical Analysis, 53 (1):386–434, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maione:2022:CFD

[MPC22] Alberto Maione, Andrea Pinamonti, and Francesco Serra Cassano. Γ-convergence for functionals depending on vector fields. II. Convergence of minimizers. SIAM Journal on Mathematical Analysis, 54 (6):5761–5791, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1432466.

Ma:2023:FAI

[MPMS23] Shiqi Ma, Leyter Potenciano-Machado, and Mikko Salo. Fixed angle inverse scattering for sound speeds close to constant. SIAM Journal on Mathematical Analysis, 55 (4):3420–3456, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M147640X.

Munoz:2021:LTB

[MPS21] Claudio Muñoz, Gustavo Ponce, and Jean-Claude Saut. On the long time behavior of solutions to the intermediate long wave equation. SIAM Journal on Mathematical Analysis, 53(1):1029–1048, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mazzoleni:2023:SAO

[MPV23] Dario Mazzoleni, Benedetta Pellacci, and Gianmaria Verzini. Singular analysis of the optimizers of the principal eigenvalue in indefinite weighted Neumann problems. SIAM Journal on Mathematical Analysis, 55(4):4162–4192, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1490600.

Miao:2022:QSR

[MPZ22] Qianyun Miao, Fa Peng, and Yuan Zhou. A quantitative Sobolev regularity for absolute minimizers involving Hamiltonian H(p) in plane. SIAM Journal on Mathematical Analysis, 54(6):5792–5853, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443601.

Mooney:2024:SRO

[MR24] Connor Mooney and Arghya Rakshit. Sobolev regularity for optimal transport maps of nonconvex planar domains. SIAM Journal on Mathematical Analysis, 56(4):4742–4758, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1582436.

Mazari:2022:QSE

[MRB22] Idriss Mazari and Domènec Ruiz-Balet. Quantitative stability for eigenvalues of Schrödinger operator, quantitative bathtub principle, and application to the turnpike property for a bilinear optimal control problem. SIAM Journal on Mathematical Analysis, 54(3):3848–3883, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1393121.

Maity:2020:MTE

[MRR20] Debayan Maity, Jean-Pierre Raymond, and Arnab Roy. Maximal-in-time existence and uniqueness of strong solution of a 3D fluid-structure interaction model. SIAM Journal on Mathematical Analysis, 52(6): 6338–6378, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marcati:2020:ARI

[MS20a] Carlo Marcati and Christoph Schwab. Analytic regularity for the incompressible Navier– Stokes equations in polygons. SIAM Journal on Mathematical Analysis, 52(3):2945–2968, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mizumachi:2020:SBL

[MS20b] Tetsu Mizumachi and Yusuke Shimabukuro. Stability of Benney–Luke line solitary waves in 2 dimensions. SIAM Journal on Mathematical Analysis, 52(5):4238–4283, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morandotti:2020:MFA

[MS20c] Marco Morandotti and Francesco Solombrino. Mean-field analysis of multipopulation dynamics with label switching. SIAM Journal on Mathematical Analysis, 52(2):1427–1462, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Murata:2020:GWP

[MS20d] Miho Murata and Yoshihiro Shibata. The global wellposedness for the compressible fluid model of Korteweg type. SIAM Journal on Mathematical Analysis, 52(6):6313–6337, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mishra:2021:IRD

[MS21] Rohit Kumar Mishra and Suman Kumar Sahoo. Injectivity and range description of integral moment transforms over m-tensor fields in R^{n}. SIAM Journal on Mathematical Analysis, 53(1):253–278, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mindrila:2022:TPW

[MS22] Claudiu Mîndrila and Sebastian Schwarzacher. Timeperiodic weak solutions for an incompressible Newtonian fluid interacting with an elastic plate. SIAM Journal on Mathematical Analysis, 54 (4):4139–4162, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458946.

Monard:2023:SXR

[MS23a] François Monard and Plamen Stefanov. Sampling the X-ray transform on simple surfaces. SIAM Journal on Mathematical Analysis, 55(3):1707–1736, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1475272.

Mukherjee:2023:ACU

[MS23b] Arpan Mukherjee and Mourad Sini. Acoustic cavitation using resonating MicroBubbles:

Analysis in the time-domain. SIAM Journal on Mathematical Analysis, 55(5):5575–5616, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1533396.

Maehlen:2024:ATW

[MS24a] Ola Mæhlen and Douglas Svensson Seth. Asymmetric traveling wave solutions of the capillary-gravity Whitham equation. SIAM Journal on Mathematical Analysis, 56 (6):8096–8124, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1632590.

Mielke:2024:CSS

[MS24b] Alexander Mielke and Stefanie Schindler. Convergence to selfsimilar profiles in reaction– diffusion systems. SIAM Journal on Mathematical Analysis, 56(6):7108–7135, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1564298.

Mimikos-Stamatopoulos:2024:WRS

[MS24c] Nikiforos Mimikos-Stamatopoulos.

Weak and renormalized solutions to a hypoelliptic mean field games system. SIAM Journal on Mathematical Analysis, 56(2):2312–2356, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Modena:2024:LNS

[MS24d] Stefano Modena and Andre Schenke. Local nonuniqueness for stochastic transport equations with deterministic drift. SIAM Journal on Mathematical Analysis, 56(4):5209–5261, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1589104.

Mimikos-Stamatopoulos:2024:RLT

[MSM24] Nikiforos Mimikos-Stamatopoulos and Sebastian Munoz. Regularity and long time behavior of one-dimensional first-order mean field games and the planning problem. SIAM Journal on Mathematical Analysis, 56 (1):43–78, January 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Mauritz:2024:BMD

[MSW24a] Marco Mauritz, Bernhard Schmitzer, and Benedikt Wirth. A Bayesian model for dynamic mass reconstruction from PET listmode data. SIAM Journal on Mathematical Analysis, 56(5):5840–5880, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M161923X.

Moll:2024:LTB

[MSW24b] Salvador Moll, Ken Shirakawa, and Hiroshi Watanabe. Largetime behavior for a phase-field system of 3D-grain boundary motion. SIAM Journal on Mathematical Analysis, 56 (5):6885–6914, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1571988.

Monobe:2023:SES

[MSY23] Harunori Monobe, Masahiko Shimojo, and Eiji Yanagida. Spreading and extinction of solutions to the logarithmic diffusion equation with a logistic reaction. SIAM Journal on Mathematical Analysis, 55(3):2261–2287, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1492544.

Mazon:2024:TRR

[MT24] José M. Mazón and Julián Toledo. Torsional rigidity in random walk spaces. SIAM Journal on Mathematical Analysis, 56(2):1604–1642, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Marion:2023:EPM

[MV23] Martine Marion and Vitaly Volpert. Existence of pulses for monotone reaction–diffusion systems. SIAM Journal on Mathematical Analysis, 55 (2):603–627, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1453219.

Ming:2020:WWP

[MW20] Mei Ming and Chao Wang. Water waves problem with surface tension in a corner domain i: a priori estimates with constrained contact angle. SIAM Journal on Mathematical Analysis, 52(5):4861–4899, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mellet:2023:CSN

[MW23] Antoine Mellet and Yijing Wu. Γ-convergence of some nonlocal perimeters in bounded domains with general boundary conditions. SIAM Journal on Mathematical Analysis, 55(4):3226–3261, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1483505.

Ma:2023:LMN

[MWX23] Lei Ma, Tian-Yi Wang, and Chunjing Xie. Low Mach number limit and far field convergence rates of irrotational flows in multidimensional nozzles with an obstacle inside. SIAM Journal on Mathematical Analysis, 55(1):36–67, ????

2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1472553.

Nemes:2020:LPA

[ND20] Gergö Nemes and Adri B. Olde Daalhuis. Large-parameter asymptotic expansions for the Legendre and Allied functions. SIAM Journal on Mathematical Analysis, 52(1):437–470, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Neiss:2020:SRP

[Nei20] Robert Axel Neiss. Symmetry reduction and periodic solutions in Hamiltonian Vlasov systems. SIAM Journal on Mathematical Analysis, 52(2): 1844–1863, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nemes:2021:PTC

[Nem21] Gergö Nemes. Proofs of two conjectures on the real zeros of the cylinder and Airy functions. SIAM Journal on Mathematical Analysis, 53(4):4328– 4349, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2020:ILN

[Ngu20] Trinh T. Nguyen. The inviscid limit of Navier–Stokes with critical Navier-slip boundary conditions for analytic data.

SIAM Journal on Mathematical Analysis, 52(5):4616–4637, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2022:CPP

[Ngu22] Hoai-Minh Nguyen. Cloaking property of a plasmonic structure in doubly complementary media and three-sphere inequalities with partial data. SIAM Journal on Mathematical Analysis, 54(2):2040–2096, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1394722.

Neamtu:2021:RCM

[NK21] Alexandra Neamtu and Christian Kuehn. Rough center manifolds. SIAM Journal on Mathematical Analysis, 53(4): 3912–3957, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Neves:2023:FBT

[NO23] Wladimir Neves and Dionicio Orlando. On fractional Benney type systems. SIAM Journal on Mathematical Analysis, 55(6):7296–7327, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Novack:2020:NWS

[Nov20] Matthew Novack. Nonuniqueness of weak solutions to the 3 dimensional quasi-geostrophic equations. SIAM Journal on Mathematical Analysis, 52(4): 3301–3349, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novaga:2020:SOG

[NP20] Matteo Novaga and Paola Pozzi. A second order gradient flow of p-elastic planar networks. SIAM Journal on Mathematical Analysis, 52(1): 682–708, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Natalini:2023:MFL

[NP23] Roberto Natalini and Thierry Paul. The mean-field limit for hybrid models of collective motions with chemotaxis. SIAM Journal on Mathematical Analysis, 55(2):900–928, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1488004.

Nguyen:2024:ODB

[NR24] Dinh-Thi Nguyen and Julien Ricaud. On one-dimensional Bose gases with two-body and (critical) attractive three-body interactions. SIAM Journal on Mathematical Analysis, 56(3): 3203–3251, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Niethammer:2020:LVE

[NS20a] Barbara Niethammer and Richard Schubert. A local version of Einstein’s formula for the effective viscosity of suspensions. SIAM Journal on Mathematical Analysis, 52(3): 2561–2591, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Noakes:2020:LRS

[NS20b] Lyle Noakes and Luchezar Stoyanov. Lens rigidity in scattering by unions of strictly convex bodies in R^{2}. SIAM Journal on Mathematical Analysis, 52(1):471–480, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Negri:2021:EEI

[NS21a] Matteo Negri and Riccardo Scala. Existence, energy identity, and higher time regularity of solutions to a dynamic viscoelastic cohesive interface model. SIAM Journal on Mathematical Analysis, 53 (5):5682–5730, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novaes:2021:HOA

[NS21b] Douglas D. Novaes and Francisco B. Silva. Higher order analysis on the existence of periodic solutions in continuous differential equations via degree theory. SIAM Journal on Mathematical Analysis, 53(2): 2476–2490, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nik:2025:OPB

[NT25] Katerina Nik and Keisuke Takasao. On an obstacle problem for the Brakke flow with a generalized right-angle boundary condition. SIAM Journal on Mathematical Analysis, 57(1):452–494, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1652180.

Nie:2024:FBP

[NWG24] Yuanyuan Nie, Chunpeng Wang, and Guanming Gai. A free boundary problem in an unbounded domain and subsonic jet flows from divergent nozzles. SIAM Journal on Mathematical Analysis, 56(5):6337–6360, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M162301X.

Nguyen:2024:SOF

[NY24] Nhu N. Nguyen and George Yin. Second-order fast-slow stochastic systems. SIAM Journal on Mathematical Analysis, 56(4):5175–5208, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1567382.

Nersesyan:2024:EMW

[NZ24] Vahagn Nersesyan and Meng Zhao. Exponential mixing for the white-forced complex Ginzburg–Landau equation in the whole space. SIAM Journal on Mathematical Analysis, 56(3):3646–3678, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1597150.

Oliynyk:2021:FGS

[Oli21] Todd A. Oliynyk. Future global stability for relativistic perfect fluids with linear equations of state p = Kρ where 1/3 < K < 1/2. SIAM Journal on Mathematical Analysis, 53 (4):4118–4141, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Osterbrink:2020:LFA

[OP20] Frank Osterbrink and Dirk Pauly. Low frequency asymptotics and Electro-magneto-statics for time-harmonic Maxwell’s equations in exterior weak Lipschitz domains with mixed boundary conditions. SIAM Journal on Mathematical Analysis, 52(5):4971– 5000, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Osborne:2025:RSS

[OS25] Yohance A. P. Osborne and Iain Smears. Regularization of stationary second-order mean field game partial differential inclusions. SIAM Journal on Mathematical Analysis, 57(5):5189–5215, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1686401.

Ou:2021:LML

[Ou21] Yaobin Ou. Low Mach and low Froude number limit for vacuum free boundary problem of all-time classical solutions of one-dimensional compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 53(3):3265–3305, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ou:2022:ILI

[OY22] Yaobin Ou and Lu Yang.

Incompressible limit of isentropic Navier–Stokes equations with ill-prepared data in bounded domains. SIAM Journal on Mathematical Analysis, 54(3):2948–2989, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1380491.

Panov:2020:DES

[Pan20] Evgeniy Y. Panov. On decay of entropy solutions to multidimensional conservation laws. SIAM Journal on Mathematical Analysis, 52(2):1310–1317, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pankavich:2021:ELT

[Pan21]  Stephen Pankavich.  Ex-

act large time behavior of spherically symmetric plasmas. SIAM Journal on Mathematical Analysis, 53(4):4474–4512, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pankavich:2023:SAB

[Pan23] Stephen Pankavich. Scattering and asymptotic behavior of solutions to the Vlasov–Poisson system in high dimension. SIAM Journal on Mathematical Analysis, 55(5):4727–4750, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1520013.

Pan:2024:GAS

[Pan24] Xinghong Pan. Globally analytical solutions of the compressible Oldroyd-B model without retardation. SIAM Journal on Mathematical Analysis, 56(4):4854–4869, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1588974.

Pecher:2020:LRW

[Pec20] Hartmut Pecher. Low regularity well-posedness for the Yang–Mills system in Fourier– Lebesgue spaces. SIAM Journal on Mathematical Analysis, 52(4):3131–3148, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:2023:CSM

[Pen23] Shaolong Peng. Classification of solutions to mixed order elliptic system with general nonlinearity. SIAM Journal on Mathematical Analysis, 55(4):2774–2812, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1510510.

Perepelitsa:2020:CSH

[Per20] Misha Perepelitsa. Classical solutions of hyperbolic systems of PDEs with kinetic structure. SIAM Journal on Mathematical Analysis, 52(2):1864–1883, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinaud:2021:ITM

[Pin21] Olivier Pinaud. Instantaneous time mirrors and wave equations with time-singular coefficients. SIAM Journal on Mathematical Analysis, 53(4):4401– 4416, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:2023:GQN

[PL23] Yue-Jun Peng and Cunming Liu. Global quasi-neutral limit for a two-fluid Euler– Poisson system in several space dimensions. SIAM Journal on Mathematical Analysis, 55 (2):1405–1438, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1501465.

Pelinovsky:2024:SBS

[PP24] Dmitry E. Pelinovsky and Michael Plum. Stability of black solitons in optical systems with intensity-dependent dispersion. SIAM Journal on Mathematical Analysis, 56(2): 2521–2568, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Perasso:2020:ABA

[PR20] Antoine Perasso and Quentin Richard. Asymptotic behavior of age-structured and delayed Lotka–Volterra models. SIAM Journal on Mathematical Analysis, 52(5):4284–4313, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Plakhov:2025:POC

[PR25] Alexander Plakhov and Vera Roshchina. The problem of optimal camouflaging. SIAM Journal on Mathematical Analysis, 57(1):95–117, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1636976.

Protasov:2020:SDT

[Pro20] Vladimir Yu. Protasov. Surface dimension, tiles, and synchronizing automata. SIAM Journal on Mathematical Analysis, 52(4):3463–3486, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Prunier:2024:FTA

[Pru24] Raphaël Prunier. Fugledetype arguments for isoperimetric problems and applications to stability among convex shapes. SIAM Journal on Mathematical Analysis, 56(2): 1560–1603, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Pierotti:2022:GSN

[PS22] Dario Pierotti and Nicola Soave. Ground states for the NLS equation with combined nonlinearities on noncompact metric graphs. SIAM Journal on Mathematical Analysis, 54(1):768–790, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1377837.

Pitcho:2023:AEN

[PS23a] Jules Pitcho and Massimo Sorella. Almost everywhere nonuniqueness of integral curves for divergencefree Sobolev vector fields.

SIAM Journal on Mathematical Analysis, 55(5):4640–4663, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487187.

Pozza:2023:LOF

[PS23b] Marco Pozza and Antonio Siconolfi. Lax-oleinik formula on networks. SIAM Journal on Mathematical Analysis, 55(3):2211–2237, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448677.

Pistoia:2023:ESS

[PST23] Angela Pistoia, Alberto Saldaña, and Hugo Tavares. Existence of solutions to a slightly supercritical pure Neumann problem. SIAM Journal on Mathematical Analysis, 55 (4):3844–3887, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1520360.

Peng:2025:SPR

[PSW25] Rui Peng, Rachidi Salako, and Yixiang Wu. Spatial profiles of a reaction–diffusion epidemic model with nonlinear incidence mechanism and constant total population. SIAM Journal on Mathematical Analysis, 57(2):1586–1620, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1662412.

Petrova:2025:PTT

[PTE25] Yulia Petrova, Sergey Tikhomirov, and Yalchin Efendiev. Propagating terrace in a two-tubes model of gravitational fingering. SIAM Journal on Mathematical Analysis, 57(1):30–64, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1632668.

Parhi:2024:DEI

[PU24] Rahul Parhi and Michael Unser. Distributional extension and invertibility of the kplane transform and its dual. SIAM Journal on Mathematical Analysis, 56(4):4662–4686, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1556721.

Pass:2021:MMO

[PVJ21] Brendan W. Pass and Adolfo Vargas-Jiménez. Multi-marginal optimal transportation problem for cyclic costs. SIAM Journal on Mathematical Analysis, 53(4):4386–4400, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pierotti:2025:NSS

[PVY25] Dario Pierotti, Gianmaria Verzini, and Junwei Yu. Normalized solutions for Sobolev critical Schrödinger equations on bounded domains. SIAM Journal on Mathematical Analysis, 57(1):262–285, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1656281.

Peng:2021:GBL

[PW21] Rui Peng and Yixiang Wu. Global L^{∞}-bounds and longtime behavior of a diffusive epidemic system in a heterogeneous environment. SIAM Journal on Mathematical Analysis, 53(3):2776–2810, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pomponio:2025:NSF

[PW25] Alessio Pomponio and Tatsuya Watanabe. Nonlinear scalar field equation with point interaction. SIAM Journal on Mathematical Analysis, 57 (5):5243–5274, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1700892.

Peng:2024:GST

[PWZZ24] Wenshuai Peng, Zhengping Wang, Xiaoyu Zeng, and Huan-Song Zhou. Ground states of two-component Bose– Einstein condensates with Josephson junction: The endpoint cases. SIAM Journal on Mathematical Analysis, 56 (6):6963–6994, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1613323.

Pei:2022:GSJ

[PX22] Yu Pei and Wei Xiang. Global subsonic jet with strong transonic shock over a convex cornered wedge for the twodimensional steady full Euler equations. SIAM Journal on Mathematical Analysis, 54(4):4407–4451, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1405320.

Paicu:2024:GWP

[PYZ24] Marius Paicu, Tianyuan Yu, and Ning Zhu. Global wellposedness and vanishing normal stress coefficients for the hydrostatic second-grade fluid equations. SIAM Journal on Mathematical Analysis, 56(3): 3252–3274, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Peng:2023:CPD

[PZ23] Hongyun Peng and Xiaoping Zhai. The Cauchy problem for the N-dimensional compressible Navier–Stokes equations without heat conductivity.

SIAM Journal on Mathematical Analysis, 55(2):1439–1463, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1504998.

Quinn:2024:AIT

[QB24] Solomon Quinn and Guillaume Bal. Approximations of interface topological invariants. SIAM Journal on Mathematical Analysis, 56(4):5521–5582, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1568387.

Qian:2024:EED

[QRW24] Zhongmin Qian, Panpan Ren, and Feng-Yu Wang. Entropy estimate for degenerate SDEs with applications to nonlinear kinetic Fokker– Planck equations. SIAM Journal on Mathematical Analysis, 56(4):5330–5349, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1634473.

Qiu:2025:TWP

[QWYY25] Lingyun Qiu, Zhongjing Wang, Hui Yu, and Shenwen Yu. Traceability of water pollution: an inversion scheme via dynamic complex geometrical optics solutions. SIAM Journal on Mathematical Analysis, 57(1):286–305, ???? 2025.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1619605.

Qin:2025:EMS

[QX25] Zengyun Qin and Xingfei Xiang. Existence of Meissner solutions to the inhomogeneous Ginzburg–Landau system in three dimensions. SIAM Journal on Mathematical Analysis, 57(2):2037–2070, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1643852.

Qi:2022:ENP

[QZ22] Shijie Qi and Wenming Zou.

Exact number of positive solutions for the Kirchhoff equation. SIAM Journal on Mathematical Analysis, 54 (5):5424–5446, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1445879.

Rankin:2023:SGC

[Ran23]  Cale Rankin.  Strict g-

convexity for generated Jacobian equations with applications to global regularity. SIAM Journal on Mathematical Analysis, 55(5):5685–5714, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1518852.

Rege:2021:VPS

[Reg21] Alexandre Rege. The Vlasov– Poisson system with a uniform magnetic field: Propagation of moments and regularity. SIAM Journal on Mathematical Analysis, 53(2):2452–2475, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ruiz-Herrera:2020:DRD

[RH20] Alfonso Ruiz-Herrera. Delay reaction–diffusion systems via discrete dynamics. SIAM Journal on Mathematical Analysis, 52(6):6297–6312, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ricciardi:2023:CPM

[Ric23] Michele Ricciardi. The convergence problem in mean field games with Neumann boundary conditions. SIAM Journal on Mathematical Analysis, 55(4):3316–3343, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1479075.

Rosenzweig:2020:JPV

[Ros20] Matthew Rosenzweig. Justification of the point vortex approximation for modified surface quasi-geostrophic equations. SIAM Journal on Mathematical Analysis, 52(2): 1690–1728, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rosler:2021:SVC

[Rös21]  Frank Rösler.  A strange

vertex condition coming from nowhere. SIAM Journal on Mathematical Analysis, 53(3): 3098–3122, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2020:VSP

[RR20] Zhenjie Ren and Mauro Rosestolato. Viscosity solutions of path-dependent PDEs with randomized time. SIAM Journal on Mathematical Analysis, 52(2):1943–1979, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Riva:2025:RAS

[RR25] Filippo Riva and Elisabetta Rocca. A rigorous approach to the sharp interface limit for phase-field models of tumor growth. SIAM Journal on Mathematical Analysis, 57(1):65–94, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1644523.

Raita:2024:SPC

[RRTT24] Bogdan Raita, Angkana Rüland, Camillo Tissot, and Antonio Tribuzio. On scaling properties for a class of two-well problems for higher order homogeneous linear differential operators. SIAM Journal on Mathematical Analysis, 56(3):3720–3758, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1588287.

Roscani:2022:OPS

[RRV22] Sabrina D. Roscani, Katarzyna Ryszewska, and Lucas Venturato. A one-phase spacefractional Stefan problem with no liquid initial domain. SIAM Journal on Mathematical Analysis, 54(5):5489–5523, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1461599.

Rakesh:2020:FAI

[RS20a] Rakesh and Mikko Salo. Fixed angle inverse scattering for almost symmetric or controlled perturbations. SIAM Journal on Mathematical Analysis, 52 (6):5467–5499, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rindler:2020:RPC

[RS20b] Filip Rindler and Giles Shaw. Relaxation for partially coercive integral functionals with linear growth. SIAM Journal on Mathematical Analysis, 52 (5):4806–4860, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rauch:2021:EPP

[RS21] Jeffrey Rauch and L. Ridgway Scott. The electrostatic potential of periodic crystals. SIAM Journal on Mathematical Analysis, 53(2):1474–1491, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Radici:2023:ESM

[RS23a] Emanuela Radici and Federico Stra. Entropy solutions of mildly singular nonlocal scalar conservation laws with congestion via deterministic particle methods. SIAM Journal on Mathematical Analysis, 55(3):2001–2041, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462994.

Roubicek:2023:VSP

[RS23b] Tomás Roubícek and Ulisse Stefanelli. Viscoelastodynamics of swelling porous solids at large strains by an Eulerian approach. SIAM Journal on Mathematical Analysis, 55(4):2677–2703, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1474229.

Romero:2024:EBT

[RS24] José Luis Romero and Michael Speckbacher. Estimation of binary time-frequency masks from ambient noise. SIAM Journal on Mathematical Analysis, 56(3):3559–3587, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Rout:2025:MDG

[RS25] Andrew Rout and Vedran Sohinger. A microscopic derivation of Gibbs measures for the 1D focusing quintic nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 57(5):4680–4755, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1681227.

Rondi:2021:SDR

[RSS21] Luca Rondi, Eva Sincich, and Mourad Sini. Stable determination of a rigid scatterer in elastodynamics. SIAM Journal on Mathematical Analysis, 53 (2):2660–2689, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ruland:2021:SMS

[Rül21] Angkana Rüland. On single measurement stability for the fractional Calderón problem.

SIAM Journal on Mathematical Analysis, 53(5):5094–5113, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rohde:2020:HNN

[RvW20] Christian Rohde and Lars von Wolff. Homogenization of nonlocal Navier–Stokes– Korteweg equations for compressible liquid-vapor flow in porous media. SIAM Journal on Mathematical Analysis, 52 (6):6155–6179, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rybka:2023:CSC

[RW23] Piotr Rybka and Glen Wheeler.

Convergence of solutions to a convective Cahn–Hilliard-type equation of the sixth order in case of small deposition rates. SIAM Journal on Mathematical Analysis, 55(5):5823–5861, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1540429.

Ren:2021:LTB

[RWZ21] Siqi Ren, Dongyi Wei, and Zhifei Zhang. Long time behavior of Alfvén waves in flowing plasma: The destruction of the magnetic Island. SIAM Journal on Mathematical Analysis, 53(5):5548–5579, ???? 2021. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Rockner:2020:SHE

[RWZZ20] Michael Röckner, Bo Wu, Rongchan Zhu, and Xiangchan Zhu. Stochastic heat equations with values in a manifold via Dirichlet forms. SIAM Journal on Mathematical Analysis, 52 (3):2237–2274, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rundell:2020:IST

[RZ20] William Rundell and Zhidong Zhang. On the identification of source term in the heat equation from sparse data. SIAM Journal on Mathematical Analysis, 52(2):1526–1548, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2021:UDA

[RZ21] Kui Ren and Yimin Zhong. Unique determination of absorption coefficients in a semilinear transport equation. SIAM Journal on Mathematical Analysis, 53(5):5158–5184, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2021:SKF

[RZZ21] Kui Ren, Hongkai Zhao, and Yimin Zhong. Separability of the kernel function in an integral formulation for the anisotropic radiative transfer equation. SIAM Journal on Mathematical Analysis, 53(5): 5613–5645, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Salazar:2021:CTE

[SAB21] Lorena Aguirre Salazar, Stan Alama, and Lia Bronsard. Convergence of the TFDW energy to the liquid drop model. SIAM Journal on Mathematical Analysis, 53(3):3493–3519, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saffirio:2020:HEV

[Saf20] Chiara Saffirio. From the Hartree equation to the Vlasov–Poisson system: Strong convergence for a class of mixed states. SIAM Journal on Mathematical Analysis, 52 (6):5533–5553, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Said:2023:GPQ

[Sai23] Ayman Rimah Said. A geometric proof of the quasi-linearity of the water waves system. SIAM Journal on Mathematical Analysis, 55(1):508–556, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M141587X.

Sprengel:2017:TIT

[SCB17] M. Sprengel, G. Ciaramella, and A. Borzì. A theoretical investigation of timedependent Kohn–Sham equations. SIAM Journal on Mathematical Analysis, 49(3):1681– 1704, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [SCB20].

Sprengel:2020:ETI

[SCB20] M. Sprengel, G. Ciaramella, and A. Borzì. Erratum: A Theoretical Investigation of Timedependent Kohn–Sham Equations. SIAM Journal on Mathematical Analysis, 52(1):1006– 1008, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [SCB17].

Scott:2024:NPL

[SD24] James M. Scott and Qiang Du. Nonlocal problems with local boundary conditions I: Function spaces and variational principles. SIAM Journal on Mathematical Analysis, 56(3):4185–4222, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1588111.

Scott:2025:NPL

[SD25] James M. Scott and Qiang Du. Nonlocal problems with local boundary conditions II:

Green’s identities and regularity of solutions. SIAM Journal on Mathematical Analysis, 57(1):404–451, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1643505.

Selim:2024:PDI

[Sel24] Salem Selim. Partial data inverse problems for magnetic Schrödinger operators with potentials of low regularity. SIAM Journal on Mathematical Analysis, 56(4):4594–4622, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1530707.

Senik:2023:HLP

[Sen23] Nikita N. Senik. Homogenization for locally periodic elliptic problems on a domain. SIAM Journal on Mathematical Analysis, 55(2):849–881, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419337.

Seth:2024:IDP

[Set24] Douglas Svensson Seth. Internal doubly periodic gravitycapillary waves with vorticity. SIAM Journal on Mathematical Analysis, 56(6):7545–7575, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1556319.

Schulz:2022:FKB

[SH22] Erick Schulz and Ralf Hiptmair. First-kind boundary integral equations for the Dirac operator in 3-dimensional Lipschitz domains. SIAM Journal on Mathematical Analysis, 54(1):616–648, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1389224.

Staudt:2025:UKP

[SHM25] Thomas Staudt, Shayan Hundrieser, and Axel Munk. On the uniqueness of Kantorovich potentials. SIAM Journal on Mathematical Analysis, 57 (2):1452–1482, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1658966.

Shu:2020:HLF

[Shu20] Yan Shu. From Hopf–Lax formula to optimal weak transfer plan. SIAM Journal on Mathematical Analysis, 52(3): 3052–3072, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shu:2021:EAD

[Shu21] Ruiwen Shu. Equilibration of aggregation–diffusion equations with weak interaction forces. SIAM Journal on Mathematical Analysis, 53(6):6562– 6616, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Silini:2023:LES

[Sil23] Lauro Silini. Local existence of smooth solutions for the semigeostrophic equations on curved domains. SIAM Journal on Mathematical Analysis, 55(6):6554–6579, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Simon:2021:QAR

[Sim21] Theresa M. Simon. Quantitative aspects of the rigidity of branching microstructures in shape memory alloys via hmeasures. SIAM Journal on Mathematical Analysis, 53(4): 4537–4567, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sinambela:2021:LAS

[Sin21]  Daniel Sinambela.  Large-

amplitude solitary waves in two-layer density stratified water. SIAM Journal on Mathematical Analysis, 53(4):4812– 4864, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sun:2022:URH

[SL22] Weiran Sun and Qin Li.

Unique reconstruction of the heat-reflection indices at solid interfaces. SIAM Journal on Mathematical Analysis, 54 (5):5742–5759, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1438773.

Snelson:2020:GBI

[Sne20] Stanley Snelson. Gaussian bounds for the inhomogeneous Landau equation with hard potentials. SIAM Journal on Mathematical Analysis, 52(2): 2081–2097, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Solci:2024:NIV

[Sol24] Margherita Solci. Nonlocalinteraction vortices. SIAM Journal on Mathematical Analysis, 56(3):3430–3451, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Sagiv:2020:TIU

[SS20] Amir Sagiv and Stefan Steinerberger. Transport and interface: an uncertainty principle for the Wasserstein distance. SIAM Journal on Mathematical Analysis, 52(3):3039–3051, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schwarzacher:2022:WSU

[SS22a] Sebastian Schwarzacher and Matthias Sroczinski. Weakstrong uniqueness for an elastic plate interacting with the Navier–Stokes equation. SIAM Journal on Mathematical Analysis, 54(4):4104–4138, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443509.

Semoradova:2022:DED

[SS22b] Iveta Semoradova and Petr Siegl. Diverging eigenvalues in domain truncations of Schrödinger operators with complex potentials. SIAM Journal on Mathematical Analysis, 54(4):5064–5101, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1439699.

Sarrazin:2024:LOT

[SS24] Clément Sarrazin and Bernhard Schmitzer. Linearized optimal transport on manifolds. SIAM Journal on Mathematical Analysis, 56(4):4970–5016, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1564535.

Saut:2024:WBG

[SSWZ24] Jean-Claude Saut, Shihan Sun, Yuexun Wang, and Yi Zhang. Wave breaking for the generalized Fornberg–Whitham equation. SIAM Journal on Mathematical Analysis, 56 (4):4440–4465, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1603431.

Saito:2020:SFB

[SSZ20] Hirokazu Saito, Yoshihiro Shibata, and Xin Zhang. Some free boundary problem for twophase inhomogeneous incompressible flows. SIAM Journal on Mathematical Analysis, 52 (4):3397–3443, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shvydkoy:2020:TBF

[ST20a] Roman Shvydkoy and Eitan Tadmor. Topologically based fractional diffusion and emergent dynamics with short-range interactions. SIAM Journal on Mathematical Analysis, 52(6):5792–5839, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sun:2020:GMD

[ST20b] Chenmin Sun and Nikolay Tzvetkov. Gibbs measure dynamics for the fractional NLS. SIAM Journal on Mathematical Analysis, 52(5):4638–4704, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stevenson:2021:TWS

[ST21] Noah Stevenson and Ian Tice. Traveling wave solutions to the multilayer free boundary incompressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 53(6): 6370–6423, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Snelson:2024:ODE

[ST24] Stanley Snelson and Eduardo V. Teixeira. Optimal domains for elliptic eigenvalue problems with rough coefficients. SIAM Journal on Mathematical Analysis, 56(3):3412– 3429, May 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Stefanov:2020:SSD

[Ste20] Plamen Stefanov. Semiclassical sampling and discretization of certain linear inverse problems. SIAM Journal on Mathematical Analysis, 52(6): 5554–5597, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stevenson:2025:PGC

[Ste25a]  Noah Stevenson.  Periodic

gravity-capillary roll wave solutions to the inclined viscous shallow water equations in two dimensions. SIAM Journal on Mathematical Analysis, 57(2):1342–1369, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1674406.

Stewart:2025:LTD

[Ste25b] Gavin Stewart. Long time decay and asymptotics for the complex mKdV equation. SIAM Journal on Mathematical Analysis, 57(1):825–885, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1528495.

Sun:2021:LTR

[Sun21] Changzhen Sun. Long-term regularity of two-dimensional Navier–Stokes–Poisson equations. SIAM Journal on Mathematical Analysis, 53(5):5114– 5157, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Steinbach:2022:NAS

[SUT22] Olaf Steinbach and Carolina Urzúa-Torres. A new approach to space-time boundary integral equations for the wave equation. SIAM Journal on Mathematical Analysis, 54(2):1370–1392, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1420034.

Schimperna:2020:CSO

[SW20a] Giulio Schimperna and Hao Wu. On a class of sixth-order Cahn–Hilliard-type equations with logarithmic potential. SIAM Journal on Mathematical Analysis, 52(5):5155–5195, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffilani:2020:SCN

[SW20b] Gigliola Staffilani and Bobby Wilson. Stability of the cubic nonlinear Schrödinger equation on an irrational torus. SIAM Journal on Mathematical Analysis, 52(2):1318–1342, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sagiv:2022:EGC

[SW22a] Amir Sagiv and Michael I. Weinstein. Effective gaps in continuous Floquet Hamiltonians. SIAM Journal on Mathematical Analysis, 54(1):986–1021, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1417363.

Saut:2022:WBW

[SW22b] Jean-Claude Saut and Yuexun Wang. The wave breaking for Whitham-type equations revisited. SIAM Journal on Mathematical Analysis, 54(2):2295–2319, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1345207.

Sini:2022:ISP

[SW22c] Mourad Sini and Haibing Wang. The inverse source problem for the wave equation revisited: a new approach. SIAM Journal on Mathematical Analysis, 54(5):5160–5181, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1463689.

Sato:2025:SNS

[SWX25] Yohei Sato, Zhi-Qiang Wang, and Jiankang Xia. Symmetric nonradial solutions for nonlinear Schrödinger systems with mixed couplings on R^{n}. SIAM Journal on Mathematical Analysis, 57(1):306–335, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1613785.

Saut:2020:LTE

[SX20a] Jean-Claude Saut and Li Xu. Long time existence for a strongly dispersive Boussinesq system. SIAM Journal on Mathematical Analysis, 52(3): 2803–2848, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schochet:2020:MFT

[SX20b] Steven Schochet and Xin Xu. Moderately fast three-scale singular limits. SIAM Journal on Mathematical Analysis, 52(4): 3444–3462, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stefanov:2022:IBP

[SZ22] Plamen Stefanov and Yimin Zhong. Inverse boundary problem for the two photon absorption transport equation. SIAM Journal on Mathematical Analysis, 54(3):2753–2767, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1417387.

Takasao:2021:OPB

[Tak21] Keisuke Takasao. On obstacle problem for Brakke’s mean curvature flow. SIAM Journal on Mathematical Analysis, 53 (6):6355–6369, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Takeuchi:2025:SSN

[Tak25] Taiki Takeuchi. On the strong solutions to the Navier–Stokes system with excessively singular external forces. SIAM Journal on Mathematical Analysis, 57(5):5382–5419, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1687492.

Tan:2021:EDM

[Tan21] Changhui Tan. Eulerian dynamics in multidimensions with radial symmetry. SIAM Journal on Mathematical Analysis, 53(3):3040–3071, ???? 2021. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Tawri:2024:SFS

[Taw24] Krutika Tawri. A stochastic fluid-structure interaction problem with the Navier-slip boundary condition. SIAM Journal on Mathematical Analysis, 56(6):7508–7544, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M164029X.

Tesfahun:2020:SDS

[Tes20] Achenef Tesfahun. Small data scattering for cubic Dirac equation with Hartree type nonlinearity in R^{1+3}. SIAM Journal on Mathematical Analysis, 52 (3):2969–3003, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tetlow:2022:RTD

[Tet22] Alexander Tetlow. Recovery of a time-dependent Hermitian connection and potential appearing in the dynamic Schrödinger equation. SIAM Journal on Mathematical Analysis, 54(2):1347–1369, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1353265.

Tione:2023:CPD

[Tio23]  Riccardo Tione.  Critical

points of degenerate polyconvex energies. SIAM Journal on Mathematical Analysis, 55(4):3205–3225, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1519018.

Tokinaga:2020:VSS

[TR20] Tokinaga and Piotr Rybka. On viscosity solutions of spacefractional diffusion equations of Caputo type. SIAM Journal on Mathematical Analysis, 52(1):653–681, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Treton:2024:BVG

[Tré24] Samuel Tréton. Blow-up vs.

global existence for a fujitatype heat exchanger system. SIAM Journal on Mathematical Analysis, 56(2):2191–2212, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Tavares:2024:SVP

[TSMO24] Eduardo H. Gomes Tavares, Marcio A. Jorge Silva, To Fu Ma, and Higidio P. Oquendo. Shearing viscoelasticity in partially dissipative Timoshenko– Boltzmann systems. SIAM Journal on Mathematical Analysis, 56(1):1149–1178, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Tsuge:2022:GES

[Tsu22] Naoki Tsuge. Global existence of a solution for isentropic gas flow in the laval nozzle with a friction term. SIAM Journal on Mathematical Analysis, 54(2):2142–2162, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1415029.

Tadmor:2022:CTG

[TT22] Eitan Tadmor and Changhui Tan. Critical threshold for global regularity of the Euler– Monge–Ampère system with radial symmetry. SIAM Journal on Mathematical Analysis, 54(4):4277–4296, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437767.

Tang:2024:RDM

[TT24] Bao Quoc Tang and Bao-Ngoc Tran. Rigorous derivation of Michaelis–Menten kinetics in the presence of slow diffusion. SIAM Journal on Mathematical Analysis, 56(5):5995–6024, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1579406.

Tao:2021:TDF

[TW21] Youshan Tao and Michael Winkler. Taxis-driven formation of singular hotspots in a May–Nowak type model for virus infection. SIAM Journal on Mathematical Analysis, 53 (2):1411–1433, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tao:2022:ETQ

[TW22a] Youshan Tao and Michael Winkler. Existence theory and qualitative analysis for a fully cross-diffusive predatorprey system. SIAM Journal on Mathematical Analysis, 54(4):4806–4864, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1449841.

Trakhinin:2022:WPM

[TW22b] Yuri Trakhinin and Tao Wang.

Well-posedness for moving interfaces with surface tension in ideal compressible MHD. SIAM Journal on Mathematical Analysis, 54(6):5888–5921, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1488429.

Trifunovic:2023:IPB

[TW23] Srdan Trifunović and Ya-Guang Wang. On the interaction problem between a compressible viscous fluid and a nonlinear thermoelastic plate. SIAM Journal on Mathematical Analysis, 55(4):3509–3566, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1385640.

Takahashi:2025:NSF

[TW25] Tomoki Takahashi and Keiichi Watanabe. Navier–Stokes flow in the exterior of a moving obstacle with a Lipschitz boundary. SIAM Journal on Mathematical Analysis, 57 (2):2071–2113, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M163801X.

Tao:2020:ZVL

[TWZ20] Tao Tao, Wendong Wang, and Zhifei Zhang. Zero-viscosity limit of the Navier–Stokes equations with the Navier friction boundary condition. SIAM Journal on Mathematical Analysis, 52(2):1040–1095, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tang:2024:MMB

[TX24] Xingdong Tang and Guixiang Xu. Minimal mass blow-up solutions for the L^{2}-critical NLS with the delta potential for even data in one dimension. SIAM Journal on Mathematical Analysis, 56(2):1727–1769, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Uhlmann:2024:ACP

[UW24] Gunther Uhlmann and Yiran Wang. The anisotropic Calderón problem for high fixed frequency. SIAM Journal on Mathematical Analysis, 56(3):4084–4103, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1579029.

Uhlmann:2023:IBV

[UZ23] Gunther Uhlmann and Yang Zhang. An inverse boundary value problem arising in nonlinear acoustics. SIAM Journal on Mathematical Analysis, 55(2):1364–1404, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1492490.

Varholm:2020:GBW

[Var20] Kristoffer Varholm. Global bifurcation of waves with multiple critical layers. SIAM Journal on Mathematical Analysis, 52(5):5066–5089, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanMeurs:2022:BLA

[vM22] Patrick van Meurs. Boundarylayer analysis of repelling particles pushed to an impenetrable barrier. SIAM Journal on Mathematical Analysis, 54(2):1742–1774, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1420198.

vanMeurs:2024:DDD

[vMP24] Patrick van Meurs and Stefania Patrizi. Discrete dislocation dynamics with annihilation as the limit of the Peierls– Nabarro model in one dimension. SIAM Journal on Mathematical Analysis, 56(1):197– 233, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Vogelius:2021:FRC

[VX21] Michael Vogelius and Jingni Xiao. Finiteness results concerning nonscattering wave numbers for incident plane and Herglotz waves. SIAM Journal on Mathematical Analysis, 53 (5):5436–5464, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vogelius:2025:FRN

[VX25] Michael S. Vogelius and Jingni Xiao. Finiteness results for nonscattering Herglotz waves: The case of inhomogeneities obtained by very general perturbations of disks. SIAM Journal on Mathematical Analysis, 57(5):5135–5160, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1709844.

Vasseur:2023:BVE

[VY23] Alexis F. Vasseur and Jincheng Yang. Boundary vorticity estimates for Navier–Stokes and application to the inviscid limit. SIAM Journal on Mathematical Analysis, 55 (4):3081–3107, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1503567.

Wang:2020:TDI

[Wan20]  Fei Wang.  The three-

dimensional inviscid limit problem with data analytic near the boundary. SIAM Journal on Mathematical Analysis, 52(4):3520–3545, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2021:GWP

[Wan21] Wenjun Wang. Global wellposedness and decay rates of strong solutions to a P 1approximation model arising in radiation hydrodynamics. SIAM Journal on Mathematical Analysis, 53(5):5361–5404, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2022:GSS

[Wan22]  Chunpeng Wang.  Global

smooth sonic-supersonic flows in a class of critical nozzles. SIAM Journal on Mathematical Analysis, 54(2):1820–1859, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1373906.

Wang:2023:PSS

[Wan23] Teng Wang. Planar stationary solution to outflow problem for compressible heat conducting gas in R_{+}^{3}: Stability and convergence rate. SIAM Journal on Mathematical Analysis, 55(1):210–274, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1488351.

Wang:2024:INB

[Wan24] Yiran Wang. Identification of nonlinear beam-hardening effects in X-ray tomography. SIAM Journal on Mathematical Analysis, 56(6):7290–7305, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1596491.

Wang:2020:RSQ

[WdlL20] Fenfen Wang and Rafael de la Llave. Response solutions to quasi-periodically forced systems, even to possibly ill-posed PDEs, with strong dissipation and any frequency vectors. SIAM Journal on Mathematical Analysis, 52(4):3149–3191, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weber:2020:HPC

[Web20] Jörg Weber. Hot plasma in a container — an optimal control problem. SIAM Journal on Mathematical Analysis, 52 (3):2895–2929, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Webber:2025:SRR

[WHQ25] James W. Webber, Sean Holman, and Eric Todd Quinto. Surface of revolution Radon transforms with centers on generalized surfaces in R^{N}. SIAM Journal on Mathematical Analysis, 57(2):2013–2036, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1627018.

Wei:2025:SDE

[WHY25] Jinlong Wei, Junhao Hu, and Chenggui Yuan. Stochastic differential equations with low regularity growing drifts and applications. SIAM Journal on Mathematical Analysis, 57 (5):4867–4907, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1636939.

Wichmann:2023:TRS

[Wic23] Jörn Wichmann. On temporal regularity of strong solutions to stochastic p-Laplace systems.

SIAM Journal on Mathematical Analysis, 55(4):3713–3730, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1460491.

Winkler:2020:SMS

[Win20]  Michael Winkler.  Small-

mass solutions in the twodimensional Keller–Segel system coupled to the Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 52(2):2041–2080, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2022:LBS

[WJW22] Yanqing Wang, Quansen Jiu, and Wei Wei. Leray’s backward self-similar solutions to the 3D Navier–Stokes equations in Morrey spaces. SIAM Journal on Mathematical Analysis, 54(3):2768–2791, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1346055.

Wang:2023:AAR

[WKM23] Renhai Wang, Kush Kinra, and Manil T. Mohan. Asymptotically autonomous robustness in probability of random attractors for stochastic Navier–Stokes equations on unbounded Poincaré domains. SIAM Journal on Mathematical Analysis, 55(4):2644–2676, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1517111.

Wei:2021:STS

[WMZZ21] Mengmeng Wei, Ming Mei, Guojing Zhang, and Kaijun Zhang. Smooth transonic steady states of hydrodynamic model for semiconductors. SIAM Journal on Mathematical Analysis, 53(4): 4908–4932, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2022:HLE

[WO22] Lei Wu and Zhimeng Ouyang.

Hydrodynamic limit of 3dimensional evolutionary Boltzmann equation in convex domains. SIAM Journal on Mathematical Analysis, 54 (2):2508–2569, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1375735.

Wickman:2023:GFP

[WO23] Clare Wickman and Kasso A.

Okoudjou. Gradient flows for probabilistic frame potentials in the Wasserstein space. SIAM Journal on Mathematical Analysis, 55(3):2324–2346, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1425633.

Wright:2024:ACM

[Wri24] J. Douglas Wright. Approximation of Calogero–Moser lattices by Benjamin–Ono equations. SIAM Journal on Mathematical Analysis, 56 (4):5583–5603, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1629869.

Weihs:2024:CFG

[WT24] Adrien Weihs and Matthew Thorpe. Consistency of fractional graph-Laplacian regularization in semisupervised learning with finite labels. SIAM Journal on Mathematical Analysis, 56(4):4253–4295, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1559087.

Wang:2022:HAM

[WW22] Chao Wang and Yuxi Wang. On the hydrostatic approximation of the MHD equations in a thin strip. SIAM Journal on Mathematical Analysis, 54(1):1241–1269, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1425360.

Wu:2023:HSE

[WW23] Wenpei Wu and Yong Wang.

On the half-space or exterior problems of the 3D compressible elastic Navier–Stokes– Poisson equations. SIAM Journal on Mathematical Analysis, 55(4):2996–3043, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1526162.

Wang:2024:VDL

[WW24] Yinghui Wang and Huanyao Wen. The vanishing diffusion limit for an Oldroyd-B model in R_{+}^{2}. SIAM Journal on Mathematical Analysis, 56(5):6551–6612, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1604795.

Wang:2021:HLB

[WWZ21] Guanfa Wang, Yong Wang, and Jiawei Zhou. Hydrodynamic limit of the Boltzmann equation to the planar rarefaction wave in three dimensional space. SIAM Journal on Mathematical Analysis, 53(4):4692– 4726, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2023:TIM

[WWZ23a] Rong-Nian Wang, Jianhong Wu, and Jia-Cheng Zhao. Theory of invariant manifolds for infinite-dimensional nonautonomous dynamical systems and applications. SIAM Journal on Mathematical Analysis, 55(5):5386–5431, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1463392.

Wang:2023:VSV

[WWZ23b] Yinghui Wang, Huanyao Wen, and Weihao Zhang. Vanishing shear viscosity limit for the Navier–Stokes equations with cylindrical symmetry: Boundary layer and optimal convergence rate. SIAM Journal on Mathematical Analysis, 55(3):1631–1675, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1509321.

Wang:2021:VVS

[WX21] Yanjin Wang and Zhouping Xin. Vanishing viscosity and surface tension limits of incompressible viscous surface waves. SIAM Journal on Mathematical Analysis, 53(1):574–648, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weng:2021:SST

[WXX21] Shangkun Weng, Chunjing Xie, and Zhouping Xin. Structural stability of the transonic shock problem in a divergent three-dimensional axisymmetric perturbed nozzle. SIAM Journal on Mathematical Analysis, 53(1):279–308, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2024:EDA

[WXZ24] Yin Wang, Qingxiang Xu, and Peng Zhou. Evolution of dispersal in advective homogeneous environments: Inflow versus outflow. SIAM Journal on Mathematical Analysis, 56(2):1643–1671, March 2024. CODEN SJMAAH. ISSN 1095-7154.

Wang:2021:WPE

[WY21] Zhan Wang and Jiaqi Yang.

Well-posedness of electrohydrodynamic interfacial waves under tangential electric field. SIAM Journal on Mathematical Analysis, 53(2):2567–2594, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2024:ESE

[WY24] Shuaikun Wang and Tong Yang. Existence and smoothing effect of measure valued solution to the homogeneous Boltzmann equation with debye-Yukawa potential. SIAM Journal on Mathematical Analysis, 56(5):5969–5994, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1562123.

Wu:2023:GSN

[WYZ23] Guochun Wu, Lei Yao, and Yinghui Zhang. Global stability and nonvanishing vacuum states of 3D compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 55(2):882–899, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1478859.

Wang:2020:BFT

[WZ20a] Ya-Guang Wang and Shi-Yong Zhu. Back flow of the two-dimensional unsteady Prandtl boundary layer under an adverse pressure gradient. SIAM Journal on Mathematical Analysis, 52(1):954–966, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2020:EBS

[WZ20b] Lina Wu and Lei Zhang. Estimates of bubbling solutions of SU(3) Toda systems at critical parameters. Part 1. SIAM Journal on Mathematical Analysis, 52(4):3738–3776, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2022:GEW

[WZ22] Guodong Wang and Bijun Zuo. Global existence of weak solutions to the compressible Navier–Stokes equations with temperature-depending viscosity coefficients. SIAM Journal on Mathematical Analysis, 54(3):3724–3756, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1405915.

Wang:2023:PRI

[WZ23] Haitao Wang and Xiongtao Zhang. Propagation of rough initial data for Navier–Stokes equation. SIAM Journal on Mathematical Analysis, 55 (2):966–1006, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1474540.

Wu:2024:LDS

[WZ24] Weina Wu and Jianliang Zhai.

Large deviations for stochastic generalized porous media equations driven by lévy noise. SIAM Journal on Mathematical Analysis, 56(1):1–42, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Wen:2025:GEL

[WZ25] Huanyao Wen and Xingyang Zhang. Global existence of large solution to the isothermal compressible Navier– Stokes equations with degenerate viscosity and vacuum. SIAM Journal on Mathematical Analysis, 57(2):1280–1314, ???? 2025. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1671803.

Wu:2025:DLO

[WZGX25] Weijun Wu, Fujun Zhou, Weihua Gong, and Yuan Xu. Diffusive limit of the one-species Vlasov–Maxwell– Boltzmann system for cutoff hard potentials. SIAM Journal on Mathematical Analysis, 57(2):1526–1585, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1665325.

Wu:2020:OLT

[WZZ20] Guochun Wu, Yinghui Zhang, and Lan Zou. Optimal largetime behavior of the two-phase fluid model in the whole space. SIAM Journal on Mathematical Analysis, 52(6):5748–5774, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wei:2024:GCL

[WZZ24] Jing Wei, Minyi Zhang, and Changjiang Zhu. Global classical large solutions for the radiation hydrodynamics model in unbounded domains. SIAM Journal on Mathematical Analysis, 56(4):4811–4833, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1598581.

Xu:2022:LTB

[XC22] Jiaohui Xu and Tomás Caraballo. Long time behavior of stochastic nonlocal partial differential equations and Wong–Zakai approximations. SIAM Journal on Mathematical Analysis, 54(3):2792–2844, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1412645.

Xu:2024:DLD

[XCV24] Jiaohui Xu, Tomás Caraballo, and José Valero. Dynamics and large deviations for fractional stochastic partial differential equations with lévy noise. SIAM Journal on Mathematical Analysis, 56(1):1016– 1067, January 2024. CODEN SJMAAH. ISSN 1095-7154.

Xiang:2020:EME

[Xia20] Xingfei Xiang. On L^{r} estimates for Maxwell’s equations with complex coefficients in Lipschitz domains. SIAM Journal on Mathematical Analysis, 52(6):6140–6154, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xie:2025:UTE

[Xie25]  Pengzhi Xie.  Uniform-in-

time estimates on the size of chaos for interacting particle systems. SIAM Journal on Mathematical Analysis, 57(5):5358–5381, ???? 2025.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1714411.

Xu:2023:SSR

[XMN23] Jianing Xu, Ming Mei, and Shinya Nishibata. Structural stability of radial interior subsonic steady-states to n-D Euler–Poisson system of semiconductor models with sonic boundary. SIAM Journal on Mathematical Analysis, 55(6):7741–7761, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Xia:2023:ROT

[XX23a] Qinglan Xia and Shaofeng Xu. Ramified optimal transportation with payoff on the boundary. SIAM Journal on Mathematical Analysis, 55(1):186–209, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1367714.

Xu:2023:URC

[XX23b] Jiali Xu and Xiang Xu. On uniqueness of recovering coefficients from localized Dirichletto-Neumann map for piecewise homogeneous piezoelectricity. SIAM Journal on Mathematical Analysis, 55(2):571–602, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1447957.

Xu:2021:TDG

[XY21] Gang Xu and Huicheng Yin.

Three-dimensional global supersonic Euler flows in the infinitely long divergent nozzles. SIAM Journal on Mathematical Analysis, 53(1):133–180, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xia:2024:MSC

[XZ24] Jiankang Xia and Xu Zhang.

Multibump solutions for critical Choquard equation. SIAM Journal on Mathematical Analysis, 56(3):3832–3860, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1581820.

Xin:2024:GWP

[XZZ24] Zhouping Xin, Liqun Zhang, and Junning Zhao. Global well-posedness and regularity of weak solutions to the Prandtl’s system. SIAM Journal on Mathematical Analysis, 56(3):3042–3081, May 2024. CODEN SJMAAH. ISSN 1095-7154.

Yamazaki:2022:NLT

[Yam22] Kazuo Yamazaki. NonUniqueness in law for two-dimensional Navier–Stokes equations with diffusion weaker than a full Laplacian. SIAM Journal on Mathematical Analysis, 54 (4):3997–4042, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1451087.

Yan:2021:IBP

[Yan21a]  Lili Yan.  Inverse bound-

ary problems for biharmonic operators in transversally anisotropic geometries. SIAM Journal on Mathematical Analysis, 53(6):6617–6653, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2021:SFB

[Yan21b] Ruoxuan Yang. Shock formation of the Burgers–Hilbert equation. SIAM Journal on Mathematical Analysis, 53(5): 5756–5802, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2025:VTA

[Yan25] Haocheng Yang. Virial theorem and its applications in instability of two-phase water-wave. SIAM Journal on Mathematical Analysis, 57 (5):5420–5453, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1660553.

Yao:2022:NSL

[YCRS22] Shuai Yao, Haibo Chen, Vicentiu D. Radulescu, and Juntao Sun. Normalized solutions for lower critical Choquard equations with critical Sobolev perturbation. SIAM Journal on Mathematical Analysis, 54(3):3696–3723, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1463136.

Yang:2021:PMS

[YGD21] Jiaqi Yang, Joan Gimeno, and Rafael De la Llave. Parameterization method for statedependent delay perturbation of an ordinary differential equation. SIAM Journal on Mathematical Analysis, 53(4): 4031–4067, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yan:2022:IVP

[YL22] Kan Yan and Yue Liu.

The initial-value problem to the modified two-component Euler–Poincaré equations. SIAM Journal on Mathematical Analysis, 54(2):2006–2039, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M138274X.

Yoshizawa:2021:REG

[Yos21] Kensuke Yoshizawa. A remark on elastic graphs with the symmetric cone obstacle. SIAM Journal on Mathematical Analysis, 53(2):1857–1885, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yuan:2023:PVS

[Yua23] Qian Yuan. Planar viscous shocks with periodic perturbations for scalar multidimensional viscous conservation laws. SIAM Journal on Mathematical Analysis, 55 (3):1499–1523, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462453.

Yuan:2022:PPC

[YY22] Qian Yuan and Yuan Yuan. Periodic perturbations of a composite wave of two viscous shocks for 1-d full compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 54(3):2876–2905, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421489.

Yu:2024:GWPb

[YY24] Xueying Yu and Haitian Yue. On the global well-posedness for the periodic quintic nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 56(2):1851–1902, March 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Yu:2021:GWP

[YYZ21] Xueying Yu, Haitian Yue, and Zehua Zhao. Global wellposedness for the focusing cubic NLS on the product space R × T^{3}. SIAM Journal on Mathematical Analysis, 53(2): 2243–2274, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yu:2024:GWPa

[YYZ24] Xueying Yu, Haitian Yue, and Zehua Zhao. Global wellposedness and scattering for fourth-order Schrödinger equations on waveguide manifolds. SIAM Journal on Mathematical Analysis, 56(1):1427–1458, February 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Yao:2021:ASS

[YZ21] Huancheng Yao and Changjiang Zhu. Asymptotic stability of the superposition of viscous contact wave with rarefaction waves for the compressible Navier–Stokes–Maxwell equations. SIAM Journal on Mathematical Analysis, 53(5):6122– 6163, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2024:DDK

[YZ24a] Xiongfeng Yang and Lixian Zhao. Derivation of decoupled Korteweg–de Vries/ Zakharov–Kuznetsov equations from Vlasov–Poisson system in the torus. SIAM Journal on Mathematical Analysis, 56(6):7350–7389, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1643025.

Yao:2024:AHE

[YZ24b] Luming Yao and Lun Zhang. Asymptotics of the hard edge Pearcey determinant. SIAM Journal on Mathematical Analysis, 56(1):137–172, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Yao:2025:NSV

[YZ25] Huancheng Yao and Changjiang Zhu. Nonlinear stability of viscous contact wave for the isentropic MHD equations with free boundary. SIAM Journal on Mathematical Analysis, 57(1):1021–1038, ???? 2025. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1688126.

Yang:2022:DTS

[YZZ22] Kai Yang, Chongchun Zeng, and Xiaoyi Zhang. Dynamics of threshold solutions for energy critical NLS with inverse square potential. SIAM Journal on Mathematical Analysis, 54(1):173–219, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1406003.

Zeng:2023:GSP

[Zen23] Huihui Zeng. Global solution to the physical vacuum problem of compressible Euler equations with damping and gravity. SIAM Journal on Mathematical Analysis, 55(6):6375–6424, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Zhang:2021:ATS

[Zha21] Yiping Zhang. Approximate two-sphere one-cylinder inequality in parabolic periodic homogenization. SIAM Journal on Mathematical Analysis, 53(5):5835–5852, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2022:SLS

[Zha22] Deng Zhang. Strichartz and local smoothing estimates for stochastic dispersive equations with linear multiplicative noise. SIAM Journal on Mathematical Analysis, 54 (6):5981–6017, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1426304.

Zhao:2023:MIP

[Zha23]  Jiefeng Zhao.  Mild ill-

posedness for 2D magnetomicropolar fluid equations near a background magnetic field. SIAM Journal on Mathematical Analysis, 55(5):5967–5992, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469432.

Zhou:2022:GSV

[Zho22] Wenshu Zhou. Global solvability and vanishing shear viscosity limit for a simplified compressible Navier–Stokes system with temperature-dependent viscosity. SIAM Journal on Mathematical Analysis, 54 (4):4865–4890, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1477441.

Zhu:2021:VAH

[Zhu21] Yuzhe Zhu. Velocity averaging and Hölder regularity for kinetic Fokker–Planck equations with general transport operators and rough coefficients. SIAM Journal on Mathematical Analysis, 53(3):2746–2775, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2024:SAE

[ZHY24] Ruixi Zhang, Qian Huang, and Wen-An Yong. Stability analysis of an extended quadrature method of moments for kinetic equations. SIAM Journal on Mathematical Analysis, 56(4):4687–4711, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M157911X.

Zeng:2024:NDP

[ZiDRdW24] Shengda Zeng, Vicen iu D. R. dulescu, and Patrick Winkert. Nonlocal double phase implicit obstacle problems with multivalued boundary conditions. SIAM Journal on Mathematical Analysis, 56(1):877–912, January 2024. CODEN SJ-MAAH. ISSN 1095-7154.

Zillinger:2023:ECL

[Zil23] Christian Zillinger. On echo chains in the linearized Boussinesq equations around traveling waves. SIAM Journal on Mathematical Analysis, 55 (5):5127–5188, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458053.

Zhai:2020:GLS

[ZLZ20] Xiaoping Zhai, Yongsheng Li, and Fujun Zhou. Global large solutions to the three dimensional compressible Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 52(2):1806–1843, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ze:2024:GDP

[ZQ24] Li Ze and Changzheng Qu. Global dynamics to the periodic ferromagnetic spin chain system. SIAM Journal on Mathematical Analysis, 56(2): 1672–1726, March 2024. CO-DEN SJMAAH. ISSN 1095-7154.

Zeng:2022:DPI

[ZRW22] Shengda Zeng, Vicentiu D. Radulescu, and Patrick Winkert. Double phase implicit obstacle problems with convection and multivalued mixed boundary value conditions. SIAM Journal on Mathematical Analysis, 54(2):1898–1926, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1441195.

Zhang:2021:ABB

[ZZ21a] Lei Zhang and Xiao-Qiang Zhao. Asymptotic behavior of the basic reproduction ratio for periodic reaction–diffusion systems. SIAM Journal on Mathematical Analysis, 53(6): 6873–6909, ???? 2021. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhao:2021:LTD

[ZZ21b] Zehua Zhao and Jiqiang Zheng. Long time dynamics for defocusing cubic nonlinear Schrödinger equations on three dimensional product space. SIAM Journal on Mathematical Analysis, 53(3):3644– 3660, ???? 2021. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhai:2023:STC

[ZZ23a] Cuili Zhai and Weiren Zhao.

Stability threshold of the Couette flow for Navier–Stokes Boussinesq system with large Richardson number γ^{2} > 1/4. SIAM Journal on Mathematical Analysis, 55(2):1284–1318, ???? 2023. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1495160.

Zhang:2023:GRR

[ZZ23b] Junhao Zhang and Huijiang Zhao. Global regularity for a radiation hydrodynamics model with viscosity and thermal conductivity. SIAM Journal on Mathematical Analysis, 55(6):6229–6261, November 2023. CODEN SJMAAH. ISSN 1095-7154.

Zhang:2024:QNL

[ZZ24] Ping Zhang and Yibin Zhang.

Quasi-neutral limit of Nernst– Planck–Navier–Stokes system. SIAM Journal on Mathematical Analysis, 56(6):7250–7289, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/24M1635843.

Zeng:2022:LSC

[ZZZ22a] Lan Zeng, Zhifei Zhang, and Ruizhao Zi. Linear stability of the Couette flow in the 3D isentropic compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 54(5):5698–5741, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1438499.

Zhang:2022:SRW

[ZZZ22b] Lan Zhang, Huijiang Zhao, and Qingsong Zhao. Stability of rarefaction waves of the compressible Navier–Stokes– Poisson system with large initial perturbation. SIAM Journal on Mathematical Analysis, 54(4):4536–4571, ???? 2022. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1473534.

Zhang:2024:CND

[ZZZ24] Nangao Zhang, Changjiang Zhu, and Limei Zhu. Convergence to nonlinear diffusion waves for solutions of blood flow model. SIAM Journal on Mathematical Analysis, 56 (5):6768–6797, ???? 2024. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1586367.